### v1

# Robust Learning of Multi-index Models via Iterative Subspace Approximation

Abstract
We study the task of learning Multi-Index Models (MIMs) in the presence of label noise under the Gaussian distribution. A $K$-MIM on $\mathbb{R}^{d}$ is any function $f$ that only depends on a $K$-dimensional subspace, i.e., $f(\mathbf{x})=g(\mathbf{W}\mathbf{x})$ for a link function $g$ on $\mathbb{R}^{K}$ and a $K\times d$ matrix $\mathbf{W}$. We consider a class of well-behaved MIMs with finite ranges that satisfy certain regularity properties. Our main contribution is a general noise-tolerant learning algorithm for this class that is qualitatively optimal in the Statistical Query (SQ) model. At a high-level, our algorithm attempts to iteratively construct better approximations to the defining subspace by computing low-degree moments of our function conditional on its projection to the subspace computed thus far, and adding directions with relatively large empirical moments. For our class of well-behaved MIMs, we show that this procedure efficiently finds a subspace $V$ so that $f(\mathbf{x})$ is close to a function of the projection of $\mathbf{x}$ onto $V$, which can then be found by brute-force. Conversely, for functions for which these conditional moments do not necessarily help in finding better subspaces, we prove an SQ lower bound providing evidence that no efficient algorithm exists.
As concrete applications of our general algorithm, we provide significantly faster noise-tolerant learners for two well-studied concept classes:
[…]
Furthermore, we show that in the presence of random classification noise, the complexity of our algorithm is significantly better, scaling polynomially with $1/\epsilon$.
1 Introduction
In this paper, we develop a new algorithm for learning a broad family of multi-index models (MIMs) under the Gaussian distribution in the presence of label noise. Before we describe our general setting and algorithm, we start with a specific basic model that was the original motivation for this work.
Motivation: Multiclass Linear Classification.
A multiclass linear classifier is a function $f:\mathbb{R}^{d}\to[K]$ of the form $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})\;,$ where $\mathbf{w}^{(i)}\in\mathbb{R}^{d}$ and $t_{i}\in\mathbb{R}$. Where ties are resolved by selecting the smallest index. Multiclass linear classification (MLC) is a textbook machine learning problem [SB14], that has also been extensively studied in the online learning literature; see, e.g., [KSST08] and references therein. In the realizable PAC setting (i.e., with clean labels), this task is efficiently solvable to 0-1 error $\epsilon$ with sample size $O(dK/\epsilon)$ via the multiclass Perceptron algorithm. On the other hand, despite extensive effort, see, e.g., [NDRT13, WLT18, LWS18, SKP+20, GWBL20], our understanding of the computational complexity of learning is poor in the noisy setting—even for random label noise.
The special case of binary classification of Linear Threshold Functions (LTFs) or halfspaces—corresponding to $K=2$—is fairly well-understood. Specifically, sample and computationally efficient learning algorithms are known, even in the presence of noisy labels [ABL17, DKS18, DGT19, DKTZ20a, DKTZ20b, DKK+21b, DKK+22]. Unfortunately, these prior approaches are specific to the binary case. In particular, it is not clear whether they can be generalized to give non-trivial efficient algorithms for the noisy multiclass setting.
Perhaps the simplest model of random label noise is that of Random Classification Noise (RCN) [AL88]. In the binary classification setting, the RCN model posits that the label of each example is independently flipped with probability exactly $\eta$, for some noise parameter $20<\eta<1/$ 2. A slight generalization of the noise model, that is essentially equivalent, allows for different probabilities $\eta_{+},\eta_{-}\in(0,1/2)$ of flipping the label of positive and negative examples. A classical result [Kea98] shows that, in the binary classification setting, any efficient Statistical Query (SQ) algorithm can be transformed into an efficient RCN-tolerant PAC learner (with a polynomial dependence in all relevant parameters). Unfortunately, no such result is known for the multiclass setting.
The natural (and well-studied) generalization of the RCN model to the multiclass setting is the following [PRKM+17, ZLA21, VRW17]. Let $D$ be the distribution on noisy samples $(\mathbf{x},y)\in\mathbb{R}^{d}\times[K]$, where in our setting the marginal distribution $D_{\mathbf{x}}$ is the standard Gaussian, and $f:\mathbb{R}^{d}\to[K]$ be the ground truth classifier. Then there exists a row-stochastic matrix $\mathbf{H}\in\mathbb{R}_{+}^{K\times K}$ such that $\mathbf{H}_{i,j}=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y=i\mid f(\mathbf{x})=j]$. Moreover, it is additionally assumed that there exists $\gamma>0$ such that $\mathbf{H}_{i,i}\geq\mathbf{H}_{i,j}+\gamma$, for all $i\neq j\in[K]$ [^1] Roughly speaking, the parameter $\gamma>0$ plays the role of $1-2\eta$ in the binary case.. Under these assumptions, MLC is information-theoretically solvable using $O(dK/(\gamma\epsilon))$ samples. On the other hand, no $\mathrm{poly}(d,K)$ time learning algorithm is known. Prior algorithmic work using gradient-based methods to achieve sample and time complexity that scales polynomially with $1/\sigma$, where $\sigma$ is the minimum singular value of $\mathbf{H}$ (see [LWS18]). This quantity could be arbitrarily small or even zero, hence such approaches do not in general lead to efficient algorithms. This discussion leads to the following question:
Is there a $\mathrm{poly}(d,K)$ time algorithm for multiclass linear classification
in the presence of random classification noise?
Perhaps surprisingly, we provide strong evidence that the answer to the above question is negative. Specifically, we establish a Statistical Query (SQ) lower bound showing that the complexity of any SQ algorithm for this task is $d^{\Omega(K)}$ (see Theorem 1.7).
Interestingly, our negative result applies for learning algorithms that achieve optimal error guarantees (or, equivalently, are able to approximate the target classifier to any desired accuracy). On the other hand, more efficient algorithms (e.g., with a fixed $\mathrm{poly}(d)$ dependence on the runtime) may be possible if one is allowed to achieve near-optimal error, e.g., up to a constant factor. As an application of our new methodology, we show that this is indeed possible (see Theorem 1.8).
A related question is to understand the algorithmic possibilities and limitations of learning MLCs in the presence of adversarial label noise. In the adversarial label noise model, an adversary is allowed to arbitrarily corrupt an unknown $\mathrm{OPT}$ fraction of the labels, for some parameter $\mathrm{OPT}<1/2$, and the goal of the learner is to compute a hypothesis whose 0-1 error is competitive with $\mathrm{OPT}$ (see Definition 1.3). The information-theoretically optimal error guarantee in this model is $\mathrm{OPT}+\epsilon$. Unfortunately, even for $K=2$, known computational lower bounds rule out the existence of such algorithms with complexity better than $d^{\mathrm{poly}(1/\epsilon)}$ [DKPZ21, DKR23]. To achieve a fixed polynomial runtime in $d$, one needs to relax the error guarantee. In particular, for $K=2$, efficient algorithms are known with complexity $\mathrm{poly}(d/\epsilon)$ and error guarantee $O(\mathrm{OPT})+\epsilon$. A natural question is whether qualitatively similar algorithms exist for the multiclass setting. Specifically, a first goal is to establish a fixed-parameter tractability guarantee:
Is there a constant-factor approximate learner for MLC with complexity $\mathrm{poly}(d)f(\epsilon,K)$?
As a corollary of our approach, we answer this question in the affirmative (see Theorem 1.6).
Broader View: Learning Multi-Index Models
The class of multiclass linear classifiers discussed above is a (very) special case of a much broader class of functions, commonly known as Multi-Index Models (MIMs). A function is called a MIM if it depends only on the projection onto a “low-dimensional” subspace. Here we focus on MIMs whose range $\cal Y$ is a finite set. Formally, we have:
Definition 1.1 (Multi-Index Model).
A function $f:\mathbb{R}^{d}\to\cal Y$ is called a $K$-MIM if there exists a subspace $U\subseteq\mathbb{R}^{d}$ of dimension at most $K$ such that $f$ depends only on the projection onto $U$, i.e., $f(\mathbf{x})=f(\mathrm{proj}_{U}\mathbf{x})$, where $\mathrm{proj}_{U}\mathbf{x}$ is the projection of $\mathbf{x}$ on $U$. Equivalently, $f$ is of the form $f(\mathbf{x})=g(\mathbf{W}\mathbf{x})$ for a link function $g:\mathbb{R}^{K}\to\cal{Y}$ and a $K\times d$ weight matrix $\mathbf{W}$.
The link function $g$ could be either a priori known or unknown to the learner. In the latter case, it is typically assumed to satisfy certain regularity conditions. Moreover, the dimension $K$ of the hidden subspace is usually significantly smaller than the ambient dimension $d$.
Multi-index models [FJS81, Hub85, Li91, HL93, XTLZ02, Xia08] capture the common assumption that real-world data are inherently structured, in the sense that the relationship between the features is of a (hidden) lower-dimensional nature. MIMs can be viewed as a lens for studying a range of ubiquitous learning models, including neural networks, intersections of halfspaces, and multiclass classifiers. Despite extensive prior investigation, there is still a vast gap in our understanding of the complexity of learning MIMs, in particular in the presence of noisy data. This discussion motivates the general direction addressed by this work:
What is the complexity of learning general classes of MIMs in the presence of noisy data?
Our main contribution makes progress towards this research direction. Specifically, we propose a condition that defines a fairly general class of MIMs—that in particular includes multiclass linear classifiers and intersections of halfspaces—and qualitatively characterize its efficient learnability in the presence of noise. That is, we give an efficient noise-tolerant learner for this class of MIMs and a qualitatively matching SQ lower bound. As a corollary, our general algorithm yields significantly faster noise-tolerant learners for both aforementioned concept classes.
1.1 Our Results
1.1.1 Learning Well-behaved Multi-Index Models with Noise
Our general algorithmic result applies to MIMs with finite range that satisfy appropriate regularity properties. Throughout this section, we will use $d$ for the dimension of the feature space and $K$ for the dimension of the hidden low-dimensional subspace. A standard assumption in the literature on MIMs, satisfied by a wide range of natural concept classes, is that the underlying function has bounded Gaussian surface area (GSA)—a complexity measure of the decision boundary (see Definition 3.1 for the definition in the multiclass case). We denote by $\Gamma$ the upper bound on the GSA of the underlying concept class.
The key structural property that we will assume for every MIM function $f$ in our concept class is, roughly speaking, the following: it is not possible to add noise to $f$ to conceal crucial directions so that they remain invisible to moment-based detection. Specifically, for any subspace $V$, one of the following must hold:
  • (a)
    $V$ already captures all essential information of $f$, in which case there is a function depending solely on the projection onto $V$ that approximates $f$ well.
  • (b)
    Otherwise, there is at least one point in $V$ such that when we condition on $\mathbf{x}^{V}$ near that point, the conditional distribution of a “noisy” version of $f$ exhibits a distinguishing moment of degree at most $m$. That is, there exists a low-degree polynomial whose conditional expectation, conditioned on $\mathbf{x}^{V}$, cannot remain hidden; and hence reveals a dependence on directions outside the subspace $V$.
We now present an informal definition of this property and refer the reader to Definition 3.2 for a more precise formulation.
Definition 1.2 (Well-Behaved Multi-Index Model).
Fix $m\in\mathbb{Z}_{+},\zeta,\tau\in(0,1)$ and $\mathcal{Y}\subseteq\mathbb{Z}$ of finite cardinality. We say that a $K$-MIM $f:\mathbb{R}^{d}\to\mathcal{Y}$ with GSA at most $\Gamma$ is is $(m,\zeta,\tau,\Gamma)$-well-behaved if for any random variable $y(\mathbf{x})$ supported on $\mathcal{Y}$ with $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq y(\mathbf{x})]\leq\zeta$ and any subspace $V\subseteq\mathbb{R}^{d}$, one of the following conditions holds:
  1. a)
    there exists a function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$, or
  2. b)
    with non-negligible probability [^2] By this we mean that the probability is not-too-small, specifically it is independent of the dimension $d$. over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, there exists a mean-zero, variance-one polynomial $p:U\to\mathbb{R}$ of degree at most $m$ and $z\in\mathcal{Y}$ such that
    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[p(\mathbf{x}^{U})\mathds{1}(y(\mathbf{x})=z)\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\geq\mathrm{poly}(\zeta/m)\;,$
    where $U=(V+W)\cap V^{\perp}$ and $W$ is the hidden subspace defining $f$.
We denote by $\mathcal{F}$ the class of $(m,\zeta,\tau,\Gamma)$-well-behaved $K$-MIMs over $\mathbb{R}^{d}$.
The aforementioned property enforces that no direction crucial to $f$ can be fully hidden when we look at a suitable low-degree moment. Intuitively, the parameter $\zeta$ in Definition 1.2 will determine the noise-tolerance of our learning algorithm, while the parameter $\tau$ will determine the final error guarantee of the algorithm. The parameter $m$ is the minimum degree of a non-zero moment on the subspace $U$. It is worth pointing out that the sample and computational complexity of our algorithm will scale with $d^{m}$.
As we will show in the next subsection, fundamental and well- studied concept classes—specifically, multiclass linear classifiers and intersections of halfspaces—satisfy Definition 1.2 for reasonable values of the underlying parameters (in particular, with $m\leq 2$).
Before we state our main algorithmic result, we recall the definition of PAC learning with adversarial label noise (aka agnostic learning) under the Gaussian distribution.
Definition 1.3 (Agnostic Multiclass Classification).
Let $\mathcal{Y}$ be finite and $\cal{C}$ a class of functions from $\mathbb{R}^{d}$ to $\mathcal{Y}$. Given $\epsilon>0$ and i.i.d. samples $(\mathbf{x},y)$ from a distribution $D$ on $\mathbb{R}^{d}\times\mathcal{Y}$ such that the marginal $D_{\mathbf{x}}$ is the standard normal and no assumptions are made on the labels, the goal is to output a hypothesis $h:\mathbb{R}^{d}\to\mathcal{Y}$ that, with high probability, has small 0-1 error $\mathrm{err}_{0-1}^{D}(h)\coloneqq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]$, as compared to the optimal 0-1 error of the class ${\cal C}$, i.e., $\mathrm{OPT}\coloneqq\inf_{c\in\mathcal{C}}\mathrm{err}_{0-1}^{D}(c)$.
Our main result is a general algorithm for learning well-behaved MIMs in the presence of adversarial label noise (see Theorem 3.3).
Theorem 1.4 (Learning Well-Behaved MIMs with Adversarial Label Noise).
[…]
At a high-level, the main advantage of the above algorithm is that—as long as $m=O(1)$—its complexity is bounded by a constant-degree polynomial in $d$. Such a bound can be viewed as a significant qualitative improvement over $d^{K}$, which may be attainable using simpler methods. In other words, we give a fixed-parameter tractability result for the task of learning well-behaved MIMs. While the complexity guarantee we obtain is new even in the realizable setting, our algorithm is additionally robust to adversarial label noise—sometimes achieving near-optimal noise-tolerance (within constant factors). In addition to its own merits, the significance of our general algorithm is illustrated by its downstream applications (given in the following subsection).
An interesting feature of Theorem 1.4 is that the complexity upper bound it establishes is qualitatively tight in the Statistical Query (SQ) model, as a function of the dimension $d$. Roughly speaking, if a function class does not satisfy the assumptions of Definition 1.2 for a specific value of $m$, then any SQ algorithm that achieves error competitive with Theorem 1.4 must have SQ complexity $d^{\Omega(m)}$. In more detail, we show (see Theorem 5.1 for a more detailed formal statement).
Theorem 1.5 (SQ Lower Bound for Learning Not-Well-behaved MIMs).
Suppose that a function class $\mathcal{G}$ is rotationally invariant and contains a bounded surface area MIM not contained in the function class $\mathcal{F}$ of Definition 1.2 for parameters $m,\tau$. Then, any SQ algorithm that learns $\mathcal{G}$ within error smaller than $O(\tau)$ has complexity at least $d^{\Omega(m)}$.
Finally, we note that a variant of our general algorithm learns well-behaved MIMs in the presence of RCN with significantly improved complexity. Specifically, for the RCN setting, our algorithm has complexity scaling polynomially in $1/\epsilon$. See Theorem 4.9 for the formal statement. By an application of the latter result, in the following subsection, we obtain the fastest known algorithm for learning multiclass linear classifiers in the presence of RCN.
1.1.2 Concrete Learning Applications to Well-studied Concept Classes
In this subsection, we present algorithmic applications of our general MIM learner to two fundamental and well-studied concept classes.
Multiclass Linear Classifiers.
We denote by $\mathcal{L}_{d,K}$ the family of $K$-class linear classifiers in $\mathbb{R}^{d}$. For the adversarial noise setting, we obtain the following result (see Theorem 2.1):
Theorem 1.6 (Learning $\mathcal{L}_{d,K}$ with Adversarial Label Noise).
[…]
Theorem 1.6 is obtained from Theorem 1.4 via a new structural result that we establish, roughly showing that the class $\mathcal{L}_{d,K}$ is $(1,\zeta,O(\zeta),O(K))$-well behaved for any $\zeta\in(0,1)$ (see Proposition 2.4).
Theorem 1.6 gives the first constant-factor approximate learner for multiclass linear classification with adversarially corrupted labels, whose complexity is a fixed-degree polynomial in $d$. We provide some remarks to interpret the statement of Theorem 1.6. First, even for $K=2$, known hardness results imply that achieving error $\mathrm{OPT}+\epsilon$ requires $d^{\mathrm{poly}(1/\epsilon)}$ time [DKPZ21, DKR23]. Hence, to achieve a fixed-degree $\mathrm{poly}(d)$ complexity guarantee, it is necessary to relax the final accuracy. The best thing to hope for here is a constant-factor approximation, i.e., a final error of $O(\mathrm{OPT})+\epsilon$ (like the one we achieve). For the $K=2$ case, prior work achieves such an error bound [DKS18] [^3] Earlier work [ABL17] achieves such a constant-factor approximation for the special case of homogeneous halfspaces.; and in fact with complexity $\mathrm{poly}(d/\epsilon)$. On the other hand, no non-trivial approximation algorithm was previously known for the multiclass case. Importantly, no polynomial-time learner with error $o(d)\,\mathrm{OPT}+\epsilon$ is known even for $K=3.$ The crucial difference is that the approximation ratio of our algorithm is an absolute constant, independent of the dimension and the number of classes.
We also investigate the complexity of learning multiclass linear classifiers in the presence of RCN. We start by establishing an SQ lower bound of $d^{\Omega(K)}$ for achieving $\mathrm{OPT}+\epsilon$ error. Specifically, we show (see Theorem 5.3):
Theorem 1.7 (SQ Lower Bound for Learning $\mathcal{L}_{d,K}$ with RCN).
For any $\epsilon\leq 1/K^{C}$, where $C$ a sufficiently large universal constant, any SQ algorithm that learns $\mathcal{L}_{d,K}$ under the Gaussian distribution in the presence of RCN to error $\mathrm{OPT}+\epsilon$, has complexity at least $d^{\Omega(K)}$.
Interestingly, using a variant of our general MIM algorithm, we obtain an RCN-tolerant learner with $O(\mathrm{OPT})+\epsilon$ error, whose complexity is qualitatively better than our agnostic learner. Specifically, we show the following (see Theorem 4.9 for the detailed statement):
Theorem 1.8 (Learning $\mathcal{L}_{d,K}$ with RCN).
Assume that the labels follow a distribution obtained by introducing RCN to a function within the class $\mathcal{L}_{d,K}$. There is an algorithm that draws $N=O(d)\,(1/\epsilon)^{\mathrm{poly}(K)}$ i.i.d. labeled samples, runs in time $\mathrm{poly}(d,N)$, and returns a hypothesis $h$ such that with high probability $\mathrm{err}_{0-1}^{D}(h)\leq O(\mathrm{OPT})+\epsilon$.
Intersections of Halfspaces.
An intersection of $K$ halfspaces is any function $f:\mathbb{R}^{d}\to\{\pm 1\}$ for which there exist $K$ halfspaces $h_{i}(\mathbf{x})=\mathrm{sign}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$ such that $f(\mathbf{x})=1$ if and only if $h_{i}(\mathbf{x})=1$ for all $i\in[K]$. We denote by $\mathcal{F}_{K,d}$ the class of intersections of $K$ halfspaces on $\mathbb{R}^{d}$.
The family of intersections of halfspaces is a fundamental concept class that has been extensively studied in many contexts for at least the past five decades. We show that our general MIM algorithm can be applied to $\mathcal{F}_{K,d}$, leading to a novel accuracy-complexity tradeoff in the agnostic setting. Specifically, we obtain the following result (see Theorem 4.1 for a more detailed statement):
Theorem 1.9 (Learning $\mathcal{F}_{K,d}$ with Adversarial Label Noise).
[…]
Theorem 1.9 is obtained from Theorem 1.4 via a new structural result, roughly showing that the class $\mathcal{F}_{d,K}$ is $(2,\zeta,K\tilde{O}(\zeta),O(\sqrt{\log(K)}))$-well behaved for any $\zeta\in(0,1)$ (see Proposition 4.2).
Prior work [KKMS05, KOS08] demonstrated that the class $\mathcal{F}_{d,K}$ is agnostically learnable under the Gaussian distribution, within optimal error of $\mathrm{OPT}+\epsilon$, using $d^{O({\log K}/{\epsilon^{4}})}$ samples and runtime. This complexity bound was recently shown to be near-optimal in the SQ model [DKPZ21, HSSV22]. A line of subsequent work has focused on understanding the tradeoff between accuracy and computational complexity, and in particular on designing faster algorithms with relaxed error guarantees. Specifically, [DKS18] developed a $\mathrm{poly}(d)(1/\epsilon)^{\mathrm{poly}(K)}$ time algorithm with error of $\mathrm{poly}(K)\tilde{O}(\mathrm{OPT}^{1/11})$. Our algorithm achieves a different tradeoff: its error guarantee is significantly better compared to [DKS18]—namely nearly-optimal as a function of $\mathrm{OPT}$—while its computational complexity incurs a $2^{\mathrm{poly}(K/\epsilon)}$ dependence (as opposed to $(1/\epsilon)^{\mathrm{poly}(K)})$).
1.2 Overview of Techniques
Prior Techniques and Obstacles for Learning Noisy Multi-Index Models.
Recall that our goal is, given access to noisy samples from well-behaved MIM with finite label set, to learn the underlying function $f$ to high accuracy. We will use $W$ to denote the hidden $K$-dimensional subspace to whose projection $f$ depends. The standard approach for solving such learning problems (see, e.g., [Vem10b]) is to first learn some suitable approximation $V$ to $W$, and then use exhaustive search to learn a function of the projection onto $V$. Computing such an approximation to $W$ is usually achieved by estimating the low-order moments of the level sets of $f$. In particular, for each output value $i$ and each low-degree polynomial $p$, we want to approximate the expectation of $p(\mathbf{x})\mathds{1}(f(\mathbf{x})=i)$. In the absence of label noise, these moments should depend solely on how $p$ interacts with the directions aligned with $W$. Using this property, one can extract an approximation of a subspace of $W$, by choosing $V$ to be the space spanned by all directions with reasonably large moments. This approach yields an algorithm with runtime approximately $d^{O(m)}$, where $m$ is the number of moments being computed, plus some (often exponential) function of $\dim(V)$—which is typically regarded as “constant”, since $d$ is much larger than $\dim(V)$.
Unfortunately, there may exist directions within $W$ where all low-degree moments vanish. An innocuous such example occurs when $f$ depends only on the projection of $\mathbf{x}$ onto a subspace $W^{\prime}\subseteq W$. In this case, all moments would vanish for directions in $W\cap W^{\prime\perp}$. However, if such moments vanish for different reasons —for example, due to label noise—the aforementioned approach could be problematic. Therefore, in order to analyze this class of algorithms, one needs to prove that every direction $\mathbf{v}\in W$ for which the moments of $f$ vanish is a direction where $f$ approximately does not depend on—i.e., $f$ is close to a function of the projection onto $\mathbf{v}^{\perp}$. Unfortunately, showing that the moments of $f$ are non-vanishing for all relevant directions can be challenging, which limits the applicability of this approach.
New Dimensionality Reduction Technique.
The main algorithmic contribution of this work is the design of a new, more general, approach that leads to significant improvements on the aforementioned dimensionality reduction technique. At a high-level, we design an iterative dimensionality reduction procedure (Algorithms 3 and 4) working as follows. At each step/iteration, after learning an approximating subspace $V$, we partition $V$ into small rectangular regions $S$, each of probability mass at least $\alpha$, where $\alpha$ is a carefully selected parameter. We then compute the first $m$ moments of $y(\mathbf{x})$, where $y(\mathbf{x})$ is the noisy version of $f(\mathbf{x})$, conditioned on the projection of $\mathbf{x}$ to $V$ lying in each region $S$; this step requires roughly $d^{m}(1/\alpha)$ samples and runtime (see Algorithm 4). In the noiseless case (i.e., with clean labels), non-vanishing moments should still align with $W$, providing us with a better approximation. We then find all directions with non-vanishing moments across all regions $S$, add them to $V$, and repeat the process. We terminate this procedure once $y$ is sufficiently close to a function $g$ of the projection of $\mathbf{x}$ onto $V$. Finally, using exhaustive search methods, we can recover an $\epsilon$-approximation to $g$, using roughly $(1/\epsilon)^{\dim(V)}$ samples and runtime.
The main advantage of our new method is that it succeeds under much weaker conditions, as compared to previous approaches. As mentioned earlier, prior techniques require that every direction in $W$ to be either unimportant or have a non-trivial low-degree moment. In contrast, our approach only requires that for each subspace $V$, either $f$ is approximately a function of the projection onto $V$, or that there exists some region $S$ with a non-trivial low-degree moment in a direction in $V^{\perp}$ correlated with $W$ (see Definition 3.2). If we can establish this structural property, it follows that unless $f$ is already close to a function of the projection onto $V$ (in which case we can apply a brute-force method), one iteration of the algorithm will find a direction in $V^{\perp}$ that is at least slightly correlated with $W$. As a result, in the next iteration, the distance between $V$ and $W$ decreases; hence, our algorithm will eventually terminate.
One disadvantage of our method as described above is that, while the runtime is typically polynomial in the dimension $d$, the dependence on other parameters can be problematic. Specifically, the number of regions $S$ is usually exponential in $\dim(V)$; hence, at each iteration, $\dim(V)$ can increase in the worst-case by an exponential amount. This can lead to power-tower dependencies on some parameters. We circumvent this issue by assuming that our function has a small fraction of regions $S$ with non-vanishing moments. With this assumption, we can only add directions with large moments in a reasonable fraction of regions using a filtering step (see Line 6), keeping the increase in $\dim(V)$ polynomial. This bound can be further improved if the noisy version of $f(\mathbf{x})$ that we observe is actually a function of the projection onto $W$, which holds in both the realizable and the RCN settings (see Section 4.2). In this case, non-vanishing moments will exactly align with directions along $W$, up to a measurement error, ensuring that $\dim(V)$ never exceeds $\dim(W)$ (see Theorem 4.9).
SQ Lower Bounds.
An appealing feature of our algorithmic technique is that it is provably qualitatively tight in the Statistical Query (SQ) model (see Section 5). In particular, suppose that a function $f$ has a subspace $V$ so that on the preimage of the projection onto any point $\mathbf{x}$ in $V$, the low-degree moments of $f$ are nearly vanishing. Then we prove an SQ lower bound showing that it is SQ-hard to distinguish between (i) the distribution $(\mathbf{x},h(\mathbf{x}))$, where $h$ is a random rotation of the function $f$ about $V$, and (ii) a product distribution $(\mathbf{x},y)$, where $\mathbf{x}$ and $y$ are conditionally independent on $\mathbf{x}^{V}$ (see Theorem 5.1). This means that unless $f$ is close to some function of the projection onto $V$, it is hard to learn $f$ in the SQ model. In more detail, unless $f$ has the property that for each $V$, either $f$ is $\tau$-close to a function of the projection onto $V$, or there is some preimage with a non-trivial moment of degree at most $m$, then any SQ algorithm that learns $f$ to accuracy better than roughly $\tau$ either uses exponentially many queries or at least one query of accuracy smaller than $d^{-\Omega(m)}$. In summary, our algorithm learns $f$ if and only if there is an SQ algorithm that does so with comparable parameters as a function of $d$.
Finally, we give an overview of our SQ lower bound for learning multiclass linear classifiers with RCN (see Theorem 5.3 in Section 5.2). This result leverages the fact that statistical queries correspond to evaluations of the confusion matrix (see Definition 4.8). If this confusion matrix is close to being non-invertible, then there is missing information between the noisy and noiseless queries. To prove this, we find a low-dimensional concept, $f$, and a corresponding confusion matrix, $\mathbf{H}$, for which the first $\Omega(K)$ moments of the noisy label sets match the moments of the standard Gaussian distribution. As a result, we establish SQ-hardness in distinguishing between the distribution of random rotations of $f$ in the $d$-dimensional space and a product distribution.
Applications to Well-studied Concept Classes.
We show that our iterative dimension-reduction technique yields significantly improved learners for multiclass linear classifiers and intersections of halfspaces in the presence of noise. In the regime of agnostically learning multiclass classifiers with $K$ classes (see Section 2), we show that our algorithm achieves an error of $O(\mathrm{OPT})+\epsilon$, using roughly $d2^{\mathrm{poly}(K/\epsilon)}$ samples and runtime (see Theorem 2.1). We show that if the noisy labels $y$ are at least $\Theta(\mathrm{OPT})+\epsilon$ far from any function depending on the projection onto a subspace $V$ (otherwise, an exhaustive search in $V$ would yield a function $g$ with the desired error bound), then at least an $\epsilon$ fraction of our partition $S$ of $V$, will have the optimal $f$ more than $\Theta(\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y\mid S])$ far from any constant function. In this case, we establish a new structural result (see Proposition 2.8) showing that the first moments of the sub-level sets of $y$ align with $W$ on a non-trivial fraction of the partition sets $S$. By using this structural result, we conclude that the class of multiclass linear classifiers is $(1,\zeta,O(\zeta),O(K))$-well behaved, as in Definition 1.2, for any $\zeta\in(0,1)$. Therefore, by applying our general algorithm we can obtain a learner with the aforementioned efficiency (Theorem 1.4)
Furthermore, we obtain an algorithm with improved time-accuracy tradeoff for agnostically learning intersections of $K$ halfspaces under the Gaussian distribution (see Section 4). Specifically, we show that an error of $K\widetilde{O}(\mathrm{OPT})$ can be achieved with complexity $d^{2}2^{\mathrm{poly}(K/\epsilon)}$ (see Theorem 4.1). An analogous structural result to the one we showed for multiclass linear classifiers can be established to prove that unless the optimal classifier is $K\widetilde{O}(\mathrm{OPT})$-close to a constant function, there exists a low-order moment (of degree at most two) that correlates non-trivially with the unknown subspace $W$ (see Proposition 4.4). Using this result, we deduce that the class of intersections of $K$ halfspaces is $(2,\zeta,K\tilde{O}(\zeta),O(\sqrt{\log(K)}))$-well behaved, as in Definition 1.2, for any $\zeta\in(0,1)$. Hence, our learner again follows by applying our general iterative dimension-reduction technique.
1.3 Prior Work
The task of learning MIMs is fundamental in machine learning, with a long line of research in both statistics and computer science [FJS81, Hub85, Li91, HL93, XTLZ02, Xia08]. The majority of the literature focuses on learning various structured families of MIMs under natural distributional assumptions with clean labels [GLM18, JSA15, DH18, BJW19, GKLW19, DKKZ20, DLS22, BBSS22, CDG+23]. A recent line of work has shown that achieving optimal error in the adversarial label setting under the Gaussian distribution cannot be achieved in fixed-degree polynomial time, even for very basic MIM classes [DKPZ21, DKMR22, Tie23]. These hardness results motivate the design of faster algorithms with relaxed error guarantees, which is the focus of this paper. Despite this being a natural direction, no general algorithmic framework was previously known in this setting.
The recent work [DKK+24] developed agnostic learners for MIMs under the Gaussian distribution with fixed-parameter tractability guarantees similar to ours. The major difference is that, while our framework relies on learning from random examples, the latter work [DKK+24] assumes query access to the target function—a much stronger access model.
The task of efficiently learning intersections of halfspaces under Gaussian marginals is one of the most well-studied problems in the computational learning theory literature. A long line of work focused on achieving strong provable guarantees for this problem in the realizable setting [Vem10b, AV99, KLT09, KOS04]. However, the aforementioned work either fails in the presence of noise or scales superpolynomially with respect to the dimension (i.e., does not achieve fixed-degree polynomial runtime). An exception is the recent work [DKS18] that sacrifices the optimal error guarantee for a polynomial dependence in the input dimension.
Binary linear classification in the presence of noise under well-behaved distributions is a fairly well-understood problem, with a substantial body of work; see, e.g., [KKMS05, ABL17, DKS18, DKTZ20a, DKTZ20b, DKK+21b, DKK+22] and references therein. However, in the multiclass setting, most known results focus on online or bandit classification [KSST08, BPS+19]. That is, the subject of designing efficient algorithms in the multiclass setting that work in the presence of noise is rather poorly understood. It is worth pointing out that the multiclass regime offers several new interesting models of noisy information that are challenging, even if the noise is random [PRKM+17, VRW17, FKKT21, CST11].
1.4 Notation and Preliminaries
Basic Notation For $n\in\mathbb{Z}_{+}$, let $[n]:=\{1,\ldots,n\}$. We use small boldface characters for vectors and capital bold characters for matrices. For $\mathbf{x}\in\mathbb{R}^{d}$ and $i\in[d]$, $\mathbf{x}_{i}$ denotes the $i$-th coordinate of $\mathbf{x}$, and $\|\mathbf{x}\|_{2}:=(\mathop{\textstyle\sum}_{i=1}^{d}\mathbf{x}_{i}^{2})^{1/2}$ denotes the $\ell_{2}$-norm of $\mathbf{x}$. Throughout this text, we will often omit the subscript and simply write $\|\mathbf{x}\|$ for the $\ell_{2}$-norm of $\mathbf{x}$. We will use $\mathbf{x}\cdot\mathbf{y}$ for the inner product of $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$ and $\theta(\mathbf{x},\mathbf{y})$ for the angle between $\mathbf{x},\mathbf{y}$. For vectors $\mathbf{x},\mathbf{v}\in\mathbb{R}^{d}$ and a subspace $V\subseteq\mathbb{R}^{d}$ denote by $\mathbf{x}^{V}$ the projection of $\mathbf{x}$ onto $V$, and by $\mathbf{x}^{\mathbf{v}}$ the projection of $\mathbf{x}$ onto the line spanned by $\mathbf{v}$. For two subspaces $V,W\subseteq\mathbb{R}^{d}$, we denote by $W^{V}=\{\mathbf{w}^{V}:\mathbf{w}\in W\}$, which it itself a subspace. Furthermore, for a set of vectors $L\subseteq\mathbb{R}^{d}$, we denote by $\mathrm{span}(L)$ the subspace of $\mathbb{R}^{d}$ defined by their span. We slightly abuse notation and denote by $\mathbf{e}_{i}$ the $i$-th standard basis vector in $\mathbb{R}^{d}$. For a matrix $A\in\mathbb{R}^{m\times n}$, the Frobenius norm of $A$ is denoted by $\left\|A\right\|_{F}=(\sum_{i=1}^{m}\sum_{j=1}^{n}A_{i,j}^{2})^{1/2}$. For a set $A\subseteq\mathbb{R}^{d}$ and a point $\mathbf{x}\in\mathbb{R}^{d}$, we define the distance from $\mathbf{x}$ to $A$ as $\mathrm{dist}(\mathbf{x},A)=\min_{\mathbf{y}\in A}\left\|\mathbf{x}-\mathbf{y}\right\|$.
We use the standard asymptotic notation, where $\widetilde{O}(\cdot)$ is used to omit polylogarithmic factors. Furthermore, we use $a\lesssim b$ to denote that there exists an absolute universal constant $C>0$ (independent of the variables or parameters on which $a$ and $b$ depend) such that $a\leq Cb$, $\gtrsim$ is defined similarly. We use the notation $g(t)\leq\mathrm{poly}(t)$ for a quantity $t\geq 1$ to indicate that there exists constants $c,C>0$ such that $g(t)\leq Ct^{c}$. Similarly we use $g(t)\geq\mathrm{poly}(t)$ for a quantity $t<1$ to denote that there exists constants $c,C>0$ such that $g(t)\geq Ct^{c}$.
Probability Notation We use $\operatorname*{\mathbf{E}}_{x\sim D}[x]$ for the expectation of the random variable $x$ according to the distribution $D$ and $\operatorname*{\mathbf{Pr}}[\mathcal{E}]$ for the probability of event $\mathcal{E}$. For simplicity of notation, we may omit the distribution when it is clear from the context. For $(\mathbf{x},y)$ distributed according to $D$, we denote by $D_{\mathbf{x}}$ to be the distribution of $\mathbf{x}$ and by $D_{y}$ to be the distribution of $y$. Let $\mathcal{N}(\boldsymbol{\mu},\mathbf{\Sigma})$ denote the $d$-dimensional Gaussian distribution with mean $\boldsymbol{\mu}\in\mathbb{R}^{d}$ and covariance $\mathbf{\Sigma}\in\mathbb{R}^{d\times d}$. We denote by $\phi_{d}(\cdot)$ the pdf of the $d$-dimensional standard normal and we use $\phi(\cdot)$ for the pdf of the $1$-dimensional standard normal.
For additional preliminaries, we refer the reader to Appendix A.
1.5 Organization
The paper is structured as follows: In Section 2, we present a special case of our subspace approximation procedure for learning multiclass linear classifiers under agnostic noise, establishing Theorem 1.6. In Section 3, we generalize this approach, stating the necessary conditions (see Definition 1.2) for our algorithm to work for general classes of MIMs, establishing Theorem 1.4. In Section 4, we consider applications of our algorithmic approach. First in Section 4.1, we apply our general theorem to the class of intersections of halfspaces, establishing Theorem 1.9. Moreover, in section Section 4.2, we consider how our algorithm performs under random noise, particularly in the Random Classification Noise (RCN) model, and demonstrate that we can get even better runtime guarantees (Theorem 1.8). Finally, Section 5 provides SQ lower bounds. First, we demonstrate that the complexity of our algorithm is qualitatively optimal with respect to the dimension, establishing Theorem 1.5. Second, we establish an SQ lower bound for learning multiclass linear classifiers under random classification noise, showing that achieving optimal error with a polynomial number of samples is impossible in the SQ model, Theorem 1.7.
2 Learning Multiclass Linear Classifiers with Adversarial Label Noise
In this section, we establish Theorem 1.6. Note that we introduce this special case as a preliminary step to build intuition for our general algorithm in Section 3. First, we state the formal version of the theorem and provide the algorithm pseudocode along with the necessary definitions.
Theorem 2.1 (Learning Multiclass Linear Classifiers with Agnostic Noise).
Let $\mathcal{L}_{d,K}$ be the class of $K$-class linear classifiers in $\mathbb{R}^{d}$. Let $D$ be a distribution over $\mathbb{R}^{d}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, let $\mathrm{OPT}=\inf_{f\in\mathcal{L}_{d,K}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$ and let $\epsilon,\delta\in(0,1)$. There exists a sufficiently large universal constant $C>0$ such that Algorithm 1 draws $N=d2^{\mathrm{poly}(1/\epsilon,K)}\log(1/\delta)$ i.i.d. samples from $D$, runs in $\mathrm{poly}(d,N)$ time, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq{C}\cdot\mathrm{OPT}+\epsilon$.
Before we present the algorithm, we provide the following definitions, which characterize how we discretize a low-dimensional subspace and approximate distributions over it.
To effectively approximate the distribution over a specific subspace $V$, we aim to preserve a substantial portion of its mass and partition it into fine regions of non-negligible mass. By selecting an orthonormal basis for $V$ and excluding regions where any coordinate exceeds $\sqrt{\log(k/\epsilon)}$, we ensure that at least a $1-\epsilon$ fraction of the mass remains. The remaining space is then partitioned into cubes aligned with $V$, each with edges of length $\epsilon$ and containing approximately $\epsilon^{\dim(V)}$ mass.
Definition 2.2 ($\epsilon$-Approximating Partition).
Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$ with an orthonormal basis $\mathbf{v}^{(1)},\dots,\mathbf{v}^{(k)}$, and let $\epsilon\in(0,1)$. We call an $\epsilon$-approximating partition with respect to $V$ a family of sets $\mathcal{S}$ of the form $S_{\mathbf{j}}=\{\mathbf{x}\in\mathbb{R}^{d}:z_{\mathbf{j}_{i}-1}\leq\mathbf{v}^{(i)}\cdot\mathbf{x}\leq z_{\mathbf{j}_{i}},i\in[k]\}$, where $\mathbf{j}\in[M_{\epsilon}]^{k}$, $z_{i}=-\sqrt{2\log(k/\epsilon)}+i\epsilon+t$, for $i\in[M_{\epsilon}]$, $t\in(0,\epsilon/2)$ and $M_{\epsilon}=\left\lceil(2\sqrt{2\log(k/\epsilon)})/\epsilon\right\rceil$. Furthermore, for each set $S_{\mathbf{j}}\in\mathcal{S}$, we associate a set of linear equation $\mathcal{G}_{S_{\mathbf{j}}}=\{\mathbf{v}^{(1)}\cdot\mathbf{x}=t_{1},\dots,\mathbf{v}^{(k)}\cdot\mathbf{x}=t_{k}\}$, where $z_{\mathbf{j}_{i}-1}\leq t_{i}\leq z_{\mathbf{j}_{i}}$ for $i\in[k]$.
Furthermore, we define the piecewise constant approximation of a distribution as the function that minimizes the 0-1 loss with respect to the labels in each subset of a partition. Our algorithm will first estimate the hidden subspace of the multiclass linear classifier and then return a piecewise constant approximation over it.
Definition 2.3 (Piecewise Constant Approximation).
Let $D$ be a distribution over $\mathbb{R}^{d}\times[K]$, let $V$ be a subspace of $\mathbb{R}^{d}$, let $\epsilon\in(0,1)$, and let $\mathcal{S}$ be an $\epsilon$-approximating partition with respect to $V$. We call a piecewise constant approximation of the distribution $D$, with respect to a partition $\mathcal{S}$, any classifier $h_{{\mathcal{S}}}:\mathbb{R^{d}}\to[K]$ of the form $h_{{\mathcal{S}}}(\mathbf{x})=\operatorname*{argmin}_{i\in[K]}(\operatorname*{\mathbf{Pr}}_{{(\mathbf{x},y)\sim D}}[i\neq y\mid\mathbf{x}\in S])$, if $\mathbf{x}\in S$ for all $S\in\mathcal{S}$, where $\mathcal{S}$ is an $\epsilon$-approximating partition with respect to $V$ according to Definition 2.2.
We now present our subspace approximation method for learning multiclass linear classifiers under agnostic noise, Algorithm 1.
Input: Accuracy $\epsilon>0$, failure probability $\delta>0$ sample access to a distribution $D$ over $\mathbb{R}^{d}\times[K]$.
Output: A hypothesis $h$ such that, with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[h(\mathbf{x})\neq y]\leq C\cdot\mathrm{OPT}{+}\epsilon$, for a universal constant $C$.
  1. 1.
    Let $c$ be a sufficiently large constant, and let $T{,p}$ be sufficiently large, constant-degree polynomials of $1/\epsilon$ and $K$.
  2. 2.
    Let $N\leftarrow d2^{T^{c}}\log(1/\delta)$, $\epsilon^{\prime}\leftarrow 1/p$, $L_{1}\leftarrow\emptyset$ and $t\leftarrow 1$.
  3. 3.
    While $t\leq T$
    1. (a)
      Draw a set $S$ of $N$ i.i.d. samples from $D$.
    2. (b)
      $\mathcal{E}_{t}\leftarrow$ Algorithm 2($\epsilon,\epsilon^{\prime},\delta,\mathrm{span}(L_{t}),S$).
    3. (c)
      $L_{t+1}\leftarrow L_{t}\cup\mathcal{E}_{t}$.
    4. (d)
      $t\leftarrow t+1$.
  4. 4.
    Construct $\mathcal{S}$ an $\epsilon^{\prime}$-approximating partition with respect to $\mathrm{span}(L_{t})$, as defined in Definition 2.2.
  5. 5.
    Return a piecewise constant approximating classifier $h_{\mathcal{S}}$, as described in Definition 2.3, using $N$ i.i.d. samples from $D$, by finding the most frequent label at any set in $\mathcal{S}$.
Algorithm 1 Learning Multiclass Linear Classifiers under Agnostic Noise
Input: $\epsilon>0$, $\epsilon^{\prime}>0$, $\delta>0$, a subspace $V$ of $\mathbb{R}^{d}$, and a set of samples $\{(\mathbf{x}^{(1)},y_{1}),\dots,(\mathbf{x}^{(N)},y_{N})\}$ from a distribution $D$ over $\mathbb{R}^{d}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$.
Output: A set of unit vectors $\mathcal{E}$.
  1. 1.
    Let $\sigma$ be a sufficiently small, constant-degree polynomial in $\epsilon,1/K$.
  2. 2.
    Construct the empirical distribution $\widehat{D}$ of $\{(\mathbf{x}^{(1)},y_{1}),\dots,(\mathbf{x}^{(N)},y_{N})\}$.
  3. 3.
    Construct an $\epsilon^{\prime}$-approximating partition, $\mathcal{S}$, with respect to $V$ (see Definition 2.2).
  4. 4.
    Compute: $\widehat{\mathbf{u}}^{(S,i)}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\widehat{D}}[\mathbf{x}\mathds{1}(y=i)\mid\mathbf{x}\in S]$ for $S\in\mathcal{S},i\in[K]$.
  5. 5.
    Let $\widehat{\mathbf{U}}=\sum_{(S,i)\in\mathcal{S}\times[K]}(\widehat{\mathbf{u}}^{(S,i)})^{\perp V}((\widehat{\mathbf{u}}^{(S,i)})^{\perp V})^{\top}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim{\widehat{D}}}[S]$.
  6. 6.
    Return the set $\mathcal{E}$ of unit eigenvectors of $\widehat{\mathbf{U}}$ corresponding to eigenvalues at least $\sigma$.
Algorithm 2 Estimating a relevant direction
Let $f^{*}$ be a multiclass linear classifier with error $\mathrm{OPT}$, i.e., $\mathbf{Pr}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$. Algorithm 1 operates as follows: In each iteration $t$, the algorithm updates the current list of unit vectors $L_{t}$ so that its span, denoted by $V_{t}$, provides a better approximation to the span of the weight vectors of $f^{*}$, denoted by $W^{*}$. To this end, Algorithm 1 uses Algorithm 2 to construct a new list $\mathcal{E}_{t}$ that contains vectors with non-trivial correlation to $(W^{*})^{\perp V_{t}}$. Our main contribution is the design and analysis of Algorithm 2, which efficiently returns a non-trivially correlating vector.
First in Section 2.1, we show that if an optimal multiclass classifier is not $C\cdot\mathrm{OPT}$-close to a constant classifier, then at least one of the first moments of the labels regions, i.e., $\operatorname*{\mathbf{E}}[\mathbf{x}\mathds{1}(y=i)]$ for some $i\in[K]$, will correlate non-trivially (by at least $\mathrm{poly}(\epsilon)$) with a vector in $W^{*}$. This, in fact, suffices to initialize a list of vectors that correlate with $W^{*}$. In order to expand this list and at the same time improve the approximation of $W^{*}$, we need to show that, we can find vectors that not only correlate with $W^{*}$ but are also orthogonal to the current list of vectors $L_{t}$.
To this end, in Section 2.2, we use the current list $L_{t}$ of weight vectors to partition $\mathbb{R}^{d}$ into sufficiently small regions $S$. In each such region $S$, the classifier $f^{*}$ has a limited dependence on the vectors in $L_{t}$. Specifically, we show that $f^{*}$ is $\epsilon$-close to a function $f_{S}^{*}$ that does not depend on weight vectors contained in $L_{t}$. Thus, within each region $S$, we obtain a multiclass classifier that depends only on vectors in $(W^{*})^{\perp V_{t}}$. Therefore, from the above discussion, in each region $S$, either the classifier can be accurately approximated by a constant classifier, or the first moments will correlate with a vector in $(W^{*})^{\perp V_{t}}$, enabling further improvement.
Furthermore, for $f^{*}$ to depend mostly on directions in $(W^{*})^{\perp V_{t}}$, the cubes $S$ in the partition must have width on the order of $\mathrm{poly}(\epsilon/K)$. Such a fine partitioning would typically yield a number of cubes that is exponential in $\dim(V_{t})$ and, hence could lead to a power-tower update in the dimension of $V_{t}$ when updating $L_{t}$. In order to avoid this power-tower dependence, we leverage the fact that the total probability mass of cubes with non-trivial correlation is at least $\mathrm{poly}(\epsilon/K)$ (this will be proved in the analysis of Algorithm 2 in Proposition 2.8). By weighting each moment vector by the probability mass of its corresponding cube (see Line 5), we are guaranteed to find a vector with non-trivial correlation by examining only a $\mathrm{poly}(K/\epsilon)$-sized list of candidate vectors.
We note that our algorithm is not proper; that is, it does not return a multiclass linear classifier. Instead, it produces a piecewise constant function over a sufficiently fine partition of $\mathbb{R}^{d}$.
The remaining part of the section is structured as follows: In Section 2.1, we state and prove the key structural result about the class of multiclass linear classifiers, Proposition 2.4, which will be exploited by our algorithm. In Section 2.2, we provide the analysis of Algorithm 2 with Proposition 2.8. In Section 2.3, we analyze Algorithm 1, providing the full proof of Theorem 2.1. Finally, in Section 2.4, we provide the proof of Lemma 2.19, a supporting lemma that is used in the proof of Proposition 2.4. The main result of this section is Theorem 2.1.
2.1 Structural Result for Multiclass Linear Classification
In this section, we prove our main structural result, Proposition 2.4, which enables us to learn multiclass linear classifiers efficiently under adversarial noise. In particular, we show that for a multiclass linear classifier $f$, if $f$ is far from any constant function by more that a constant multiple of the error of $f$ with respect to the noisy label $y$, then the first moment of at least one of the noisy label regions has a positive correlation with a vector in the subspace of weights.
We prove this by finding carefully chosen vectors $\mathbf{v}^{(i)}{\in}\mathbb{R}^{d}$, $i{\in}[K]$, such that $\sum_{i=1}^{K}\operatorname*{\mathbf{E}}[\mathbf{v}^{(i)}\cdot\mathbf{x}\mathds{1}(y=i)]$ is guaranteed to be positive. Note that this sum is equal to $\sum_{i=1}^{K}\operatorname*{\mathbf{E}}[\mathbf{v}^{(i)}\cdot\mathbf{x}\mathds{1}(f(\mathbf{x})=i)]$ plus an error term for instances when $y$ is not equal to $f(\mathbf{x})$. By carefully selecting these vectors, we can ensure that a small probability of error cannot corrupt the sum by much. Formally, we show the following:
Proposition 2.4 (Structural Result).
Let $d\in\mathbb{Z}_{+}$, $\epsilon\in(0,1)$, $D$ be a distribution on labeled examples $(\mathbf{x},y)\in\mathbb{R}^{d}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$. Also, let $f$ be a multiclass linear classifier, $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$ with $\mathbf{w}^{(1)},\ldots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$, and let $C$ be a sufficiently large positive universal constant. If $\epsilon\leq 1/C$ and $f(\mathbf{x})$ is not close to any constant classifier, i.e., ${\min_{i\in[K]}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq i]\geq C\cdot{\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]}+\epsilon$, then for $j=\operatorname*{argmax}_{i\in[K]}t_{i}$ there exists $i\in[K]$ with $\mathbf{w}^{(i)}-\mathbf{w}^{(j)}\neq\mathbf{0}$ such that
$\frac{\mathbf{w}^{(i)}-\mathbf{w}^{(j)}}{\left\|\mathbf{w}^{(i)}-\mathbf{w}^{(j)}\right\|}\cdot\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}[y=i]]\geq\mathrm{poly}(1/K,\epsilon)\;.$
Proof.
Define $\mathrm{OPT}\coloneqq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$. First note that without loss of generality we can assume that $\mathbf{w}^{(1)}=\mathbf{0},t_{1}=0$. This can be seen by noting that the output label does not change if we subtract the same function from all affine functions that define the multiclass linear classifier. Specifically, by subtracting $\mathbf{w}^{(j)}\cdot\mathbf{x}+t_{j}$, where $j=\operatorname*{argmax}_{i\in[K]}t_{i}$ from all affine functions and renaming the labels, we ensure that one of the affine functions becomes $\mathbf{0}$ and that $t_{i}\leq 0$ for all $i\in[K]$.
We define $L_{i}:\mathbb{R}^{d}\to[K],L_{i}(\mathbf{x})=\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}$ the corresponding affine functions such that $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}L_{i}(\mathbf{x})$ and we define the function $s(\mathbf{x})\coloneqq\max_{i\in[K]}L_{i}(\mathbf{x})$.
For some parameter $\gamma>0$ which we will assign later, let $\mathbf{v}^{(i)}\coloneqq\mathbf{w}^{(i)}/(\gamma-t_{i})$ (this is well-defined as $\gamma>0$ and $t_{i}\leq 0$) for all $i\in[K]$. Note that for all nonzero vectors $\mathbf{v}^{(i)}$, we have that $L_{i}(\mathbf{x})\geq\gamma$ if and only if $\mathbf{v}^{(i)}\cdot\mathbf{x}\geq 1$. We can decompose the first moment of label $i$ as follows:
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(y=i)]$ $=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(f(\mathbf{x})=i)]-\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(y\neq i,f(\mathbf{x})=i)]+\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(y=i,f(\mathbf{x})\neq i)]$
$=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{x}\mathds{1}(f(\mathbf{x})=i)]-\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}S_{i}(\mathbf{x},y)]+\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}T_{i}(\mathbf{x},y)]\;,$ (1)
where $S_{i}(\mathbf{x},y)=\mathds{1}(y\neq i,f(\mathbf{x})=i)$ and $T_{i}(\mathbf{x},y)=\mathds{1}(y=i,f(\mathbf{x})\neq i)$ for $i\in[K]$. Taking the inner product with $\mathbf{v}^{(i)}$ and summing Equation 1 for all $i\in[K]$, we have that
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}^{(y)}\cdot\mathbf{x}]$ $=\underbrace{\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}]}_{I_{1}}\underbrace{\;-\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}^{(i)}\cdot\mathbf{x}S_{i}(\mathbf{x},y)]+\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(x,y)\sim D}[\mathbf{v}^{(i)}\cdot\mathbf{x}T_{i}(\mathbf{x},y)]}_{I_{2}}\;.$
Note that $I_{1}$ is the first moment of $\mathbf{v}^{(f(\mathbf{x}))}$, i.e., the first moment of the noiseless multiclass linear classifier, and $I_{2}$ is the contribution that corresponds to the noise. Let $g(\mathbf{x})\coloneqq\max_{j\in[K]}\mathbf{v}^{(j)}\cdot\mathbf{x}$. We first bound $I_{1}$ from below.
Claim 2.5.
It holds that $I_{1}\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]$.
Proof of Claim 2.5.
Note that it holds $\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}\geq 0$ because $L_{f(\mathbf{x})}(\mathbf{x})=\max_{i}L_{i}(\mathbf{x})\geq L_{1}(\mathbf{x})=0$ and therefore $\mathbf{w}^{(f(\mathbf{x}))}\cdot\mathbf{x}\geq|t_{i}|$ which gives that $\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}\geq 0$. Hence, by applying Markov’s inequality, we have
$I_{1}=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}]\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}\geq 1]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[L_{f(\mathbf{x})}(\mathbf{x})\geq\gamma]\;,$
where the last equality holds by the definition of $\mathbf{v}^{(i)}$. Note that the event $L_{f(\mathbf{x})}(\mathbf{x})\geq\gamma$ is a superset of the event $g(\mathbf{x})\geq 1$. Indeed, if $g(\mathbf{x})\geq 1$, there exists some $i\in[K]$ such that $\mathbf{v}^{(i)}\cdot\mathbf{x}\geq 1$, which implies that $L_{i}(\mathbf{x})\geq\gamma$, and therefore $L_{f(\mathbf{x})}(\mathbf{x})\geq\gamma$. This gives that $I_{1}\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]$ and completes the proof of Claim 2.5. ∎
It remains to bound the contribution from the noise, i.e., the quantity $I_{2}$.
Claim 2.6.
There exists a $\gamma>0$ such that $I_{2}\geq-C\left(\sqrt{\mathrm{OPT}\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]}+\mathrm{OPT}\right)$, where $C>0$ is a universal constant, and $961/48\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]/$ 96.
Proof.
We show how to bound the contribution of the terms containing $S_{i}$, as similar arguments can be used to bound the contribution of the terms involving $T_{i}$. In order to bound the contribution of the $S_{i}$ terms from above, notice that
$\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}^{(i)}\cdot\mathbf{x}S_{i}(\mathbf{x},y)]$ $\leq\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}\left[\max_{j\in[K]}(\mathbf{v}^{(j)}\cdot\mathbf{x})S_{i}(\mathbf{x},y)\right]$
$=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}\left[\max_{j\in[K]}(\mathbf{v}^{(j)}\cdot\mathbf{x})S(\mathbf{x},y)\right]$
$=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\max_{j\in[K]}(\mathbf{v}^{(j)}\cdot\mathbf{x})\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]\right]\;,$
where $S\coloneqq\sum_{i=1}^{K}S_{i}$. From the assumptions of Proposition 2.4, we have that $f$ has a missclassification error at most $\mathrm{OPT}$, and therefore we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)]\leq\mathrm{OPT}$. Furthermore, only one of $S_{i}$ and $T_{i}$ for all $i\in[K]$ can be non-zero for each point $(\mathbf{x},y)$, which gives that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]\leq 1$. Let $R(\mathbf{x})\coloneqq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]$ and note that from the above we have that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)]\leq\mathrm{OPT}$ and $R(\mathbf{x})\in[0,1]$.
By the definition of the $\mathbf{v}^{(i)}$’s, we have $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\max_{i\in[K]}L_{i}(\mathbf{x})\geq\gamma]$. Note that $g(\mathbf{x})\geq 0$ by construction, therefore for any $t>0$, we have that
$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})R(\mathbf{x})]$ $\leq t\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})\mathds{1}(g(\mathbf{x})\leq t)]+\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})R(\mathbf{x})\mathds{1}(g(\mathbf{x})>t)]$
$\leq t\mathrm{OPT}+\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\mathds{1}(g(\mathbf{x})>t)]\;,$
where we use the fact that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]\leq\mathrm{OPT}$ and $R(\mathbf{x})\leq 1$. We next bound the term $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\mathds{1}(g(\mathbf{x})>t)]$. Using that for any positive function $G(\mathbf{x})$ it holds that $G(\mathbf{x})=\int_{0}^{\infty}\mathds{1}(G(\mathbf{x})\geq z)\mathrm{d}z$, we have that
$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\mathds{1}(g(\mathbf{x})>t)]=\int_{t}^{\infty}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathds{1}(g(\mathbf{x})\geq y)]\mathrm{d}y$ $\leq\int_{t}^{\infty}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathds{1}(g(\mathbf{x})\geq\lfloor y\rfloor)]\mathrm{d}y$
$=\sum_{n=\lfloor t\rfloor}^{\infty}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[n\leq g(\mathbf{x})]\;,$
where in the first equality, we exchange the order of the integrals using Tonelli’s theorem and in the inequality, we used that $\mathds{1}(g(\mathbf{x})\geq y)$ is a decreasing function of $y$. In order to bound the terms $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[n\leq g(\mathbf{x})]$, we prove that the tails of $g$ are at least quadratically decreasing. We show the following:
Lemma 2.7.
Let $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$, $t_{1},\dots,t_{K}\in\mathbb{R}_{-}$ and $s:\mathbb{R}^{d}\to\mathbb{R}$ such that $s(\mathbf{x})=\max_{i\in[K]}\max(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i},0)$. There exists $\gamma>0$ such that
$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]/{96}\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]\leq 1/{48}$
and
$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq n\gamma]\lesssim(1/n^{2})\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]$
for all $n\in\mathbb{Z}_{+}$.
For the proof of Lemma 2.7, we refer the reader to Section 2.4. Now, we leverage Lemma 2.7 to complete the proof. Note that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma].$ Furthermore, since $t_{i}\leq 0$ for all $i\in[K]$, it follows that if $\mathbf{w}^{(i)}\geq n(\gamma-t_{i})$, then $\mathbf{w}^{(i)}\geq n\gamma-t_{i}$. This, in turn, implies that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq n]\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq n\gamma].$ Therefore, using Lemma 2.7, we have that
$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\mathds{1}(g(\mathbf{x})>t)]\lesssim\sum_{n=\lfloor t\rfloor}^{\infty}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[1\leq g(\mathbf{x})]/n^{2}\lesssim\frac{\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[1\leq g(\mathbf{x})]}{t-2}\;,$
where in the last inequality, we used the approximation of the sum by an integral. Therefore, by choosing $t=2+(\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[1\leq g(\mathbf{x})]/\mathrm{OPT})^{1/2}$, we have that
$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})R(\mathbf{x})]\lesssim t\mathrm{OPT}+\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[1\leq g(\mathbf{x})]/(t-2)\lesssim\sqrt{\mathrm{OPT}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[1\leq g(\mathbf{x})]}+\mathrm{OPT}.$
For the noisy terms defined by $T_{i}$, we have that
$\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(x,y)\sim D}[\mathbf{v}^{(i)}\cdot\mathbf{x}T_{i}(\mathbf{x},y)]$ $=-\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(x,y)\sim D}[\mathbf{v}^{(i)}\cdot(-\mathbf{x})T_{i}(\mathbf{x},y)]=-\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(x,y)\sim D^{\prime}}[\mathbf{v}^{(i)}\cdot\mathbf{x}T_{i}(\mathbf{x},y)]\;,$
where $D^{\prime}$ is a reflection of distribution $D$ with respect to $\mathbf{0}$. As a result, we can arrive at the same bound as for the term defined by the $S_{i}$’s. Therefore, we conclude that the total contribution of outliers is at most $O(\sqrt{\mathrm{OPT}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[1\leq g(\mathbf{x})]}+\mathrm{OPT})$, which completes the proof of Claim 2.6. ∎
By the choice of $\gamma$ in Claim 2.6 and the assumption that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq i]\geq C\cdot\mathrm{OPT}+\epsilon$ we have that $\mathrm{OPT}\leq({96}/C)\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]$. Therefore, for a sufficiently large constant $C>0$, we get that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}^{(y)}\cdot\mathbf{x}]\gtrsim\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]$ and this gives us
$\sum_{i:\mathbf{w}^{(i)}\neq\mathbf{0}}\mathbf{w}^{(i)}/(\gamma-t_{i})\cdot\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(y=i)]\gtrsim\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]\;.$ (2)
Moreover, if for some $i\in[K]$, we have that $\gamma-t_{i}\leq\epsilon\left\|\mathbf{w}^{(i)}\right\|$ then $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[L_{i}(\mathbf{x})\geq\gamma]=\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(0,1)}[x\geq(\gamma-t_{i})/\left\|\mathbf{w}^{(i)}\right\|]\geq\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(0,1)}[x\geq\epsilon]$. However, for $\epsilon<2\sqrt{\log({48})}$ from A.1 we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\max_{i\in[K]}L_{i}(\mathbf{x})\geq\gamma]>1/{48}$ which leads to a contradiction. Hence, we have that $\gamma-t_{i}\geq\epsilon\left\|\mathbf{w}^{(i)}\right\|$ for all $i\in[K]$. Now from the fact that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]\geq C\cdot\mathrm{OPT}+\epsilon$ we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]\gtrsim C\cdot\mathrm{OPT}+\epsilon\gtrsim\epsilon$.
Concluding, from (2) we have that there exists $i\in[K]$ such that $\mathbf{w}^{(i)}/(\gamma-t_{i})\cdot\operatorname*{\mathbf{E}}[\mathbf{x}\mathds{1}(y=i)]\gtrsim\epsilon$. Now since $\gamma-t_{i}\geq 0$, we have that $\mathbf{w}^{(i)}\cdot\operatorname*{\mathbf{E}}[\mathbf{x}\mathds{1}(y=i)]\geq 0$ and hence $\left(\mathbf{w}^{(i)}/\left\|\mathbf{w}^{(i)}\right\|\right)\cdot\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(y=i)]\gtrsim\epsilon^{2}$. Using the fact that we centered the affine functions defining $f$ by subtracting $\mathbf{w}^{(1)}\cdot\mathbf{x}+t_{1}$ from each of them, gives us the statement. This completes the proof of Proposition 2.4.
2.2 Finding a Relevant Direction
This section contains the proof of Proposition 2.8, which states that if our current subspace approximation does not lead to a good enough error, then Algorithm 2 would return at least one vector with non-trivial projection onto the hidden subspace.
Proposition 2.8.
Let $d,k,K\in\mathbb{Z}_{+}$, $\epsilon,\delta\in(0,1)$ and $\epsilon^{\prime}\leq c\epsilon^{2}/({K^{2}\sqrt{k}})$, for $c$ a sufficiently small constant. Let $D$ be a distribution supported on $\mathbb{R}^{d}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ and $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$. Let $f$ be a multiclass linear classifier with weight vectors $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$ and $\mathrm{OPT}=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$. Additionally, let $h_{\mathcal{S}}$ be a piecewise constant approximation of $D$, with respect to an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$, according $V$, as defined in Definitions 2.2 and 2.3.
There exists $N=d\mathrm{poly}(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}\log(1/\delta)$ and a universal constant $C$ such that if $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{{\mathcal{S}}}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+\epsilon$, then Algorithm 2 with probability at least $1-\delta$ returns a list of unit vectors $\mathcal{E}$ of size $|\mathcal{E}|=\mathrm{poly}(1/\epsilon,K)$, such that: There exists $\mathbf{v}\in\mathcal{E}$ and $i,j\in[K],i\neq j$ for which
$\frac{(\mathbf{w}^{(i)}-\mathbf{w}^{(j)})^{\perp V}}{\left\|(\mathbf{w}^{(i)}-\mathbf{w}^{(j)})^{\perp V}\right\|}\cdot\mathbf{v}\geq\mathrm{poly}(1/K,\epsilon)\;.$
Before proving Proposition 2.8, we develop the necessary tools for working with multiclass linear classifiers restricted to cubes extending along the orthogonal complement of a subspace. First, we define the notion of a projected classifier onto a set of linear equations, Definition 2.9.
The intuition behind this definition is that when we restrict our classifier to sufficiently small cubes on $V$, its behavior mainly depends on its components in the orthogonal complement of $V$. In other words, the classifier is nearly equivalent to one that fixes its component on $V$ to a specific point. We denote by the projected classifier of $f$ with respect to a set of linear equations $\mathcal{G}$ the classifier obtained by fixing the coordinates of $\mathbf{x}^{V}$ according to $\mathcal{G}$.
Definition 2.9 (Projected Classifier).
Let $f(\mathbf{x})=\operatorname*{argmax}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$ with $\mathbf{w}^{(i)}\in\mathbb{R}^{d},t_{i}\in\mathbb{R}$ for $i\in[K]$ be a multiclass linear classifier, let $\mathcal{G}=\{\mathbf{v}^{(1)}\cdot\mathbf{x}=z_{1},\ldots\mathbf{v}^{(k)}\cdot\mathbf{x}=z_{k}\}$ be a set of $k$ linear equations and let $V$ be the subspace spanned by orthonormal vectors $\mathbf{v}^{(1)},\ldots,\mathbf{v}^{(k)}\in\mathbb{R}^{d}$. We denote by $f_{\mathcal{G}}(\mathbf{x})$ the linear classifier $f$ projected onto $\mathcal{G}$, i.e., $f_{\mathcal{G}}(\mathbf{x})=\operatorname*{argmax}((\mathbf{w}^{(i)})^{\perp_{V}}\cdot\mathbf{x}+t_{i}+t_{i,\mathcal{G}}^{\prime})$, where $t_{i,\mathcal{G}}^{\prime}=\sum_{j=1}^{k}z_{j}a_{j}$ where $a_{j}=\mathbf{w}^{(i)}\cdot\mathbf{v}^{(j)}$.
Now we prove Lemma 2.10, which states that for any partition of a subspace into sufficiently small cubes, the target function is well-approximated by a piecewise function constructed using a projected classifier for each region. This will be our main supporting result for proving Proposition 2.8, as it allows us to reduce the problem to instances on the orthogonal complement and use Proposition 2.4.
Lemma 2.10 (Projection Lemma).
Let $d,K\in\mathbb{Z}_{+}{,\epsilon>0}$, $f:\mathbb{R}^{d}\to[K]$ be a multiclass linear classifier, $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d},{k\geq 1}$ and $\epsilon>0$. Let $\mathcal{S}$ an $\epsilon$-approximating partition of $\mathbb{R}^{d}$ with respect to $V$ as in Definition 2.2. Then there exists a set $\mathcal{T}\subseteq\mathcal{S}$ such that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathcal{N}(\mathbf{0},\mathbf{I})}[S]\geq 1-\epsilon$ and for any $S\in\mathcal{T}$, it holds that
$\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]\leq K^{2}{\sqrt{k}}\epsilon\;,$
where $f_{\mathcal{G}_{S}}$ denotes the projected classifier onto a set of linear equations $\mathcal{G}_{S}$ that is associated with $S$, as defined in Definition 2.9.
Proof.
Let $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$ and $t_{1},\dots,t_{K}\in\mathbb{R}$ such as $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$. Denote by $\mathbf{v}^{(1)},\dots,\mathbf{v}^{(k)}$ an orthonormal basis of $V$. Fix a set of the partition $S\in\mathcal{S}$ and denote by $f_{\mathcal{G}_{S}}=\operatorname*{argmax}((\mathbf{w}^{(i)})^{\perp{V}}\cdot\mathbf{x}+t_{i}+t_{i}^{\prime})$ the projected classifier that corresponds to the set of linear equations $\mathcal{G}_{S}=\{\mathbf{v}^{(1)}\cdot\mathbf{x}=z_{1},\ldots\mathbf{v}^{(k)}\cdot\mathbf{x}=z_{k}\}$ (see Definition 2.2). Let $\mathbf{w}^{(i,j)}=\mathbf{w}^{(i)}-\mathbf{w}^{(j)}$, $t_{i,j}=t_{i}-t_{j}$ and $t_{i,j}^{\prime}=t_{i}^{\prime}-t_{j}^{\prime}$. By the union bound, we have that
$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\leq$
$\sum_{i\neq j}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathrm{sign}(\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j})\neq\mathrm{sign}((\mathbf{w}^{(i,j)})^{\perp V}\cdot\mathbf{x}+t_{i,j}+t^{\prime}_{i,j})\mid\mathbf{x}\in S]\;.$
Furthermore, by the definition of $f_{\mathcal{G}_{S}}$, if $\mathbf{x}\in S$ then we have that for any $i,j\in[K]$
$\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-t_{i,j}^{\prime}\rvert$ $=\left|\sum_{l=1}^{k}(\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)})(\mathbf{x}\cdot\mathbf{v}^{(l)}-z_{l})\right|$
$\leq\sum_{l=1}^{k}\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)}\rvert\lvert\mathbf{x}\cdot\mathbf{v}^{(l)}-z_{l}\rvert$
$\leq\sum_{l=1}^{k}\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)}\rvert\epsilon$
$\leq\|(\mathbf{w}^{(i,j)})^{V}\|{\sqrt{k}}\epsilon\;,$
where we used the fact that the vector $\sum_{i=1}^{k}z_{i}\mathbf{v}^{(i)}$ belongs in $S$ by the definition of $\mathcal{G}$.
Notice that in order for $\mathrm{sign}(\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j})\neq\mathrm{sign}((\mathbf{w}^{(i,j)})^{\perp V}\cdot\mathbf{x}+t_{i,j}+t^{\prime}_{i,j})$, we need to have that $\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\rvert\leq\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-t_{i,j}^{\prime}\rvert$, because we can rewrite $t^{\prime}_{i,j}=\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}+t_{i,j}^{\prime}$.
Averaging over $S\in\mathcal{S}$ gives us
$\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]$ $\leq\sum_{i\neq j\in[K]}\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}[\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\rvert\leq\|(\mathbf{w}^{(i,j)})^{V}\|{\sqrt{k}}\epsilon\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]$
$\leq\sum_{i\neq j\in[K]}\operatorname*{\mathbf{Pr}}[\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\rvert\leq\|(\mathbf{w}^{(i,j)})^{V}\|{\sqrt{k}}\epsilon]$
$\leq\sum_{i\neq j\in[K]}\frac{\|(\mathbf{w}^{(i,j)})^{V}\|{\sqrt{k}}\epsilon}{\|\mathbf{w}^{(i,j)}\|}\leq K^{2}{\sqrt{k}}\epsilon\;.$
We now proceed with the proof of Proposition 2.8.
Proof of Proposition 2.8.
Let $\mathbf{w}^{(1)},\ldots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$ and $t_{1},\ldots,t_{K}\in\mathbb{R}$ be the parameters defining $f$, i.e., $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$. Denote by $\mathbf{w}^{(i,j)}=\mathbf{w}^{(i)}-\mathbf{w}^{(j)}$, by $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}=(\mathbf{w}^{(i,j)})^{\perp V}/\left\|(\mathbf{w}^{(i,j)})^{\perp V}\right\|$, and let $C$ be a constant that satisfies the statement of Proposition 2.4. Let $\mathcal{G}_{S}$ be the set of linear equations associated with each cube $S\in\mathcal{S}$ as defined in Definition 2.2, and let $f_{\mathcal{G}_{S}}$ be the projection of $f$ onto $\mathcal{G}_{S}$ as defined in Definition 2.9. Furthermore, denote by $\mathbf{u}^{(S,i)}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(y=i)\mid\mathbf{x}\in S],S\in\mathcal{S},i\in[K]$ the population moments of each discretization set and by $\widehat{\mathbf{u}}^{(S,i)}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\widehat{D}}[\mathbf{x}\mathds{1}(y=i)\mid\mathbf{x}\in S],S\in\mathcal{S},i\in[K]$ its empirical counterpart, computed in Line 4. Also, let $\widehat{\mathbf{U}}$ be the matrix computed at Line 5 of Algorithm 2.
We first show that if the current subspace $V$ does not lead to a good approximation of $f$, then by splitting the space into small cubes with respect to $V$, there will be a nontrivial fraction of them for which $f$ is far from constant.
Claim 2.11 (Large Fraction of Good Cubes).
If $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{{\mathcal{S}}}(\mathbf{x})\neq y]>(C+1)\mathrm{OPT}+5\epsilon$, then there exists a subset $\mathcal{T}\subseteq\mathcal{S}$ such that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]\geq\epsilon$; and for all $i\in[K],S\in\mathcal{T}$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]+\epsilon$.
Proof of Claim 2.11.
Define $\mathcal{T}$ to be all sets $S\subseteq\mathcal{S}$ such that $\operatorname*{\mathbf{Pr}}[i\neq y\mid\mathbf{x}\in S]>(C+1)\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y\mid\mathbf{x}\in S]+2\epsilon$, for all $i\in[K],S\in\mathcal{T}$. Assume for the sake of contradiction that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]\leq 2\epsilon$. We have that
$\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{{\mathcal{S}}}(\mathbf{x})\neq y]$ $\leq\sum_{S\in\mathcal{S}}\min_{\widehat{y}\in[K]}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[\widehat{y}\neq y\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]+\epsilon$
$\leq\sum_{S\not\in\mathcal{T}}\min_{\widehat{y}\in[K]}\left((C+1)\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}\in S]+2\epsilon\right)\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]+3\epsilon$
$\leq(C+1)\mathrm{OPT}+5\epsilon\;,$
where in the first equality, we used the fact that at most an $\epsilon$ fraction of the mass is outside the discretization (see A.1). This leads to a contradiction and hence $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]>2\epsilon$.
Furthermore, note that by Lemma 2.10 there exists $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ such that $\sum_{S\in\mathcal{T}^{\prime}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]>\epsilon$ and for all $i\in[K],S\in\mathcal{T}^{\prime}$, it holds that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[i\neq y\mid\mathbf{x}\in S]>(C+1)\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(x)\neq y\mid\mathbf{x}\in S]+\epsilon$. Since $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[i\neq y\mid\mathbf{x}\in S]\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]+\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]$ we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]+\epsilon$ for all $S\in\mathcal{T}^{\prime}$, which completes the proof. ∎
In the next claim, we show that if there exists a cube where the target function is far from being constant, then we can find a vector that correlates with the unknown parameters.
Claim 2.12 (Correlating Moments).
Let $S\in\mathcal{S}$ be a subset of the partition such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]+\epsilon$. Then there exists some $i\in[K]$ and a unit vector $\mathbf{w}\in W$ for which $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{u}^{(S,i)}\geq\mathrm{poly}(1/K,\epsilon)$.
Proof of Claim 2.12.
First, notice that the output label of $f_{\mathcal{G}_{S}}$ is only determined by the projection at the orthogonal complement of $V$, i.e.,
$f_{\mathcal{G}_{S}}(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}((\mathbf{w}^{(i)})^{\perp V}\cdot\mathbf{x}+t_{i}+t_{\mathcal{G}_{S},i}^{\prime})=\operatorname*{argmax}_{i\in[K]}((\mathbf{w}^{(i)})^{\perp V}\cdot\mathbf{x}^{\perp V}+t_{i}+t_{\mathcal{G}_{S},i}^{\prime})=f_{\mathcal{G}_{S}}(\mathbf{x}^{\perp{V}}).$
Furthermore, note that the condition whether $\mathbf{x}$ belongs to $S$ depends only on the projection of $\mathbf{x}^{V}$, hence without loss of generality it can written as $\mathds{1}(\mathbf{x}^{V}\in S)$. Thus, we have that
$\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]$ $=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x}^{\perp V})\neq y\mid\mathbf{x}\in S]$
[…]
$=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y]\;,$
where $D_{V^{\perp}}$ is defined to be the marginal obtained by averaging $D$ over $V$ conditioned on $S$. Similarly, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y]+\epsilon$ for all $i\in[K]$. Moreover, by the properties of the standard Gaussian distribution, we have that the $\mathbf{x}$-marginal of $D_{V^{\perp}}$ is a standard Gaussian.
As a result, by applying Proposition 2.4 for $f_{\mathcal{G}_{S}}$, we have that there exist $i,j\in[K],i\neq j$ such that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[\mathbf{x}\mathds{1}[y=i]]\geq\mathrm{poly}(1/K,\epsilon)$. Furthermore, note that by averaging over $V$ conditioning on $S$ we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}^{\perp V}\mathds{1}[y=i]\mid\mathbf{x}\in S]=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[\mathbf{x}\mathds{1}[y=i]]$. Therefore, there exist $i,j\in[K],i\neq j$, such that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{u}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$ for some $l\in[K]$.
In the next claim, we show that by taking enough samples, the empirical moments are close to the population ones.
Claim 2.13 (Concentration of Empirical Distribution).
Let $p(\epsilon,1/K)$ be a sufficiently small polynomial in $\epsilon$ and $1/K$. If $N=d(k/\epsilon^{\prime})^{Ck}(K/\epsilon)^{C}\log(1/\delta)$ for a sufficiently large universal constant $C>0$, then with probability $1-\delta$ we have that $\left\|\widehat{\mathbf{u}}^{(S,i)}-{\mathbf{u}}^{(S,i)}\right\|\leq{p(\epsilon,1/K)}$ and $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S},i\in[K]$.
Proof of Claim 2.13.
First note that by the definition of an approximating partition (see Definition 2.2) and A.1, for all $S\in\mathcal{S}$ we have that $\operatorname*{\mathbf{Pr}}_{D}[S]{=(\epsilon^{\prime}/k)^{\Omega(k)}}$ and $\lvert\mathcal{S}\rvert{=(k/\epsilon^{\prime})^{O(k)}}$. Therefore, by union bound and Hoeffding’s inequality, it holds that if $N\geq{(k/\epsilon^{\prime})^{Ck}}\log(\lvert\mathcal{S}\rvert/\delta)={(k/\epsilon^{\prime})^{O(k)}}\log(1/\delta)$ for a sufficiently large constant $C>0$, then with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]-\operatorname*{\mathbf{Pr}}_{D}[S]\rvert\leq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S}$. Hence, it is true that $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$, i.e., the number of samples that fall in each set $S\in\mathcal{S}$ is at least $N\operatorname*{\mathbf{Pr}}_{D}[S]/2{=N(\epsilon^{\prime}/k)^{\Omega(k)}}$.
Furthermore, note that conditioning on any $S$, the random vector $\mathbf{x}\mathds{1}(y=i)$ with $(\mathbf{x},y)\sim D$ is subgaussian in any unit direction with subgaussian norm at most a constant. Because, we can decompose $\mathbf{v}\cdot\mathbf{x}=\mathbf{v}\cdot\mathbf{x}^{V}+\mathbf{v}\cdot\mathbf{x}^{\perp V}$ with $\mathbf{v}\cdot\mathbf{x}^{V},\mathbf{v}\cdot\mathbf{x}^{\perp V}$ being independent, where the zero mean centering $\mathbf{v}\cdot\mathbf{x}^{V}$ is bounded in any direction by $\epsilon^{\prime}\sqrt{k}\leq\epsilon$ since we have conditioned on $S$ and $\mathbf{v}\cdot\mathbf{x}^{\perp V}$ is a standard gaussian since $S$ depends only on $\mathbf{x}^{V}$. Therefore, by the use of a standard covering argument it follows that if the number of samples that falls in each discretization set $S$ is at least $d\mathrm{poly}(K,1/\epsilon)\log(K/\delta)$ for a sufficiently large polynomial, then $\left\|\widehat{\mathbf{u}}^{(S,i)}-{\mathbf{u}}^{(S,i)}\right\|\leq{p(\epsilon,1/K)}$ for all $i\in[K]$.
Combining the above results with the allong with union bound gives us that there exists an $N=d(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}\log(1/\delta)$ that suffices.
Applying Claims 2.11 and 2.12, we have that if $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{{\mathcal{S}}}(\mathbf{x})\neq y]>(C+1)\mathrm{OPT}+\epsilon$ then there exists a subset $\mathcal{T}\subseteq\mathcal{S}$ of the $\epsilon^{\prime}$-approximating partition with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{D}[S]=\Omega(\epsilon)$ such that for all $S\in\mathcal{T}$ we have that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{u}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$ for some $i,j,l\in[K],i\neq j$. Moreover, if $N=d(k/\epsilon^{\prime})^{C^{\prime}k}(K/\epsilon)^{C^{\prime}}\log(1/\delta)$ for a sufficiently large universal constant $C^{\prime}>0$, using the concentration result of Claim 2.13, we have that the empirical moments of each set satisfy the same guarantees, i.e., $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\widehat{\mathbf{u}}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$, for some $i,j,l\in[K],i\neq j,$ for all $S\in\mathcal{T}$. Furthermore, note that by Claim 2.13 we also have that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]=\Omega(\epsilon)$.
Therefore, there exists a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ and $i,j\in[K]$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\Omega(\epsilon/K^{2})$ such that for all $S\in\mathcal{T}$ we have that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\widehat{\mathbf{u}}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$ for some $l\in[K]$. Hence, for some $i\neq j\in[K]$ we have that
$\left((\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\right)^{\top}\widehat{\mathbf{U}}(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\geq\frac{\epsilon}{K^{2}}\mathrm{poly}(\epsilon,1/K)=\mathrm{poly}(1/K,\epsilon)\;,$
where we used the fact that $\mathbf{x}\mathbf{x}^{\top}$ is PSD for any $\mathbf{x}\in\mathbb{R}^{d}$.
Moreover, note that $(\mathbf{u}^{(S,i)})^{\perp V}=\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp V}}[\mathbf{x}^{\perp V}\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[\mathds{1}(y=i)\mid\mathbf{x}\in S]]$ as $S$ depends only on $\mathbf{x}^{V}$ and
$\left\|(\mathbf{u}^{(S,i)})^{\perp V}\right\|$ $\leq\max_{\left\|\mathbf{u}\right\|=1}\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp V}}[\mathbf{u}\cdot\mathbf{x}^{\perp V}\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[\mathds{1}(y=i)\mid\mathbf{x}\in S]]$
$\leq\max_{\left\|\mathbf{u}\right\|=1}(\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp V}}[(\mathbf{u}\cdot\mathbf{x}^{\perp V})^{2}]\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathds{1}(y=i)\mid\mathbf{x}\in S]])^{1/2}\leq 1\;,$
where we used the Cauchy-Schwarz inequality and the dual characterization of the $\ell_{2}$ norm. Therefore, by the triangle inequality, it follows that
$\left\|\widehat{\mathbf{U}}\right\|_{F}\leq\sum_{S\in\mathcal{S},i\in[K]}\|(\widehat{\mathbf{u}}^{(S,i)})^{\perp V}\|\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\leq\sum_{S\in\mathcal{S},i\in[K]}(1+\mathrm{poly}(\epsilon,1/K))\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\leq 2K\;.$
Hence, by Claim B.1, we have that there exists a unit eigenvector $\mathbf{v}$ of $\widehat{\mathbf{U}}$ with eigenvalue at least ${\mathrm{poly}(\epsilon/K)}$ for a sufficiently small polynomial, such that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{v}\geq\mathrm{poly}(1/K,\epsilon)$, for some $i\neq j\in[K]$. Moreover, the number of such eigenvectors is at most $2K/\mathrm{poly}(1/K,\epsilon)=\mathrm{poly}(K,1/\epsilon)$, which completes the proof of Proposition 2.8.
Before proving our main result of this section, Theorem 2.1, we need to show that if we have learned a direction of our target model to sufficiently high accuracy, then we can assume, without loss of generality, that the model is defined only on the remaining directions. Specifically, we establish the following:
Lemma 2.14.
Let $K\geq 2,\epsilon\in(0,1)$ and $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$ with $\mathbf{w}^{(i)}\in\mathbb{R}^{d}$ for $i\in[K]$ be a multiclass linear classifier. Let $V$ be a subspace such that for all $i,j\in[K]$ with $i\neq j$ it holds that $\|(\mathbf{w}^{(i)}-\mathbf{w}^{(j)})^{V}\|\leq(\epsilon^{2}/K^{4})\|\mathbf{w}^{(i)}-\mathbf{w}^{(j)}\|$. Then, it holds that
$\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f^{\perp V}(\mathbf{x})]\lesssim\epsilon\;,$
where $f^{\perp V}(\mathbf{x})=\operatorname*{argmax}((\mathbf{w}^{(i)})^{\perp V}\cdot\mathbf{x}+t_{i})$.
Proof.
For simplicity, let $\mathbf{w}^{(i,j)}=\mathbf{w}^{(i)}-\mathbf{w}^{(j)}$ and $t_{i,j}=t_{i}-t_{j}$. We have that
$\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f^{\perp V}(\mathbf{x})]\leq\sum_{i,j\in[K]}\operatorname*{\mathbf{Pr}}[\mathrm{sign}(\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j})\neq\mathrm{sign}((\mathbf{w}^{(i,j)})^{\perp V}\cdot\mathbf{x}+t_{i,j})]\;.$
We can write $\mathbf{w}^{(i,j)}=(\mathbf{w}^{(i,j)})^{V}+(\mathbf{w}^{(i,j)})^{\perp V}$, with $\|(\mathbf{w}^{(i,j)})^{V}\|\leq(\epsilon^{2}/K^{4})\|\mathbf{w}^{(i,j)}\|$ and $\|(\mathbf{w}^{(i,j)})^{\perp V}\|\geq\sqrt{1-(\epsilon^{4}/K^{8})}\|\mathbf{w}^{(i,j)}\|$. Therefore, using the independence of any two orthogonal directions of a standard Gaussian, along with A.1, we have that
$\operatorname*{\mathbf{Pr}}[\mathrm{sign}(\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j})\neq\mathrm{sign}((\mathbf{w}^{(i,j)})^{\perp V}\cdot\mathbf{x}+t_{i,j})]$ $\leq\operatorname*{\mathbf{Pr}}[|(\mathbf{w}^{(i,j)})^{V}\cdot\mathbf{x}|\geq\lvert(\mathbf{w}^{(i,j)})^{\perp V}\cdot\mathbf{x}+t_{i,j}\rvert]$
$\leq\operatorname*{\mathbf{Pr}}_{z\sim\mathcal{N}(0,1)}[|z|\geq\frac{\|(\mathbf{w}^{(i,j)})^{\perp V}\|(\epsilon/K^{2})}{\|(\mathbf{w}^{(i,j)})^{V}\|}]+\epsilon/K^{2}$
$\leq\operatorname*{\mathbf{Pr}}_{z\sim\mathcal{N}(0,1)}[|z|\geq\frac{\sqrt{1-\epsilon^{4}/K^{8}}}{\epsilon/K^{2}}]+\epsilon/K^{2}\leq 2\epsilon/K^{2}\;.$
Therefore, we have that
$\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f^{\perp V}(\mathbf{x})]\leq 2\epsilon\;.$
2.3 Proof of Theorem 2.1
In this section, given Proposition 2.4, we proceed to the proof of Theorem 2.1.
Proof of Theorem 2.1.
We show that Algorithm 1, with high probability, returns a hypothesis $h$ with 0-1 error $O(\mathrm{OPT})+\epsilon$. Let $\mathbf{w}^{*(1)},\ldots,\mathbf{w}^{*(K)}\in\mathbb{R}^{d}$ and $t_{1}^{*},\ldots,t_{K}^{*}\in\mathbb{R}$ be the parameters of the affine functions of an optimal classifier $f^{*}$, i.e., $f^{*}(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{*(i)}\cdot\mathbf{x}+t_{i}^{*})$ and denote by $W^{*}=\mathrm{span}(\mathbf{w}^{*(1)},\ldots,\mathbf{w}^{*(K)})$. Let $C>0$ be an absolute constant that satisfies the statement of Proposition 2.8. Let $L_{t}$ be the list of vectors updated by the algorithm (Line 3c) and $V_{t}=\mathrm{span}(L_{t})$, $\dim(V_{t})=k_{t}$. Also let $\epsilon^{\prime}=\mathrm{poly}(\epsilon/K)$ for a sufficiently small polynomial, and for $t\in[T]$ let $\mathcal{S}_{t}$ be arbitrary $\epsilon^{\prime}$-approximating partitions with respect to $V_{t}$ (see Definition 2.2). Let $h_{t}:\mathbb{R}^{d}\to[K]$ be piecewise constant classifiers, defined as $h_{t}=h_{\mathcal{S}_{t}}$ according to Definition 2.3 for the distribution $D$.
To prove the correctness of Algorithm 1, we need to show that if $h_{t}$ has significant error, then the algorithm improves its approximation $V_{t}$ to $W^{*}$. For quantifying the improvement at every step, we consider the following potential function
$\Phi_{t}=\sum_{i=1}^{K}\left\|(\mathbf{b}^{(i)})^{\perp V_{t}}\right\|^{2}\;,$
where $\mathbf{b}^{(1)},\dots,\mathbf{b}^{(K)}$ is an orthonormal basis of $W^{*}$.
Our first step is to show that as long as the algorithm has not terminated, then with high probability we can find a weight vector that correlates non-trivially with the unknown subspace $W^{*}$.
Claim 2.15 (Large Error Implies Good Correlation).
For all $k_{t}=\mathrm{poly}(K/\epsilon)$, if $\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+2\epsilon$ and $N\geq d\mathrm{poly}(K,k_{t},1/\epsilon)^{k_{t}}\log(1/\delta)$, then with probability at least $1-\delta$ there exists a unit vector $\mathbf{v}^{(t)}\in V_{t+1}$ such that $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\mathrm{poly}(1/K,\epsilon)$ for some unit vector $\mathbf{w}\in W^{*}$.
Proof of Claim 2.15.
First, note that using a sufficiently small polynomial for $\epsilon^{\prime}$, we have that $\epsilon^{\prime}\leq c\epsilon^{2}/(K^{2}\sqrt{k_{t}})$ for all $k_{t}=\mathrm{poly}(K/\epsilon)$ for a sufficiently small constant $c$. Denote by $\mathbf{u}^{(1)},\dots,\mathbf{u}^{(k^{\prime})}$ the unit vectors return by Algorithm 2 at iteration $t$, i.e., the vectors in $\mathcal{E}_{t}$. For the sake of contradiction assume that $\lvert\mathbf{u}^{(i)}\cdot\mathbf{w}\rvert\leq\epsilon^{2}/(C^{2}K^{4})$, for all unit vectors $\mathbf{w}\in W^{*}$.
Let $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}$ be vectors in $\mathbb{R}^{d}$. For simplicity, define $\mathbf{w}^{(i,j)}\coloneqq\mathbf{w}^{(i)}-\mathbf{w}^{(j)},\mathbf{w}^{*(i,j)}\coloneqq\mathbf{w}^{*(i)}-\mathbf{w}^{*(j)}$. By Lemma 2.14, there exists a classifier $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}^{*})$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq f^{*}(\mathbf{x})]\leq\epsilon/C$, and $\mathbf{w}^{(i)}=(\mathbf{w}^{*(i)})^{\perp U},$ for all $i\in[K]$, where $U=\mathrm{span}(\mathbf{u}^{(1)},\dots,\mathbf{u}^{(k^{\prime})})$.
Since $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]\leq\mathrm{OPT}+\epsilon/C,$ we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]+\epsilon$. Therefore, by applying Proposition 2.8, for $f$, with probability $1-\delta$ there exists a vector $\mathbf{u}^{(l)}\in\mathcal{E}_{t}$ such that $\mathbf{u}^{(l)}\cdot\mathbf{w}^{(i,j)}\geq\mathrm{poly}(1/K,\epsilon)\left\|(\mathbf{w}^{(i,j)})^{\perp V_{t}}\right\|$ for some $i,j\in[K]$ with $(\mathbf{w}^{(i,j)})^{\perp V_{t}}\neq\mathbf{0}$. However, since $\mathbf{w}^{(i,j)}\in U^{\perp}$ we have a contradiction, which concludes the proof. ∎
Our next claim quantifies the decrease of the potential function $\Phi_{t}$ after adding a new weight vector.
Claim 2.16 (Potential Decrease).
Let $\beta\geq 0$. If there exists a unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ and $\mathbf{w}\in W^{*}$ such that
$\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\beta\;,$
then $\Phi_{t+1}\leq\Phi_{t}-\beta^{2}/2$.
Proof of Claim 2.16.
First note that $\sum_{1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})(\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}})\geq\beta$ because we can write $\mathbf{w}$ as $\sum_{1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})\mathbf{b}^{(i)}$. Also note that $\lvert\mathbf{b}^{(i)}\cdot\mathbf{w}\rvert\leq\left\|\mathbf{w}\right\|=1$ for all $i\in[K]$. Hence, by the Pythagorean theorem and the dual characterization of norms, we have that
$\Phi_{t+1}=\sum_{i=1}^{K}\|(\mathbf{b}^{(i)})^{\perp V_{t}}\|^{2}-\|(\mathbf{b}^{(i)})^{V_{t+1}\cap\perp V_{t}}\|^{2}$ $\leq\Phi_{t}-\sum_{i=1}^{K}\left\lvert\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\right\rvert^{2}$
$\leq\Phi_{t}-\left\lvert\sum_{i=1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})(\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}})\right\rvert^{2}/2$
$\leq\Phi_{t}-\beta^{2}/2\;.$
This completes the proof of Claim 2.16. ∎
Note that by Claims 2.15 and 2.16, if $h_{t}$ has large misclassification error, we can decrease the potential $\Phi_{t}$ by a non-trivial amount. However, since the potential function is nonnegative, this implies that at some iteration $t$, $h_{t}$ must have a small error. We aim to show that once this happens, the subsequent hypotheses $h_{t}$ also maintain small error. In the following claim, we show that by adding new vectors to the subspace $V_{t}$, we decrease the error of our piecewise constant approximation.
Claim 2.17 (Error Decrease).
For each $t\in[T]$, it holds $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t+1}(\mathbf{x})\neq y]\leq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]$.
Proof of Claim 2.17.
Since the statement of Proposition 2.8 holds for any $\epsilon^{\prime}$-partitions $\mathcal{S}_{t},t\in[T]$, independent of the choice of basis and threshold points, we can assume that all the classifiers $h_{t}$ are computed with respect to extensions of a common orthonormal basis and that all threshold points are aligned. Hence, $\mathcal{S}_{t+1}$ is a subdivision of $\mathcal{S}_{t}$. Thus, each set $S\in\mathcal{S}_{t}$ is equal to a union of sets $S_{1},\dots,S_{l}\in\mathcal{S}_{t+1}$, i.e., $S=\cup_{i=1}^{l}S_{i}$. Therefore, by the definition of $h_{t}$, we have that
$\operatorname*{\mathbf{Pr}}[h_{t+1}(\mathbf{x})\neq y,\mathbf{x}\in S]$ $=\sum_{i=1}^{l}\operatorname*{\mathbf{Pr}}[h_{t+1}(\mathbf{x})\neq y\mid\mathbf{x}\in S_{i}]\operatorname*{\mathbf{Pr}}[S_{i}]=\sum_{i=1}^{l}\min_{\widehat{y}\in[K]}\operatorname*{\mathbf{Pr}}[y\neq\widehat{y}\mid\mathbf{x}\in S_{i}]\operatorname*{\mathbf{Pr}}[S_{i}]$
$\leq\min_{\widehat{y}\in[K]}\sum_{i=1}^{l}\operatorname*{\mathbf{Pr}}[y\neq\widehat{y}\mid\mathbf{x}\in S_{i}]\operatorname*{\mathbf{Pr}}[S_{i}]=\min_{\widehat{y}\in[K]}\operatorname*{\mathbf{Pr}}[\widehat{y}\neq y,\mathbf{x}\in\cup_{i=1}^{l}S_{i}]$
$\leq\operatorname*{\mathbf{Pr}}[h_{t}(\mathbf{x})\neq y,\mathbf{x}\in S]\;.$
Summing over all $S\in\mathcal{S}$ we conclude the statement. ∎
In the following claim, we show that using a sufficient number of samples, we can obtain a good estimate for the population piecewise constant approximation function.
Claim 2.18.
Let $\epsilon,\epsilon^{\prime},\delta\in(0,1),k,K\in\mathbb{Z}_{+}$ with $\epsilon^{\prime}\leq\epsilon/2$. Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$, and let $h$ and $\widehat{h}$ be the piecewise constant approximating functions defined according to the same partition, with width $\epsilon^{\prime}$ and subspace $V$ with respect to distributions $D$ and $\widehat{D}$, respectively (Definition 2.3), where $\widehat{D}$ is the empirical distribution of $N$ i.i.d. samples from $D$. There exists $N={(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}}\log(1/\delta)$ such that with probability $1-\delta$ we have that $\operatorname*{\mathbf{Pr}}[\widehat{h}(\mathbf{x})\neq y]\leq\operatorname*{\mathbf{Pr}}[h(\mathbf{x})\neq y]+\epsilon\;.$
Proof.
Denote by $\mathcal{S}$ the approximating partition for the functions $h,\widehat{h}$. First, note that by the definition of an approximating partition (Definition 2.2) and by A.1, we have that $\operatorname*{\mathbf{Pr}}_{D}[S]={(\epsilon^{\prime}/k)^{\Omega(k)}}$ and hence $\lvert\mathcal{S}\rvert={(k/\epsilon^{\prime})^{O(k)}}$. Therefore, by the union bound and Hoeffding’s inequality it holds that if $N\geq C\log(\lvert\mathcal{S}\rvert K/\delta)/(\epsilon/\lvert S\rvert)^{2}={(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}}\log(1/\delta)$ for a sufficiently large universal constant $C>0$, then with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[\mathbf{x}\in S,y=i]-\operatorname*{\mathbf{Pr}}_{D}[\mathbf{x}\in S,y=i]\rvert\leq\epsilon/(4\lvert\mathcal{S}\rvert)$ for all $S\in\mathcal{S},i\in[K]$. Hence, with probability $1-\delta$ we have that for all $S\in\mathcal{S}$ it holds $\operatorname*{\mathbf{Pr}}[\widehat{h}(\mathbf{x})\neq y]\leq\operatorname*{\mathbf{Pr}}[h(\mathbf{x})\neq y]+\epsilon/(2\lvert\mathcal{S}\rvert)$. The claim follows, as the total error is simply the summation of the error over cubes and the error outside the partition which is at most $\epsilon^{\prime}$ from Definition 2.2. ∎
Note that from Lines 1, 3 of Algorithm 1 we perform at most $\mathrm{poly}(K,1/\epsilon)$ iterations. Furthermore, in each iteration, we update the vector set with at most $\mathrm{poly}(K,1/\epsilon)$ vectors. Hence, it follows that $k_{t}\leq\mathrm{poly}(K,1/\epsilon)$ for all $t=1,\dots,T$.
Assume that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+\epsilon$ for all $t=1,\dots,T$. Using the fact that $N=d2^{\mathrm{poly}(K,1/\epsilon)}\log(1/\delta)$ for a sufficiently large polynomial (Line 2 of Algorithm 1), we can apply Claim 2.15 and conclude that with probability $1-\delta$ there exist unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ for $t=[T]$ such that $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\mathrm{poly}(1/K,\epsilon)$. Thus, from Claim 2.16, we have that with probability $1-\delta$, for all $t\in[T]$, $\Phi_{t}\leq\Phi_{t-1}-\mathrm{poly}(\epsilon,1/K)$. After $T$ iterations, it follows that $\Phi_{T}\leq\Phi_{0}-T\mathrm{poly}(\epsilon,1/K)=1-T\mathrm{poly}(\epsilon,1/K)$. However, if $T$ is set to be a sufficiently large polynomial of $\epsilon$ and $K$, we would arrive at a contradiction, since $\Phi_{T}\geq 0$. Hence, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]\leq C\cdot\mathrm{OPT}+\epsilon$ for some $t\in\{1,\dots,T\}$. Since the error of $h_{t}$ can only be decreasing by Claim 2.17 and $h_{t}$ is close to its sample variant by Claim 2.18, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]<C\cdot\mathrm{OPT}+\epsilon$.
Sample and Computational Complexity: From the analysis above, we have that the algorithm terminates in $\mathrm{poly}(1/\epsilon,K)$ iterations and at each iteration we draw $d2^{\mathrm{poly}(K,1/\epsilon)}\log(1/\delta)$ samples. Hence, we have that the total number of samples is of the order of $d2^{\mathrm{poly}(K,1/\epsilon)}\log(1/\delta)$. Moreover, the algorithm runs in $\mathrm{poly}(N)$ time as all operations can be implemented in polynomial time. ∎
2.4 Proof of Lemma 2.7
This section proves Lemma 2.7, one of the supporting lemmas used in the proof of Proposition 2.4. For completeness we restate the lemma. See 2.7 The proof of the lemma immediately follows from Lemma 2.19 and Claim 2.20. First, we establish Lemma 2.19, which provides a bound on the tails of a piecewise affine function. Then, using Claim 2.20, we show that there is indeed a suitable $\gamma$ to which this tail bound can be applied.
Lemma 2.19 (Tail Bounds).
Let $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$, $t_{1},\dots,t_{K}\in\mathbb{R}$ and $s:\mathbb{R}^{d}\to\mathbb{R}$ such that $s(\mathbf{x})=\max_{i\in[K]}\max(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i},0)$. For any $\gamma>0$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]\leq 1/{48}$, we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq n\gamma]\lesssim(1/n^{2})\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]$ for all $n\in\mathbb{Z}_{+}$.
Proof.
By the properties of addition and scalar multiplication of independent Gaussian random variables, we have that for any $n\in\mathbb{Z}_{+}$ a standard Gaussian random variable $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ can be written as $\mathbf{x}=\frac{1}{n}\sum_{i=1}^{n^{2}}\mathbf{z}^{(i)}$ where $\mathbf{z}^{(i)}$ are drawn i.i.d. from $\mathcal{N}(\mathbf{0},\mathbf{I})$. Now if for some $i\in[n^{2}]$ we have that $s(\mathbf{z}^{(i)})\geq n\gamma$, then there exists a $j\in[K]$ such that $\mathbf{w}^{(j)}\cdot\mathbf{z}^{(i)}+t_{j}\geq n\gamma$ and, furthermore, we have that
$\mathbf{w}^{(j)}\cdot\mathbf{x}+t_{j}$ $=\frac{1}{n}\mathbf{w}^{(j)}\cdot\mathbf{z}^{(i)}+t_{j}+\frac{1}{n}\sum_{k\neq i}\mathbf{w}^{(j)}\cdot\mathbf{z}^{(k)}\geq\gamma+\mathbf{w}^{(j)}\cdot\bigg{(}\frac{1}{n}\sum_{k\neq i}\mathbf{z}^{(k)}\bigg{)}.$
Note that $\frac{1}{n}\sum_{k\neq i}\mathbf{z}^{(k)}$ is distributed according to $\mathcal{N}(\mathbf{0},\frac{n^{2}-1}{n^{2}}\mathbf{I})$; hence, by symmetry we have that $\operatorname*{\mathbf{Pr}}[\mathbf{w}^{(j)}\cdot(\frac{1}{n}\sum_{k\neq i}\mathbf{z}^{(k)})\geq 0]=1/2$. As a result we have that $\operatorname*{\mathbf{Pr}}[s(\mathbf{x})\geq\gamma\mid s(\mathbf{z}^{(i)})\geq n\gamma]\geq 1/2$, for all $i\in[n^{2}]$.
Denote by $p_{n}=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq n\gamma]$. To prove Lemma 2.19, it suffices to show that $p_{n}\leq 4p_{1}/n^{2}$, for all $n\geq 1$. Note that this holds trivially true for $n=1$. Assume as an inductive hypothesis that $p_{n-1}\leq 4p_{1}/(n-1)^{2}$ for some $n\geq 2$. We will prove that $p_{n}\leq 4p_{1}/n^{2}$ for $n\geq 2$.
Fix $i\in[n^{2}]$ and denote by $A$ the event that $s(\mathbf{x})\geq\gamma$ and by $B_{i}^{(n)}$ the event that $s(\mathbf{z}^{(i)})\geq n\gamma$. Note that
$\operatorname*{\mathbf{Pr}}\left[A,B_{i}^{(n)},\bigcap_{j\neq i}B_{j}^{(n)}\right]$ $\geq\operatorname*{\mathbf{Pr}}[A,B_{i}^{(n)}]-\sum_{\begin{subarray}{c}j=1\\ j\neq i\end{subarray}}^{n^{2}}\operatorname*{\mathbf{Pr}}[B_{i}^{(n)},B_{j}^{(n)}]$
$\geq p_{n}/2-(n^{2}-1)p_{n}^{2}\;,$
where in the first inequality we used the union bound, and in the last inequality we used the independence of the $\mathbf{z}^{(i)}$’s and that $\operatorname*{\mathbf{Pr}}[s(\mathbf{x})\geq\gamma\mid s(\mathbf{z}^{(i)})\geq n\gamma]\geq 1/2$, for all $i\in[n^{2}]$.
Note that it holds $p_{n}\leq p_{n-1}$, since $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq n\gamma]\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq(n-1)\gamma]$. Consequently, by the inductive hypothesis we have that $(n^{2}-1)p_{n}^{2}\leq 4p_{1}(n^{2}-1)/(n-1)^{2}p_{n}$. Hence, since $(n^{2}-1)/(n-1)^{2}\leq 3$ for all $n\geq 2$ and $p_{1}\leq 1/48$, we have that $(n^{2}-1)p_{n}^{2}\leq p_{n}/4$.
Summing over all $i\in[n^{2}]$ we have
$p_{1}=\operatorname*{\mathbf{Pr}}[A]\geq\operatorname*{\mathbf{Pr}}\left[A,\bigcup_{i\in[n^{2}]}\left(B_{i}^{(n)},\bigcap_{j\neq i}B_{j}^{(n)}\right)\right]\geq n^{2}p_{n}/4\;.$
Therefore, we have that $p_{n}\leq 4p_{1}/n^{2}$, which by the principle of mathematical induction completes the proof of Lemma 2.19. ∎
In order to apply Lemma 2.19, we need to find a threshold such that the probability that $s$ is greater than this threshold is bounded from above by $1/{48}$. We show the following
Claim 2.20.
Fix $c\in(0,1)$, there exists $\gamma>0$ such that
$c\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]\geq\frac{c}{2}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]\;.$
Proof of Claim 2.20.
First notice that the quantity $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]$ is continuous with respect to $\gamma$ for any $\gamma>0$. Indeed, $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\gamma\leq s(\mathbf{x})\leq\gamma+\epsilon]\leq\sum_{i=1}^{K}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[-t_{i}+\gamma\leq\mathbf{w}^{(i)}\cdot\mathbf{x}\leq-t_{i}+\gamma+\epsilon]$ where each term such that $\mathbf{w}^{(i)}\neq\mathbf{0}$ tends to 00 as $\epsilon$ tends to 00 and each term such that $\mathbf{w}^{(i)}=\mathbf{0}$ is equal to 00 from the fact that $-t_{i}+\gamma>0$. Also note that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]$ tends to $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]$ as $\gamma$ approaches 00; hence, there exists $\gamma^{\prime}$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]-\epsilon/2<\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma^{\prime}]\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]$. Therefore, if $c\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]$ we have that $\gamma^{\prime}$ satisfies the claim, because $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]\geq\epsilon$.
If $c<\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]$, assume that for all $\gamma$ such that $c\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]$ we have that $c/2\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]$. Let $\gamma_{0}=\sup\{\gamma:\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]\geq c\}$. Notice that this set is non-empty as $\gamma^{\prime}$ belongs in this set; hence $\gamma_{0}\geq\gamma^{\prime}>0$. We have that for all $\delta>0$ sufficiently small $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma_{0}-\delta]\geq c$ while $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma_{0}+\delta]\leq c/2$, which is a contradiction as $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]$ is continuous for all $\gamma>0$. Thus, there exists $\gamma$ such that $c\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]\geq c/2\geq(c/2)(\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0])$. ∎
Combining the the above two statements completes the proof of Lemma 2.7.
3 Learning Multi-Index Models with Agnostic Noise
In this section, we extend the algorithm of Section 2 for a general class of functions that satisfy specific regularity conditions (Definition 1.2), establishing Theorem 1.4. First, we state the formal versions of the conditions and theorem, and then provide pseudocode for the algorithm along with the necessary definitions.
As we will discuss later, we need a continuity condition on the function in order to be able to approximate it over small regions. To this end, we define the Gaussian Surface Area of a multiclass classifier. A bound on this quantity will be crucial for the analysis of our algorithm.
Definition 3.1 (Gaussian Surface Area).
For a Borel set $A\subseteq\mathbb{R}^{d}$, its Gaussian surface area is defined by $\Gamma(A):=\liminf_{\delta\to 0}\frac{\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{x}\in A_{\delta}\setminus A]}{\delta},$ where $A_{\delta}=\{\mathbf{x}:\mathrm{dist}(\mathbf{x},A)\leq\delta\}$. Let $\mathcal{Y}$ be a set of finite cardinality. For a multiclass function $f:\mathbb{R}^{d}\mapsto\mathcal{Y}$, we define its Gaussian surface area to be the total surface area of each of the label sets $K_{z}=\{\mathbf{x}\in\mathbb{R}^{d}:f(\mathbf{x})=z\}$, i.e., $\Gamma(f)=\sum_{z\in\mathcal{Y}}\Gamma(K_{z})$. For a class of multiclass concepts $\mathcal{C}$, we define $\Gamma(\mathcal{C}):=\sup_{f\in\mathcal{C}}\Gamma(f)$.
Now we proceed to define the regularity conditions that we will assume in order to get efficient algorithms. Formally, we define a class of well-behaved functions as follows:
Definition 3.2 (Well-Behaved Multi-Index Model).
Fix $d,K,m\in\mathbb{Z}_{+},\delta,\epsilon,\sigma,\tau\in(0,1)$ and $\mathcal{Y}\subseteq\mathbb{Z}$ of finite cardinality. We say that a class of functions $\mathcal{F}$ from $\mathbb{R}^{d}\mapsto\mathcal{Y}$ is a class of $(m,\epsilon,\delta,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models if, for every $f\in\mathcal{F}$, the following conditions hold:
  1. 1.
    There exists a subspace $W\subseteq\mathbb{R}^{d}$ of dimension at most $K$ such that $f$ depends only on the projection onto $W$, i.e., for every $\mathbf{x}\in\mathbb{R}^{d}$ it holds $f(\mathbf{x})=f(\mathbf{x}^{W})$.
  2. 2.
    The Gaussian surface area of $f$ is bounded by $\Gamma$, i.e., $\Gamma(f)\leq\Gamma$.
  3. 3.
    Let $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, for any random variable $y$ supported on $\mathcal{Y}$ with $\operatorname*{\mathbf{Pr}}_{\mathbf{x},y}[f(\mathbf{x})\neq y]\leq\epsilon$ and any subspace $V\subseteq\mathbb{R}^{d}$,
    1. (a)
      either there exists a function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x},y}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$.
    2. (b)
      or with probability at least $\delta$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, where $\mathbf{x}_{0}$ is independent of $\mathbf{x}$, there exists a polynomial $p:U\to\mathbb{R}$ of degree at most $m$ and a label $z\in\mathcal{Y}$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x}}[p(\mathbf{x}^{U})]=0$ and $\operatorname*{\mathbf{E}}_{\mathbf{x},y}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>\sigma\|p(\mathbf{x}^{U})\|_{2}$, where $U=(V+W)\cap V^{\perp}$.
We can now state the main theorem of this section.
Theorem 3.3 (Approximately Learning Multi-Index Models with Agnostic Noise).
Fix $d,K,m\in\mathbb{Z}_{+}$, $\zeta,\alpha,\tau,\sigma\in(0,1)$ and $\Gamma\in\mathbb{R}_{+}$. Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ and let $\epsilon,\delta\in(0,1)$. Let $\mathcal{F}$ a class of $(m,\zeta,\alpha,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models according to Definition 3.2 and let $\mathrm{OPT}=\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$. If $\zeta\geq\mathrm{OPT}+\epsilon$, then, Algorithm 3 draws $N={d}^{O(m)}2^{\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))}\log(1/\delta)$ i.i.d. samples from $D$, runs in time $\mathrm{poly}(d,N,\lvert\mathcal{Y}\rvert)$, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon\;.$
Definition 3.2 outlines the crucial properties necessary for our algorithm to produce a satisfactory solution. We first assume that our concept class consists of Multi-Index Models with a bounded Gaussian surface area. This condition bounds the complexity of the decision boundary of the multi-index models or equivalently it bounds how fast the function changes labels. We use this condition in order to show that if we partition our space to sufficiently small cubes then the function should be almost constant. This is a condition that holds for many classes of interest, such as multiclass linear classifiers and intersections of halfspaces.
In addition, we impose a structural condition on the class to ensure progress at each step of the algorithm. Intuitively, if $V$ is the subspace we have currently identified, we require that either it already provides a good approximation Condition 3(a) or we can use moments to uncover further directions in the hidden subspace $W$ Condition 3(b). Specifically, we assume that for any perturbation $y$ of a function $f$ from the class that we are attempting to learn, and for any given subspace $V$, one of the following must hold: (i) there exists a good approximation to $y$ using a function defined over $V$ (Condition 3(a) of Definition 3.2), or (ii) with some probability over orthogonal cylinders to $V$, there exists a low-degree, zero-mean, variance-one polynomial along directions that correlate with the ground-truth subspace $W$, which non-trivially correlates with the level sets of $y$ over that orthogonal cylinder (Condition 3(b) of Definition 3.2). In this manner, extracting these directions and adding them to $V$ enables an improved approximation of the hidden-subspace $W$. Although Condition 3(b) is defined over specific points with a measure of zero, we demonstrate in Lemma 3.5 that, by leveraging our assumptions, there exists a $\delta$ fraction of finite width cylinders that exhibit this property.
Condition 3(a) of Definition 3.2 further provides an approximation factor that the optimal function within the subspace $V$ can achieve. This function can be approximated using a non-parametric regression algorithm. In fact, in Lemma 3.4, we formally show that a piece-wise constant approximation is enough to approximate the unknown function $g$.
We prove this by leveraging the Gaussian noise sensitivity of $f$ to show that a noisy version of $g$ is nearly constant on sufficiently thin cubes defined over $V$. This way using the noisy version of $g$ we can construct a good enough piecewise constant approximation. Consequently, the function that minimizes the disagreement over these cubes will also exhibit low error.
Input: Accuracy $\epsilon\in(0,1)$, failure probability $\delta\in(0,1)$, and sample access to a distribution $D$ over $\mathbb{R}^{d}\times\mathcal{Y}$.
Output: A hypothesis $h$ such that, with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$.
  1. 1.
    Let $T,p$ be sufficiently large polynomials in $1/\alpha,1/\epsilon,1/\sigma,K,\Gamma,\lvert\mathcal{Y}\rvert,m$, and let $C$ be a sufficiently large universal constant..
  2. 2.
    Initialize $L_{1}=\emptyset$, $t\leftarrow 1$, $\epsilon^{\prime}\leftarrow 1/p$, $N\leftarrow{d}^{Cm}2^{T^{C}}\log(1/\delta)$.
  3. 3.
    While $t\leq T$
    1. (a)
      Draw a set $S$ of $N$ i.i.d. samples from $D$.
    2. (b)
      $\mathcal{E}_{t}\leftarrow$ Algorithm 4($\epsilon,\epsilon^{\prime},\delta,\mathrm{span}(L_{t}),S$).
    3. (c)
      $L_{t+1}\leftarrow L_{t}\cup\mathcal{E}_{t}$.
    4. (d)
      $t\leftarrow t+1$.
  4. 4.
    Construct an $\epsilon^{\prime}$-approximating partition, $\mathcal{S}$, with respect to $\mathrm{span}(L_{t})$, as defined in Definition 2.2.
  5. 5.
    Return a piecewise constant approximating classifier $h_{\mathcal{S}}$, as defined in Definition 2.3, using $N$ i.i.d. samples from $D$, by finding the most frequent label at any subset of $\mathcal{S}$.
Algorithm 3 Learning Multi-Index Models with Agnostic Noise
Input: $\epsilon>0$, $\epsilon^{\prime}>0$, $\delta>0$, a subspace $V$ of $\mathbb{R}^{d}$ and a set of samples $\{(\mathbf{x}^{(1)},y_{1}),\dots,(\mathbf{x}^{(N)},y_{N})\},\mathbf{x}^{(i)}\in\mathbb{R}^{d},y_{i}\in\mathcal{Y}$.
Output: A set of unit vectors $\mathcal{E}$.
  1. 1.
    Let $t,\eta$ be sufficiently small polynomials in $\epsilon,\alpha,\sigma,1/K,1/\Gamma,1/m,1/\lvert\mathcal{Y}\rvert$.
  2. 2.
    Construct the empirical distribution $\widehat{D}$ of $\{(\mathbf{x}^{(1)},y_{1}),\dots,(\mathbf{x}^{(N)},y_{N})\}$.
  3. 3.
    Construct $\mathcal{S}$, an $\epsilon^{\prime}$-approximating partition with respect to $V$ (Definition 2.2).
  4. 4.
    For each $S\in\mathcal{S}$ and each $i\in\mathcal{Y}$, find a polynomial $p_{S,i}(\mathbf{x})$ such that
    $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[$ $(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}^{\perp V}))^{2}\mid\mathbf{x}\in S]$
    $\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}^{\perp V}))^{2}\mid\mathbf{x}\in S]+\eta^{2}\;.$
  5. 5.
    For each $S\in\mathcal{S}$, compute $U_{S,i}$ as the set of unit eigenvectors of $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim D_{\mathbf{x}}}[\nabla p_{S,i}(\mathbf{x}^{\perp V})\nabla p_{S,i}(\mathbf{x}^{\perp V})^{\top}\mid\mathbf{x}\in S]$ with eigenvalues at least $\sigma^{2}/(4K)$ and let $U_{S}=\cup_{i\in\mathcal{Y}}U_{S,i}$.
  6. 6.
    Let $\widehat{\mathbf{U}}=\sum_{S\in\mathcal{S},\mathbf{u}\in U_{S}}(\mathbf{u})^{\perp V}((\mathbf{u})^{\perp V})^{\top}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim{\widehat{D}}}[S]$.
  7. 7.
    Return the set $\mathcal{E}$ of unit eigenvectors of $\widehat{\mathbf{U}}$ with corresponding eigenvalues at least $t$.
Algorithm 4 Estimating a relevant direction
Lemma 3.4 (Piecewise Constant Approximation Suffices).
Let $d\in\mathbb{Z}_{+},\epsilon,\epsilon^{\prime}\in(0,1)$ and let $\mathcal{Y}$ be a set of finite cardinality. Let $D$ be a distribution supported on $\mathbb{R}^{d}\times\mathcal{Y}$, whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, and let $V\subseteq\mathbb{R}^{d}$ be a $k$-dimensional subspace, with $k\geq 1$. Suppose that for a function $f\in\mathcal{F}$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\epsilon$. Additionally, assume that there exists a function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$. Let $\eta=\epsilon^{\prime 2}/(C\Gamma(f)\lvert\mathcal{Y}\rvert\sqrt{k})$, for $C>0$ a sufficiently large universal constant, and $\mathcal{S}$ be an $\eta$-piecewise approximating partition with respect to $V$, as defined in Definition 2.2. Then, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]\leq\tau+\epsilon+\epsilon^{\prime}$, where $h_{\mathcal{S}}$ is an $\eta$-piecewise constant approximation function as defined in Definition 2.3.
Proof.
Fix $i\in\mathcal{Y}$ and consider the boolean functions $f_{i}(\mathbf{x})\coloneqq\mathds{1}(f(\mathbf{x})=i)$, $g_{i}(\mathbf{x})\coloneqq\mathds{1}(g(\mathbf{x}^{V})=i)$.
By $T_{\rho}$ we denote the Gaussian noise operator defined in Definition A.3. By A.4, we have that $\left\|f_{i}-g_{i}\right\|_{L_{1}}\geq\left\|T_{\rho}f_{i}-T_{\rho}g_{i}\right\|_{L_{1}}$. Hence, from the triangle inequality we also have that $\left\|f_{i}-g_{i}\right\|_{L_{1}}+\left\|T_{\rho}f_{i}-f_{i}\right\|_{L_{1}}\geq\left\|f_{i}-T_{\rho}g_{i}\right\|_{L_{1}}$. Furthermore, by A.5, we have that $\left\|T_{\rho}f_{i}-f_{i}\right\|_{L_{1}}\leq\sqrt{\pi}\Gamma(f_{i})\rho$. Therefore, we have that
$\left\|f_{i}-T_{\rho}g_{i}\right\|_{L_{1}}\leq\left\|f_{i}-g_{i}\right\|_{L_{1}}+\sqrt{\pi}\Gamma(f_{i})\rho\;.$
Moreover, note that for any continuously differentiable function $F:\mathbb{R}^{d}\mapsto\mathbb{R}$, it holds that $F(\mathbf{a})-F(\mathbf{b})=\int_{0}^{1}\nabla F(\mathbf{b}+t(\mathbf{a}-\mathbf{b}))\cdot(\mathbf{a}-\mathbf{b})\mathrm{d}t$ and from the Cauchy-Schwarz inequality, we have that $|F(\mathbf{a})-F(\mathbf{b})|\leq\max_{t\in(0,1)}\|\nabla F(\mathbf{b}+t(\mathbf{a}-\mathbf{b}))\|\|\mathbf{a}-\mathbf{b}\|$. By taking $F=T_{\rho}g_{i}$ and using A.4, we get that
$|T_{\rho}g_{i}(\mathbf{x})-T_{\rho}g_{i}(\mathbf{y})|\leq\frac{\|\mathbf{x}-\mathbf{y}\|}{\rho}\;,$
for any $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$. For each $S\in\mathcal{S}$, pick a point $\mathbf{x}_{S}\in S$. By averaging over the partition, we have that
$\sum_{S\in\mathcal{S}}\left\|(T_{\rho}g_{i}(\mathbf{x}_{S})-T_{\rho}g_{i})\mathds{1}(\mathbf{x}\in S)\right\|_{L_{1}}\leq\sum_{S\in\mathcal{S}}\frac{\sqrt{k}\eta}{\rho}\operatorname*{\mathbf{Pr}}[S]\leq\frac{\sqrt{k}\eta}{\rho}\;.$
This in turn gives that
$\sum_{S\in\mathcal{S}}\left\|(T_{\rho}g_{i}(\mathbf{x}_{S})-f_{i})\mathds{1}(\mathbf{x}\in S)\right\|_{L_{1}}\leq\left\|f_{i}-g_{i}\right\|_{L_{1}}+\sqrt{\pi}\Gamma(f_{i})\rho+\frac{\sqrt{k}\eta}{\rho}\;.$
Furthermore, observe that by the linearity of expectation and the fact that $\sum_{i\in\mathcal{Y}}g_{i}(\mathbf{x})=1$ we have that $\sum_{i\in\mathcal{Y}}T_{\rho}g_{i}(\mathbf{x})=1$ for $\mathbf{x}\in\mathbb{R}^{d}$. Moreover, $T_{\rho}g_{i}(\mathbf{x})\in[0,1]$ for all $\mathbf{x}\in\mathbb{R}^{d},i\in\mathcal{Y}$. Therefore, we can consider the multiclass function $G:\mathbb{R}^{d}\to\mathcal{Y}$ such that $G(\mathbf{x})$ is a random label sampled from the probability distribution induced by $T_{\rho}g_{i}(\mathbf{x}_{S})$, where $S\in\mathcal{S}$ such that $\mathbf{x}\in S$. Note that for every probability vector $\mathbf{p}\in\mathbb{R}^{|\mathcal{Y}|}$ and label $z\in\mathcal{Y}$ we have that
$\|\mathbf{p}-\mathbf{e}_{z}\|_{1}=\lvert\mathbf{p}_{z}-1\rvert+\sum_{i\neq z,i\in\mathcal{Y}}\lvert\mathbf{p}_{i}\rvert=2\operatorname*{\mathbf{Pr}}_{i\sim\mathbf{p}}[i\neq z]\;.$
Denote by $A=\bigcup_{S\in\mathcal{S}}S$ the set of points that lie inside the regions of $\mathcal{S}$. Considering the disagreement between $G$ and $f$ in $A$, we obtain
$\operatorname*{\mathbf{Pr}}[G(\mathbf{x})\neq f(\mathbf{x}),\mathbf{x}\in A]$ $\leq\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}[G(\mathbf{x})\neq f(\mathbf{x}),\mathbf{x}\in S]=\sum_{i\in\mathcal{Y}}\sum_{S\in\mathcal{S}}\left\|(T_{\rho}g_{i}(\mathbf{x}_{S})-f_{i})\mathds{1}(\mathbf{x}\in S)\right\|_{L_{1}}/2$
$\leq\sum_{i\in\mathcal{Y}}\|f_{i}-g_{i}\|_{L_{1}}/2+\sqrt{\pi}\Gamma(f)\rho+\lvert\mathcal{Y}\rvert\sqrt{k}\eta/\rho\;.$
Noting that $\sum_{i\in\mathcal{Y}}\|f_{i}-g_{i}\|=2\tau$ set $\rho=\epsilon^{\prime}/(C{\Gamma(f)})$ and $\eta={\epsilon^{\prime}\rho}/{(C|\mathcal{Y}|\sqrt{k}})$, for a sufficiently large constant $C>0$. With these choices, we obtain $\operatorname*{\mathbf{Pr}}[G(\mathbf{x})\neq f(\mathbf{x}),\mathbf{x}\in A]\leq\tau+\epsilon^{\prime}/2$. By the triangle inequality we have that $\operatorname*{\mathbf{Pr}}[G(\mathbf{x})\neq y,\mathbf{x}\in A]]\leq\tau+\epsilon+\epsilon^{\prime}/2$. Since $h_{\mathcal{S}}$ chooses the label that minimizes the error at each $S\in\mathcal{S}$, we have that $\operatorname*{\mathbf{Pr}}[h_{\mathcal{S}}(\mathbf{x})\neq y,\mathbf{x}\in A]\leq\operatorname*{\mathbf{Pr}}[G(\mathbf{x})\neq y,\mathbf{x}\in A]]$. Finally, considering that $\operatorname*{\mathbf{Pr}}[\mathbf{x}\not\in A]\leq\eta$ completes the proof of Lemma 3.4.
In the following lemma, we show that, due to the bound on the Gaussian surface area of $f$, if there exist level sets of $f$ for which the conditional moments on a subspace $V$ are large with non-trivial probability over $V$, then the same property holds on sufficiently small cubes. That is, when $V$ is partitioned into sufficiently thin cubes, the conditional moments of the level sets of $f$ on these cubes will be large with non-trivial probability over the choice of cube. The intuition behind the proof is that, since the function is smooth (by the GSA bound), re-randomizing over sufficiently small cubes barely affects its output. However, by re-randomizing over cubes, the moments of the level-sets become independent of which point in the cube we condition on. As a result, if we observe distinguishing moments when conditioning on a point in the cube, then these moments persist when conditioning on the entire cube.
Lemma 3.5 (Existence of Large Moments on Small Cubes).
Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ and let $\mathcal{F}$ be a class of $(m,\epsilon,\delta,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models according to Definition 3.2. Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$, with $k\geq 1$, and let $\mathcal{S}$ be an $\eta$-approximating partition of $V$ according to Definition 2.2, with $\eta=\epsilon^{\prime 3}/(C\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2})$, for $C>0$ a sufficiently large universal constant.
If for a function $f\in\mathcal{F}$ with $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\epsilon+\epsilon^{\prime}$ there exist no function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$, then there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\delta$ such that for each $S\in\mathcal{T}$ there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x},y}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}\in S]>\sigma\|p(\mathbf{x}^{U})\|_{2}$, where $U=(V+W)\cap V^{\perp}$.
Proof.
Note that, since $\mathcal{F}$ satisfies Definition 3.2, with probability at least $\delta$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, independent of $\mathbf{x}$, there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\mathcal{D}}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>\sigma\|p(\mathbf{x}^{U})\|_{2}\;,$
where $U=(V+W)\cap V^{\perp}$.
In order to prove the statement, we need to show that the above condition can be extended from a single point to encompass an entire cube $S$. To achieve this, we introduce a random variable $y^{\prime}$, which is $\epsilon$-close to $f$, such that, when conditioned on any $S\subset\mathcal{S}$, $y^{\prime}$ is distributed uniformly within $S$ along the directions defined by $V$, while the other directions remain consistent with those of $y$.
For any $\mathbf{x}\in\mathbb{R}^{d}$, denote by $\mathcal{S}_{\mathbf{x}}$ the set $S\in\mathcal{S}$ such $\mathbf{x}\in S$. We define a random vector $\mathbf{x}^{\prime}$ such that $\mathbf{x}^{\prime\perp V}=\mathbf{x}^{\perp V}$ and $\mathbf{x}^{\prime V}$ is the sampled from the standard Gaussian over $V$ conditioned on $\mathcal{S}_{\mathbf{x}}$, i.e., $\mathcal{N}(\mathbf{0},\mathbf{I})\mid\mathcal{S}_{\mathbf{x}}$. Note that $\mathbf{x}^{\prime}$ also follows the standard normal distribution over $\mathbb{R}^{d}$ because its $V$-component is resampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$.
In the following claim we will prove that $f(\mathbf{x})$ and $f(\mathbf{x}^{\prime})$ are approximately the same.
Claim 3.6.
It holds that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f(\mathbf{x}^{\prime})]=O(\epsilon^{\prime})$.
Proof.
Define the boolean function $f_{i}(\mathbf{x})\coloneqq\mathds{1}(f(\mathbf{x})=i)$ and denote by $g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})\coloneqq T_{\rho}f_{i}(\mathbf{x}_{0}^{\perp V}+\mathbf{x}^{V})$, where $T_{\rho}$ denotes the Gaussian noise operator defined in Definition A.3. Fix $\mathbf{x}_{0}\in\mathbb{R}^{d},S\in\mathcal{S}$ we first show that $\lvert g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})-g_{\mathbf{x}_{0}}^{(i)}(\mathbf{z})\rvert$ is small for $\mathbf{x}^{V},\mathbf{z}^{V}\in S$.
Note that for any continuously differentiable function $F:V\mapsto\mathbb{R},\mathbf{a},\mathbf{b}\in V$, it holds that $F(\mathbf{a})-F(\mathbf{b})=\int_{0}^{1}\nabla F(\mathbf{b}+t(\mathbf{a}-\mathbf{b}))\cdot(\mathbf{a}-\mathbf{b})\mathrm{d}t$ and from Cauchy-Schwarz inequality, we have that $|F(\mathbf{a})-F(\mathbf{b})|\leq\max_{t\in(0,1)}\|\nabla F(\mathbf{b}+t(\mathbf{a}-\mathbf{b}))\|\|\mathbf{a}-\mathbf{b}\|$. By taking $F(\mathbf{x})=g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})$ and using A.4, we get that
$|g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})-g_{\mathbf{x}_{0}}^{(i)}(\mathbf{z})|\leq\frac{\eta\sqrt{k}}{\rho}\;,$
for any $\mathbf{x}^{V},\mathbf{\mathbf{z}}^{V}\in S$.
Consequently by the triangle inequality and Claim A.7 we have
$\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{E}}_{\mathbf{x}_{0},\mathbf{x}^{V}}[\lvert(f_{i}(\mathbf{x}_{0}^{\perp V}+\mathbf{x}^{V})g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x}))\mathds{1}(\mathbf{x}\in S)\rvert]$
$\leq\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{E}}_{\mathbf{x}_{0},\mathbf{x}^{V}}[\lvert(f_{i}(\mathbf{x}_{0}^{\perp V}+\mathbf{x}^{V})-g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x}_{S}))\mathds{1}(\mathbf{x}\in S)\rvert]+\eta/\rho\lesssim\sqrt{\rho}\Gamma(f_{i})+\eta\sqrt{k}/\rho\;,$
for any collection of points $\mathbf{x}_{S}\in S$. Furthermore, observe that by the linearity of expectation and the fact that $\sum_{i\in\mathcal{Y}}f_{i}(\mathbf{x})=1$ we have that $\sum_{i\in\mathcal{Y}}g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})=1$ and $g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})\in[0,1]$ for all $\mathbf{x},\mathbf{x}_{0}\in\mathbb{R}^{d},i\in\mathcal{Y}$. Therefore, we can consider the multiclass function $G:\mathbb{R}^{d}\to\mathcal{Y}$ such that $G(\mathbf{x})$ is a random label sampled from the probability distribution induced by $g_{\mathbf{x}}^{(i)}(\mathbf{x}_{S})$, where $S\in\mathcal{S}$ such that $\mathbf{x}\in S$. Note that for every probability vector $\mathbf{p}\in\mathbb{R}^{|\mathcal{Y}|}$ and label $z\in\mathcal{Y}$ we have that
$\|\mathbf{p}-\mathbf{e}_{z}\|_{1}=\lvert\mathbf{p}_{z}-1\rvert+\sum_{i\neq z,i\in\mathcal{Y}}\lvert\mathbf{p}_{i}\rvert=2\operatorname*{\mathbf{Pr}}_{i\sim\mathbf{p}}[i\neq z]\;.$
Therefore, we have that the $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq G(\mathbf{x})]\lesssim\sqrt{\rho}\Gamma(f)+\eta\sqrt{k}\lvert\mathcal{Y}\rvert/\rho\lesssim\epsilon^{\prime}$.
Notice that since $\mathbf{x}^{\prime}$ is distributed as $\mathcal{N}(\mathbf{0},\mathbf{I})$, we have that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq G(\mathbf{x}^{\prime})]=O(\epsilon^{\prime})$. However, as $G(\mathbf{x})$ is independent $\mathbf{x}^{V}$, we have that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq G(\mathbf{x}^{\prime})]=\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq G(\mathbf{x})]=O(\epsilon^{\prime})$. Therefore, by the triangle inequality we have that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f(\mathbf{x}^{\prime})]=O(\epsilon^{\prime})$, which completes the proof of Claim 3.6.∎
Define $y^{\prime}$ such that or all $\mathbf{z}\in\mathbb{R}^{d}$, the distribution of $y^{\prime}$ given $\mathbf{x}^{\prime}=\mathbf{z}$ is the same as the distribution of $y$ given $\mathbf{x}=\mathbf{z}$. By the definition of $y^{\prime}$, it follows that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}^{\prime},y^{\prime}}[f(\mathbf{x}^{\prime})\neq y^{\prime}]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}^{\prime}}\left[\operatorname*{\mathbf{E}}_{y^{\prime}}[f(\mathbf{x}^{\prime})\neq y^{\prime}\mid\mathbf{x}^{\prime}]\right]\leq\epsilon$. Therefore, it holds that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y^{\prime}]\leq\epsilon+O(\epsilon^{\prime})$.
Therefore, with probability at least $\delta$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ independent of $\mathbf{x}$, there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that
$\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>\sigma\|p(\mathbf{x}^{U})\|_{2}\;.$
However, since the moments of $y^{\prime}$ are identical when conditioned on each $\mathbf{x}_{0}^{V}$ that belongs to the same region in $S$, we deduce that if the moment condition holds for some polynomial at a particular point $\mathbf{x}_{0}^{V}$, then it must hold for the entire cube $S$, such that $\mathbf{x}_{0}^{V}\in S$. Therefore, there exists a subset $\mathcal{T}\subseteq\mathcal{S}$ of the partition with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\delta$ satisfying the following: for all $S\in\mathcal{T}$ there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}\in S]>\sigma\|p(\mathbf{x}^{U})\|_{2}\;.$
Noticing that the moments with respect to $U$ on each $S\in\mathcal{S}$ are the same for $y$ and $y^{\prime}$, completes the proof of Lemma 3.5. ∎
The remaining part of the section is structured as follows: In Section 3.1, we provide the analysis of Algorithm 4 with Proposition 3.7. In Section 3.2, we analyze Algorithm 3 providing the full proof of Theorem 3.3, which is the main result of this section.
3.1 Finding a Relevant Direction
In this section, we prove Proposition 3.7, which states that if the class satisfies Definition 3.2 and the current approximation of the ground-truth subspace provides a suboptimal piecewise approximation, then Algorithm 4 outputs a small set of vectors, at least one of which has a non-trivial projection onto the ground-truth subspace.
Proposition 3.7.
Let $d,k\in\mathbb{Z}_{+}$, $\epsilon,\delta\in(0,1)$ and $\epsilon^{\prime}{\leq}c\epsilon^{3}/(\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2}))$, for $c$ a sufficiently small universal constant. Let $D$ be a distribution supported on $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, let $V\subseteq\mathbb{R}^{d}$ be a $k$-dimensional subspace, and let $\mathcal{F}$ be a class of $({m,\zeta+O(\epsilon)},\alpha,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models according to Definition 3.2. Let $f\in\mathcal{F}$ with $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]={\mathrm{OPT}}$ and let $W$ be a subspace such that $f(\mathbf{x})=f(\mathbf{x}^{W})$ for all $\mathbf{x}\in\mathbb{R}^{d}$, with $K=\dim(W)$. Additionally, let $h_{\mathcal{S}}$ be a piecewise constant approximation of $D$, with respect to an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$, according $V$, as defined in Definitions 2.2 and 2.3.
There exists $N=(dm)^{O(m)}{(k/\epsilon^{\prime})^{O(k)}}\log(1/\delta){(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{O(1)}}$ such that, If $\zeta\geq\mathrm{OPT}$ and $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>{\tau+\mathrm{OPT}+3\epsilon}$, then Algorithm 4 given $N$ i.i.d. samples from $D$, runs in $\mathrm{poly}(N,d,\lvert\mathcal{Y}\rvert)$ time, and with probability at least $1-\delta$, returns a list of unit vectors $\mathcal{E}$ of size $|\mathcal{E}|=\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))$, such that: For some $\mathbf{v}\in\mathcal{E}$ and unit vector $\mathbf{w}\in W$ it holds that
$\mathbf{w}^{\perp V}\cdot\mathbf{v}\geq\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))\;.$
Before proving Proposition 3.7, we need to show that if we have learned a direction of our target model to accuracy $\epsilon$, then we only observe moments that depend on the remaining directions. The intuition behind the proof is that re-randomizing over directions that have negligible projection to $W$ barely affects the function’s output. Hence, the function is close to a mixing $y^{\prime}$ of the noisy version $y$ that has these components re-randomized. As a result, if the function is far from a projection onto $V$, we have large moments for level-sets of $y^{\prime}$. Consequently, the re-randomized directions are averaged out, and we are left with distinguishing moments on the remaining directions. The proof is similar in structure to that of Lemma 3.5, the main difference being in what is being re-randomized.
Lemma 3.8.
Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, let $\mathcal{F}$ be a class of $(m,\epsilon+O(\epsilon^{\prime}),\delta,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models according to Definition 3.2. Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$, with $k\geq 1$, and let $\mathcal{S}$ be an $\eta$-approximating partition of $V$ according to Definition 2.2, with $\eta=\epsilon^{\prime 3}/(C\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2})$, for $C>0$ a sufficiently large universal constant. Let $f\in\mathcal{F}$ be a function that depends only on a $K$-dimensional subspace $W$, and let $y$ be random variable over $\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\epsilon$.
Suppose that there exists no function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$. Then, there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\delta$ such that for each $S\in\mathcal{T}$ there exists a zero-mean, variance-one polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x},y}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}\in S]>\sigma$, where $U{=}(V{+}W)\cap V^{\perp}$. Moreover, for each such polynomial $p$, we have that $\nabla p(\mathbf{x}^{U})\cdot\mathbf{v}=0$ for any $\mathbf{v}\in E$ and $\mathbf{x}\in\mathbb{R}^{d}$, where $E\subseteq{V^{\perp}}$ is a subspace satisfying $\|\mathbf{v}^{W}\|\leq\epsilon^{\prime}/\Gamma(f)$ for every unit vector $\mathbf{v}\in E$.
Proof.
Let $\mathbf{x},\mathbf{x}^{\prime}$ be random variables defined as $\mathbf{x}=\mathbf{x}^{\perp E}+\mathbf{x}^{E}$ and $\mathbf{x}^{\prime}=\mathbf{x}^{\perp E}+{\mathbf{z}}^{E}$, where $\mathbf{z}$ is an independent standard Gaussian vector. Denote by $\mathcal{S}$ an $\eta$-approximating partition with respect to $V$. Notice that for any unit direction $\mathbf{w}\in W$, both $\mathbf{w}\cdot\mathbf{x}$ and $\mathbf{w}\cdot\mathbf{x}^{\prime}$ are distributed according to a standard Gaussian. Furthermore, we have that by independence
$\operatorname*{\mathbf{E}}[(\mathbf{w}\cdot\mathbf{x})(\mathbf{w}\cdot\mathbf{x}^{\prime})]$ $=\operatorname*{\mathbf{E}}[(\mathbf{w}\cdot\mathbf{x}^{\perp E})^{2}]=\|\mathbf{w}^{\perp E}\|^{2}\;.$
Given that $E$ is the subspace of small projections, i.e., $\left\|\mathbf{w}^{E}\right\|\leq\epsilon^{\prime}/\Gamma(f)$; it follows that $\operatorname*{\mathbf{E}}[(\mathbf{w}\cdot\mathbf{x}^{\perp E})^{2}]\geq 1-(\epsilon^{\prime}/\Gamma(f))^{2}$. As a result, $\mathbf{x}$ and $\mathbf{x}^{\prime}$ are at least $1-(\epsilon^{\prime}/\Gamma(f))^{2}$ correlated in $W$. Thus by the Gaussian noise sensitivity bound A.5, we have that
$\operatorname*{\mathbf{Pr}}_{\mathbf{x},\mathbf{x}^{\prime}}[f(\mathbf{x})\neq f(\mathbf{x}^{\prime})]\leq O(\epsilon^{\prime})\;.$
Consider the random variable $y^{\prime}$ supported on $\mathcal{Y}$ that is distributed like $y$ for $\mathbf{x}^{\prime}$, i.e., $D(y^{\prime}=i\mid\mathbf{x}^{\prime}=\mathbf{z})=D(y=i\mid\mathbf{x}=\mathbf{z})$, for all $\mathbf{z}\in\mathbb{R}^{d},i\in\mathcal{Y}$. We have that
$\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y^{\prime}]\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq y^{\prime}]+O(\epsilon^{\prime})\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]+O(\epsilon^{\prime})\leq\epsilon+O(\epsilon^{\prime})\;.$
As a result, by Lemma 3.5, we have that there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\delta$ such that for each $S\in\mathcal{T}$ there exists a zero-mean, variance-one polynomial $p:U\to\mathbb{R}$ of degree at most $m$ along with a label $z\in\mathcal{Y}$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}\in S]>\sigma$, where $U=(V+W)\cap V^{\perp}$. However we have that for all $S\in\mathcal{T}$
$\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}\in S]=\operatorname*{\mathbf{E}}[\operatorname*{\mathbf{E}}_{\mathbf{x}^{E}}[p(\mathbf{x}^{U})\mid\mathbf{x}^{\perp E}]\operatorname*{\mathbf{E}}_{\mathbf{x}^{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp E}]\mid\mathbf{x}\in S]]\;.$
Notice that $p^{\prime}(\mathbf{x}^{U})\coloneqq\operatorname*{\mathbf{E}}_{\mathbf{x}^{E}}[p(\mathbf{x}^{U})\mid\mathbf{x}^{\perp E}]$ is a mean-zero polynomial of degree at most $m$ with variance at most one by Jensen’s inequality. Furthermore, as $p^{\prime}$ is independent of $\mathbf{x}^{E}$ we have that $\nabla p^{\prime}(\mathbf{x}^{U})\cdot\mathbf{v}=0$ for any $\mathbf{v}\in E$ and $\mathbf{x}\in\mathbb{R}^{d}$. ∎
We now proceed with the proof of Proposition 3.7.
Proof of Proposition 3.7.
Denote by $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}$ an orthonormal basis of the subspace $W$. Also, let $\widehat{\mathbf{U}}$ be the matrix computed at Line 6 of Algorithm 4 and let $\eta^{2}$ be the $l_{2}^{2}$ error chosen in the polynomial-regression step, i.e., Line 4 of Algorithm 4.
First note that as $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>{\tau+\mathrm{OPT}+3\epsilon}$ we have that form Lemma 3.4 for every function $g:V\to\mathcal{Y}$ it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\geq\tau$. Thus from Lemma 3.8 we have that there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\alpha$, such that for each $S\in\mathcal{T}$ there exists a $z\in\mathcal{Y}$ and a zero-mean variance-one polynomial $q_{S}:U\to\mathbb{R}$ of degree at most $m$ such that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[q_{S}(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}\in S]>\sigma$, where $U=(V+W)\cap(V^{\perp})$. Furthermore, define $E_{S}\coloneqq\mathrm{span}(\{\mathbf{v}\in U_{S}:\left\|\mathbf{v}^{W}\right\|\leq\epsilon/(\sqrt{\lvert U_{S}\rvert}\Gamma)\})$. Then, for every unit vector $\mathbf{v}\in E_{S}$, we have $\nabla q_{S}(\mathbf{x}^{U})\cdot\mathbf{v}=0$ since $\left\|\mathbf{v}^{W}\right\|\leq\epsilon/\Gamma$ holds for every unit vector $\mathbf{v}\in E_{S}$.
In the following claim, we leverage the existence of non-trivial moments over $U$ for regions $S\in\mathcal{T}$, to show that if the sample size is sufficiently large then for each region $S\in\mathcal{T}$ the associated set, $U_{S}$ (see Line 5 of Algorithm 4) contains a vector that correlates with $W$.
Claim 3.9 (Existence of Correlating Vectors).
Fix, $S\in\mathcal{T}$, if the number of samples that fall in $S$ is $N_{S}\geq(dm)^{Cm}\log(\lvert\mathcal{Y}\rvert/\delta)/{\eta}^{C}$, for a sufficiently large universal constant $C>0$, then with probability at least $1-\delta$ there exists a unit vector $\mathbf{v}\in U_{S}$ such that
$(\mathbf{w}^{(i)})^{\perp V}\cdot\mathbf{v}\geq{\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)}\;,$
for some $i\in[K]$. Moreover, it holds that $\lvert U_{S}\rvert=\mathrm{poly}(mK\lvert\mathcal{Y}\rvert/\sigma)$.
Proof of Claim 3.9.
Denote by $p_{S,i}$ for $S\in\mathcal{S},i\in\mathcal{Y}$ the polynomials of degree at most $m$ computed at Line 4 of Algorithm 4. First, notice that since whether or not $\mathbf{x}\in S$ depends only in $\mathbf{x}^{V}$, we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}^{\perp V}))^{2}\mid\mathbf{x}\in S]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}^{\perp V}))^{2}\mid\mathbf{x}\in S]+\eta^{2}$ is equivalent to $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}))^{2}]+\eta^{2}$, where $D_{V^{\perp}}$ is defined to be the marginal obtained by averaging $D$ over $V$ conditioned on $S$, i.e., $D_{V^{\perp}}(\mathbf{x}^{\perp V},y)=\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[D(\mathbf{x},y)\mid\mathbf{x}\in S]$. Moreover, by the properties of the Gaussian distribution, we have that the $\mathbf{x}$-marginal of $D_{V^{\perp}}$ is a standard Gaussian. Therefore, if $N_{S}\geq(dm)^{Cm}\log(\lvert\mathcal{Y}\rvert/\delta)/{\eta^{C}}$ for a sufficiently large universal constant $C>0$, then by the union bound and B.2, we have that with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}))^{2}]+\eta^{2}$. Moreover, by the orthogonality of Hermite polynomials we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]=\sum_{\beta\subseteq\mathbb{N}^{d}}(\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x})\mathds{1}(y=i)]-\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x})p_{S,i}(\mathbf{x})])^{2}$, it holds that $\sum_{t=1}^{m}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)^{[t]}-p_{S,i}^{[t]}(\mathbf{x}))^{2}]\leq\eta^{2}$.
From the rotational invariance of the Gaussian without loss of generality, we can denote by $\mathbf{e}_{1},\dots,\mathbf{e}_{k^{\prime}}$ an orthonormal basis of the space of $U$ projected onto $E_{S}^{\perp}$. We have that $k^{\prime}\geq 1$ otherwise we would not have the existence of $q_{S}$ that is defined over this space. Furthermore, since $U$ is a subset of $V^{\perp}$ we similarly have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[q_{S}(\mathbf{x}^{U})\mathds{1}(y=z)]>\sigma$ for some $z\in\mathcal{Y}$. Decomposing $q_{S}(\mathbf{x}^{U})$ to the basis of Hermite polynomials we have that $q_{S}(\mathbf{x}^{U})=\sum_{\beta\in\mathbb{N}^{d^{\prime}},1\leq\left\|\beta\right\|_{1}\leq m}\hat{q}_{S}(\beta)H_{\beta}(\mathbf{x}^{U})$ where $d^{\prime}=\dim(U)$. Moreover, note that since $q_{S}$ has no component outside $\mathbf{e}_{1},\dots,\mathbf{e}_{k^{\prime}}$ we have that $q_{S}(\mathbf{x}^{U})=\sum_{\beta\in J}\hat{q}_{S}(\beta)H_{\beta}(\mathbf{x}^{U})$, where $J$ denotes the set of $\beta\in\mathbb{N}^{d^{\prime}}$ such that $\beta_{i}\geq 1$ for some $i\in[k^{\prime}]$. Considering the correlation of $q_{S}$ with the label region $z$ we have that
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[q_{S}(\mathbf{x}^{U})\mathds{1}(y=z)]$ $=\sum_{\beta\in J}\hat{q}_{S}(\beta)\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x}^{U})\mathds{1}(y=z)]$
$\leq\left(\sum_{\beta\in J}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x}^{U})\mathds{1}(y=z)]^{2}\right)^{1/2}\;,$
where we used the fact that $\|q_{S}\|_{2}=1$ and the Cauchy-Schwarz inequality. Hence, we have that the sum of the squares of the Hermite coefficients for of the degree $m$ restriction the random variable $\mathds{1}(y=z)$ is at least $\sigma^{2}$.
Now evaluating the quadratic form of $\mathbf{M}=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim D_{\mathbf{x}}}[\nabla p_{S,z}(\mathbf{x}^{\perp V})\nabla p_{S,z}(\mathbf{x}^{\perp V})^{\top}\mid\mathbf{x}\in S]$ (Line 5 of Algorithm 4) using A.2 gives us
$\sum_{i=1}^{k^{\prime}}\mathbf{e}_{i}^{\top}\mathbf{M}\mathbf{e}_{i}$ $=\sum_{i=1}^{k^{\prime}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\nabla p_{S,z}(\mathbf{x})\cdot\mathbf{e}_{i})^{2}]\geq\sum_{\beta\in\mathbb{N}^{d}}(\sum_{i=1}^{k^{\prime}}\beta_{i})(\hat{p}_{S,z}(\beta))^{2}$
$\geq\sum_{\beta\in J}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x}^{U})\mathds{1}(y=z)]^{2}-2\eta$
$\geq\sigma^{2}-2\eta\geq\sigma^{2}/2$
for $\eta\leq\sigma/2$. Thus, since $\dim(U)\leq K$ we have that $\mathbf{e}_{j}^{\top}\mathbf{M}\mathbf{e}_{j}\geq\sigma^{2}/(2K)$ for some $j\in[k^{\prime}]$.
Since $\mathbf{M}$ is symmetric and therefore diagonalizable, we can denote by $\mathbf{u}^{(i)}$ the unit eigenvectors of $\mathbf{M}$ and by $\lambda_{i}$ the corresponding eigenvalues. Additionally, let $U_{S,z}$ denote the set of eigenvectors of $\mathbf{M}$ with eigenvalues greater than ${\sigma^{2}}/{4K}$ (see Line 5 of Algorithm 4). Decomposing $\mathbf{e}_{j}$ to the basis of eigenvectors gives us
$\frac{\sigma^{2}}{4K}+\sum_{\mathbf{u}^{(i)}\in U_{S,z}}\lambda_{i}(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\sum_{i=1}^{d-k}\lambda_{i}(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\frac{\sigma^{2}}{2K}\;.$
Hence, there exists an eigenvector $\mathbf{u}^{(i)}\in U_{S,z}$ such that $\lambda_{i}(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\sigma^{2}/(4K\lvert U_{S,z}\rvert)$ thus $(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\sigma^{2}/(4\lvert U_{S,z}\rvert K\text{tr}(\mathbf{M}))$ as $\mathbf{M}$ is positive definite. Therefore, by B.3 we have that $\lvert U_{S,z}\rvert=\mathrm{poly}(mK/\sigma),\text{tr}(\mathbf{M})=O(m)$ which implies that $\lvert\mathbf{e}_{j}\cdot\mathbf{u}^{(i)}\rvert\geq\mathrm{poly}(\sigma/(mK))$. Finally, note that since $\mathbf{e}_{j}\cdot\mathbf{u}^{(i)}\neq 0$, we have that $\mathbf{u}^{(i)}$ cannot belong in the space $E_{S}$. Thus, there exists a unit vector $\mathbf{w}\in W$ such that $\lvert\mathbf{u}^{(i)}\cdot\mathbf{w}\rvert\geq\epsilon/(\sqrt{\lvert U_{S}\rvert}\Gamma)\geq\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)$, which concludes the proof of Claim 3.9. ∎
First note that by the definition of an approximating partition (see Definition 2.2) and A.1, for all $S\in\mathcal{S}$ we have that $\operatorname*{\mathbf{Pr}}_{D}[S]{=(\epsilon^{\prime}/k)^{\Omega(k)}}$ and $\lvert\mathcal{S}\rvert{=(k/\epsilon^{\prime})^{O(k)}}$. Therefore, by union bound and Hoeffding’s inequality, it holds that if $N\geq{(k/\epsilon^{\prime})^{Ck}}\log(\lvert\mathcal{S}\rvert/\delta)={(k/\epsilon^{\prime})^{O(k)}}\log(1/\delta)$ for a sufficiently large constant $C>0$, then with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]-\operatorname*{\mathbf{Pr}}_{D}[S]\rvert\leq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S}$. Hence, it is true that $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$, i.e., the number of samples that fall in each set $S\in\mathcal{S}$ is at least $N\operatorname*{\mathbf{Pr}}_{D}[S]/2=N{(\epsilon^{\prime}/k)^{\Omega(k)}}$.
Therefore, for $N\geq(dm)^{Cm}{(k/\epsilon^{\prime})^{Ck}}\log(1/\delta){(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{C}}$ for a sufficiently large universal constant $C>0$, by applying Claim 3.9, we have that for all $S\in\mathcal{T}$ it holds that $\mathbf{v}\cdot(\mathbf{w}^{(i)})^{\perp V}\geq{\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)}$ for some $\mathbf{v}\in U_{S},i\in[K]$. Hence, we have that there exists a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ and $i\in[K]$ with $\sum_{S\in\mathcal{T^{\prime}}}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\sum_{S\in\mathcal{T^{\prime}}}\operatorname*{\mathbf{Pr}}_{D}[S]/2=\Omega(\alpha/K)$ such that for all $S\in\mathcal{T^{\prime}}$ we have that $\mathbf{v}\cdot(\mathbf{w}^{(i)})^{\perp V}\geq{\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)}$ for some $\mathbf{v}\in U_{S}$.
Denote by $\widehat{\mathbf{U}}$ the matrix computed at Line 6 of Algorithm 4. For some $i\in[K]$ we have that
$\left((\mathbf{w}^{(i)})^{\perp V}\right)^{\top}\widehat{\mathbf{U}}(\mathbf{w}^{(i)})^{\perp V}\gtrsim\frac{\alpha}{K}{\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)}\geq{\mathrm{poly}(\epsilon\alpha\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)}\;,$
where we used the fact that $\widehat{\mathbf{U}}$ is psd.
Moreover, as $\left\|\mathbf{v}\right\|=1$ for all $\mathbf{v}\in U_{S},S\in\mathcal{S}$, by the triangle inequality we can see that
$\left\|\widehat{\mathbf{U}}\right\|_{F}\leq\sum_{S\in\mathcal{S},\mathbf{v}\in U_{S}}\left\|\mathbf{v}\right\|\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\leq\sum_{S\in\mathcal{S},\mathbf{v}\in U_{S}}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\leq\mathrm{poly}(mK\lvert\mathcal{Y}\rvert/\sigma)\;,$
where we used the fact that $\lvert U_{S}\rvert=\mathrm{poly}(mK\lvert\mathcal{Y}\rvert/\sigma)$ by Claim 3.9.
Hence, by Claim B.1, we have that there exists a unit eigenvector $\mathbf{v}$ of $\widehat{\mathbf{U}}$ with eigenvalue at least ${\mathrm{poly}(\epsilon\alpha\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)}$ for a sufficiently small polynomial, such that $(\mathbf{w}^{(i)})^{\perp V}\cdot\mathbf{v}\geq\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))$ for some $i\in[K]$. Moreover, the number of such eigenvectors is at most $\mathrm{poly}(mK\lvert\mathcal{Y}\rvert\Gamma/(\epsilon\alpha\sigma))$, which completes the proof of Proposition 3.7. ∎
3.2 Proof of Theorem 3.3
In this section, given Proposition 3.7, we proceed to the proof of Theorem 3.3. Recall that, in Proposition 3.7, we have shown that if our current approximation to the hidden subspace is not accurate enough to produce a classifier that has sufficiently small error, then Algorithm 4 efficiently finds a direction that has non-trivial correlation to the hidden subspace. In the proof that follows, we iteratively apply this argument to show that, after a moderate number of iterations, Algorithm 3 outputs a classifier with sufficiently small error.
Proof of Theorem 3.3.
Denote by $f^{*}\in\mathcal{F}$ a classifier such that $\operatorname*{\mathbf{Pr}}[f^{*}(\mathbf{x})\neq y]\leq\mathrm{OPT}+\epsilon/4$. We show that Algorithm 3, with high probability, returns a hypothesis $h$ with 0-1 error at most $\tau+\mathrm{OPT}+\epsilon$. Let $W^{*}$ be a subspace of dimension $K$ such that $f^{*}(\mathbf{x})=f^{*}(\mathbf{x}^{W^{*}})$ and denote by $\mathbf{w}^{*(1)},\ldots,\mathbf{w}^{*(K)}\in\mathbb{R}^{d}$ an orthonormal basis of $W^{*}$. Let $L_{t}$ be the list of vectors updated by the algorithm (Line 3c of Algorithm 3) and $V_{t}=\mathrm{span}(L_{t})$, $\dim(V_{t})=k_{t}$. Also, let $\epsilon^{\prime}$ be the partition width parameter set at Line 2 of Algorithm 3, and for $t\in[T]$ let $\mathcal{S}_{t}$ be arbitrary $\epsilon^{\prime}$-approximating partitions with respect to $V_{t}$ (see Definition 2.2). Let $h_{t}:\mathbb{R}^{d}\to[K]$ be a piecewise constant classifiers, defined as $h_{t}=h_{\mathcal{S}_{t}}$ according to Definition 2.3 for the distribution $D$.
To prove the correctness of Algorithm 3, we need to show that if $h_{t}$ has significant error, then the algorithm improves its approximation $V_{t}$ to $W^{*}$. For quantifying the improvement at every step, we consider the following potential function
$\Phi_{t}=\sum_{i=1}^{K}\left\|(\mathbf{w}^{*(i)})^{\perp V_{t}}\right\|^{2}\;.$
Note that from Lines 1 and 3 of Algorithm 3, we perform at most $\mathrm{poly}(mK\Gamma{\lvert\mathcal{Y}\rvert}/(\epsilon\alpha\sigma))$ iterations. Furthermore, in each iteration, we update the vector set with at most $\mathrm{poly}(mK\Gamma{\lvert\mathcal{Y}\rvert}/(\epsilon\alpha\sigma))$ vectors. Hence, it follows that $k_{t}\leq\mathrm{poly}(mK\Gamma{\lvert\mathcal{Y}\rvert}/(\epsilon\alpha\sigma))$, for all $t=1,\dots,T$.
Assume that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+\epsilon/2$ for all $t=1,\dots,T$. Using the fact that $N={d}^{Cm}2^{(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))^{C}}\log(1/\delta)$ for a sufficiently large universal constant $C>0$ (Line 2 of Algorithm 3), we can apply Proposition 3.7 and conclude that, with probability $1-\delta$, there exists unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ for $t=[T]$ such that $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))$. Thus, by Claim 2.16, we have that with probability $1-\delta$, for all $t\in[T]$, $\Phi_{t}\leq\Phi_{t-1}-\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))$. After $T$ iterations, it follows that $\Phi_{T}\leq\Phi_{0}-T\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))=K-T\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))$. However, since $T$ is set to be a sufficiently large polynomial of $m,\Gamma,K,{\lvert\mathcal{Y}\rvert},1/\epsilon,1/\alpha$ and $1/\sigma$ we would arrive at a contradiction, since $\Phi_{T}\geq 0$. Hence, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon/2$, for some $t\in\{1,\dots,T\}$. Since the error of $h_{t}$ can only be decreasing by Claim 2.17 and $h_{t}$ is close to its sample variant by Claim 2.18, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$.
Sample and Computational Complexity: From the analysis above we have that the algorithm terminates in $\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))$ iterations and at each iteration we draw of the order of ${d}^{O(m)}2^{\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))}\log(1/\delta)$ samples. Hence, we have that the total number of samples is ${d}^{O(m)}2^{\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))}\log(1/\delta)$. Moreover, we use at most $\mathrm{poly}(N,d,\lvert\mathcal{Y}\rvert)$ time as all operations can be implemented in polynomial time. ∎
4 Algorithmic Applications of General MIM Learner
In this section, we present applications of our main result. First, in Section 4.1 we present an application of our algorithmic approach to the class of intersections of halfspaces, establishing Theorem 1.9. Second, in Section 4.2, we show that a slight modification of our algorithm can achieve improved performance in the RCN setting, establishing Theorem 1.8.
Additionally, note that our result on learning Multiclass Linear Classifiers, as presented in Theorem 2.1, can also be derived as a corollary of Theorem 3.3, using similar arguments.
4.1 Learning Intersections of Halfspaces with Adversarial Label Noise
In this section, we present Theorem 4.1, on learning intersections of halfspaces under agnostic noise, which is derived as a corollary of Theorem 3.3. The proof of this result follows from Proposition 4.2, which relies on a new structural result about the class of intersections of $K$-halfspaces, Proposition 4.4. First, we state the formal version of the theorem.
Theorem 4.1 (Learning Intersections of Halfspaces with Adversarial Label Noise).
Fix the class $\mathcal{F}_{K,d}$ of intersections of $K$ halfspaces in $\mathbb{R}^{d}$. Let $D$ be a distribution over $\mathbb{R}^{d}\times\{0,1\}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ and let $\epsilon,\delta\in(0,1),K,d\in\mathbb{Z}_{+}$. Then, Algorithm 3 draws $N=d^{2}2^{\mathrm{poly}(K/\epsilon)}\log(1/\delta)$ i.i.d. samples from $D$, runs in time $\mathrm{poly}(d,N)$, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq K\,O(\mathrm{OPT}\log(1/\mathrm{OPT}))+\epsilon$, where $\mathrm{OPT}=\inf_{f\in\mathcal{F}_{K,d}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$.
For this class of Multi-Index Models, we adapt the general algorithm Algorithm 4 to perform polynomial regression with polynomials of degree at most $2$ in Line 4 of Algorithm 4. This degree suffices for learning intersections of halfspaces, as demonstrated in Proposition 4.2. This task can be efficiently reduced to covariance estimation with respect to the Frobenius norm, as the entries of the covariance matrix correspond to the Hermite coefficients. Therefore, we have that it can be performed in polynomial time using $O(d^{2})$ samples. Consequently, the final sample complexity of the algorithm scales with $d^{2}$ rather than $d^{O(1)}$.
The following proposition shows that the class of intersections of halfspaces satisfies Definition 3.2. To prove this result, we leverage that if a function is far from any function defined over a given subspace, then it must also be far from any constant function on a non-trivial fraction of the cylinders restricted to that subspace.
Proposition 4.2.
Let $d,K\in\mathbb{Z}_{+},\zeta,\epsilon\in(0,1)$. $\mathcal{F}_{K,d}$ is a class of $(2,\zeta,\epsilon,K,KO(\zeta\log(1/\zeta))+O(\epsilon),\mathrm{poly}(\epsilon/K),O(\sqrt{\log K}))$-well-behaved Multi-Index Models according to Definition 3.2.
Proof.
Let $f\in\mathcal{F}_{K,d}$ and denote by $W$ the subspace spanned by the normal vectors of the halfspaces that define $f$, which has dimension at most $K$. From [KOS08, CCK17] we have that the Gaussian surface area of $\mathcal{F}_{K,d}$ is bounded by $O(\sqrt{\log(K)})$. It remains to show that Condition 3 of Definition 3.2 is satisfied for the class $\mathcal{F}_{K,d}$.
Let $D$ be a distribution over $\mathbb{R}^{d}\times\{0,1\}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\zeta$. Let $V$ be a subspace of $\mathbb{R}^{d}$, and let $\tau(x)=x\log(1/x)$. Assume that for all functions $g:V\to\{0,1\}$ it holds that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq g(\mathbf{x}^{V})]\geq CK\tau(\zeta)+2\epsilon$, for a sufficiently large constant $C$. Hence, the same holds true for the function $h$ defined as $h(\mathbf{x})=\operatorname*{argmin}_{i\in\{\pm 1\}}\mathbf{Pr}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}]$, since by the definition $h(\mathbf{x})=h(\mathbf{x}^{V})$. Thus, $\operatorname*{\mathbf{Pr}}[h(\mathbf{x})\neq y]\geq(C-1)K\tau(\zeta)+2\epsilon$.
We first show that if the current subspace $V$ is far from optimal, then by splitting the space into cylinders with respect to each point of $V$, the target function is far from being constant on a nontrivial fraction of these cylinders. Specifically, we show the following
Claim 4.3 (Fraction of Good Points).
With probability at least $\epsilon$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, independent of $\mathbf{x}$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>(C-2)K\tau\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\right)+\epsilon$ for all $i\in\{\pm 1\}$ and $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\leq 1/(C-2)$.
Proof of Claim 4.3.
For the sake of contradiction, assume that the set $S$ of $\mathbf{x}_{0}\in\mathbb{R}^{d}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>(C-1)K\tau\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\right)+\epsilon$ for all $i\in\{\pm 1\}$ has probability mass less than $\epsilon$. Then, we have that
$\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]$ $=\operatorname*{\mathbf{E}}_{\mathbf{x}_{0}^{V}}\left[\min_{\hat{y}\in\{\pm 1\}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}\left[\hat{y}\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\right]\right]$ (3)
$\leq(C-1)K\operatorname*{\mathbf{E}}_{\mathbf{x}_{0}^{V}}\left[\tau\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}\left[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\right]\right)\right]+2\epsilon$
$\leq(C-1)K\tau(\zeta)+2\epsilon\;,$
where in the last inequality we used Jensen’s inequality since $\tau$ is concave. This leads to a contradiction, and therefore we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{x}\in S]>\epsilon$. Observe that for each such $\mathbf{x}_{0}$ as $(C-1)K\tau(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}])+\epsilon\leq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\leq 1$, we have that, for a sufficiently large choice of $C$, the probability $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]$ is less than the constant specified in the statement for all $\mathbf{x}_{0}\in S$. Finally, noting that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\leq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]+\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]$ and $\tau(x)\geq x$ for any $x\in(0,1)$ less than a constant gives us the statement. ∎
Fix a point $\mathbf{x}_{0}\in\mathbb{R}^{d}$ that satisfies the statement of Claim 4.3. Let $D_{V^{\perp}}$ be a distribution supported on $V^{\perp}\times\{\pm 1\}$ defined as $D_{V^{\perp}}(\mathbf{x}^{\perp V},y)=D(\mathbf{x},y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V})$. By the properties of the Gaussian distribution, we have that the $\mathbf{x}$-marginal of $D_{V^{\perp}}$ is a standard Gaussian. Moreover, by the definition of $D_{V^{\perp}}$, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f^{\prime}(\mathbf{x})\neq i]>(C-2)K\tau(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f^{\prime}(\mathbf{x})\neq y])+\epsilon$ for all $i\in\{\pm 1\}$ and $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f^{\prime}(\mathbf{x})\neq y]\leq 1/(C-2)$, where $f^{\prime}(\mathbf{x})=f(\mathbf{x}^{\perp V}+\mathbf{x}_{0}^{V})$. Finally, applying Proposition 4.4 for every intersection of halfspaces $f^{\prime}(\mathbf{x})=f(\mathbf{x}^{\perp V}+\mathbf{x}_{0}^{V})$ for $\mathbf{x}_{0}^{V}$ that satisfies the statement of Claim 4.3 completes the proof of Proposition 4.2. ∎
Below we state and prove a structural result concerning the intersections of halfspaces. This result asserts that if an intersection of halfspaces is far from being constant, then either the first or the second moment exhibits a non-trivial correlation with its weight vectors. This result is crucial for the execution of our algorithm, as it enables us to improve at every iteration.
Proposition 4.4 (Structural Result for Intersections of Halfspaces).
Let $d\in\mathbb{Z}_{+}$, $\epsilon,\mathrm{OPT}\in(0,1)$, and let $D$ be a distribution over $\mathbb{R}^{d}\times\{0,1\}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$. Let $f$ be a classifier in $\mathcal{F}_{K,d}$ that satisfies $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$, and assume that $f$ can be expressed as $f(\mathbf{x})=\prod_{i\in[K]}\mathds{1}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}\geq 0)$ with $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d},t_{1},\dots,t_{K}\in\mathbb{R}$. If it holds that $\min(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq 0],\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq 1])\geq CK\mathrm{OPT}\log(1/\mathrm{OPT})+\epsilon$ and $\mathrm{OPT}\leq 1/C$ for some sufficiently large absolute constant $C$, then there exists some $i\in[K]$ such that
$\max\left(\left|\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}\left[\mathbf{w}^{(i)}\cdot\mathbf{x}y\right]\right|,\left|\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}\left[\left((\mathbf{w}^{(i)}\cdot\mathbf{x})^{2}-1\right)y\right]\right|\right)\geq\mathrm{poly}(1/K,\epsilon)\;.$
Proof.
Let $W=\mathrm{span}(\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)})$ and without loss of generality assume that $\left\|\mathbf{w}^{(i)}\right\|=1$ for all $i\in[K]$, since we can normalize each halfspace without changing the output of $f$. Assume that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]\rvert\leq p(\epsilon/K)$, for all $i\in[K]$, where $p$ is a sufficiently small universal polynomial (to be quantified later). We show that there exists $i\in[K]$ such that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[((\mathbf{w}^{(i)}\cdot\mathbf{x})^{2}-1)y]\rvert\geq\mathrm{poly}(\epsilon/K)$ for a sufficiently small universal polynomial. We can decompose the second-moment matrix as follows:
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})y]$ $=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})f(\mathbf{x})]+\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})(\mathds{1}(f(\mathbf{x})\neq y=1)-\mathds{1}(y\neq f(\mathbf{x})=1))]$
$=\underbrace{\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[({\mathbf{x}}{\mathbf{x}}^{\top}-\mathbf{I})f(\mathbf{x})]}_{{\mathbf{I}_{1}}}\underbrace{+\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[({\mathbf{x}}{\mathbf{x}}^{\top}-\mathbf{I})(T(\mathbf{x},y)-S(\mathbf{x},y))]}_{{\mathbf{I}_{2}}}\;,$
where $S(\mathbf{x},y)\coloneqq\mathds{1}(y\neq f(\mathbf{x})=1),T(\mathbf{x},y)\coloneqq\mathds{1}(f(\mathbf{x})\neq y=1)$. Note that ${\mathbf{I}_{1}}$ is the second moment of $f(\mathbf{x})$, i.e., the noiseless classifier, and ${\mathbf{I}_{2}}$ is the contribution corresponding to the noise.
First, we show the following tail bounds, which will help us bound the contribution corresponding to the noise.
Claim 4.5.
Let $\epsilon\in(0,1)$ be less than a sufficiently small universal constant, and let $R:\mathbb{R}^{d}\to[0,1]$ be a function such that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]=\epsilon$. For all unit vectors $\mathbf{v}\in\mathbb{R}^{d}$ the following holds:
  • i)
    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[(\mathbf{v}\cdot\mathbf{x})^{2}R(\mathbf{x})]\lesssim\epsilon\log(1/\epsilon)$.
  • ii)
    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{v}\cdot\mathbf{x}R(\mathbf{x})]\lesssim\epsilon\sqrt{\log(1/\epsilon)}$.
Proof.
First, we will prove (i) and then we apply it to prove (ii). Note that the expectation $\operatorname*{\mathbf{E}}[(\mathbf{v}\cdot\mathbf{x})^{2}R(\mathbf{x})]$ is minimized if $R$ puts all of its mass at the tails of the function $(\mathbf{v}\cdot x)^{2}$, i.e., $R(\mathbf{x})=\mathds{1}((\mathbf{v}\cdot\mathbf{x})^{2}\geq y_{\epsilon})$, where $y_{\epsilon}$ is the smallest number such that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathds{1}((\mathbf{v}\cdot\mathbf{x})^{2}\geq y_{\epsilon})]\leq\epsilon$. Hence, we have that
$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[(\mathbf{v}\cdot\mathbf{x})^{2}R(\mathbf{x})]$ $\leq\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[(\mathbf{v}\cdot\mathbf{x})^{2}\mathds{1}((\mathbf{v}\cdot\mathbf{x})^{2}\geq y_{\epsilon})]$
[…]
$=\frac{1}{\sqrt{2\pi}}y_{\epsilon}e^{-{y_{\epsilon}}/2}-\epsilon$ (4)
$\leq\epsilon\log(1/\epsilon)+\epsilon\lesssim\epsilon\log(1/\epsilon)\;,$ (5)
where in (4) we simply used partial integration. In (5) we used the fact that the function $xe^{-x/2}$ is decreasing when $x$ is larger than $2$, note that from A.1 we have that $y_{\epsilon}\gtrsim\log(1/\epsilon)\geq 2$, since $\epsilon$ is assumed to be less than some sufficiently small constant.
Finally, we can prove (ii) by simply using Cauchy-Schwarz
$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{v}\cdot\mathbf{x}R(\mathbf{x})]^{2}\leq\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[(\mathbf{v}\cdot\mathbf{x})^{2}R(\mathbf{x})]\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]\lesssim\epsilon^{2}\log(1/\epsilon)\;.$
Which concludes the proof of Claim 4.5. ∎
We next show that there exists a direction within $W$ along which the second moment is small.
Claim 4.6.
There exists an $i\in[K]$ such that
$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\mathbf{w}^{(i)}\lesssim-\frac{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]}{K}+p(\epsilon/K)+\mathrm{OPT}\log(1/\mathrm{OPT})\;.$
Proof.
Denote by $\mu\coloneqq\operatorname*{\mathbf{E}}[\mathbf{x}\mid f(\mathbf{x})]$, by $\bar{\mathbf{x}}\coloneqq\mathbf{x}-\mathbf{\mu}$ and by $\mathbf{M}=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\bar{\mathbf{x}}\bar{\mathbf{x}}^{\top}f(x)]$. For a defining vector $\mathbf{w}^{(i)}$, by B.4 we have that $(\mathbf{w}^{(i)})^{\top}\mathbf{M}\mathbf{w}^{(i)}\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1](1-ce^{-(\max(0,t_{i}))^{2}/2}),$ for a sufficiently small universal constant $c>0$. Furthermore, we claim that there exists $i\in[K]$ for which the bias $t_{i}$ is at most $2\sqrt{\ln(\frac{K}{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]})}$. To see that, note that if for every $i\in[K]$ we have that $t_{i}>2\sqrt{\ln(\frac{K}{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]})},$ then by the union bound
$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq 1]\leq\sum_{i=1}^{K}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}\leq 0]<\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]\;,$
which is a contradiction. Hence, it holds that $(\mathbf{w}^{(i)})^{\top}\mathbf{M}\mathbf{w}^{(i)}\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1](1-c\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]/K)$ for some $i\in[K]$.
Therefore, we have that
$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\mathbf{w}^{(i)}$ $=(\mathbf{w}^{(i)})^{\top}\mathbf{M}\mathbf{w}^{(i)}-\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=1]+\frac{\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{w}^{(i)}\cdot\mathbf{x}f(\mathbf{x})]^{2}}{\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=1]}$
$\leq-c\frac{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]}{K}+\frac{\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{w}^{(i)}\cdot\mathbf{x}f(\mathbf{x})]^{2}}{\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=1]}\;.$
Note that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{w}^{(i)}\cdot\mathbf{x}f(\mathbf{x})]=\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]+\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}(S(\mathbf{x},y)-T(\mathbf{x},y))]$. By the assumption that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]\rvert\leq\epsilon p(\epsilon/K)$ for all $i\in[K]$, and the fact that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\geq\operatorname*{\mathbf{Pr}}[y=1]-\mathrm{OPT}\geq\epsilon$, it holds that $\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]^{2}/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\leq p(\epsilon/K)$.
Moreover, note that $\lvert\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}S(\mathbf{x},y)]\rvert=\lvert\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}R(\mathbf{x})]\rvert$, for $R(\mathbf{x})=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]$. Therefore, since $R(\mathbf{x})\in[0,1]$ for all $\mathbf{x}\in\mathbb{R}^{d}$ and $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]\leq\mathrm{OPT}$, by Claim 4.5 we have that $\lvert\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}(S(\mathbf{x},y)]\rvert\lesssim\mathrm{OPT}\sqrt{\log(1/\mathrm{OPT})}$. Consequently, since $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\geq\mathrm{OPT}$ we have that $\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}S(\mathbf{x},y)]^{2}/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\lesssim\mathrm{OPT}\log(1/\mathrm{OPT})$. Using similar arguments we have that $\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}T(\mathbf{x},y)]^{2}/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\lesssim\mathrm{OPT}\log(1/\mathrm{OPT})$.
Combining the above results gives us
$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\mathbf{w}^{(i)}\lesssim-\frac{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]}{K}+p(\epsilon/K)+\mathrm{OPT}\log(1/\mathrm{OPT})\;.$
Which concludes the proof of Claim 4.6. ∎
Now we bound the contribution corresponding to the noise, using Claim 4.5.
Claim 4.7.
For every unit vector $\mathbf{v}\in\mathbb{R}^{d}$, it holds that $\lvert{\mathbf{v}^{\top}\mathbf{I}_{2}\mathbf{v}}\rvert=O(\mathrm{OPT}\log(1/\mathrm{OPT}))$.
Proof.
Let $\mathbf{v}\in\mathbb{R}^{d}$ be a unit vector, and denote by $R(\mathbf{x})=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]$. Note that $R(\mathbf{x})\in[0,1]$ for all $\mathbf{x}\in\mathbb{R}^{d}$, and moreover $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]\leq\mathrm{OPT}$. Hence, by Claim 4.5 we have that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{v}\cdot\mathbf{x})^{2}S(\mathbf{x},y)]\rvert=O(\mathrm{OPT}\log(1/\mathrm{OPT}))$. By using similar arguments we have that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{v}\cdot\mathbf{x})^{2}T(\mathbf{x},y)]\rvert=O(\log(1/\mathrm{OPT})\mathrm{OPT})$. Therefore, we have that
$\lvert{\mathbf{v}^{\top}\mathbf{I}_{2}\mathbf{v}}\rvert$ $\leq\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{v}\cdot\mathbf{x})^{2}S(\mathbf{x},y)]\rvert+\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{v}\cdot\mathbf{x})^{2}T(\mathbf{x},y)]\rvert+\operatorname*{\mathbf{E}}[S(\mathbf{x},y)+T(\mathbf{x},y)]$
$=O(\log(1/\mathrm{OPT})\mathrm{OPT})\;.$
Which concludes the proof of Claim 4.7. ∎
Denote by $t$ the quantity $t\coloneqq C^{\prime}\log(1/\mathrm{OPT})\mathrm{OPT}$, for sufficiently large constant $C^{\prime}$. Combining the Claims 4.6 and 4.7 we have that
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})y]$ $\lesssim-\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]/K-t+p(\epsilon/K)$
$\leq-\left(\min_{i\in\{0,1\}}(\operatorname*{\mathbf{Pr}}[y=i])-\mathrm{OPT}\right)/(2K)-t+p(\epsilon/K)$
$\leq-\epsilon/(2K)+p(\epsilon/K)\;,$
where we used the fact that one of $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1],\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]$ is greater than $1/2$ and the assumption that for all labels $i\in\{0,1\}$, $\operatorname*{\mathbf{Pr}}[y=i]\geq CK\mathrm{OPT}\log(1/\mathrm{OPT})+\epsilon$, for a sufficiently large constant $C$. Setting $p(\epsilon/K)\leq c\epsilon/K$ for a sufficiently small universal constant $c>0$, concludes the proof of Proposition 4.4. ∎
4.2 Learning Multi-Index Models with RCN
In this section, we demonstrate that our algorithmic approach becomes more efficient when the labeled examples are perturbed by uniform random noise. In particular, we study the performance of our approach in the RCN setting (see Definition 4.8). We show that Algorithm 3 can be modified into Algorithm 5, which achieves improved complexity in the RCN regime compared to the agnostic noise setting. The main result of this section is Theorem 4.9.
Definition 4.8 (RCN Distribution).
Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$. We say that $D$ is a RCN distribution for the class $\mathcal{F}\subseteq\{f:\mathbb{R}^{d}\to\mathcal{Y}\}$ if there exists a function $f\in\mathcal{F}$ and a row-stochastic matrix $\mathbf{H}\in\mathbb{R}^{\lvert\mathcal{Y}\rvert\times\lvert\mathcal{Y}\rvert}$ with $\mathbf{H}_{i,i}>\mathbf{H}_{i,j}+\gamma$ for all $i,j\in\mathcal{Y}$, for some $\gamma>0$, such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y=i\mid\mathbf{x}]=\mathbf{H}_{f(\mathbf{x}),i}$. We refer to $\mathbf{H}$ as the confusion matrix of the distribution.
The key distinction that allows for this improvement is that, in the RCN setting, the output label depends solely on the ground truth label. Therefore, the moments of the noisy class regions still rely only on the subspace that characterizes the classifier. Consequently, any non-zero moment results in vectors that approximately lie within that subspace (see Proposition 4.10). The main result of this section is Theorem 4.9.
Input: Accuracy $\epsilon>0$, $\delta>0$ and sample access to a distribution $D$ over $\mathbb{R}^{d}\times\mathcal{Y}$ that satisfies Definition 4.8.
Output: A hypothesis $h$ such that with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$.
  1. 1.
    Let $C$ be a sufficiently large universal constant.
  2. 2.
    Let $N\leftarrow(dm)^{{C}m}2^{K^{C}}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{{C}K}\log(1/\delta)/(\sigma\alpha)^{{C}}$, $\epsilon^{\prime}\leftarrow\epsilon^{3}/(C\lvert\mathcal{Y}\rvert\Gamma^{2}K^{2}))$, $L_{1}\leftarrow\emptyset$,$t\leftarrow 1$.
  3. 3.
    While $t\leq 2K$
    1. (a)
      Draw a set $S$ of $N$ i.i.d. samples from $D$.
    2. (b)
      $\mathcal{E}_{t}\leftarrow$ Algorithm 4($\epsilon,\epsilon^{\prime},\delta,\mathrm{span}(L_{t}),S$).
    3. (c)
      Choose any vector $\mathbf{v}\in\mathcal{E}_{t}$.
    4. (d)
      $L_{t+1}\leftarrow L_{t}\cup\{\mathbf{v}\}$.
    5. (e)
      $t\leftarrow t+1$.
  4. 4.
    Construct $\mathcal{S}$ an $\epsilon^{\prime}$-approximating partition with respect to $\mathrm{span}(L_{t})$, as defined in Definition 2.2.
  5. 5.
    Return a piecewise constant approximating classifier $h_{\mathcal{S}}$, as described in Definition 2.3, using $N$ i.i.d. samples from $D$, by finding the most frequent label at any set in $\mathcal{S}$.
Algorithm 5 Learning Multi-Index Models under RCN
Theorem 4.9 (Learning with RCN).
Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, and let $\epsilon,\delta\in(0,1)$. Let $\mathrm{OPT}=\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$, where $\mathcal{F}$ is a class of $(m,{\zeta+\epsilon},\alpha,K,\Gamma,\tau,\sigma)$-well-behaved Multi-Index Models, as in Definition 3.2. If $\zeta\geq\mathrm{OPT}$ and $D$ satisfies Definition 4.8, then Algorithm 5 draws $N={(dm)^{O(m)}2^{\mathrm{poly}(K)}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{O(K)}\log(1/\delta)/(\sigma\alpha)^{O(1)}}$ i.i.d. samples from $D$, runs in time $\mathrm{poly}(d,N,\lvert\mathcal{Y}\rvert)$, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$.
We remark that for the case of multiclass linear classifiers, similarly to Proposition 4.2, using Proposition 2.4 we can prove that the class of multiclass linear classifiers is $(1,\zeta,\epsilon,K,O(\zeta)+O(\epsilon),\mathrm{poly}(\epsilon/K),O(K\sqrt{\log(K)}))$-well behaved. As a result, we obtain an algorithm for learning under RCN by applying Theorem 4.9. Moreover, note that in this case Algorithm 4 performs linear regression under Gaussian marginals (see Line 4). This task that can be efficiently reduced to mean estimation, specially to estimating $\operatorname*{\mathbf{E}}[\mathds{1}(y=i)]$ and $\operatorname*{\mathbf{E}}[\mathbf{x}\mathds{1}(y=i)]$. Therefore, this step can be performed in polynomial time using $O(d)$ samples. Consequently, the final sample complexity of the algorithm scales with $d$ rather than $d^{O(1)}$, which establishes Theorem 1.8.
Our proof of the Theorem 4.9 relies on the following proposition, which states that, given that our distribution is an RCN distribution with respect to a function in the class, all vectors returned by Algorithm 4 have an improved correlation with the ground truth subspace compared to the agnostic case (see Proposition 3.7).
Proposition 4.10.
Let $d,k,K\in\mathbb{Z}_{+}$, $\epsilon,\delta\in(0,1)$ and $\epsilon^{\prime}\leq c\epsilon^{3}/(\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2}))$, for $c$ a sufficiently small universal constant. Let $D$ be a distribution supported on $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$ and let $\mathcal{F}\subseteq\{f:\mathbb{R}^{d}\to\mathcal{Y}\}$ be a class of $(m,{\zeta+O(\epsilon)},\alpha,K,\Gamma,\tau,\sigma)$-well-behaved Multi-Index Models according to Definition 3.2. Let $f\in\mathcal{F}$ with $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$ and let $W$ be a subspace such that $f(\mathbf{x})=f(\mathbf{x}^{W})$ for all $\mathbf{x}\in\mathbb{R}^{d}$, with $K=\dim(W)$. Additionally, let $h_{\mathcal{S}}$ be a piecewise constant approximation of $D$, with respect to an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$, according to $V$, as defined in Definitions 2.2 and 2.3.
There exists $N=(dm)^{O(m)}{(k/\epsilon^{\prime})^{O(k)}}\log(1/\delta){(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{O(1)}}$ such that, if $D$ satisfies Definition 4.8 for $f$, $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+\epsilon$ and $\zeta\geq\mathrm{OPT}$, then Algorithm 4 given $N$ i.i.d. samples from $D$, runs in $\mathrm{poly}(N,d,\lvert\mathcal{Y}\rvert)$ time, and with probability at least $1-\delta$, returns a list of unit vectors $\mathcal{E}$ of size $|\mathcal{E}|=\mathrm{poly}(mK\Gamma/(\epsilon\sigma\alpha))$, such that: for any vector $\mathbf{v}\in\mathcal{E}$, there exists a unit vector $\mathbf{w}\in W$ for which
$\mathbf{w}\cdot\mathbf{v}\geq 1-\epsilon\;.$
Proof.
Denote by $\mathbf{H}$ the stochastic matrix that perturbs the labels of $f$. Also let $\widehat{\mathbf{U}}$ be the matrix computed at Line 6 of Algorithm 5 and let $\eta^{2}=\mathrm{poly}(\epsilon\sigma\alpha/(mK\Gamma))$ be the $l_{2}^{2}$ error chosen in the polynomial-regression step, i.e., Line 4 of Algorithm 5.
We begin by showing that, for each subset of the partition, the computed vectors have only a small component in the subspace orthogonal to $W$. This, in turn, implies that any large vector must have a significant component in $W$.
Claim 4.11.
Fix $S\in\mathcal{S}$. If the number of samples that fall in $S$ is $N_{S}\geq(dm)^{Cm}\log(\lvert\mathcal{Y}/\delta\rvert)/\eta^{C}$ for a sufficiently large universal constant $C>0$, then for any $\mathbf{u}\in U_{S}$ and unit vector $\mathbf{v}\in W^{\perp}$ we have that $\lvert\mathbf{v}\cdot\mathbf{u}\rvert\leq\sqrt{2Km}\eta/\sigma$.
Proof.
Let $N_{S},S\in\mathcal{S}$, be the number of samples that fall in the set $S$. From the rotational invariance of the standard Gaussian, without loss of generality let $\mathbf{e}_{1},\dots,\mathbf{e}_{t}$ be a basis of $W^{\perp}$, $t\in\mathbb{Z}_{+}$. By B.2 and the union bound, we have that with $N_{S}=(dm)^{O(m)}\log(\lvert\mathcal{Y}\rvert/\delta)/{\eta^{O(1)}}$ i.i.d. samples and runtime $\mathrm{poly}(N_{S},d)$, we can find polynomials $p_{S,i}$ such that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}^{\perp V}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}^{\perp V}))^{2}]+\eta^{2}$. Hence, by averaging over $V$ conditioning on $S$, we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}))^{2}]+\eta^{2}$, where $D_{V^{\perp}}$ is defined to be the marginal obtained by averaging $D$ over $V$ conditioned on $S$, i.e., $D_{V^{\perp}}(\mathbf{x}^{\perp V},y)=\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[D(\mathbf{x},y)\mid\mathbf{x}\in S]$.
Define $\hat{p}_{S,i}(\beta)\coloneqq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x})p_{S,i}(\mathbf{x})],\hat{g}_{i}(\beta)\coloneqq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x})\mathds{1}(y=i)]$ for each $\beta\in\mathbb{N}^{d}$. Note that $\hat{p}_{S,i}(\beta)$ and $\hat{g}_{i}(\beta)$ are the Hermite coefficients of $p_{S,i}$ and $\mathds{1}(y=i)$, respectively. Since by the orthogonality of Hermite polynomials we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]=\sum_{\beta\in\mathbb{N}^{d}}(\hat{p}_{S,i}(\beta)-\hat{g}_{i}(\beta))^{2}$, it holds that $\sum_{\beta\in\mathbb{N}^{d},1\leq\|\beta\|_{1}\leq m}(\hat{p}_{S,i}(\beta)-\hat{g}_{i}(\beta))^{2}\leq\eta^{2}$.
Moreover, if $\beta_{i}\neq 0$, then for some $i\in[t]$
$\hat{g}_{i}(\beta)$ $=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[H_{\beta}(\mathbf{x}^{\perp V})\mathds{1}(y=i)\mid S]=\sum_{j=1}^{K}\mathbf{H}_{j,i}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[H_{\beta}(\mathbf{x}^{\perp V})\mathds{1}(f(\mathbf{x})=j)\mid S]$
$=\sum_{j=1}^{K}\mathbf{H}_{j,i}\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp V\cap W}}[H_{\beta}(\mathbf{x}^{\perp V\cap W})\mathds{1}(f(\mathbf{x}^{W})=j)]\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp V\cap\perp W}}[H_{\beta}(\mathbf{x}^{\perp V\cap\perp W})]\mid S]=0\;,$
where we used the fact that the $\mathbf{x}\in S$ depends only on $\mathbf{x}^{V}$, and the fact that non-constant Hermite polynomials have mean zero under the Gaussian distribution. Therefore, we have that the sum of the Hermite coefficients $\hat{p}_{S,i}(\beta)$, for all $\beta\in\mathbb{N}^{d}$ such that $1\leq\|\beta\|_{1}\leq m$ and $\beta_{i}\neq 0$ for some $i\in[t]$, squared is less than $\eta^{2}$.
By fact A.2, for any $S\in\mathcal{S},i\in[K]$, we have that
$\mathbf{e}_{1}^{\top}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[\nabla p_{S,i}(\mathbf{x})\nabla p_{S,i}(\mathbf{x})^{\top}]\mathbf{e}_{1}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\nabla p_{S,i}(\mathbf{x})\cdot\mathbf{e}_{1})^{2}]$ $=\sum_{\beta\in\mathbb{N}^{d}}\beta_{1}(\hat{p}_{S,i}(\beta))^{2}$
$\leq m\eta^{2}\;.$
Hence, as $U_{S,i}$ has the eigenvectors of $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[\nabla p_{S,i}(\mathbf{x})\nabla p_{S,i}(\mathbf{x})^{\top}]=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\nabla p_{S,i}(\mathbf{x}^{\perp V})\nabla p_{S,i}(\mathbf{x}^{\perp V})^{\top}\mid S]$ with eigenvalue greater than $\sigma^{2}/(2K)$, by Claim 4.12 we have the statement of the claim. ∎
Claim 4.12.
Let $\mathbf{M}\in\mathbb{R}^{d\times d}$ be a symmetric, psd matrix and let $\mathbf{v}\in\mathbb{R}^{d}$ be a vector such that $\mathbf{v}^{\top}\mathbf{M}\mathbf{v}\leq\epsilon$. Let $U$ denote the set of eigenvectors of $\mathbf{M}$ with eigenvalue greater than $\lambda$. Then, for every $\mathbf{u}\in U$, it holds that $\lvert\mathbf{u}\cdot\mathbf{v}\rvert\leq\sqrt{\epsilon/\lambda}$.
Proof.
As $\mathbf{M}$ is symmetric, by the Spectral Decomposition Theorem, we can decompose it to the basis of its eigenvectors $\mathbf{u}^{(1)},\dots\mathbf{u}^{(d)}$. Denoting by $\lambda_{1},\dots,\lambda_{d}$ the corresponding eigenvalues, we have that
$\epsilon\geq\mathbf{v}^{\top}\mathbf{M}\mathbf{v}=\sum_{i=1}^{d}\lambda_{i}(\mathbf{u}^{(i)}\cdot\mathbf{v})^{2}\geq\sum_{\mathbf{u}\in U}\lambda(\mathbf{u}\cdot\mathbf{v})^{2}\;,$
where we used the fact that $\mathbf{M}$ is psd. Which completes the proof of Claim 4.12. ∎
First note that by the definition of an approximating partition (see Definition 2.2) and A.1, for all $S\in\mathcal{S}$ we have that $\operatorname*{\mathbf{Pr}}_{D}[S]{=(\epsilon^{\prime}/k)^{\Omega(k)}}$ and $\lvert\mathcal{S}\rvert{=(k/\epsilon^{\prime})^{O(k)}}$. Therefore, by union bound and Hoeffding’s inequality, it holds that if $N\geq{(k/\epsilon^{\prime})^{Ck}}\log(\lvert\mathcal{S}\rvert/\delta)={(k/\epsilon^{\prime})^{O(k)}}\log(1/\delta)$ for a sufficiently large constant $C>0$, then with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]-\operatorname*{\mathbf{Pr}}_{D}[S]\rvert\leq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S}$. Hence, it is true that $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$, i.e., the number of samples that fall in each set $S\in\mathcal{S}$ is at least $N\operatorname*{\mathbf{Pr}}_{D}[S]/2=N{(\epsilon^{\prime}/k)^{\Omega(k)}}$.
Therefore, if $N\geq(dm)^{Cm}{(k/\epsilon^{\prime})^{Ck}}\log(1/\delta){(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{C}}$ for a sufficiently large universal constant $C>0$, then by applying Claim 4.11 we have that with probability $1-\delta$ for any unit vector $\mathbf{v}\in W^{\perp}$
$\mathbf{v}^{\top}\widehat{\mathbf{U}}\mathbf{v}\leq\sum_{S\in\mathcal{S},\mathbf{u}\in U_{S}}(\mathbf{v}\cdot\mathbf{u})^{2}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\leq\eta^{2}\mathrm{poly}(mK\lvert\mathcal{Y}\rvert/\sigma)\;.$
where we used the fact that $\lvert U_{S}\rvert=\mathrm{poly}(mK\lvert\mathcal{Y}\rvert/\sigma)$ by Claim 3.9. Hence, since $\widehat{\mathbf{U}}$ is a symmetric psd matrix, applying Claim 4.12 we have that for all $\mathbf{u}\in\mathcal{E}$ and unit vectors $\mathbf{v}\in W^{\perp}$ it holds that $\lvert\mathbf{u}\cdot\mathbf{v}\rvert\leq\eta/\sqrt{\mathrm{poly}(\epsilon\sigma\alpha/(mK\Gamma\lvert\mathcal{Y}\rvert)}$. As a result, we have that if $\eta$ is chosen to be a sufficiently small polynomial of $\epsilon,\sigma,\alpha,1/\Gamma,1/m,1/K,1/\lvert\mathcal{Y}\rvert$, then for all $\mathbf{u}^{(i)}\in\mathcal{E}$ and any $\mathbf{v}\in W^{\perp}$ we have that $\lvert\mathbf{u}^{(i)}\cdot\mathbf{v}\rvert\leq\epsilon$.
Moreover, from Proposition 3.7 we have that if $N\geq(dm)^{Cm}{(k/\epsilon^{\prime})^{Ck}}\log(1/\delta){(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{C}}$ for a sufficiently large universal constant $C>0$, then there exists a $\mathbf{v}\in\mathcal{E}$. Moreover, as $\mathbf{v}$ is a unit vector and $\left\|\mathbf{v}^{\perp W}\right\|\leq\epsilon$, we have that $\left\|\mathbf{v}^{W}\right\|\geq\sqrt{1-\epsilon^{2}}\geq 1-\epsilon$ which proves Proposition 4.10. ∎
Now we prove our main result for learning with RCN. Our proof similar to the proof of Theorem 3.3. The distinction is that we are going to use Proposition 4.10 which provides improved correlation when compared to Proposition 3.7.
Proof of Theorem 4.9.
We show that Algorithm 5, with high probability, returns a hypothesis $h$ with 0-1 error at most $\tau+\mathrm{OPT}+\epsilon$. Let $f^{*}$ be an optimal classifier and denote by $W^{*}$ the subspace of its weight vectors. Let $L_{t}$ be the list of vectors updated by the algorithm (Line 3d) and $V_{t}=\mathrm{span}(L_{t})$, $\dim(V_{t})=k_{t}$. Denote by $T=2K$ the total number of iterations. Also let $\epsilon^{\prime}$ be the parameter set at Line 2 of Algorithm 5, and for $t\in[T]$ let $\mathcal{S}_{t}$ be arbitrary $\epsilon^{\prime}$-approximating partitions with respect to $V_{t}$ (see Definition 2.2). Let $h_{t}:\mathbb{R}^{d}\to[K]$ be piecewise constant classifiers, defined as $h_{t}=h_{\mathcal{S}_{t}}$ according to Definition 2.3 for the distribution $D$.
Note that from Lines 2, 3 of Algorithm 5, we perform $O(K)$ iterations. Furthermore, in each iteration, we update the vector set by adding one vector. Hence, $k_{t}=O(K)$ for all $t=1,\dots,T$.
Assume that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+\epsilon$ for all $t=1,\dots,T$. Using the fact that $N={(dm)^{Cm}2^{K^{C}}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{CK}\log(1/\delta)/(\sigma\alpha)^{C}}$ for a sufficiently large universal constant $C>0$ (Line 2 of Algorithm 5), we can apply Proposition 4.10 and conclude that, with probability $1-\delta$, there exist unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ for $t\in[T]$ such that $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq 1-\epsilon\geq 1/2$. Thus, from Claim 2.16, we have that with probability $1-\delta$, for all $t\in[T]$, $\Phi_{t}\leq\Phi_{t-1}-1/2$. After $T$ iterations, it follows that $\Phi_{T}\leq\Phi_{0}-T/2=K-T/2$. However, if $T$ is set to be $2K$ we would arrive at a contradiction, since $\Phi_{T}\geq 0$. Hence, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$ for some $t\in\{1,\dots,T\}$. Since the error of $h_{t}$ can only be decreasing by Claim 2.17, and $h_{t}$ is close to its sample variant by Claim 2.18, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+2\epsilon$.
Sample and Computational Complexity: Note that the algorithm terminates in $O(K)$ iterations and at each iteration we draw $N={(dm)^{O(m)}2^{\mathrm{poly}(K)}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{O(K)}\log(1/\delta)/(\sigma\alpha)^{O(1)}}$ samples. Hence, the total sample size is ${(dm)^{O(m)}2^{\mathrm{poly}(K)}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{O(K)}\log(1/\delta)/(\sigma\alpha)^{O(1)}}$. Moreover we use at most $\mathrm{poly}(N,d,\lvert\mathcal{Y}\rvert)$ time as all operations can be implemented in polynomial time in $N,d$ and $\lvert\mathcal{Y}\rvert$. ∎
5 Statistical Query Lower Bounds
Here we prove Theorems 1.5 and 1.7. First, in Section 5.1 we present SQ lower bounds about general classes of Multi-Index Models that do not satisfy Definition 1.2. Essentially showing that the algorithm of Section 3 is quantitatively optimal in the SQ model. Second, in Section 5.2, we present an SQ lower bound for the problem of learning multiclass linear classifiers under RCN.
5.1 SQ Lower Bounds for Learning Multi-Index Models
Here we show that if a Multi-Index Model class $\mathcal{F}$ does not satisfy Definition 3.2, any SQ algorithm that learns $\mathcal{F}$ to an accuracy comparable to that of Theorem 3.3 requires super-polynomial complexity. Specifically, we establish the following theorem:
Theorem 5.1.
(SQ Optimality of General Algorithm) Let $d,K,m\in\mathbb{Z}_{+}$ with $d^{c}\geq\max(m,K)$ for a sufficiently small absolute constant $c>0$, and let $\mathcal{Y}$ be a set of finite cardinality. Additionally, let $\tau,\mathrm{OPT}\in(0,1)$ with $\tau,\mathrm{OPT}\geq d^{-O(m)}$. Assume the existence of (i) a function $f:\mathbb{R}^{d}\mapsto\mathcal{Y}$ (ii) a random variable $y$ over $\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]=\mathrm{OPT}$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, (iii) a subspace $W$ of dimension at most $K$, and iv) a subspace $V$ such that:
  1. 1.
    The function $f$ depends only on the projection onto $W$, i.e., $f(\mathbf{x})=f(\mathbf{x}^{W})$.
  2. 2.
    For every function $g:V\to\mathcal{Y}$ we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\geq\tau$.
  3. 3.
    For any $\mathbf{x}_{0}\in\mathbb{R}^{d},i\in[K]$, and for any zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$, it holds that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[p(\mathbf{x}^{U})\mathds{1}(y=i)\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]=0$, where $U=(V+W)\cap V^{\perp}$.
Let $\mathcal{F}$ denote the class of functions defined as =$\mathcal{F}=\{f(\mathbf{x}^{\perp U},\mathbf{U}\mathbf{x}^{U}):\mathrm{for\ all\ }\mathbf{U}\in\mathbb{R}^{k\times d}:\mathbf{U}\mathbf{U}^{\top}=\mathbf{I}_{k}\}$, where $k=\dim(U)$. Then, any SQ algorithm that learns $\mathcal{F}$ up to an error of smaller than $\tau-2\mathrm{OPT}$ with probability at least $2/3$ requires either $2^{d^{\Omega(1)}}$ queries or at least one query to $\text{VSTAT}\left(d^{\Omega(m)}/|\mathcal{Y}|\right)$.
Proof.
Let $U=(V+W)\cap V^{\perp}$ with $K^{\prime}=\dim(U)$ and let $U^{\perp}$ be the orthogonal complement of $U$.
Let $\mathcal{Q}\subseteq\mathbb{R}^{d\times K^{\prime}}$ be the set of matrices from C.4 for which it holds that $\lvert\mathcal{Q}\rvert=2^{\Omega(d^{c})},\mathbf{Q}\mathbf{Q}^{\top}=\mathbf{I}_{K^{\prime}}$ and $\left\|\mathbf{Q}\mathbf{P}^{\top}\right\|_{F}\leq O(d^{2c-1+2\alpha})$ for all $\mathbf{Q},\mathbf{P}\in\mathcal{Q}$, where $\alpha=\log K^{\prime}/\log d$. For each $\mathbf{Q}\in\mathcal{Q}$, we define the function $f_{\mathbf{Q}}(\mathbf{x}):=f(\mathbf{x}^{\perp U},\mathbf{Q}\mathbf{x}^{U})$ along with the coupled random variables $y_{\mathbf{Q}}$ satisfying $\operatorname*{\mathbf{Pr}}[y_{\mathbf{Q}}=i|\mathbf{x}=\mathbf{z}]=\operatorname*{\mathbf{Pr}}[y=i|\mathbf{x}=\mathbf{z}^{\perp U}+\mathbf{Q}\,\mathbf{z}^{U}]$, for all $i\in\mathcal{Y},\mathbf{z}\in\mathbb{R}^{d}$. We consider the following class of functions $\mathcal{F}_{\mathcal{Q}}=\{f_{\mathbf{Q}}(\mathbf{x}):\mathbf{Q}\in\mathcal{Q}\}$ and note that by definition $\mathcal{F}_{\mathcal{Q}}\subseteq\mathcal{F}$.
For each $\mathbf{Q}\in\mathcal{Q}$, let $D_{\mathbf{Q}}$ be the joint distribution of $(\mathbf{x},y_{\mathbf{Q}})$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. Denote by $\mathcal{D}=\{D_{\mathbf{Q}}:\mathbf{Q}\in\mathcal{Q}\}$ the class of these distributions. Denote by $D_{0}$ the distribution of pairs $(\mathbf{x},y_{0})$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $y_{0}$ a random variable supported on $\mathcal{Y}$ that depends only on $W^{V}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{0}}[y_{0}=i\mid\mathbf{x}^{\perp U}=\mathbf{z}^{\perp U}]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[y=i\mid\mathbf{x}^{\perp U}=\mathbf{z}^{\perp U}]$, for all $i\in\mathcal{Y},\mathbf{z}\in\mathbb{R}^{d}$. Denote by $y_{i}(\mathbf{x}^{\perp U})=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{0}}[y_{0}=i\mid\mathbf{x}^{\perp U}]$.
We first show that for all $D_{\mathbf{P}},D_{\mathbf{Q}}\in\mathcal{D}$ with $\mathbf{P}\neq\mathbf{Q}$, the correlation $\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{Q}})$ is small. Using the law of total expectation, we have that
$\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{Q}})$ $=\sum_{i\in\mathcal{Y}}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\frac{1}{y_{i}(\mathbf{x}^{\perp U})}(\mathds{1}(y_{\mathbf{P}}=i)-y_{i}(\mathbf{x}^{\perp U}))(\mathds{1}(y_{\mathbf{Q}}=i)-y_{i}(\mathbf{x}^{\perp U}))\right]$
$=\sum_{i\in\mathcal{Y}}\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp U}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\bigg{[}\frac{1}{y_{i}(\mathbf{x}^{\perp U})}\operatorname*{\mathbf{E}}_{\mathbf{x}^{U}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathds{1}(y_{\mathbf{P}}=i)-y_{i}(\mathbf{x}^{\perp U}))(\mathds{1}(y_{\mathbf{Q}}=i)-y_{i}(\mathbf{x}^{\perp U}))\mid\mathbf{x}^{\perp U}]\bigg{]}\;.$
Considering each term separately, by C.5, we have that
$\operatorname*{\mathbf{E}}_{\mathbf{x}^{U}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[(\mathds{1}(y_{\mathbf{P}}=i)-$ $y_{i}(\mathbf{x}^{\perp U}))(\mathds{1}(y_{\mathbf{Q}}=i)-y_{i}(\mathbf{x}^{\perp U}))\mid\mathbf{x}^{\perp U}]$
$\leq\sum_{t=m}^{\infty}\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{t}\operatorname*{\mathbf{E}}[\mathds{1}(y=i)^{2}\mid\mathbf{x}^{\perp U}]\leq\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{m}\operatorname*{\mathbf{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp U}]\;,$
where we used Parseval’s identity and Assumption (3) of the theorem statement. Therefore, we have that
$\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{Q}})$ $\leq\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{m}\sum_{i\in\mathcal{Y}}\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp U}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\frac{\operatorname*{\mathbf{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp U}]}{y_{i}(\mathbf{x}^{\perp U})}\right]\leq\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{m}|\mathcal{Y}|\;.$
Hence, $\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{Q}})=d^{-\Omega(m)}\lvert\mathcal{Y}\rvert$. Similarly, we can show that $\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{P}})\leq\lvert\mathcal{Y}\rvert$ for all $D_{\mathbf{P}}\in\mathcal{D}$. Therefore, we have that $\mathcal{D}$ is $(d^{-\Omega(m)}\lvert\mathcal{Y}\rvert,\lvert\mathcal{Y}\rvert)$-correlated relative to $D_{0}$ (see Definition C.1), and thus from C.3 we have that any SQ algorithm that solves the decision problem $\mathcal{B}(\mathcal{D},D_{0})$ with probability at least $2/3$ requires $2^{d^{\Omega(1)}}$ queries to the $\text{VSTAT}(d^{\Omega(m)}/\lvert\mathcal{Y}\rvert)$ oracle.
Hence, in order to complete the statement of the theorem, we need to show the following standard reduction from the testing problem to the learning problem (see, e.g. Chapter $8$ of [DK23]).
Claim 5.2.
Any SQ algorithm that learns $\mathcal{F}$ with agnostic noise with error better than $\tau-2\mathrm{OPT}$ can be used to solve the decision problem $\mathcal{B}(\mathcal{D},D_{0})$ using one more query of accuracy $\mathrm{OPT}$.
Proof.
Assume the existence of an SQ algorithm $\mathcal{A}$ that given access to a distribution $D^{\prime}$ over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, returns a hypothesis $h$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D^{\prime}}[h(\mathbf{x})\neq y]<\tau-2\mathrm{OPT}$, with $\mathrm{OPT}=\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D^{\prime}}[f(\mathbf{x})\neq y]$. Using $\mathcal{A}$, we can reduce the testing problem to the learning problem.
Note that by running algorithm $\mathcal{A}$ to distribution $D_{\mathbf{P}}\in\mathcal{D}$, $\mathcal{A}$ returns a hypothesis $h$ with error $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{\mathbf{P}}}[h(\mathbf{x})\neq y]<\tau-2\mathrm{OPT}$. However, running a $\mathcal{A}$ with SQ oracle access to $D_{0}$ would result in a hypothesis $h$ that cannot perform better than $\tau-\mathrm{OPT}$, as $D_{0}$ is a product distribution for each fixed projection of $\mathbf{x}$ in $V$. As $\tau-\mathrm{OPT}>\tau-2\mathrm{OPT}$, we have that with an additional query of accuracy less than $\mathrm{OPT}$, we can solve $\mathcal{B}(\mathcal{D},D_{0})$. ∎
5.2 SQ Lower Bounds for Learning Multiclass Linear Classifiers with RCN
In this section, we present SQ hardness results for learning Multiclass Linear Classifiers in the presence of RCN under the Gaussian distribution. In particular, we show that achieving $\mathrm{OPT}+\epsilon$ error in the SQ model requires, either a super-polynomial number of queries, or queries with accuracy that depends on the spectral properties of the confusion matrix (see Definition 4.8), rather than on the amount of noise. The main result of this section is Theorem 5.3.
Theorem 5.3 (SQ Lower Bound for Learning under RCN).
Any SQ algorithm that learns $K$-Multiclass Linear Classifiers under the $d$-dimensional standard normal in the presence of RCN to error $\mathrm{OPT}+\mathrm{poly}(1/K)$, requires either $2^{d^{\Omega(1)}}$ queries or at least one query to $\text{VSTAT}\left(d^{\Omega(K)}\right)$.
SQ-Hard Instances
To construct our family of hard instances, we first define a two-dimensional model $f:\mathbb{R}^{2}\to[K]$ as follows:
$\mathbf{w}^{(k)}=\begin{bmatrix}\cos\left(\frac{2\pi k}{K}\right)\\ \sin\left(\frac{2\pi k}{K}\right)\\ \end{bmatrix},k\in[K],\text{ and }f(\mathbf{x})=\operatorname*{argmax}_{k\in[K]}(\mathbf{w}^{(k)}\cdot\mathbf{x})\;.$ (6)
Geometrically, the classifier $f$ partitions the plane into $K$ equal angular sectors, assigning each sector to a different class. Now, given $\mathcal{Q}$, a set of $2\times d$ matrices, we can define a class of functions from $\mathbb{R}^{d}$ to $[K]$ as
$\mathcal{F}_{\mathcal{Q}}=\{f(\mathbf{Q}\mathbf{x}):\mathbf{Q}\in\mathcal{Q}\}\;.$ (7)
Before proving Theorem 5.3, we first need to show that we can match the appropriate number of moments with the standard Gaussian.
Lemma 5.4 (Moment matching under RCN).
Let $K\in\mathbb{Z}_{+}$ with $K\geq 8$ such that $K$ is divisible by $4$. There exists a $K\times K$ doubly stochastic matrix $\mathbf{H}$ such that the following hold:
  1. i)
    $\min_{i\neq j}\mathbf{H}_{i,i}-\mathbf{H}_{i,j}=\Omega(1/K^{3}).$
  2. ii)
    For every polynomial $p:\mathbb{R}^{2}\to\mathbb{R}$ of degree at most $K/2-2$, it holds that
    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}\left[y=i\mid\mathbf{x}\right]-{1}/{K}\right)p(\mathbf{x})\right]=0$ for all $i\in[K]$ ,
where $D$ is the distribution over $\mathbb{R}^{2}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ that satisfies Definition 4.8 for the confusion matrix $\mathbf{H}$ and the function $f$ as defined in Equation 6.
Proof.
Let $\mathbf{H}\in\mathbb{R}^{K\times K}$ be a doubly stochastic matrix and define the functions $\tilde{f}_{i}$ for $i\in[K]$ as $\tilde{f}_{i}(\mathbf{x})=\sum_{j=1}^{K}\mathbf{H}_{j,i}\mathds{1}(f(\mathbf{x})=j)$. For any function $p:\mathbb{R}^{2}\to\mathbb{R}$, we denote by $\bar{p}$ the zero mean centering of $p$, i.e., $\bar{p}(\mathbf{x})=p(\mathbf{x})-\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[p(\mathbf{x})\right]$, and by $\mathbf{v}^{(p)}\in\mathbb{R}^{K}$ the vector defined as $\mathbf{v}^{(p)}_{i}=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\bar{p}(\mathbf{x})\mathds{1}(f(\mathbf{x})=i)\right]$. We will refer to $\mathbf{v}^{{(p)}}$ as the vector of expectations of $p$. Also denote by $\boldsymbol{\omega}^{(k)}$ the root of unity vector of rate $k$, i.e. ${\boldsymbol{\omega}}^{(k)}\in\mathbb{C}^{K}$ such that $\boldsymbol{\omega}_{j}^{(k)}=e^{\frac{2\pi kji}{K}}$.
First, we show that a function $p$ has zero correlation with $\tilde{f}_{i}-1/K$ for all $i\in[K]$ under the standard Gaussian if and only if the vector of expectations of $p$ over the level sets belongs in the kernel of $\mathbf{H}^{\top}$.
Claim 5.5.
Let $p:\mathbb{R}^{2}\to\mathbb{R}$, then we have that $\mathbf{v}^{(p)}\in\ker(\mathbf{H}^{\top})\text{ if and only if }\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}[(\tilde{f}_{i}(\mathbf{x})-1/K)p(\mathbf{x})]=0\text{ for all $i\in[K]$}\;.$
Proof of Claim 5.5.
First note that from the rotational invariance of the standard Gaussian distribution we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}^{2}}[f(\mathbf{x})=i]=1/K$, for all $i\in[K]$. Hence, we have that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}[\tilde{f}_{i}(\mathbf{x})]=\sum_{j=1}^{K}\mathbf{H}_{ji}\operatorname*{\mathbf{E}}[\mathds{1}(f(\mathbf{x})=i)]=1/K$, from the assumption that $\mathbf{H}$ is column stochastic. Thus, for all $i\in[K]$ we have that
$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\left(\tilde{f}_{i}(\mathbf{x})-\frac{1}{K}\right)p(\mathbf{x})\right]$ $=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\tilde{f}_{i}(\mathbf{x})p(\mathbf{x})\right]-\frac{1}{K}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[p(\mathbf{x})\right]$
$=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\tilde{f}_{i}(\mathbf{x})p(\mathbf{x})\right]-\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[p(\mathbf{x})\right]\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\tilde{f}_{i}(\mathbf{x})\right]$
$=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\tilde{f}_{i}(\mathbf{x})\bar{p}(\mathbf{x})\right]\;.$
Therefore, substituting the definition of $\tilde{f}_{i},i\in[K],$ we have that
$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\tilde{f}_{i}(\mathbf{x})\bar{p}(\mathbf{x})\right]=\sum_{j=1}^{K}\mathbf{H}_{j,i}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\bar{p}(\mathbf{x})\mathds{1}(f(\mathbf{x})=j)\right]=(\mathbf{H}^{\top})_{i}\mathbf{v}^{(p)}\;,$
which concludes the proof. ∎
We next note that that all polynomials of degree at most $m$ can be expressed as a linear combination of polynomials of the form $(x+iy)^{a}(x-iy)^{b}$, where $a,b>0,a+b\leq m$. In the following claim, we show that the vectors associated with each polynomial of this basis are precisely the root of unity vectors.
Claim 5.6.
For $p(x,y)=(x+iy)^{a}(x-iy)^{b},a,b\in\mathbb{Z}_{+}$. If $a\neq b$, then there exists a constant $c\in\mathbb{C}$ such that $\mathbf{v}^{(p)}=c\boldsymbol{\omega}^{(a-b)}$; and if $a=b$, then $\mathbf{v}=\mathbf{0}$.
Proof of Claim 5.6.
First, note that by the definition of $f$ the region $\mathds{1}(f(x,y)=j)$ is equivalent to $\theta\in\left[{2\pi j}/{K}-{\pi}/{K},{2\pi j}/{K}+{\pi}/{K}\right]$, where $\theta=\arctan\left({y}/{x}\right)$. Fix $a,b\in\mathbb{Z}_{+},a\neq b$, and let $\theta_{1,j}={2\pi j}/{K}-{\pi}/{K}$, $\theta_{2,j}={2\pi j}/{K}+{\pi}/{K}$, for $j\in[K]$, and $p(x,y)=(x+iy)^{a}(x+iy)^{b}$. We have that
$\mathbf{v}_{j}^{(p)}=\operatorname*{\mathbf{E}}_{(x,y)\sim\mathcal{N}^{2}}[p(x,y)\mathds{1}(f(x,y)=j)]$ $=\operatorname*{\mathbf{E}}_{(x,y)\sim\mathcal{N}^{2}}[\left(x^{2}+y^{2}\right)^{\frac{a+b}{2}}e^{i\theta(a-b)}\mathds{1}(f(x,y)=j)]$
[…]
$=c\frac{1}{i(a-b)}\left(e^{i\theta_{2,j}(a-b)}-e^{i\theta_{1,j}(a-b)}\right)\;,$
[…]
Therefore, note that $\mathbf{v}_{(j+1)\mod K}^{(p)}=\mathbf{v}_{j}^{(p)}e^{\frac{2\pi i}{K}(a-b)}$ for all $j\in\{0,\dots,K-1\}$, since $\theta_{1,(j+1)}=\theta_{1,j}+{2\pi}/{K}$ and $\theta_{2,(j+1)}=\theta_{2,j}+{2\pi}/{K}$. Moreover, if $a\neq b$, we observe that the integral of $p(x,y)$ over the entire region is zero, which concludes the proof. ∎
Hence, by Claim 5.6, in order to have 00 correlation with all polynomials of degree at most $m$, we need to have that $\boldsymbol{\omega}^{(k)}\in\ker(\mathbf{H}^{\top})$ for all $k\in\{1,\dots,m,K-m,\dots,K-1\}$, as the roots of unity are symmetric, i.e., $\boldsymbol{\omega}^{(-k)}=\boldsymbol{\omega}^{(K-k)}$.
We first construct a stochastic vector $\mathbf{h}\in\mathbb{R}^{K}$ with $\mathbf{h}_{0}-\mathbf{h}_{j}=\Omega(1/K^{3})$, for all $j\in\{1,\dots,K-1\}$ that is also orthogonal to the aforementioned root of unity vectors for $m=\Omega(K)$. Next, we will extend this vector to a matrix. Define the vectors $\mathbf{h}^{(1)}=\boldsymbol{\omega}^{(0)}/K$, $\mathbf{h}^{(2)}=(\boldsymbol{\omega}^{(K/2-1)}+\boldsymbol{\omega}^{(K/2+1)})/K$ and $\mathbf{h}^{\prime}=2\mathbf{h}^{(1)}+\mathbf{h}^{(2)}$. Note that $\mathbf{h}^{(1)}_{j}=1/K$ and $\mathbf{h}_{j}^{(2)}=({2}/{K})(-1)^{j}\cos\left({2\pi j}/{K}\right)$ for all $j\in\{0,\dots,K-1\}$. Hence, we have that all the coordinates of $\mathbf{h}^{\prime}$ are nonnegative. Moreover, $\mathbf{h}^{\prime}$ attains its maximum value at $\mathbf{h}_{0}^{\prime}={4}/{K}$, since $\mathbf{h}_{K/2}^{\prime}=0$, because $(-1)^{j}\cos\left({2\pi j}/{K}\right)=-1$ when $j=K/2$ and $K$ is divisible by $4$. Furthermore, it holds that $\max_{j\in[K-1]}\mathbf{h}_{j}^{\prime}\leq 2/K+\max_{j\in[K-1]}\lvert\mathbf{h}_{j}^{(2)}\rvert=2/K+2/K\cos\left({2\pi}/{K}\right)$. As a result
$\mathbf{h}^{\prime}_{0}-\max_{j\in[K-1]}\mathbf{h}_{j}^{\prime}\geq\frac{2}{K}\left(1-\cos\left(\frac{2\pi}{K}\right)\right)\geq\frac{2}{K}\left(\left(\frac{2\pi}{K}\right)^{2}/2-\left(\frac{2\pi}{K}\right)^{4}/24\right)\gtrsim\frac{1}{K^{3}}\;,$
where the second inequality follows from the Taylor series $\cos(x)=1-\frac{x^{2}}{2!}+\frac{x^{4}}{4!}-\frac{x^{6}}{6!}+\cdots$, which is an alternating decreasing series when $x={2\pi}/{K}\leq 1$, thus it can be upper bounded by the first three terms. Let $\mathbf{h}$ be the normalization of $\mathbf{h}^{\prime}$ such that the sum of its coordinates equals $1$. Given that each element of $\mathbf{h}^{\prime}$ is at most $4/K$, we have that the normalization factor is at most $4$. Hence, $\mathbf{h}_{0}-\max_{j\in[K-1]}\mathbf{h}_{j}\gtrsim{1}/{K^{3}}$. Moreover, from the well-known fact that $\boldsymbol{\omega}^{(0)},\dots,\boldsymbol{\omega}^{(K-1)}$ form an orthogonal basis of $\mathbb{C}^{K}$, we have that $\mathbf{h}$ is orthogonal to $\boldsymbol{\omega}^{(k)}$ for all $k\in\{1,\dots,K/2-2,K/2,K/2+2,\dots,K-1\}$, since it is a linear combination of $\boldsymbol{\omega}^{(0)},\boldsymbol{\omega}^{(K/2-1)},\boldsymbol{\omega}^{(K/2+1)}$.
Now consider $\mathbf{H}$, the circulant matrix generated by row shifts of the vector $h$. By the definition of $\mathbf{H}$, we have that $\mathbf{H}$ is a row stochastic matrix and that $\min_{i\neq j}\mathbf{H}_{ii}-\mathbf{H}_{i,j}=\Omega(1/K^{3})$. Moreover, $\mathbf{H}$ is symmetric, because it holds that $\mathbf{H}_{i,j}=\mathbf{h}_{j-i\mod K}=\mathbf{h}_{i-j\mod K}=\mathbf{H}_{j,i}$ and $\mathbf{h}_{j}=\mathbf{h}_{K-j},j\in[K-1]$. This equality is justified by the identity $(-1)^{j}\cos\left({2\pi j}/{K}\right)=(-1)^{K-j}\cos\left({2\pi(K-j)}/{K}\right)$, which holds since $K$ is an even integer. Hence $\mathbf{H}$ is also column stochastic.
Finally, by A.6 we have that the eigenvectors of any circulant matrix are exactly $\boldsymbol{\omega}^{(k)}$ with eigenvalues $\mathbf{h}\cdot\boldsymbol{\omega}^{(k)},k\in\{0,\dots,K-1\}$. Hence, $\boldsymbol{\omega}^{(k)}\in\ker(\mathbf{H})=\ker(\mathbf{H}^{\top})$ for all $k\in\{1,\dots,K/2-2,K/2,K/2+2,\dots,K-1\}$. Therefore, by applying Claim 5.5, we conclude the proof of Lemma 5.4. ∎
5.2.1 Proof of Theorem 5.3
Proof of Theorem 5.3.
Consider the case where $K$ is divisible by $4$ and let $\mathbf{H}$ be a matrix that satisfies the statement of Lemma 5.4 Let $f:\mathbb{R}^{2}\to[K]$ be the linear classifier defined in Equation 6, and let $\mathcal{F}_{\mathcal{Q}}$ be the class of function defined in Equation 7, where $\mathcal{Q}\subseteq\mathbb{R}^{d\times 2}$ the family that satisfies the statement of C.4. We define the family of distributions $\mathcal{D}$ such that $D_{\mathbf{Q}}\in\mathcal{D}$ if and only if $D_{\mathbf{Q}}$ is a joint distribution $(\mathbf{x},y_{\mathbf{Q}})$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $y_{\mathbf{Q}}$ is a random variable supported on $[K]$ defined as $\operatorname*{\mathbf{Pr}}[y_{\mathbf{Q}}=i\mid\mathbf{x}]=\mathbf{H}_{f(\mathbf{Q}\mathbf{x}),i}$ for some $\mathbf{Q}\in\mathbf{\mathcal{missing}}{Q}$. Note that all distributions in $\mathcal{D}$ satisfy Definition 4.8 for confusion matrix $\mathbf{H}$ and some function in $\mathcal{F}_{\mathcal{Q}}$.
Denote by $D$ the product distribution $\mathcal{N}^{d}\times\text{Unif}([K])$. For $D_{\mathbf{Q}},D_{\mathbf{P}}\in\mathcal{D}$ with $\mathbf{Q}\neq\mathbf{P}$, we have that
$\chi_{D}(D_{\mathbf{Q}},D_{\mathbf{P}})$ $=K\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{d}}\left[(\operatorname*{\mathbf{Pr}}[y_{\mathbf{Q}}=i\mid\mathbf{x}]-\frac{1}{K})(\operatorname*{\mathbf{Pr}}[y_{\mathbf{P}}=i\mid\mathbf{x}]-\frac{1}{K})\right]$
$=K\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{d}}\left[(\mathbf{H}_{f(\mathbf{Q}\mathbf{x}),i}-\frac{1}{K})(\mathbf{H}_{f(\mathbf{P}\mathbf{x}),i}-\frac{1}{K})\right]$
$=K\sum_{i=1}^{K}\sum_{j=1}^{\infty}\left\|\mathbf{P}\mathbf{Q}^{\top}\right\|_{2}^{j}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[((\mathbf{H}_{f(\mathbf{x}),i}-\frac{1}{K})^{[j]})^{2}\right]$
$=\left\|\mathbf{P}\mathbf{Q}^{\top}\right\|_{2}^{K/2-2}K\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[(\mathbf{H}_{f(\mathbf{x}),i}-\frac{1}{K})^{2}\right]\leq d^{\Omega(K)(c-1/2)}K\;,$
where we used C.5, Parseval’s identity and the fact that $\mathbf{H}$ satisfies the statement of Lemma 5.4. Similarly, we have that $\chi_{D}(D_{\mathbf{Q}},D_{\mathbf{Q}})\leq K$ for all $D_{\mathbf{Q}}\in\mathcal{D}$.
Consequently, we have that there exists a family of distributions of size $2^{\Omega(d^{c})}$ that is $(d^{\Omega(K)(c-1/2)}K,K)$ correlated. Hence, by C.3 we have that any SQ algorithm that solves the decision problem $\mathcal{B}(\mathcal{D},D)$ requires $2^{\Omega(d^{c})}d^{\Omega(K)(c-1/2)}$ queries to the $\text{VSTAT}(1/(d^{\Omega(K)(c-1/2)}))$.
In order to complete the statement of the theorem, we need to show the following standard reduction from the testing problem to the learning problem.
Claim 5.7.
Let $\gamma=\min_{i,j}\mathbf{H}_{i,i}-\mathbf{H}_{i,j}$. Any SQ algorithm that learns multiclass linear classifiers under the standard normal in the presence of RCN with error better than $\mathrm{OPT}+\gamma/2$ can be used to solve the decision problem $\mathcal{B}(\mathcal{D},D)$ using one more query of accuracy $\gamma/4$.
Proof.
Assume that there exists an SQ algorithm $\mathcal{A}$ that if given access to a distribution $D^{\prime}$ over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is standard normal returns a hypothesis $h$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D^{\prime}}[h(\mathbf{x})\neq y]\leq\mathrm{OPT}+\gamma/4$. Using $\mathcal{A}$, we can reduce the testing problem to the learning problem.
Note that by running $\mathcal{A}$ to distribution $D_{\mathbf{P}}\in\mathcal{D}$, $\mathcal{A}$ returns a hypothesis $h$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{\mathbf{P}}}[h(\mathbf{x})\neq y]\leq\mathrm{OPT}+\gamma/4$. However, running $\mathcal{A}$ with SQ oracle access to $D$ would result in a hypothesis $h$ that cannot perform better than $1-1/K,$ as $D$ is a product distribution with the uniform distribution over the labels. Since $\mathrm{OPT}<1-1/K-\gamma/2$, because $\mathbf{H}_{i,i}\geq 1/K+\gamma/2$ as $\sum_{j}\mathbf{H}_{i,j}=1,i\in[K]$, we have that with an additional query of accuracy less than $\gamma/4$, we can solve $\mathcal{B}(\mathcal{D},D)$. ∎
For $K$ not divisible by $4$, we add $r=4\mod K$ additional classes whose weight vectors are identical to $\mathbf{w}^{(1)}$, and with no noise, i.e., for these added classes $j$, we have that $\mathbf{H}_{jj}=1$ and $\mathbf{H}_{j,i}=\mathbf{H}_{i,j}=0$ for all $j\neq i$. For the remaining classes, i.e., for $i,j\in[K-r]$, we consider a matrix $\mathbf{H}_{i,j}$ that satisfies the conditions of Lemma 5.4. Note that since these additional classes have zero probability (as ties are broken by the smallest index), our SQ hardness result extends to the general case of any $K$. ∎
Appendix
Appendix A Additional Preliminaries
Gaussian Space
We denote by $L^{2}(\mathcal{N})$ the vector space of all functions $f:\mathbb{R}^{d}\to\mathbb{R}$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f^{2}(x)]<\infty$. We define the standard $L^{p}$ norms with respect to the Gaussian measure, i.e., $\|g\|_{L^{p}}=(\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[|g(\mathbf{x})|^{p})^{1/p}$. The usual inner product for this space is $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})g(\mathbf{x})]$. While usually one considers the probabilists’ or physicists’ Hermite polynomials, in this work we define the normalized Hermite polynomial of degree $i$ to be $H_{0}(x)=1,H_{1}(x)=x,H_{2}(x)=\frac{x^{2}-1}{\sqrt{2}},\ldots,H_{i}(x)=\frac{He_{i}(x)}{\sqrt{i!}},\ldots$ where by $He_{i}(x)$ we denote the probabilists’ Hermite polynomial of degree $i$. These normalized Hermite polynomials form a complete orthonormal basis for the single-dimensional version of the inner product space defined above. To get an orthonormal basis for $L^{2}(\mathcal{N})$, we use a multi-index $V\in\mathbb{N}^{d}$ to define the $d$-variate normalized Hermite polynomial as $H_{V}(\mathbf{x})=\prod_{i=1}^{d}H_{v_{i}}(x_{i})$. The total degree of $H_{V}$ is $|V|=\sum_{v_{i}\in V}v_{i}$.
Fact A.1 (Gaussian Density Properties).
Let $\mathcal{N}$ be the standard one-dimensional normal distribution. Then, the following properties hold:
  1. 1.
    For any $t>0$, it holds $e^{-t^{2}/2}/4\leq\operatorname*{\mathbf{Pr}}_{z\sim\mathcal{N}}[z>t]\leq e^{-t^{2}/2}/2$.
  2. 2.
    For any $a,b\in\mathbb{R}$ with $a\leq b$, it holds $\operatorname*{\mathbf{Pr}}_{z\sim\mathcal{N}}[a\leq z\leq b]\leq(b-a)/\sqrt{2\pi}$.
Fact A.2 (see, e.g., Lemma 6 in [KTZ19]).
Let $f\in L^{2}(\mathbb{R}^{d},\mathcal{N}(0,I))$ with its $k$-degree Hermite expansion $f(\mathbf{x})=\sum_{\alpha\in\mathbb{N}^{d},\|\alpha\|_{1}\leq k}\widehat{f}(\alpha)H_{\alpha}(\mathbf{x})$. It holds that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[(\nabla f(\mathbf{x})\cdot\mathbf{e}_{i})^{2}\right]=\sum_{\alpha\in\mathbb{N}^{d}\|\alpha\|_{1}\leq k}\alpha_{i}(\widehat{f}(\alpha))^{2}$.
Definition A.3 (Ornstein-Uhlenbeck Noise Operator).
Let $k\in\mathbb{N}$ and $\rho\in[0,1]$. We define the Ornstein-Uhlenbeck operator $T_{\rho}:\{\mathbb{R}^{d}\to\mathbb{R}\}\to\{\mathbb{R}^{d}\to\mathbb{R}\}$ that maps $f:\mathbb{R}^{d}\to\mathbb{R}$ to the function $T_{\rho}f:\mathbb{R}^{d}\to\mathbb{R}$ with
$T_{\rho}f(x)=\operatorname*{\mathbf{E}}_{z\sim\mathcal{N}}\left[f\left(\sqrt{1-\rho^{2}}\cdot x+\rho\cdot z\right)\right].$
Fact A.4 (see, e.g., [Bog98]).
The following hold:
  1. 1.
    For any $f,g\in L_{1}$ and $\rho\in(0,1)$, it holds that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim{\cal N}}[(T_{\rho}f)g]=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim{\cal N}}[(T_{\rho}g(\mathbf{x}))f(\mathbf{x})]\;.$
  2. 2.
    For any $g\in L_{1}$, it holds:
    1. (a)
      For any $\rho\in(0,1)$, $T_{\rho}g(\mathbf{x})$ is differentiable at every point $\mathbf{x}$.
    2. (b)
      For any $\rho\in(0,1)$ the $T_{\rho}g(\mathbf{x})$ is $\|g\|_{\infty}/\rho$-Lipschitz, i.e., $\|\nabla T_{\rho}g(\mathbf{x})\|\leq\|g\|_{\infty}/\rho$ for all $\mathbf{x}\in\mathbb{R}^{d}$.
    3. (c)
      For any $p\geq 1$, $T_{\rho}$ is a contraction with respect to the norm $\|\cdot\|_{L_{p}}$, i.e., it holds $\|T_{\rho}g\|_{L^{p}}\leq\|g\|_{L^{p}}$.
Fact A.5 (Ledoux-Pisier [Pis86, Led94]).
Let $f:\mathbb{R}^{d}\mapsto\{\pm 1\}$ be a Boolean function. It holds
$\operatorname*{\mathbf{E}}_{x\sim\mathcal{N}}[f(x)T_{\rho}f(x)]\geq 1-2\sqrt{\pi}\,\Gamma(f)\,\rho.$
Fact A.6.
Let $\mathbf{c}\in\mathbb{R}^{n}$. Then the matrix
$\mathbf{M}=\begin{bmatrix}c_{0}&c_{1}&c_{2}&\cdots&c_{n-1}\\ c_{n-1}&c_{0}&c_{1}&\cdots&c_{n-2}\\ c_{n-2}&c_{n-1}&c_{0}&\cdots&c_{n-3}\\ \vdots&\vdots&\vdots&\ddots&\vdots\\ c_{1}&c_{2}&c_{3}&\cdots&c_{0}\end{bmatrix}$
given by $\mathbf{M}_{j,k}=c_{(k-j)\mod n}$ has eigenvalues
$\lambda_{j}=\mathbf{c}\cdot\boldsymbol{\omega}^{(j)}\quad j=0,1,\ldots,n-1,$
where $\boldsymbol{\omega}^{(j)}\in\mathbb{C}^{n}$ a complex vector such that $\boldsymbol{\omega}^{(j)}_{k}=e^{\frac{2\pi i}{n}jk},k=0,1,\ldots,n-1$.
The following claim bounds the Gaussian noise sensitivity of a function when we re-randomize its input over a subspace rather than over the entire space $\mathbb{R}^{d}$. This claim is used in the proof of Claim 3.6 to show that a function with bounded Gaussian surface area remains nearly invariant when its input is re-randomized over sufficiently small cubes.
Claim A.7 (Gaussian Noise Sensitivity over a Subspace).
Let $f:\mathbb{R}^{d}\to[K]$ be a a $K$-wise function and $V$ be a subspace of $\mathbb{R}^{d}$
$\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f(\mathbf{x}^{\perp V}+\mathbf{x}^{\prime V})]\lesssim\sqrt{\epsilon}\Gamma(f)\;,$
where $\mathbf{x}^{V}$ and $\mathbf{x}^{\prime V}$ are $(1-\epsilon)$-correlated standard Gaussians.
Proof.
Without loss of generality, we assume that $f$ is a Boolean function, as we can use the union bound over the level sets to reduce to this case. Let the Hermite expansion of $f$ be $f(\mathbf{x})=\sum_{S\in\mathbb{N}^{d}}\widehat{f}(S)H_{S}(\mathbf{x})$. Then we can express $f^{\prime}(\mathbf{x})=f(\mathbf{x}^{\perp V}+\mathbf{x}^{\prime V})$ in the same basis as
$f^{\prime}(\mathbf{x})=\sum_{S\in\mathbb{N}^{d}}\widehat{f}(S)H_{S}(\mathbf{x}^{\perp V}+\mathbf{x}^{\prime V})=$ $\sum_{S\in\mathbb{N}^{d}}\widehat{f}(S)H_{S}(\mathbf{x}^{\perp V})T_{\epsilon}[H_{S}(\mathbf{x}^{\prime V})]=\sum_{S\in\mathbb{N}^{d}}\epsilon^{\lvert S_{V}\rvert}\widehat{f}(S)H_{S}(\mathbf{x})\;,$
where $S_{V}$ denotes subset of $S$ that corresponds to the coordinates in $V$. Hence the sensitivity over the subspace $V$ can be written as
$\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f(\mathbf{x}^{\perp V}+\mathbf{x}^{\prime V})]$ $=\frac{1}{2}-\frac{1}{2}\operatorname*{\mathbf{E}}[f(\mathbf{x})f^{\prime}(\mathbf{x})]=\frac{1}{2}-\frac{1}{2}\sum_{S\in\mathbb{N}^{d}}\epsilon^{\lvert S_{V}\rvert}(\widehat{f}(S))^{2}$
$\leq\frac{1}{2}-\frac{1}{2}\sum_{S\in\mathbb{N}^{d}}\epsilon^{\lvert S\rvert}(\widehat{f}(S))^{2}=\frac{1}{2}-\frac{1}{2}\operatorname*{\mathbf{E}}[f(\mathbf{x})T_{\epsilon}[f(\mathbf{x})]]\;.$
Noting that the final expression is the Gaussian sensitivity, the result follows. ∎
Appendix B Omitted Content from Sections 2, 3 and 4
The following claim shows that if a matrix has a small Frobenius norm but exhibits a large quadratic form in some direction, then by selecting those eigenvectors whose eigenvalues exceed a given threshold, one obtains a short list of vectors among which at least one correlates well with that direction. This claim is used for proving Propositions 2.8 and 3.7.
Claim B.1.
Let $\mathbf{M}\in\mathbb{R}^{d\times d}$ a symmetric positive semi-definite (PSD) matrix and let $\mathbf{v}\in\mathbb{R}^{d}$ with $\left\|\mathbf{v}\right\|\leq 1$ such that $\mathbf{v}^{\top}\mathbf{M}\mathbf{v}\geq\alpha$. Then there exists a unit eigenvector $\mathbf{u}$ of $\mathbf{M}$ with eigenvalue at least $\alpha/2$ such that $\lvert\mathbf{u}\cdot\mathbf{v}\rvert\gtrsim({\alpha}/{\|\mathbf{M}\|_{F}})^{3/2}\;.$ Moreover, the number of eigenvectors of $\mathbf{M}$ with eigenvalue greater than $\alpha/2$ is at most ${4\|\mathbf{M}\|_{F}}/{\alpha^{2}}\;.$
Proof.
From the spectral theorem, let $\lambda_{1}\geq\lambda_{2}\geq\cdots\geq\lambda_{d}\geq 0$ be the eigenvalues of $\mathbf{M}$ with their corresponding orthonormal eigenvectors $\mathbf{x}^{(1)},\mathbf{x}^{(2)},\ldots,\mathbf{x}^{(d)}\in\mathbb{R}^{d}$. Since $\mathbf{M}$ is symmetric and psd, it can be decomposed as $\mathbf{M}=\sum_{i=1}^{d}\lambda_{i}\mathbf{x}^{(i)}\mathbf{(}\mathbf{x}^{(i)})^{\top}.$ Moreover, we can write $\mathbf{v}$ in the eigenbasis of $\mathbf{M}$, i.e., $\mathbf{v}=\sum_{i=1}^{d}(\mathbf{v}\cdot\mathbf{x}^{(i)})\mathbf{x}^{(i)}.$ Therefore, we have that $\mathbf{v}^{\top}\mathbf{M}\mathbf{v}=\sum_{i=1}^{d}\lambda_{i}(\mathbf{v}\cdot\mathbf{x}^{(i)})^{2}$, and using that $\mathbf{v}^{\top}\mathbf{M}\mathbf{v}\geq\alpha$, we get that $\sum_{i=1}^{n}\lambda_{i}(\mathbf{v}\cdot\mathbf{x}^{(i)})^{2}\geq\alpha$.
Let $S=\{i:\lambda_{i}\geq\frac{\alpha}{2}\}$. Using the fact that $\sum_{i=1}^{d}(\mathbf{v}\cdot\mathbf{x}^{(i)})^{2}\leq 1$, we have that
$\sum_{i=1}^{d}\lambda_{i}(\mathbf{v}\cdot\mathbf{x}^{(i)})^{2}=\sum_{i\not\in S}\lambda_{i}(\mathbf{v}\cdot\mathbf{x}^{(i)})^{2}+\sum_{i\in S}\lambda_{i}(\mathbf{v}\cdot\mathbf{x}^{(i)})^{2}\leq\frac{\alpha}{2}+\sum_{i\in S}\lambda_{i}(\mathbf{v}\cdot\mathbf{x}^{(i)})^{2}\;.$
Hence, we have that $\sum_{i\in S}\lambda_{i}(\mathbf{v}\cdot\mathbf{x}^{(i)})^{2}\geq\frac{a}{2}.$ Furthermore, note that there must be at least one index $j\in S$ such that $\lambda_{j}(\mathbf{v}\cdot\mathbf{x}^{(j)})^{2}\geq\frac{a}{2|S|}.$ Also, from the fact that $\lambda_{j}\leq\|\mathbf{M}\|_{F}$, we have that $(\mathbf{v}\cdot\mathbf{x}^{(j)})^{2}\geq\frac{a}{2|S|\|\mathbf{M}\|_{F}}.$ Taking the sum of the squares of the eigenvalues in $S$, we get that
$|S|\cdot\left(\frac{\alpha}{2}\right)^{2}\leq\sum_{i\in S}\lambda_{i}^{2}\leq\|\mathbf{M}\|_{F}^{2}\Rightarrow|S|\leq\frac{4\|\mathbf{M}\|_{F}^{2}}{\alpha^{2}}\;.$
Combining the above, we have that $\mathbf{v}\cdot\mathbf{x}^{(j)}\gtrsim(\alpha/\|\mathbf{M}\|_{F})^{3/2}$, for some $j\in S$, and $\lvert S\rvert\leq 4\|\mathbf{M}\|_{F}/\alpha^{2}$, which completes the proof of Claim B.1. ∎
The following two facts are important properties of polynomial regression which are be used in Proposition 3.7 for the analysis of our general algorithm. B.2 shows that by using a sufficiently small number of samples in polynomial time, one can efficiently obtain a low-degree polynomial whose $L_{2}$ approximation error is within an additive $\epsilon$ of the best achievable by any polynomial of the same degree. B.3 shows that the subspace spanned by the eigenvectors of a low-degree regression polynomial contains only a few eigenvectors corresponding to large eigenvalues.
Fact B.2 (see, e.g., Lemma 3.3 in [DKK+21a]).
Let $\mathcal{D}$ be a distribution on $\mathbb{R}^{d}\times\{\pm 1\}$ whose x-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$. Let $k\in\mathbb{Z}_{+}$ and $\epsilon,\delta>0$. There is an algorithm that draws $N=(dk)^{O(k)}\log(1/\delta)/\epsilon^{2}$ samples from $\mathcal{D}$, runs in time $\text{poly}(N,d)$, and outputs a polynomial $P(\mathbf{x})$ of degree at most $k$ such that
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\mathcal{D}}[(y-P(\mathbf{x}))^{2}]\leq\min_{P^{\prime}\in\mathcal{P}_{k}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\mathcal{D}}[(y-P^{\prime}(\mathbf{x}))^{2}]+\epsilon,$
with probability $1-\delta$.
Fact B.3 (see, e.g., Lemma 3.3 in [DKK+21a]).
Fix $\epsilon\in(0,1)$ and let $P(\mathbf{x})$ be a degree-$k$ polynomial, such that
$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\mathcal{D}}[(y-P(\mathbf{x}))^{2}]\leq\min_{P^{\prime}\in\mathcal{P}_{k}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\mathcal{D}}[(y-P^{\prime}(\mathbf{x}))^{2}]+O(\epsilon).$
Let $\mathbf{M}=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{D}_{\mathbf{x}}}[\nabla P(\mathbf{x})\nabla P(\mathbf{x})^{\top}]$ and $V$ be the subspace spanned by the eigenvectors of $\mathbf{M}$ with eigenvalues larger than $\eta$. Then the dimension of the subspace $V$ is $\dim(V)=O(k/\eta)$ and moreover $\mathrm{tr}(\mathbf{M})=O(k)$.
B.4 shows that for an intersection of halfspaces, the variance in any defining direction is reduced from one by an amount that depends on how compressed the function is along that direction. This fact is the key ingredient in our structural result on intersections of halfspaces (see Proposition 4.4).
Fact B.4 (Gaussian Monotonicity, [Vem10a]).
Let $\mathcal{K}\subseteq\mathbb{R}^{d}$ be an intersection of halfspaces and let $\mathcal{N}_{d}|_{\mathcal{K}}$ be the truncation of the standard Gaussian distribution in $d$ dimensions $\mathcal{N}_{d}$ to $\mathcal{K}$. For any $u\in\mathbb{S}^{d-1}$, we have $\operatorname{Var}_{\mathbf{x}\sim\mathcal{N}_{d}|_{\mathcal{K}}}(\mathbf{u}\cdot\mathbf{x})\leq 1$. Moreover, if for some $T\in\mathbb{R}$ the halfspace $\{\mathbf{x}:\mathbf{u}\cdot\mathbf{x}+T\geq 0\}$ is one of the defining halfspaces of the intersection then, we have variance reduction along $u$, i.e., $\operatorname{Var}_{\mathbf{x}\sim\mathcal{N}_{d}|_{\mathcal{K}}}(\mathbf{u}\cdot\mathbf{x})\leq 1-\frac{1}{2}e^{-\frac{1}{2}\max\{0,T\}^{2}}$ for a sufficiently large universal constant $C>0$.
Appendix C Basics on Statistical Query Model
In the following definition, we state the notion of pairwise correlation between two distributions with respect to a third distribution, and then we introduce the concept of a $(\gamma,\beta)$-correlated family.
Definition C.1.
[…]
The following notion of statistical dimension, introduced in [FGR+13], effectively characterizes the difficulty of the decision problem.
Definition C.2.
For $\gamma,\beta>0$, a decision problem $\mathcal{B}(\mathcal{D},\mathcal{D})$, where $D$ is fixed and $\mathcal{D}$ is a family of distributions over $\mathcal{X}$, let $s$ be the maximum integer such that there exists $\mathcal{D}\subseteq\mathcal{D}$ such that $\mathcal{D}$ is $(\gamma,\beta)$-correlated relative to $D$ and $|\mathcal{D}|\geq s$. We define the Statistical Query dimension with pairwise correlations $(\gamma,\beta)$ of $\mathcal{B}$ to be $s$ and denote it by $SD(\mathcal{B},\gamma,\beta)$.
The connection between SQ dimension and lower bounds is captured by the following lemma.
Fact C.3 ([FGR+13]).
Let $\mathcal{B}(\mathcal{D},D)$ be a decision problem, where $D$ is the reference distribution and $\mathcal{D}$ is a class of distributions over $\mathcal{X}$. For $\gamma,\beta>0$, let $s=SD(\mathcal{B},\gamma,\beta)$. Any SQ algorithm that solves $\mathcal{B}$ with probability at least $2/3$ requires at least $s\cdot\gamma/\beta$ queries to the VSTAT(1/$\gamma$) oracle.
In order to construct a large set of nearly uncorrelated hypotheses, we need the following fact:
Fact C.4 (see, e.g., Lemma 2.5 in [DKPZ21]).
Let $0<a,c<\frac{1}{2}$ and $m,n\in\mathbb{Z}_{+}$ such that $m\leq n^{a}$. There exists a set $S$ of $\Omega(n^{c})$ matrices in $\mathbb{R}^{m\times n}$ such that every $\mathbf{U}\in S$ satisfies $\mathbf{U}\mathbf{U}^{\top}=\mathbf{I}_{m}$ and every pair $\mathbf{U},\mathbf{V}\in S$ with $\mathbf{U}\neq\mathbf{V}$ satisfies $\|\mathbf{UV}^{\top}\|_{F}\leq O(n^{2c-1+2a})$.
Furthermore, to bound the correlation between two rotated versions of the same function, we will use the following fact, which relates this correlation to the Hermite coefficients of the function and the angle between the rotations.
Fact C.5 (Lemma 2.3 in [DKPZ21]).
Let $g:\mathbb{R}^{m}\to\mathbb{R}$ and $\mathbf{U},\mathbf{V}\in\mathbb{R}^{m\times n}$ be linear maps such that $\mathbf{U}\mathbf{U}^{\top}=\mathbf{V}\mathbf{V}^{\top}=\mathbf{I}_{m}$. Then, we have that
$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}_{n}}\left[g(\mathbf{U}\mathbf{x})g(\mathbf{V}\mathbf{x})\right]\leq\sum_{t=0}^{\infty}\|\mathbf{UV}^{\top}\|_{2}^{t}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}_{m}}\left[(g^{[t]}(\mathbf{x}))^{2}\right],$
where $g^{[t]}$ denotes the degree-$t$ Hermite part of $g$.

### v2

@@
 We consider a class of well-behaved MIMs with finite ranges that satisfy certain regularity properties.
-Our main contribution is a general noise-tolerant learning algorithm for this class that is qualitatively optimal in the Statistical Query (SQ) model.
+Our main contribution is a general noise-tolerant learning algorithm for this class whose complexity is qualitatively optimal in the Statistical Query (SQ) model.
 At a high-level, our algorithm attempts to iteratively construct better approximations to the defining subspace by computing low-degree moments of our function conditional on its projection to the subspace computed thus far, and adding directions with relatively large empirical moments.
-For our class of well-behaved MIMs, we show that this procedure efficiently finds a subspace $V$ so that $f(\mathbf{x})$ is close to a function of the projection of $\mathbf{x}$ onto $V$, which can then be found by brute-force.
+For well-behaved MIMs, we show that this procedure efficiently finds a subspace $V$ so that $f(\mathbf{x})$ is close to a function of the projection of $\mathbf{x}$ onto $V$, which can then be found by brute-force.
 Conversely, for functions for which these conditional moments do not necessarily help in finding better subspaces, we prove an SQ lower bound providing evidence that no efficient algorithm exists.
@@
 Before we describe our general setting and algorithm, we start with a specific basic model that was the original motivation for this work.
-Motivation: Multiclass Linear Classification.
-A multiclass linear classifier is a function $f:\mathbb{R}^{d}\to[K]$ of the form $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})\;,$ where $\mathbf{w}^{(i)}\in\mathbb{R}^{d}$ and $t_{i}\in\mathbb{R}$.
-Where ties are resolved by selecting the smallest index.
+Motivation: Multiclass Linear Classification
+A multiclass linear classifier is a function $f:\mathbb{R}^{d}\to[K]$ of the form $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})\;,$ where $\mathbf{w}^{(i)}\in\mathbb{R}^{d}$ and $t_{i}\in\mathbb{R}$ (ties are resolved by selecting the smallest index).
 Multiclass linear classification (MLC) is a textbook machine learning problem [SB14], that has also been extensively studied in the online learning literature; see, e.g., [KSST08] and references therein.
-In the realizable PAC setting (i.e., with clean labels), this task is efficiently solvable to 0-1 error $\epsilon$ with sample size $O(dK/\epsilon)$ via the multiclass Perceptron algorithm.
-On the other hand, despite extensive effort, see, e.g., [NDRT13, WLT18, LWS18, SKP+20, GWBL20], our understanding of the computational complexity of learning is poor in the noisy setting—even for random label noise.
+In the realizable PAC setting (i.e., with clean labels), this task is efficiently solvable to any desired error $\epsilon$ with sample size $O(dK/\epsilon)$ via linear programming (or via the multiclass Perceptron algorithm for well-behaved marginal distributions, including the Gaussian).
+On the other hand, despite extensive effort, see, e.g., [NDRT13, WLT18, LWS18, SKP+20, GWBL20], our understanding of the computational complexity of learning is poor in the noisy setting—even for the simplest case of random label noise.
 The special case of binary classification of Linear Threshold Functions (LTFs) or halfspaces—corresponding to $K=2$—is fairly well-understood.
-Specifically, sample and computationally efficient learning algorithms are known, even in the presence of noisy labels [ABL17, DKS18, DGT19, DKTZ20a, DKTZ20b, DKK+21b, DKK+22].
+Specifically, sample and computationally efficient learning algorithms are known, even in the presence of partially corrupted labels [ABL17, DKS18, DGT19, DKTZ20a, DKTZ20b, DKK+21b, DKK+22].
 Unfortunately, these prior approaches are specific to the binary case.
-In particular, it is not clear whether they can be generalized to give non-trivial efficient algorithms for the noisy multiclass setting.
+In particular, it is unclear whether they can be generalized to give any non-trivial guarantees for the noisy multiclass setting.
 Perhaps the simplest model of random label noise is that of Random Classification Noise (RCN) [AL88].
@@
 A classical result [Kea98] shows that, in the binary classification setting, any efficient Statistical Query (SQ) algorithm can be transformed into an efficient RCN-tolerant PAC learner (with a polynomial dependence in all relevant parameters).
-Unfortunately, no such result is known for the multiclass setting.
-The natural (and well-studied) generalization of the RCN model to the multiclass setting is the following [PRKM+17, ZLA21, VRW17].
+Unfortunately, no such result is known for the multiclass setting [^1] In fact, Theorem 1.9 implies that such a generalization is impossible without strong assumptions on the noise matrix..
+The standard generalization of the RCN model to the multiclass setting is the following [PRKM+17, ZLA21, VRW17].
 Let $D$ be the distribution on noisy samples $(\mathbf{x},y)\in\mathbb{R}^{d}\times[K]$, where in our setting the marginal distribution $D_{\mathbf{x}}$ is the standard Gaussian, and $f:\mathbb{R}^{d}\to[K]$ be the ground truth classifier.
@@
 Prior algorithmic work using gradient-based methods to achieve sample and time complexity that scales polynomially with $1/\sigma$, where $\sigma$ is the minimum singular value of $\mathbf{H}$ (see [LWS18]).
-This quantity could be arbitrarily small or even zero, hence such approaches do not in general lead to efficient algorithms.
+This quantity could be arbitrarily small or even zero, hence such approaches do not in general lead to a computationally efficient algorithm.
 This discussion leads to the following question:
@@
 As an application of our new methodology, we show that this is indeed possible (see Theorem 1.11).
-A related question is to understand the algorithmic possibilities and limitations of learning MLCs in the presence of adversarial label noise.
+An even more challenging goal is to understand the algorithmic possibilities and limitations of learning MLCs in the presence of adversarial label noise.
 In the adversarial label noise model, an adversary is allowed to arbitrarily corrupt an unknown $\mathrm{OPT}$ fraction of the labels, for some parameter $\mathrm{OPT}<1/2$, and the goal of the learner is to compute a hypothesis whose 0-1 error is competitive with $\mathrm{OPT}$ (see Definition 1.4).
@@
 Unfortunately, even for $K=2$, known computational lower bounds rule out the existence of such algorithms with complexity better than $d^{\mathrm{poly}(1/\epsilon)}$ [DKPZ21, DKR23].
-To achieve a fixed polynomial runtime in $d$, one needs to relax the error guarantee.
-In particular, for $K=2$, efficient algorithms are known with complexity $\mathrm{poly}(d/\epsilon)$ and error guarantee $O(\mathrm{OPT})+\epsilon$.
-A natural question is whether qualitatively similar algorithms exist for the multiclass setting.
-Specifically, a first goal is to establish a fixed-parameter tractability guarantee:
+Consequently, to achieve a fixed polynomial runtime in $d$, one needs to relax the error guarantee.
+For the special case $K=2$, such algorithms are known.
+In particular, [DKS18] gave a $\mathrm{poly}(d/\epsilon)$ time algorithm with error guarantee $C\cdot\mathrm{OPT}+\epsilon$, where $C$ is a universal constant.
+A natural question is whether qualitatively similar algorithms exist for the multiclass setting. Specifically, a first goal is to establish a fixed-parameter tractability guarantee:
 Is there a constant-factor approximate learner for MLC with complexity $\mathrm{poly}(d)f(\epsilon,K)$?
@@
 Equivalently, $f$ is of the form $f(\mathbf{x})=g(\mathbf{W}\mathbf{x})$ for a link function $g:\mathbb{R}^{K}\to\cal{Y}$ and a $K\times d$ weight matrix $\mathbf{W}$.
-The link function $g$ could be either a priori known or unknown to the learner.
-In the latter case, it is typically assumed to satisfy certain regularity conditions.
-Moreover, the dimension $K$ of the hidden subspace is usually significantly smaller than the ambient dimension $d$.
+In typical applications, the dimension $K$ of the hidden subspace is usually significantly smaller than the ambient dimension $d$.
+In regards to the problem of learning MIMs from labeled examples, there two qualitatively different versions: the link function $g$ in Definition 1.1 could be a priori known or unknown to the learner.
+In the case of unknown $g$, it is typically assumed that it satisfies certain regularity conditions; as without any assumptions on $g$, learning is information-theoretically impossible.
+In this work, we focus on the setting where $g$ is unknown.
 Multi-index models [FJS81, Hub85, Li91, HL93, XTLZ02, Xia08] capture the common assumption that real-world data are inherently structured, in the sense that the relationship between the features is of a (hidden) lower-dimensional nature.
@@
 1.1 Our Results
+In this section, we present our main results. For a concise overview of our algorithmic contributions, see Table 1.
+Function Class Prior Work Our Work
+Runtime Error Runtime Error
+[…]
+$K$-MLC with RCN $d^{\mathrm{poly}(K/\epsilon)}$ $\mathrm{OPT}+\epsilon$ (Theorem 1.11) $\mathrm{poly}(d)(1/\epsilon)^{\mathrm{poly}(K)}$ $O(\mathrm{OPT})+\epsilon$
+[…]
+[…]
+Table 1: Comparison of runtime and error guarantees between state-of-the-art efficient algorithms (see [KOS08, DKS18] for intersections of $K$ halfspaces) and our subspace approximation procedure.
 1.1.1 Learning Well-behaved Multi-Index Models with Noise
-Our general algorithmic result applies to MIMs with finite range that satisfy appropriate regularity properties.
+Our general algorithmic result (Theorem 1.5) applies to MIMs with finite range that satisfy a collection of regularity properties (Definition 1.2).
 Throughout this section, we will use $d$ for the dimension of the feature space and $K$ for the dimension of the hidden low-dimensional subspace.
-A standard assumption in the literature on MIMs, satisfied by a wide range of natural concept classes, is that the underlying function has bounded Gaussian surface area (GSA)—a complexity measure of the decision boundary (see Definition 3.1 for the definition in the multiclass case).
+A standard assumption in the literature on MIMs (see, e.g., [KOS08, DKK+23]), satisfied by a wide range of natural concept classes, is that the underlying function has bounded Gaussian surface area (GSA)—a complexity measure of the decision boundary (see Definition 3.1 for the definition in the multiclass case).
 We denote by $\Gamma$ the upper bound on the GSA of the underlying concept class.
-The key structural property that we will assume for every MIM function $f$ in our concept class is, roughly speaking, the following: it is not possible to add noise to $f$ to conceal crucial directions so that they remain invisible to moment-based detection.
+The structural property that we will assume for every MIM function $f$ in our class is, roughly speaking, the following: it is not possible to add noise to $f$ to conceal crucial directions so that they remain invisible to moment-based detection.
 Specifically, for any subspace $V$, one of the following must hold:
@@
     Otherwise, there is at least one point in $V$ such that when we condition on $\mathbf{x}^{V}$ near that point, the conditional distribution of a “noisy” version of $f$ exhibits a distinguishing moment of degree at most $m$.
-That is, there exists a low-degree polynomial whose conditional expectation, conditioned on $\mathbf{x}^{V}$, cannot remain hidden; and hence reveals a dependence on directions outside the subspace $V$.
+That is, there exists a degree-$m$ polynomial whose conditional expectation, conditioned on $\mathbf{x}^{V}$ (the projection of $\mathbf{x}$ on $V$), cannot remain hidden; and hence reveals a dependence on directions outside the subspace $V$.
+The parameter $m$ above turns out to essentially characterize the complexity of the learning problem.
+We show, qualitatively speaking, that there exists an algorithm with complexity $d^{m}$; and that this bound is essentially best possible in the SQ model.
 We now present an informal definition of this property and refer the reader to Definition 3.2 for a more precise formulation.
 Definition 1.2 (Well-Behaved Multi-Index Model).
-Fix $m\in\mathbb{Z}_{+},\zeta,\tau\in(0,1)$ and $\mathcal{Y}\subseteq\mathbb{Z}$ of finite cardinality.
-We say that a $K$-MIM $f:\mathbb{R}^{d}\to\mathcal{Y}$ with GSA at most $\Gamma$ is is $(m,\zeta,\tau,\Gamma)$-well-behaved if for any random variable $y(\mathbf{x})$ supported on $\mathcal{Y}$ with $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq y(\mathbf{x})]\leq\zeta$ and any subspace $V\subseteq\mathbb{R}^{d}$, one of the following conditions holds:
+Fix $m\in\mathbb{Z}_{+},\zeta,\tau\in(0,1)$, $\Gamma>0$, and $\mathcal{Y}\subseteq\mathbb{Z}$ a set of finite cardinality.
+We say that a $K$-MIM $f:\mathbb{R}^{d}\to\mathcal{Y}$ is $(m,\zeta,\tau,\Gamma)$-well-behaved if $f$ has Gaussian surface area at most $\Gamma$, and for any random variable $y(\mathbf{x})$ supported on $\mathcal{Y}$ with $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq y(\mathbf{x})]\leq\zeta$ and any subspace $V\subseteq\mathbb{R}^{d}$, one of the following conditions holds:
   1. a)
@@
     $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[p(\mathbf{x}^{U})\mathds{1}(y(\mathbf{x})=z)\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\geq\mathrm{poly}(\zeta/m)\;,$
-    where $U=(V+W)\cap V^{\perp}$ and $W$ is the hidden subspace defining $f$.
+    where $U=(V+W)\cap V^{\perp}$ is the projection of $W$ onto $V^{\perp}$ and $W$ is the subspace defining $f$.
 We denote by $\mathcal{F}$ the class of $(m,\zeta,\tau,\Gamma)$-well-behaved $K$-MIMs over $\mathbb{R}^{d}$.
@@
 It is worth pointing out that the sample and computational complexity of our algorithm will scale with $d^{m}$.
+Remark 1.3.
+Definition 1.2 implicitly introduces a complexity measure for learning a family of discrete MIMs.
+Let $\mathcal{C}$ be a family of MIMs with Gaussian Surface Area bounded by $\Gamma$, and let $\zeta,\tau>0$ be parameters.
+Define $m^{\ast}$ to be the smallest integer for which every $f\in\mathcal{C}$ is $(m^{\ast},\zeta,\tau,\Gamma)$-well-behaved.
+It turns out that the quantity $m^{\ast}$ is (essentially) a generalization of the generative exponent [DPVLB24], a measure characterizing the SQ complexity of learning Single-Index Models (SIMs).
+For further details, see Appendix D.
 As we will show in the next subsection, fundamental and well- studied concept classes—specifically, multiclass linear classifiers and intersections of halfspaces—satisfy Definition 1.2 for reasonable values of the underlying parameters (in particular, with $m\leq 2$).
@@
 Let $\mathcal{Y}$ be finite and $\cal{C}$ a class of functions from $\mathbb{R}^{d}$ to $\mathcal{Y}$.
-Given $\epsilon>0$ and i.i.d.
+Given i.i.d.
 samples $(\mathbf{x},y)$ from a distribution $D$ on $\mathbb{R}^{d}\times\mathcal{Y}$ such that the marginal $D_{\mathbf{x}}$ is the standard normal and no assumptions are made on the labels, the goal is to output a hypothesis $h:\mathbb{R}^{d}\to\mathcal{Y}$ that, with high probability, has small 0-1 error $\mathrm{err}_{0-1}^{D}(h)\coloneqq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]$, as compared to the optimal 0-1 error of the class ${\cal C}$, i.e., $\mathrm{OPT}\coloneqq\inf_{c\in\mathcal{C}}\mathrm{err}_{0-1}^{D}(c)$.
@@
 At a high-level, the main advantage of the above algorithm is that—as long as $m=O(1)$—its complexity is bounded by a constant-degree polynomial in $d$.
-Such a bound can be viewed as a significant qualitative improvement over $d^{K}$, which may be attainable using simpler methods.
 In other words, we give a fixed-parameter tractability result for the task of learning well-behaved MIMs.
@@
 In more detail, we show (see Theorem 5.2 for a more detailed formal statement).
-Theorem 1.5 (SQ Lower Bound for Learning Not-Well-behaved MIMs).
-Suppose that a function class $\mathcal{G}$ is rotationally invariant and contains a bounded surface area MIM not contained in the function class $\mathcal{F}$ of Definition 1.2 for parameters $m,\tau$.
-Then, any SQ algorithm that learns $\mathcal{G}$ within error smaller than $O(\tau)$ has complexity at least $d^{\Omega(m)}$.
+Theorem 1.6 (SQ Lower Bound for Learning Not-Well-behaved MIMs, Informal).
+Suppose that a function class $\mathcal{G}$ is rotationally invariant and contains a bounded surface area MIM not contained in the function class $\mathcal{F}$ of Definition 1.2 for parameters $m,\tau,\zeta$.
+Then, any SQ algorithm that learns $\mathcal{G}$ in the presence of $\zeta$ label noise to error smaller than $\tau-O(\zeta)$ has complexity at least $d^{\Omega(m)}$.
+Remark 1.7 (Characterization).
+Theorems 1.5 and 1.6 together essentially characterize the complexity of learning a MIM family.
+In particular, let $\mathcal{C}$ be a rotationally invariant family of MIMs with Gaussian Surface Area bounded by $\Gamma$, and let $\zeta,\tau>0$ be parameters.
+Define $m^{\ast}$ to be the smallest integer for which every $f\in\mathcal{C}$ is $(m^{\ast},\zeta,\tau,\Gamma)$-well-behaved.
+Then there is an SQ algorithm that learns a (slightly less than) $\zeta$-noisy function from $\mathcal{C}$ to error $\tau+O(\zeta)$ with resources $d^{m^{\ast}}$ (times some function of the other parameters), but not an SQ algorithm that learns a $\zeta$-noisy function from $\mathcal{C}$ to error $\tau-O(\zeta)$ with resources $d^{o(m^{\ast})}$.
 Finally, we note that a variant of our general algorithm learns well-behaved MIMs in the presence of RCN with significantly improved complexity.
@@
 In this subsection, we present algorithmic applications of our general MIM learner to two fundamental and well-studied concept classes.
-Multiclass Linear Classifiers.
+Multiclass Linear Classifiers
 We denote by $\mathcal{L}_{d,K}$ the family of $K$-class linear classifiers in $\mathbb{R}^{d}$. For the adversarial noise setting, we obtain the following result (see Theorem 2.1):
@@
 […]
-Theorem 1.6 is obtained from Theorem 1.4 via a new structural result that we establish, roughly showing that the class $\mathcal{L}_{d,K}$ is $(1,\zeta,O(\zeta),O(K))$-well behaved for any $\zeta\in(0,1)$ (see Proposition 2.4).
+Theorem 1.8 can be derived from Theorem 1.5 via a new structural result that we establish, roughly showing that the class $\mathcal{L}_{d,K}$ is $(1,\zeta,O(\zeta),O(K))$-well behaved for any $\zeta\in(0,1)$ (see Proposition 2.4).
+For the sake of exposition, in Section 2 we give a direct algorithm establishing Theorem 1.8.
 Theorem 1.8 gives the first constant-factor approximate learner for multiclass linear classification with adversarially corrupted labels, whose complexity is a fixed-degree polynomial in $d$.
-We provide some remarks to interpret the statement of Theorem 1.6.
+We provide some remarks to interpret this statement.
 First, even for $K=2$, known hardness results imply that achieving error $\mathrm{OPT}+\epsilon$ requires $d^{\mathrm{poly}(1/\epsilon)}$ time [DKPZ21, DKR23].
@@
 For any $\epsilon\leq 1/K^{C}$, where $C$ a sufficiently large universal constant, any SQ algorithm that learns $\mathcal{L}_{d,K}$ under the Gaussian distribution in the presence of RCN to error $\mathrm{OPT}+\epsilon$, has complexity at least $d^{\Omega(K)}$.
+Digression: SQ Lower Bound for MLC with Partial Labels
+As an additional implication of our SQ lower bound technique establishing Theorem 1.9, we establish the first super-polynomial SQ lower bound for the well-studied task of learning multiclass linear classifiers with partial (or coarse) labels.
+This is a classical and widely studied model of misspecification in multi-class classification [CST11, LD14, VRW17, FKKT21].
+In this model, instead of observing instance-label pairs $(\mathbf{x}^{(i)},f(\mathbf{x}^{(i)}))_{i=1}^{m}$, we observe instance-label subsets $(\mathbf{x}^{(i)},S_{i})_{i=1}^{m}$, where each $S_{i}\subseteq\mathcal{Y}$ and the true label always belongs to the observed subset (i.e., $f(\mathbf{x}^{(i)})\in S_{i}$).
+The goal is to find a hypothesis with error $\epsilon$ with respect to the original labels.
+A necessary assumption for identifiability in this model is that any incorrect label does not appear too frequently in the observed subset (as quantified by the parameter $\gamma$ in the theorem statement below); see Definition E.1 for a formal definition.
+Under this assumption, the sample complexity of ERM for this task is $O(dK/(\gamma\epsilon))$ [LD14].
+On the computational front, we establish the following SQ lower bound (see Theorem E.2):
+Theorem 1.10 (SQ Lower Bound for MLC with Partial Labels).
+Any SQ algorithm that learns $\mathcal{L}_{d,K}$ given partial labels, with $\gamma=\mathrm{poly}(1/K)$, under the Gaussian distribution to error $O(1/K)$, has complexity $d^{\Omega(K)}$.
+The details are deferred to Appendix E.
+We now return to our algorithmic contributions.
 Interestingly, using a variant of our general MIM algorithm, we obtain an RCN-tolerant learner with $O(\mathrm{OPT})+\epsilon$ error, whose complexity is qualitatively better than our agnostic learner.
-Specifically, we show the following (see Theorem 4.9 for the detailed statement):
+Specifically, we show the following (see Theorem 4.9):
 Theorem 1.11 (Learning $\mathcal{L}_{d,K}$ with RCN).
@@
 labeled samples, runs in time $\mathrm{poly}(d,N)$, and returns a hypothesis $h$ such that with high probability $\mathrm{err}_{0-1}^{D}(h)\leq O(\mathrm{OPT})+\epsilon$.
-Intersections of Halfspaces.
+Remark 1.12.
+Recent work [DMRT25] has shown that the above error guarantee is not possible for efficient SQ algorithms in the distribution-free setting.
+Specifically, any SQ algorithm for MLC with RCN that achieves any constant factor approximation requires super-polynomial in $d$ time.
+This hardness result further motivates the study of distribution-specific learning in the multiclass setting, even with random noise only.
+Intersections of Halfspaces
 An intersection of $K$ halfspaces is any function $f:\mathbb{R}^{d}\to\{\pm 1\}$ for which there exist $K$ halfspaces $h_{i}(\mathbf{x})=\mathrm{sign}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$ such that $f(\mathbf{x})=1$ if and only if $h_{i}(\mathbf{x})=1$ for all $i\in[K]$.
@@
 Theorem 1.13 is obtained from Theorem 1.5 via a new structural result, roughly showing that the class $\mathcal{F}_{d,K}$ is $(2,\zeta,K\tilde{O}(\zeta),O(\sqrt{\log(K)}))$-well behaved for any $\zeta\in(0,1)$ (see Proposition 4.2).
-Prior work [KKMS05, KOS08] demonstrated that the class $\mathcal{F}_{d,K}$ is agnostically learnable under the Gaussian distribution, within optimal error of $\mathrm{OPT}+\epsilon$, using $d^{O({\log K}/{\epsilon^{4}})}$ samples and runtime.
-This complexity bound was recently shown to be near-optimal in the SQ model [DKPZ21, HSSV22].
-A line of subsequent work has focused on understanding the tradeoff between accuracy and computational complexity, and in particular on designing faster algorithms with relaxed error guarantees.
-Specifically, [DKS18] developed a $\mathrm{poly}(d)(1/\epsilon)^{\mathrm{poly}(K)}$ time algorithm with error of $\mathrm{poly}(K)\tilde{O}(\mathrm{OPT}^{1/11})$.
-Our algorithm achieves a different tradeoff: its error guarantee is significantly better compared to [DKS18]—namely nearly-optimal as a function of $\mathrm{OPT}$—while its computational complexity incurs a $2^{\mathrm{poly}(K/\epsilon)}$ dependence (as opposed to $(1/\epsilon)^{\mathrm{poly}(K)})$).
+To highlight the significance of Theorem 1.13, some remarks are in order.
+Note that the information-theoretically optimal error in our setting is $\mathrm{OPT}+\epsilon$.
+The only known algorithm achieving this guarantee is the $L_{1}$ polynomial regression [KKMS05, KOS08], which has complexity $d^{O({\log K}/{\epsilon^{4}})}$ for the class $\mathcal{F}_{d,K}$.
+(This complexity bound was recently shown to be near-optimal in the SQ model [DKPZ21, HSSV22].)
+A fundamental open question in the literature has been whether one can achieve near-optimal error guarantees with qualitatively improved sample and computational complexity.
+Theorem 1.13 answers this question in the affirmative.
+Specifically, for $K=O(1)$, it achieves near-optimal error guarantee (up to a $O(\log(1/\mathrm{OPT}))$ factor), and its complexity has fixed-polynomial dependence in the dimension (namely, $d^{2}$ as opposed to $d^{f(k,1/\epsilon)}$).
+In comparison, [DKS18] developed a $\mathrm{poly}(d)(1/\epsilon)^{\mathrm{poly}(K)}$ time algorithm, achieving an error guarantee of $\mathrm{poly}(K)\tilde{O}(\mathrm{OPT}^{1/11})$.
+The key distinction is that the error guarantee achieved by the [DKS18] algorithm is highly suboptimal as a function of $\mathrm{OPT}$ (by a large polynomial factor).
+Additionally, the dependence on $K$ in our Theorem 1.13 is significantly better by a polynomial factor.
 1.2 Overview of Techniques
-Prior Techniques and Obstacles for Learning Noisy Multi-Index Models.
-Recall that our goal is, given access to noisy samples from well-behaved MIM with finite label set, to learn the underlying function $f$ to high accuracy.
+Prior Techniques and Obstacles for Learning Multi-Index Models
+Recall that our goal is, given access to noisy samples from a well-behaved MIM with finite label set, to learn the underlying function $f$ to high accuracy.
 We will use $W$ to denote the hidden $K$-dimensional subspace to whose projection $f$ depends.
@@
 Unfortunately, showing that the moments of $f$ are non-vanishing for all relevant directions can be challenging, which limits the applicability of this approach.
-New Dimensionality Reduction Technique.
+New Dimensionality Reduction Technique
 The main algorithmic contribution of this work is the design of a new, more general, approach that leads to significant improvements on the aforementioned dimensionality reduction technique.
@@
 As a result, in the next iteration, the distance between $V$ and $W$ decreases; hence, our algorithm will eventually terminate.
+This generalization of the previously used one-shot dimension reduction technique has two main advantages.
+Practically speaking, the more relaxed condition is often easier to prove.
+This allows, for example, our qualitative improvements in the error for robustly learning intersections of halfspaces.
+As far as we know, the one-shot dimension reduction technique may suffice to obtain error $\tilde{O}(\mathrm{OPT})$ in this case.
+However, the difficulty of proving such results means that the best analysis we know of gives substantially worse bounds (and only after a quite involved analysis).
+On the other hand, the condition needed for our iterative dimension reduction is somewhat easier to obtain, giving us a substantially better $\tilde{O}(K\mathrm{OPT})$ error with a simpler analysis.
+Furthermore, our new technique allows us to efficiently learn a broader family of MIMs compared to what was previously possible with one-shot dimension reduction techniques, even in principle.
+In particular, as we show in Appendix F, there exist Multi-Index Models that cannot be learned to better than constant error with any constant number of moments using the one-shot technique; and on the other hand, can be efficiently learned to arbitrary accuracy using only second moments via our iterative technique.
 One disadvantage of our method as described above is that, while the runtime is typically polynomial in the dimension $d$, the dependence on other parameters can be problematic.
@@
 In this case, non-vanishing moments will exactly align with directions along $W$, up to a measurement error, ensuring that $\dim(V)$ never exceeds $\dim(W)$ (see Theorem 4.9).
-SQ Lower Bounds.
+For an illustration of our iterative procedure for the special case of multiclass linear classification, see Figure 1 in Section 2.
+SQ Lower Bounds
 An appealing feature of our algorithmic technique is that it is provably qualitatively tight in the Statistical Query (SQ) model (see Section 5).
@@
 As a result, we establish SQ-hardness in distinguishing between the distribution of random rotations of $f$ in the $d$-dimensional space and a product distribution.
-Applications to Well-studied Concept Classes.
+Applications to Well-studied Concept Classes
 We show that our iterative dimension-reduction technique yields significantly improved learners for multiclass linear classifiers and intersections of halfspaces in the presence of noise.
@@
 Despite this being a natural direction, no general algorithmic framework was previously known in this setting.
-The recent work [DKK+24] developed agnostic learners for MIMs under the Gaussian distribution with fixed-parameter tractability guarantees similar to ours.
-The major difference is that, while our framework relies on learning from random examples, the latter work [DKK+24] assumes query access to the target function—a much stronger access model.
+The recent work [DKK+23] developed agnostic learners for MIMs under the Gaussian distribution with fixed-parameter tractability guarantees similar to ours.
+The major difference is that, while our framework relies on learning from random examples, the latter work [DKK+23] assumes query access to the target function—a much stronger access model.
 The task of efficiently learning intersections of halfspaces under Gaussian marginals is one of the most well-studied problems in the computational learning theory literature.
@@
 1.4 Notation and Preliminaries
-Basic Notation For $n\in\mathbb{Z}_{+}$, let $[n]:=\{1,\ldots,n\}$.
+Basic Notation For $n\in\mathbb{Z}_{+}$, let $[n]:=\{1,\ldots,n\}$ and $\overline{\mathbb{R}}\coloneqq\mathbb{R}\cup\{\pm\infty\}$.
+We denote by $2^{S}=\{T\mid T\subseteq S\}$ the power set of the set $S$.
 We use small boldface characters for vectors and capital bold characters for matrices.
@@
 We will use $\mathbf{x}\cdot\mathbf{y}$ for the inner product of $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$ and $\theta(\mathbf{x},\mathbf{y})$ for the angle between $\mathbf{x},\mathbf{y}$.
-For vectors $\mathbf{x},\mathbf{v}\in\mathbb{R}^{d}$ and a subspace $V\subseteq\mathbb{R}^{d}$ denote by $\mathbf{x}^{V}$ the projection of $\mathbf{x}$ onto $V$, and by $\mathbf{x}^{\mathbf{v}}$ the projection of $\mathbf{x}$ onto the line spanned by $\mathbf{v}$.
-For two subspaces $V,W\subseteq\mathbb{R}^{d}$, we denote by $W^{V}=\{\mathbf{w}^{V}:\mathbf{w}\in W\}$, which it itself a subspace.
+For vectors $\mathbf{x},\mathbf{v}\in\mathbb{R}^{d}$ and a subspace $V\subseteq\mathbb{R}^{d}$ denote by $\mathbf{x}^{V}$ the projection of $\mathbf{x}$ onto $V$, $\mathbf{x}^{\perp V}$ the projection of $\mathbf{x}$ onto the orthogonal complement of $V$ and by $\mathbf{x}^{\mathbf{v}}$ the projection of $\mathbf{x}$ onto the line spanned by $\mathbf{v}$.
+For two subspaces $V,W\subseteq\mathbb{R}^{d}$, we denote by $W^{V}=\{\mathbf{w}^{V}:\mathbf{w}\in W\}$ and by $V+W=\{\mathbf{w}+\mathbf{v}:\mathbf{w}\in W,\mathbf{v}\in V\}$, note that both $W^{V}$ and $V+W$ are subspaces.
 Furthermore, for a set of vectors $L\subseteq\mathbb{R}^{d}$, we denote by $\mathrm{span}(L)$ the subspace of $\mathbb{R}^{d}$ defined by their span.
@@
 In Section 3, we generalize this approach, stating the necessary conditions (see Definition 1.2) for our algorithm to work for general classes of MIMs, establishing Theorem 1.5.
-In Section 4, we consider applications of our algorithmic approach.
-First in Section 4.1, we apply our general theorem to the class of intersections of halfspaces, establishing Theorem 1.9.
-Moreover, in section Section 4.2, we consider how our algorithm performs under random noise, particularly in the Random Classification Noise (RCN) model, and demonstrate that we can get even better runtime guarantees (Theorem 1.8).
-Finally, Section 5 provides SQ lower bounds.
-First, we demonstrate that the complexity of our algorithm is qualitatively optimal with respect to the dimension, establishing Theorem 1.5.
-Second, we establish an SQ lower bound for learning multiclass linear classifiers under random classification noise, showing that achieving optimal error with a polynomial number of samples is impossible in the SQ model, Theorem 1.7.
+In Section 4, we present additional applications of our algorithmic approach.
+In Section 4.1, we apply our general theorem to the class of intersections of halfspaces, establishing Theorem 1.13.
+In Section 4.2, we adapt our general algorithm for the case of random noise, and demonstrate that it yields qualitative stronger complexity guarantees (Theorem 1.11).
+Finally, Section 5 gives our SQ lower bounds.
+First, we demonstrate that the complexity of our algorithm is qualitatively optimal as a function of the dimension, establishing Theorem 1.6.
+Second, we give an SQ lower bound (Theorem 1.9) for learning multiclass linear classifiers under random classification noise (and in the presence of coarse labels).
 2 Learning Multiclass Linear Classifiers with Adversarial Label Noise
@@
 There exists a sufficiently large universal constant $C>0$ such that Algorithm 1 draws $N=d2^{\mathrm{poly}(1/\epsilon,K)}\log(1/\delta)$ i.i.d.
-samples from $D$, runs in $\mathrm{poly}(d,N)$ time, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq{C}\cdot\mathrm{OPT}+\epsilon$.
+samples from $D$, runs in $\mathrm{poly}(N)$ time, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq C\cdot\mathrm{OPT}+\epsilon$.
 Before we present the algorithm, we provide the following definitions, which characterize how we discretize a low-dimensional subspace and approximate distributions over it.
@@
 Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$ with an orthonormal basis $\mathbf{v}^{(1)},\dots,\mathbf{v}^{(k)}$, and let $\epsilon\in(0,1)$.
-We call an $\epsilon$-approximating partition with respect to $V$ a family of sets $\mathcal{S}$ of the form $S_{\mathbf{j}}=\{\mathbf{x}\in\mathbb{R}^{d}:z_{\mathbf{j}_{i}-1}\leq\mathbf{v}^{(i)}\cdot\mathbf{x}\leq z_{\mathbf{j}_{i}},i\in[k]\}$, where $\mathbf{j}\in[M_{\epsilon}]^{k}$, $z_{i}=-\sqrt{2\log(k/\epsilon)}+i\epsilon+t$, for $i\in[M_{\epsilon}]$, $t\in(0,\epsilon/2)$ and $M_{\epsilon}=\left\lceil(2\sqrt{2\log(k/\epsilon)})/\epsilon\right\rceil$.
-Furthermore, for each set $S_{\mathbf{j}}\in\mathcal{S}$, we associate a set of linear equation $\mathcal{G}_{S_{\mathbf{j}}}=\{\mathbf{v}^{(1)}\cdot\mathbf{x}=t_{1},\dots,\mathbf{v}^{(k)}\cdot\mathbf{x}=t_{k}\}$, where $z_{\mathbf{j}_{i}-1}\leq t_{i}\leq z_{\mathbf{j}_{i}}$ for $i\in[k]$.
-Furthermore, we define the piecewise constant approximation of a distribution as the function that minimizes the 0-1 loss with respect to the labels in each subset of a partition.
+An $\epsilon$-approximating partition with respect to $V$ is a collection of sets $\mathcal{S}$ defined as follows: For each multi-index $\mathbf{j}=(\mathbf{j}_{1},\dots,\mathbf{j}_{k})\in[M_{\epsilon}]^{k}$, define $S_{\mathbf{j}}=\{\mathbf{x}\in\mathbb{R}^{d}:z_{\mathbf{j}_{i}-1}\leq\mathbf{v}^{(i)}\cdot\mathbf{x}\leq z_{\mathbf{j}_{i}},i\in[k]\}$ where $z_{i}$’s are defined as $z_{i}=-\sqrt{2\log(k/\epsilon)}+i\epsilon+t$, for $i\in\{0,\dots,M_{\epsilon}\},t\in(0,\epsilon/2)$ and $M_{\epsilon}=\left\lceil(2\sqrt{2\log(k/\epsilon)})/\epsilon\right\rceil$.
+Furthermore, to each set $S_{\mathbf{j}}\in\mathcal{S}$, we associate a set of linear equations $\mathcal{G}_{S_{\mathbf{j}}}=\{\mathbf{v}^{(1)}\cdot\mathbf{x}=t_{1},\dots,\mathbf{v}^{(k)}\cdot\mathbf{x}=t_{k}\}$, where for each $i\in[k]$, $t_{i}$ is chosen so that $z_{\mathbf{j}_{i}-1}\leq t_{i}\leq z_{\mathbf{j}_{i}}$.
+Furthermore, we define the piecewise constant approximation of a distribution over labeled examples as the function that minimizes the 0-1 loss with respect to the labels in each subset of a partition.
 Our algorithm will first estimate the hidden subspace of the multiclass linear classifier and then return a piecewise constant approximation over it.
 Definition 2.3 (Piecewise Constant Approximation).
-Let $D$ be a distribution over $\mathbb{R}^{d}\times[K]$, let $V$ be a subspace of $\mathbb{R}^{d}$, let $\epsilon\in(0,1)$, and let $\mathcal{S}$ be an $\epsilon$-approximating partition with respect to $V$.
-We call a piecewise constant approximation of the distribution $D$, with respect to a partition $\mathcal{S}$, any classifier $h_{{\mathcal{S}}}:\mathbb{R^{d}}\to[K]$ of the form $h_{{\mathcal{S}}}(\mathbf{x})=\operatorname*{argmin}_{i\in[K]}(\operatorname*{\mathbf{Pr}}_{{(\mathbf{x},y)\sim D}}[i\neq y\mid\mathbf{x}\in S])$, if $\mathbf{x}\in S$ for all $S\in\mathcal{S}$, where $\mathcal{S}$ is an $\epsilon$-approximating partition with respect to $V$ according to Definition 2.2.
-We now present our subspace approximation method for learning multiclass linear classifiers under agnostic noise, Algorithm 1.
-Input: Accuracy $\epsilon>0$, failure probability $\delta>0$ sample access to a distribution $D$ over $\mathbb{R}^{d}\times[K]$.
-Output: A hypothesis $h$ such that, with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[h(\mathbf{x})\neq y]\leq C\cdot\mathrm{OPT}{+}\epsilon$, for a universal constant $C$.
+Let $D$ be a distribution over $\mathbb{R}^{d}\times[K]$, let $V$ be a subspace of $\mathbb{R}^{d}$ and let $\epsilon\in(0,1)$.
+Let $\mathcal{S}$ be an $\epsilon$-approximating partition with respect to $V$ (see Definition 2.2).
+A piecewise constant approximation of the distribution $D$, with respect to $\mathcal{S}$, is any classifier $h_{\mathcal{S}}:\mathbb{R}^{d}\to[K]$ such that for each $S\in\mathcal{S}$ and $\mathbf{x}\in S$, $h_{\mathcal{S}}$ is defined as follows $h_{\mathcal{S}}(\mathbf{x})\in\operatorname*{argmin}_{i\in[K]}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[i\neq y\mid\mathbf{x}\in S]$.
+We now present our subspace approximation method for learning multiclass linear classifiers under agnostic noise (see Algorithm 1).
+Input: Accuracy $\epsilon>0$, failure probability $\delta>0$ sample access to a distribution $D$ over $\mathbb{R}^{d}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$.
+Output: A hypothesis $h$ such that with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[h(\mathbf{x})\neq y]\leq C\cdot\mathrm{OPT}{+}\epsilon$, where $C$ is a universal constant and $\mathrm{OPT}\coloneqq\inf_{f\in\mathcal{L}_{d,K}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$.
   1. 1.
-    Let $c$ be a sufficiently large constant, and let $T{,p}$ be sufficiently large, constant-degree polynomials of $1/\epsilon$ and $K$.
+    Let $c$ be a sufficiently large constant, and let $T$ and $p$ be sufficiently large, constant-degree polynomials in $1/\epsilon$ and $K$.
   2. 2.
@@
     1. (a)
-      Draw a set $S$ of $N$ i.i.d. samples from $D$.
+      Draw a set $S_{t}$ of $N$ i.i.d. samples from $D$.
     2. (b)
-      $\mathcal{E}_{t}\leftarrow$ Algorithm 2($\epsilon,\epsilon^{\prime},\delta,\mathrm{span}(L_{t}),S$).
+      $\mathcal{E}_{t}\leftarrow$ Algorithm 2($\epsilon,\epsilon^{\prime},\delta,\mathrm{span}(L_{t}),S_{t}$).
     3. (c)
@@
   4. 4.
-    Construct $\mathcal{S}$ an $\epsilon^{\prime}$-approximating partition with respect to $\mathrm{span}(L_{t})$, as defined in Definition 2.2.
+    Construct $\mathcal{S}$, an $\epsilon^{\prime}$-approximating partition with respect to $\mathrm{span}(L_{t})$ (see Definition 2.2).
   5. 5.
-    Return a piecewise constant approximating classifier $h_{\mathcal{S}}$, as described in Definition 2.3, using $N$ i.i.d.
-samples from $D$, by finding the most frequent label at any set in $\mathcal{S}$.
+    Draw $N$ i.i.d.
+samples from $D$ and construct the piecewise constant classifier $h_{\mathcal{S}}$ as follows: For each $S\in\mathcal{S}$, assign a label that appears most frequently among the samples falling in $S$.
+Return $h_{\mathcal{S}}$.
 Algorithm 1 Learning Multiclass Linear Classifiers under Agnostic Noise
-Input: $\epsilon>0$, $\epsilon^{\prime}>0$, $\delta>0$, a subspace $V$ of $\mathbb{R}^{d}$, and a set of samples $\{(\mathbf{x}^{(1)},y_{1}),\dots,(\mathbf{x}^{(N)},y_{N})\}$ from a distribution $D$ over $\mathbb{R}^{d}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$.
+Input: $\epsilon,\epsilon^{\prime},\delta>0$, a subspace $V$ of $\mathbb{R}^{d}$, and a set of samples $\{(\mathbf{x}^{(1)},y_{1}),\dots,(\mathbf{x}^{(N)},y_{N})\}$ from a distribution $D$ over $\mathbb{R}^{d}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$.
 Output: A set of unit vectors $\mathcal{E}$.
@@
   3. 3.
-    Construct an $\epsilon^{\prime}$-approximating partition, $\mathcal{S}$, with respect to $V$ (see Definition 2.2).
+    Construct an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$ with respect to $V$ (see Definition 2.2).
   4. 4.
-    Compute: $\widehat{\mathbf{u}}^{(S,i)}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\widehat{D}}[\mathbf{x}\mathds{1}(y=i)\mid\mathbf{x}\in S]$ for $S\in\mathcal{S},i\in[K]$.
+    For each $S\in\mathcal{S}$ and for each label $i\in[K]$ compute $\widehat{\mathbf{u}}^{(S,i)}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\widehat{D}}[\mathbf{x}\mathds{1}(y=i)\mid\mathbf{x}\in S]$.
   5. 5.
@@
 Algorithm 2 Estimating a relevant direction
-Let $f^{*}$ be a multiclass linear classifier with error $\mathrm{OPT}$, i.e., $\mathbf{Pr}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$.
-Algorithm 1 operates as follows: In each iteration $t$, the algorithm updates the current list of unit vectors $L_{t}$ so that its span, denoted by $V_{t}$, provides a better approximation to the span of the weight vectors of $f^{*}$, denoted by $W^{*}$.
+Let $f^{*}$ be a $K$-multiclass linear classifier with error $\mathrm{OPT}$, i.e., $\mathbf{Pr}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$.
+Algorithm 1 proceeds in a sequence of iterations: In each iteration $t$, it maintains and updates a set of unit vectors $L_{t}$ so that its span $V_{t}$ provides a better approximation to the span of the weight vectors of $f^{*}$, denoted by $W^{*}$.
 To this end, Algorithm 1 uses Algorithm 2 to construct a new list $\mathcal{E}_{t}$ that contains vectors with non-trivial correlation to $(W^{*})^{\perp V_{t}}$.
-Our main contribution is the design and analysis of Algorithm 2, which efficiently returns a non-trivially correlating vector.
-First in Section 2.1, we show that if an optimal multiclass classifier is not $C\cdot\mathrm{OPT}$-close to a constant classifier, then at least one of the first moments of the labels regions, i.e., $\operatorname*{\mathbf{E}}[\mathbf{x}\mathds{1}(y=i)]$ for some $i\in[K]$, will correlate non-trivially (by at least $\mathrm{poly}(\epsilon)$) with a vector in $W^{*}$.
-This, in fact, suffices to initialize a list of vectors that correlate with $W^{*}$.
+First, in Section 2.1, we show that if $f^{*}$ is not $C\cdot\mathrm{OPT}$-close to any constant classifier, then there exists at least one label $i\in[K]$ for which the first moment $\operatorname*{\mathbf{E}}[\mathbf{x}\mathds{1}(y=i)]$ correlates non-trivially (by at least $\mathrm{poly}(\epsilon/K)$) with a vector in $W^{*}$.
+This, in fact, suffices to initialize a list of vectors that correlate with a vector in $W^{*}$.
 In order to expand this list and at the same time improve the approximation of $W^{*}$, we need to show that, we can find vectors that not only correlate with $W^{*}$ but are also orthogonal to the current list of vectors $L_{t}$.
-To this end, in Section 2.2, we use the current list $L_{t}$ of weight vectors to partition $\mathbb{R}^{d}$ into sufficiently small regions $S$.
+To this end, Algorithm 2 uses the current set $L_{t}$ of weight vectors to partition $\mathbb{R}^{d}$ into sufficiently small regions $S$.
 In each such region $S$, the classifier $f^{*}$ has a limited dependence on the vectors in $L_{t}$.
@@
 Thus, within each region $S$, we obtain a multiclass classifier that depends only on vectors in $(W^{*})^{\perp V_{t}}$.
-Therefore, from the above discussion, in each region $S$, either the classifier can be accurately approximated by a constant classifier, or the first moments will correlate with a vector in $(W^{*})^{\perp V_{t}}$, enabling further improvement.
+Therefore, in every region $S$, either the classifier can be well-approximated by a constant function or the first moments will correlate non-trivially with a vector in $(W^{*})^{\perp V_{t}}$, thus enabling further improvement of the approximation.
 Furthermore, for $f^{*}$ to depend mostly on directions in $(W^{*})^{\perp V_{t}}$, the cubes $S$ in the partition must have width on the order of $\mathrm{poly}(\epsilon/K)$.
@@
 In order to avoid this power-tower dependence, we leverage the fact that the total probability mass of cubes with non-trivial correlation is at least $\mathrm{poly}(\epsilon/K)$ (this will be proved in the analysis of Algorithm 2 in Proposition 2.8).
-By weighting each moment vector by the probability mass of its corresponding cube (see Line 5), we are guaranteed to find a vector with non-trivial correlation by examining only a $\mathrm{poly}(K/\epsilon)$-sized list of candidate vectors.
-We note that our algorithm is not proper; that is, it does not return a multiclass linear classifier.
+By weighting each moment vector by the probability mass of its corresponding cube (see Line 5 of Algorithm 2), we are guaranteed to find a vector with non-trivial correlation by examining only a $\mathrm{poly}(K/\epsilon)$-sized list of candidate vectors.
+As follows from the above, our algorithm is not proper; that is, it does not return a multiclass linear classifier.
 Instead, it produces a piecewise constant function over a sufficiently fine partition of $\mathbb{R}^{d}$.
-The remaining part of the section is structured as follows: In Section 2.1, we state and prove the key structural result about the class of multiclass linear classifiers, Proposition 2.4, which will be exploited by our algorithm.
-In Section 2.2, we provide the analysis of Algorithm 2 with Proposition 2.8.
+See Figure 1 for an illustration.
+Organization of Section 2
+The body of this section is structured as follows: In Section 2.1, we state and prove the key structural result about the class of multiclass linear classifiers, Proposition 2.4, which will be exploited by our algorithm.
+In Section 2.2, we provide the analysis of Algorithm 2 in Proposition 2.8.
 In Section 2.3, we analyze Algorithm 1, providing the full proof of Theorem 2.1.
-Finally, in Section 2.4, we provide the proof of Lemma 2.19, a supporting lemma that is used in the proof of Proposition 2.4.
-The main result of this section is Theorem 2.1.
+Finally, in Section 2.4, we provide the proof of Claim 2.19, a supporting lemma that is used in the proof of Proposition 2.4.
+$\mathbf{v}^{(1)}$
+$\mathbf{v}^{(2)}$$\mathbf{v}^{(1)}$
+Figure 1: Illustration of two consecutive steps of our iterative subspace approximation procedure for 4-class Multiclass Linear Classification.
+In the left figure, the space $\mathbb{R}^{d}$ is partitioned according to one already discovered direction.
+The highlighted rectangular regions correspond to areas where a constant function over the current subspace $V$ yields an insufficient approximation, indicating that additional relevant directions can be found.
+In contrast, the non-highlighted regions already permit a sufficiently good constant approximation.
+The right figure illustrates that we discovered the second hidden direction so that partitioning the space to sufficiently small cubes leads to a good approximation.
 2.1 Structural Result for Multiclass Linear Classification
-In this section, we prove our main structural result, Proposition 2.4, which enables us to learn multiclass linear classifiers efficiently under adversarial noise.
-In particular, we show that for a multiclass linear classifier $f$, if $f$ is far from any constant function by more that a constant multiple of the error of $f$ with respect to the noisy label $y$, then the first moment of at least one of the noisy label regions has a positive correlation with a vector in the subspace of weights.
-We prove this by finding carefully chosen vectors $\mathbf{v}^{(i)}{\in}\mathbb{R}^{d}$, $i{\in}[K]$, such that $\sum_{i=1}^{K}\operatorname*{\mathbf{E}}[\mathbf{v}^{(i)}\cdot\mathbf{x}\mathds{1}(y=i)]$ is guaranteed to be positive.
+In this section, we prove our main structural result, Proposition 2.4, which enables us to efficiently learn multiclass linear classifiers under adversarial noise.
+In particular, we show that for a multiclass linear classifier $f$, if $f$ is far from any constant function by more than a constant multiple of its error with respect to the noisy label $y$, then the first moment of at least one of the noisy label regions has a non-trivial correlation with a vector in the subspace of weights defining $f$.
+We prove this by finding carefully chosen vectors $\mathbf{v}^{(i)}{\in}\mathbb{R}^{d}$, $i{\in}[K]$, such that $\sum_{i=1}^{K}\operatorname*{\mathbf{E}}[\mathbf{v}^{(i)}\cdot\mathbf{x}\mathds{1}(y=i)]$ is guaranteed to be non-trivially large.
 Note that this sum is equal to $\sum_{i=1}^{K}\operatorname*{\mathbf{E}}[\mathbf{v}^{(i)}\cdot\mathbf{x}\mathds{1}(f(\mathbf{x})=i)]$ plus an error term for instances when $y$ is not equal to $f(\mathbf{x})$.
-By carefully selecting these vectors, we can ensure that a small probability of error cannot corrupt the sum by much.
+By carefully selecting these vectors, we ensure that even a small error probability does not significantly reduce the overall sum.
 Formally, we show the following:
@@
 Define $\mathrm{OPT}\coloneqq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$.
-First note that without loss of generality we can assume that $\mathbf{w}^{(1)}=\mathbf{0},t_{1}=0$.
-This can be seen by noting that the output label does not change if we subtract the same function from all affine functions that define the multiclass linear classifier.
-Specifically, by subtracting $\mathbf{w}^{(j)}\cdot\mathbf{x}+t_{j}$, where $j=\operatorname*{argmax}_{i\in[K]}t_{i}$ from all affine functions and renaming the labels, we ensure that one of the affine functions becomes $\mathbf{0}$ and that $t_{i}\leq 0$ for all $i\in[K]$.
-We define $L_{i}:\mathbb{R}^{d}\to[K],L_{i}(\mathbf{x})=\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}$ the corresponding affine functions such that $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}L_{i}(\mathbf{x})$ and we define the function $s(\mathbf{x})\coloneqq\max_{i\in[K]}L_{i}(\mathbf{x})$.
-For some parameter $\gamma>0$ which we will assign later, let $\mathbf{v}^{(i)}\coloneqq\mathbf{w}^{(i)}/(\gamma-t_{i})$ (this is well-defined as $\gamma>0$ and $t_{i}\leq 0$) for all $i\in[K]$.
-Note that for all nonzero vectors $\mathbf{v}^{(i)}$, we have that $L_{i}(\mathbf{x})\geq\gamma$ if and only if $\mathbf{v}^{(i)}\cdot\mathbf{x}\geq 1$.
+Without loss of generality, we may assume that $\mathbf{w}^{(1)}=\mathbf{0}$ and $t_{1}=0$.
+This follows from the observation that subtracting the same function from every classifier does not change the output label.
+In particular, if we subtract $\mathbf{w}^{(j)}\cdot\mathbf{x}+t_{j}$, where $j\in\operatorname*{argmax}_{i\in[K]}t_{i}$, from all affine functions and then rename the labels accordingly, one of the resulting affine functions becomes identically $\mathbf{0}$ and we obtain $t_{i}\leq 0$ for all $i\in[K]$.
+We define $L_{i}:\mathbb{R}^{d}\to\mathbb{R},L_{i}(\mathbf{x})=\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}$ the corresponding affine functions such that $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}L_{i}(\mathbf{x})$ and we define the function $s(\mathbf{x})\coloneqq\max_{i\in[K]}L_{i}(\mathbf{x})$.
+For some parameter $\gamma>0$ which we will chose appropriately below, let $\mathbf{v}^{(i)}\coloneqq\mathbf{w}^{(i)}/(\gamma-t_{i})$ (this is well-defined as $\gamma>0$ and $t_{i}\leq 0$) for all $i\in[K]$.
+Let $g(\mathbf{x})\coloneqq\max_{i\in[K]}\mathbf{v}^{(i)}\cdot\mathbf{x}$.
+Note that for all nonzero vectors $\mathbf{v}^{(i)}$, we have that $L_{i}(\mathbf{x})\geq\gamma$ if and only if $\mathbf{v}^{(i)}\cdot\mathbf{x}\geq 1$ which means that $g(\mathbf{x})\geq 1$.
+Furthermore, because $\mathbf{w}^{(1)}=\mathbf{0}$ and $t_{1}=0$, we have that $g(\mathbf{x})\geq 0$ for all $\mathbf{x}\in\mathbb{R}^{d}$.
 We can decompose the first moment of label $i$ as follows:
@@
 where $S_{i}(\mathbf{x},y)=\mathds{1}(y\neq i,f(\mathbf{x})=i)$ and $T_{i}(\mathbf{x},y)=\mathds{1}(y=i,f(\mathbf{x})\neq i)$ for $i\in[K]$.
-Taking the inner product with $\mathbf{v}^{(i)}$ and summing Equation 1 for all $i\in[K]$, we have that
-$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}^{(y)}\cdot\mathbf{x}]$ $=\underbrace{\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}]}_{I_{1}}\underbrace{\;-\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}^{(i)}\cdot\mathbf{x}S_{i}(\mathbf{x},y)]+\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(x,y)\sim D}[\mathbf{v}^{(i)}\cdot\mathbf{x}T_{i}(\mathbf{x},y)]}_{I_{2}}\;.$
-Note that $I_{1}$ is the first moment of $\mathbf{v}^{(f(\mathbf{x}))}$, i.e., the first moment of the noiseless multiclass linear classifier, and $I_{2}$ is the contribution that corresponds to the noise.
-Let $g(\mathbf{x})\coloneqq\max_{j\in[K]}\mathbf{v}^{(j)}\cdot\mathbf{x}$.
+Taking the inner product of $\operatorname*{\mathbf{E}}[x\mathds{1}(y=i)]$ with $\mathbf{v}^{(i)}$ and summing Equation 1 over $i\in[K]$, we have that
+$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}^{(y)}\cdot\mathbf{x}]$ $=\underbrace{\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}]}_{I_{1}}\underbrace{\;-\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}^{(i)}\cdot\mathbf{x}S_{i}(\mathbf{x},y)]+\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(x,y)\sim D}[\mathbf{v}^{(i)}\cdot\mathbf{x}T_{i}(\mathbf{x},y)]}_{I_{2}}\;.$ (2)
+Note that $I_{1}$ is the the contribution of each $\mathbf{v}^{(i)}$ over $\mathbf{x}$ in the region where $f(\mathbf{x})=i$, and $I_{2}$ is the contribution that corresponds to the noise.
 We first bound $I_{1}$ from below.
@@
 Proof of Claim 2.5.
-Note that it holds $\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}\geq 0$ because $L_{f(\mathbf{x})}(\mathbf{x})=\max_{i}L_{i}(\mathbf{x})\geq L_{1}(\mathbf{x})=0$ and therefore $\mathbf{w}^{(f(\mathbf{x}))}\cdot\mathbf{x}\geq|t_{i}|$ which gives that $\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}\geq 0$.
+Note that the random variable $\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}$ is non-negative since $L_{f(\mathbf{x})}(\mathbf{x})=\max_{i}L_{i}(\mathbf{x})\geq L_{1}(\mathbf{x})=0$ which implies that $\mathbf{w}^{(f(\mathbf{x}))}\cdot\mathbf{x}\geq|t_{i}|$ and consequently, $\mathbf{v}^{(f(\mathbf{x}))}\cdot\mathbf{x}\geq 0$.
 Hence, by applying Markov’s inequality, we have
@@
 where the last equality holds by the definition of $\mathbf{v}^{(i)}$.
-Note that the event $L_{f(\mathbf{x})}(\mathbf{x})\geq\gamma$ is a superset of the event $g(\mathbf{x})\geq 1$.
-Indeed, if $g(\mathbf{x})\geq 1$, there exists some $i\in[K]$ such that $\mathbf{v}^{(i)}\cdot\mathbf{x}\geq 1$, which implies that $L_{i}(\mathbf{x})\geq\gamma$, and therefore $L_{f(\mathbf{x})}(\mathbf{x})\geq\gamma$.
-This gives that $I_{1}\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]$ and completes the proof of Claim 2.5.
+It remains to show that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[L_{f(\mathbf{x})}(\mathbf{x})\geq\gamma]\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]$.
+To see this, note that the event $L_{f(\mathbf{x})}(\mathbf{x})\geq\gamma$ contains the event $g(\mathbf{x})\geq 1$.
+If $g(\mathbf{x})\geq 1$ then by definition there exists some $i\in[K]$ such that $\mathbf{v}^{(i)}\cdot\mathbf{x}\geq 1$, which implies that $L_{i}(\mathbf{x})\geq\gamma$ (recall that $\mathbf{v}^{(i)}=\mathbf{w}^{(i)}/(\gamma-t_{i})$), and consequently, $L_{f(\mathbf{x})}(\mathbf{x})\geq\gamma$.
+This gives that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[L_{f(\mathbf{x})}(\mathbf{x})\geq\gamma]\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]$ and completes the proof of Claim 2.5.
 ∎
@@
 Claim 2.6.
-There exists a $\gamma>0$ such that $I_{2}\geq-C\left(\sqrt{\mathrm{OPT}\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]}+\mathrm{OPT}\right)$, where $C>0$ is a universal constant, and $961/48\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]/$ 96.
+There exists an absolute constant $C^{\prime}>0$ and $\gamma>0$ such that
+$I_{2}\geq-C^{\prime}\left(\sqrt{\mathrm{OPT}\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]}+\mathrm{OPT}\right)\;,$
+and $961/48\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]/$ 96.
 Proof.
-We show how to bound the contribution of the terms containing $S_{i}$, as similar arguments can be used to bound the contribution of the terms involving $T_{i}$.
-In order to bound the contribution of the $S_{i}$ terms from above, notice that
+We show how to bound the contribution of the terms involving $S_{i}$; similar arguments can be used to bound the contribution of the terms involving $T_{i}$.
+In order to bound the contribution of the $S_{i}$ terms we can write
 $\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}^{(i)}\cdot\mathbf{x}S_{i}(\mathbf{x},y)]$ $\leq\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}\left[\max_{j\in[K]}(\mathbf{v}^{(j)}\cdot\mathbf{x})S_{i}(\mathbf{x},y)\right]$
 $=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}\left[\max_{j\in[K]}(\mathbf{v}^{(j)}\cdot\mathbf{x})S(\mathbf{x},y)\right]$
-$=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\max_{j\in[K]}(\mathbf{v}^{(j)}\cdot\mathbf{x})\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]\right]\;,$
-where $S\coloneqq\sum_{i=1}^{K}S_{i}$.
-From the assumptions of Proposition 2.4, we have that $f$ has a missclassification error at most $\mathrm{OPT}$, and therefore we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)]\leq\mathrm{OPT}$.
-Furthermore, only one of $S_{i}$ and $T_{i}$ for all $i\in[K]$ can be non-zero for each point $(\mathbf{x},y)$, which gives that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]\leq 1$.
-Let $R(\mathbf{x})\coloneqq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]$ and note that from the above we have that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)]\leq\mathrm{OPT}$ and $R(\mathbf{x})\in[0,1]$.
-By the definition of the $\mathbf{v}^{(i)}$’s, we have $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\max_{i\in[K]}L_{i}(\mathbf{x})\geq\gamma]$.
-Note that $g(\mathbf{x})\geq 0$ by construction, therefore for any $t>0$, we have that
+$=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\max_{j\in[K]}(\mathbf{v}^{(j)}\cdot\mathbf{x})\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]\right]$
+$=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[g(\mathbf{x})\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]\right]\;,$
+where $S=\sum_{i=1}^{K}S_{i}$.
+Since $f$ has misclassification error $\mathrm{OPT}$, we have $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)]\leq\mathrm{OPT}$.
+Moreover, for each point $(\mathbf{x},y)$, at most one among $\{S_{i},T_{i}:i\in[K]\}$ can be nonzero which implies that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]\leq 1$.
+Define $R(\mathbf{x})\coloneqq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]$.
+Then, by the law of total expectation, we have that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)]\leq\mathrm{OPT}$ and $R(\mathbf{x})\in[0,1]$.
+Recall that $g(\mathbf{x})\geq 0$ by construction, hence for any $t>0$,
 $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})R(\mathbf{x})]$ $\leq t\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})\mathds{1}(g(\mathbf{x})\leq t)]+\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})R(\mathbf{x})\mathds{1}(g(\mathbf{x})>t)]$
@@
 $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\mathds{1}(g(\mathbf{x})>t)]=\int_{t}^{\infty}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathds{1}(g(\mathbf{x})\geq y)]\mathrm{d}y$ $\leq\int_{t}^{\infty}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathds{1}(g(\mathbf{x})\geq\lfloor y\rfloor)]\mathrm{d}y$
-$=\sum_{n=\lfloor t\rfloor}^{\infty}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[n\leq g(\mathbf{x})]\;,$
+$=\sum_{n=\lfloor t\rfloor}^{\infty}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[n\leq g(\mathbf{x})]\;,$ (3)
 where in the first equality, we exchange the order of the integrals using Tonelli’s theorem and in the inequality, we used that $\mathds{1}(g(\mathbf{x})\geq y)$ is a decreasing function of $y$.
-In order to bound the terms $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[n\leq g(\mathbf{x})]$, we prove that the tails of $g$ are at least quadratically decreasing.
+In order to bound the terms $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[n\leq g(\mathbf{x})]$, we show that the tails of $g$ decay at least quadratically.
 We show the following:
@@
 Now, we leverage Lemma 2.7 to complete the proof.
-Note that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma].$ Furthermore, since $t_{i}\leq 0$ for all $i\in[K]$, it follows that if $\mathbf{w}^{(i)}\geq n(\gamma-t_{i})$, then $\mathbf{w}^{(i)}\geq n\gamma-t_{i}$.
-This, in turn, implies that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq n]\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq n\gamma].$ Therefore, using Lemma 2.7, we have that
+Recall that the function $s(\mathbf{x})$ of Lemma 2.7 is defined as $s(\mathbf{x})=\max_{i}L_{i}(\mathbf{x})$.
+We need to show that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq n]\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq n\gamma]$.
+First, note that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma],$ by construction.
+Furthermore, recall that $g(\mathbf{x})=\max_{i\in[K]}\mathbf{v}^{(i)}\cdot\mathbf{x}$ where $\mathbf{v}^{(i)}=\mathbf{w}^{(i)}/(\gamma-t_{i})$ and that $L_{i}(\mathbf{x})=\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}$.
+Since $t_{i}\leq 0$ for all $i\in[K]$, it follows that if $\mathbf{w}^{(i)}\cdot\mathbf{x}\geq n(\gamma-t_{i})$, then $\mathbf{w}^{(i)}\cdot\mathbf{x}\geq n\gamma-t_{i}$.
+This, in turn, implies that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq n]\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq n\gamma].$ Therefore, using Lemma 2.7 into Equation 3, we have that
 $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\mathds{1}(g(\mathbf{x})>t)]\lesssim\sum_{n=\lfloor t\rfloor}^{\infty}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[1\leq g(\mathbf{x})]/n^{2}\lesssim\frac{\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[1\leq g(\mathbf{x})]}{t-2}\;,$
-where in the last inequality, we used the approximation of the sum by an integral.
+where in the last inequality we used the approximation of the sum by an integral.
 Therefore, by choosing $t=2+(\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[1\leq g(\mathbf{x})]/\mathrm{OPT})^{1/2}$, we have that
@@
 As a result, we can arrive at the same bound as for the term defined by the $S_{i}$’s.
-Therefore, we conclude that the total contribution of outliers is at most $O(\sqrt{\mathrm{OPT}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[1\leq g(\mathbf{x})]}+\mathrm{OPT})$, which completes the proof of Claim 2.6.
+Therefore, we conclude that the total contribution of the noise is at most $O(\sqrt{\mathrm{OPT}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[1\leq g(\mathbf{x})]}+\mathrm{OPT})$, which completes the proof of Claim 2.6.
 ∎
-By the choice of $\gamma$ in Claim 2.6 and the assumption that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq i]\geq C\cdot\mathrm{OPT}+\epsilon$ we have that $\mathrm{OPT}\leq({96}/C)\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]$.
-Therefore, for a sufficiently large constant $C>0$, we get that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}^{(y)}\cdot\mathbf{x}]\gtrsim\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]$ and this gives us
-$\sum_{i:\mathbf{w}^{(i)}\neq\mathbf{0}}\mathbf{w}^{(i)}/(\gamma-t_{i})\cdot\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(y=i)]\gtrsim\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]\;.$ (2)
-Moreover, if for some $i\in[K]$, we have that $\gamma-t_{i}\leq\epsilon\left\|\mathbf{w}^{(i)}\right\|$ then $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[L_{i}(\mathbf{x})\geq\gamma]=\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(0,1)}[x\geq(\gamma-t_{i})/\left\|\mathbf{w}^{(i)}\right\|]\geq\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(0,1)}[x\geq\epsilon]$.
-However, for $\epsilon<2\sqrt{\log({48})}$ from A.1 we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\max_{i\in[K]}L_{i}(\mathbf{x})\geq\gamma]>1/{48}$ which leads to a contradiction.
-Hence, we have that $\gamma-t_{i}\geq\epsilon\left\|\mathbf{w}^{(i)}\right\|$ for all $i\in[K]$.
-Now from the fact that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]\geq C\cdot\mathrm{OPT}+\epsilon$ we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]\gtrsim C\cdot\mathrm{OPT}+\epsilon\gtrsim\epsilon$.
-Concluding, from (2) we have that there exists $i\in[K]$ such that $\mathbf{w}^{(i)}/(\gamma-t_{i})\cdot\operatorname*{\mathbf{E}}[\mathbf{x}\mathds{1}(y=i)]\gtrsim\epsilon$.
-Now since $\gamma-t_{i}\geq 0$, we have that $\mathbf{w}^{(i)}\cdot\operatorname*{\mathbf{E}}[\mathbf{x}\mathds{1}(y=i)]\geq 0$ and hence $\left(\mathbf{w}^{(i)}/\left\|\mathbf{w}^{(i)}\right\|\right)\cdot\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(y=i)]\gtrsim\epsilon^{2}$.
-Using the fact that we centered the affine functions defining $f$ by subtracting $\mathbf{w}^{(1)}\cdot\mathbf{x}+t_{1}$ from each of them, gives us the statement.
+Now we show that $I_{2}\geq-\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]/2$. Recall that by the assumptions of Proposition 2.4 we have
+$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}f(\mathbf{x})\neq i\big{]}\geq C\,\mathrm{OPT}+\epsilon\quad\text{for all }i\in[K].$
+We show that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}s(\mathbf{x})\neq 0\big{]}\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}f(\mathbf{x})\neq 1\big{]}$.
+To see this, observe that for any $i\in[K]$ for which $L_{i}$ is not identically zero, we have that $L_{i}(\mathbf{x})\neq 0$ almost surely (with respect to the Gaussian measure).
+Moreover, if for some $i\in[K]$ the function $L_{i}$ is identically zero and $i\neq 1$, then without loss of generality we may assume that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y=i]=0$, since we could set $1$ to be the label with minimum index in the set $\operatorname*{argmax}_{i\in[K]}t_{i}$ (for the original biases $t_{i}$).
+Hence, in every case for any $i\neq 1$ we have $L_{i}(\mathbf{x})\neq 0$ almost surely.
+Therefore, if $f(\mathbf{x})\neq 1$ then, except for a measure zero set it holds that $L_{i}(\mathbf{x})>0$ for some $i\neq 1$, which implies that $s(\mathbf{x})\neq 0$.
+As a result, it holds that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}s(\mathbf{x})\neq 0\big{]}\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}f(\mathbf{x})\neq 1\big{]}$.
+Hence, it follows that
+$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}s(\mathbf{x})\neq 0\big{]}\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}f(\mathbf{x})\neq 1\big{]}\geq C\,\mathrm{OPT}+\epsilon.$
+By the choice of $\gamma$ in Claim 2.6, we also have
+$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}s(\mathbf{x})\geq\gamma\big{]}\geq\frac{1}{96}\,\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}s(\mathbf{x})\neq 0\big{]}.$
+Since $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}g(\mathbf{x})\geq 1\big{]}=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}s(\mathbf{x})\geq\gamma\big{]}$, we get that that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]\geq({C\,\mathrm{OPT}+\epsilon})/{96}$.
+Finally, by choosing $C$ sufficiently large (in particular, ensuring that $C\geq 256/((C^{\prime})^{2}+1)$), from Claim 2.6, we conclude that
+$I_{2}\geq-\frac{1}{2}\,\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}g(\mathbf{x})\geq 1\big{]}\;.$
+and that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\big{[}g(\mathbf{x})\geq 1\big{]}\gtrsim\epsilon$. Combining the above with Equation 2 and Claim 2.5, we obtain
+$\sum_{i:\mathbf{w}^{(i)}\neq\mathbf{0}}\mathbf{w}^{(i)}/(\gamma-t_{i})\cdot\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(y=i)]\gtrsim\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[g(\mathbf{x})\geq 1]\gtrsim\epsilon\;.$
+This implies that there exists some $j\in[K]$ such that
+[…]
+To conclude the proof, we must show that $\gamma-t_{j}\geq\epsilon\,\left\|\mathbf{w}^{(j)}\right\|$.
+Suppose, for the sake of contradiction, that for some $i\in[K]$ we have $\gamma-t_{i}\leq\epsilon\left\|\mathbf{w}^{(i)}\right\|$.
+Then, by the definition of $L_{i}$, we have
+$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\Big{[}L_{i}(\mathbf{x})\geq\gamma\Big{]}=\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(0,1)}\Bigg{[}x\geq\frac{\gamma-t_{i}}{\left\|\mathbf{w}^{(i)}\right\|}\Bigg{]}\geq\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(0,1)}[x\geq\epsilon]\;.$
+However, for $\epsilon<2\sqrt{\log(48)}$, A.1 implies that
+$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\Big{[}\max_{i\in[K]}L_{i}(\mathbf{x})\geq\gamma\Big{]}>\frac{1}{48},$
+which leads to a contradiction from Lemma 2.7 (and the choice of $\gamma$). Thus, we must have $\gamma-t_{i}\geq\epsilon\,\left\|\mathbf{w}^{(i)}\right\|$, for all $i\in[K]$. As a result, we obtain
+[…]
+Finally, using the fact that we centered the affine functions defining $f$ by subtracting $\mathbf{w}^{(1)}\cdot\mathbf{x}+t_{1}$ from each of them, we obtain the desired statement.
 This completes the proof of Proposition 2.4.
@@
 Proposition 2.8.
-Let $d,k,K\in\mathbb{Z}_{+}$, $\epsilon,\delta\in(0,1)$ and $\epsilon^{\prime}\leq c\epsilon^{2}/({K^{2}\sqrt{k}})$, for $c$ a sufficiently small constant.
+Let $d,k,K\in\mathbb{Z}_{+}$, $\epsilon,\delta\in(0,1)$ and $\epsilon^{\prime}\leq c\epsilon^{2}/(K^{2}\sqrt{k})$, for some sufficiently small absolute constant $c>0$.
 Let $D$ be a distribution supported on $\mathbb{R}^{d}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ and $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$.
-Let $f$ be a multiclass linear classifier with weight vectors $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$ and $\mathrm{OPT}=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$.
-Additionally, let $h_{\mathcal{S}}$ be a piecewise constant approximation of $D$, with respect to an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$, according $V$, as defined in Definitions 2.2 and 2.3.
-There exists $N=d\mathrm{poly}(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}\log(1/\delta)$ and a universal constant $C$ such that if $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{{\mathcal{S}}}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+\epsilon$, then Algorithm 2 with probability at least $1-\delta$ returns a list of unit vectors $\mathcal{E}$ of size $|\mathcal{E}|=\mathrm{poly}(1/\epsilon,K)$, such that: There exists $\mathbf{v}\in\mathcal{E}$ and $i,j\in[K],i\neq j$ for which
+Let $f$ be a multiclass linear classifier with weight vectors $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$ and define $\mathrm{OPT}=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$.
+Furthermore, let $h_{\mathcal{S}}$ be a piecewise constant approximation of $D$, with respect to an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$ defined relative to the subspace $V$ (as in Definitions 2.2 and 2.3).
+There exists a sample size $N=d(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}\log(1/\delta)$ and a universal constant $C>0$, such that if $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+\epsilon$, then Algorithm 2 given $N$ samples, runs in $\mathrm{poly}(N)$ time and with probability at least $1-\delta$ returns a set of unit vectors $\mathcal{E}$ of size $|\mathcal{E}|=\mathrm{poly}(1/\epsilon,K)$ with the following property: there exists a $\mathbf{v}\in\mathcal{E}$ and $i,j\in[K]$, $i\neq j$, such that
 $\frac{(\mathbf{w}^{(i)}-\mathbf{w}^{(j)})^{\perp V}}{\left\|(\mathbf{w}^{(i)}-\mathbf{w}^{(j)})^{\perp V}\right\|}\cdot\mathbf{v}\geq\mathrm{poly}(1/K,\epsilon)\;.$
 Before proving Proposition 2.8, we develop the necessary tools for working with multiclass linear classifiers restricted to cubes extending along the orthogonal complement of a subspace.
-First, we define the notion of a projected classifier onto a set of linear equations, Definition 2.9.
+First, we define the notion of a projected classifier onto a set of linear equations (Definition 2.9).
 The intuition behind this definition is that when we restrict our classifier to sufficiently small cubes on $V$, its behavior mainly depends on its components in the orthogonal complement of $V$.
@@
 Definition 2.9 (Projected Classifier).
-Let $f(\mathbf{x})=\operatorname*{argmax}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$ with $\mathbf{w}^{(i)}\in\mathbb{R}^{d},t_{i}\in\mathbb{R}$ for $i\in[K]$ be a multiclass linear classifier, let $\mathcal{G}=\{\mathbf{v}^{(1)}\cdot\mathbf{x}=z_{1},\ldots\mathbf{v}^{(k)}\cdot\mathbf{x}=z_{k}\}$ be a set of $k$ linear equations and let $V$ be the subspace spanned by orthonormal vectors $\mathbf{v}^{(1)},\ldots,\mathbf{v}^{(k)}\in\mathbb{R}^{d}$.
-We denote by $f_{\mathcal{G}}(\mathbf{x})$ the linear classifier $f$ projected onto $\mathcal{G}$, i.e., $f_{\mathcal{G}}(\mathbf{x})=\operatorname*{argmax}((\mathbf{w}^{(i)})^{\perp_{V}}\cdot\mathbf{x}+t_{i}+t_{i,\mathcal{G}}^{\prime})$, where $t_{i,\mathcal{G}}^{\prime}=\sum_{j=1}^{k}z_{j}a_{j}$ where $a_{j}=\mathbf{w}^{(i)}\cdot\mathbf{v}^{(j)}$.
+Let $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$ with $\mathbf{w}^{(i)}\in\mathbb{R}^{d},t_{i}\in\mathbb{R}$ for $i\in[K]$, let $\mathcal{G}=\{\mathbf{v}^{(1)}\cdot\mathbf{x}=z_{1},\ldots,\mathbf{v}^{(k)}\cdot\mathbf{x}=z_{k}\}$ be a set of $k$ linear equations and let $V$ be the subspace spanned by orthonormal vectors $\mathbf{v}^{(1)},\ldots,\mathbf{v}^{(k)}\in\mathbb{R}^{d}$.
+We denote by $f_{\mathcal{G}}(\mathbf{x})$ the linear classifier $f$ projected onto $\mathcal{G}$, i.e., $f_{\mathcal{G}}(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}((\mathbf{w}^{(i)})^{\perp_{V}}\cdot\mathbf{x}+t_{i}+t_{i,\mathcal{G}}^{\prime})$, where $t_{i,\mathcal{G}}^{\prime}=\sum_{j=1}^{k}z_{j}a_{ij}$ where $a_{ij}=\mathbf{w}^{(i)}\cdot\mathbf{v}^{(j)}$.
 Now we prove Lemma 2.10, which states that for any partition of a subspace into sufficiently small cubes, the target function is well-approximated by a piecewise function constructed using a projected classifier for each region.
@@
 Lemma 2.10 (Projection Lemma).
-Let $d,K\in\mathbb{Z}_{+}{,\epsilon>0}$, $f:\mathbb{R}^{d}\to[K]$ be a multiclass linear classifier, $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d},{k\geq 1}$ and $\epsilon>0$.
-Let $\mathcal{S}$ an $\epsilon$-approximating partition of $\mathbb{R}^{d}$ with respect to $V$ as in Definition 2.2.
+Let $d,K,k\in\mathbb{Z}_{+},\epsilon,\epsilon^{\prime}>0$, $f:\mathbb{R}^{d}\to[K]$ be a multiclass linear classifier, $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d},k\geq 1$.
+Let $\mathcal{S}$ be an $\epsilon^{\prime}$-approximating partition of $\mathbb{R}^{d}$ with respect to $V$ as in Definition 2.2, with $\epsilon^{\prime}\leq\epsilon^{2}/(K^{2}\sqrt{k})$.
 Then there exists a set $\mathcal{T}\subseteq\mathcal{S}$ such that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathcal{N}(\mathbf{0},\mathbf{I})}[S]\geq 1-\epsilon$ and for any $S\in\mathcal{T}$, it holds that
-$\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]\leq K^{2}{\sqrt{k}}\epsilon\;,$
+$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\leq\epsilon\;,$
 where $f_{\mathcal{G}_{S}}$ denotes the projected classifier onto a set of linear equations $\mathcal{G}_{S}$ that is associated with $S$, as defined in Definition 2.9.
 Proof.
-Let $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$ and $t_{1},\dots,t_{K}\in\mathbb{R}$ such as $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$.
-Denote by $\mathbf{v}^{(1)},\dots,\mathbf{v}^{(k)}$ an orthonormal basis of $V$.
-Fix a set of the partition $S\in\mathcal{S}$ and denote by $f_{\mathcal{G}_{S}}=\operatorname*{argmax}((\mathbf{w}^{(i)})^{\perp{V}}\cdot\mathbf{x}+t_{i}+t_{i}^{\prime})$ the projected classifier that corresponds to the set of linear equations $\mathcal{G}_{S}=\{\mathbf{v}^{(1)}\cdot\mathbf{x}=z_{1},\ldots\mathbf{v}^{(k)}\cdot\mathbf{x}=z_{k}\}$ (see Definition 2.2).
+Let $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$ and $t_{1},\dots,t_{K}\in\mathbb{R}$ such that $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$.
+Let $\mathbf{v}^{(1)},\dots,\mathbf{v}^{(k)}$ be an orthonormal basis of $V$.
+Fix a set of the partition $S\in\mathcal{S}$ and denote by $f_{\mathcal{G}_{S}}=\operatorname*{argmax}((\mathbf{w}^{(i)})^{\perp{V}}\cdot\mathbf{x}+t_{i}+t_{i}^{\prime})$ (see Definition 2.9 for the definition of $t_{i}^{\prime}$) the projected classifier that corresponds to the set of linear equations $\mathcal{G}_{S}=\{\mathbf{v}^{(1)}\cdot\mathbf{x}=z_{1},\ldots,\mathbf{v}^{(k)}\cdot\mathbf{x}=z_{k}\}$ (see Definition 2.2 for the definition of $\mathcal{G}_{S}$).
 Let $\mathbf{w}^{(i,j)}=\mathbf{w}^{(i)}-\mathbf{w}^{(j)}$, $t_{i,j}=t_{i}-t_{j}$ and $t_{i,j}^{\prime}=t_{i}^{\prime}-t_{j}^{\prime}$.
@@
 $\sum_{i\neq j}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathrm{sign}(\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j})\neq\mathrm{sign}((\mathbf{w}^{(i,j)})^{\perp V}\cdot\mathbf{x}+t_{i,j}+t^{\prime}_{i,j})\mid\mathbf{x}\in S]\;.$
-Furthermore, by the definition of $f_{\mathcal{G}_{S}}$, if $\mathbf{x}\in S$ then we have that for any $i,j\in[K]$
-$\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-t_{i,j}^{\prime}\rvert$ $=\left|\sum_{l=1}^{k}(\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)})(\mathbf{x}\cdot\mathbf{v}^{(l)}-z_{l})\right|$
-$\leq\sum_{l=1}^{k}\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)}\rvert\lvert\mathbf{x}\cdot\mathbf{v}^{(l)}-z_{l}\rvert$
-$\leq\sum_{l=1}^{k}\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)}\rvert\epsilon$
-$\leq\|(\mathbf{w}^{(i,j)})^{V}\|{\sqrt{k}}\epsilon\;,$
-where we used the fact that the vector $\sum_{i=1}^{k}z_{i}\mathbf{v}^{(i)}$ belongs in $S$ by the definition of $\mathcal{G}$.
-Notice that in order for $\mathrm{sign}(\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j})\neq\mathrm{sign}((\mathbf{w}^{(i,j)})^{\perp V}\cdot\mathbf{x}+t_{i,j}+t^{\prime}_{i,j})$, we need to have that $\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\rvert\leq\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-t_{i,j}^{\prime}\rvert$, because we can rewrite $t^{\prime}_{i,j}=\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}+t_{i,j}^{\prime}$.
-Averaging over $S\in\mathcal{S}$ gives us
-$\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]$ $\leq\sum_{i\neq j\in[K]}\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}[\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\rvert\leq\|(\mathbf{w}^{(i,j)})^{V}\|{\sqrt{k}}\epsilon\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]$
-$\leq\sum_{i\neq j\in[K]}\operatorname*{\mathbf{Pr}}[\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\rvert\leq\|(\mathbf{w}^{(i,j)})^{V}\|{\sqrt{k}}\epsilon]$
-$\leq\sum_{i\neq j\in[K]}\frac{\|(\mathbf{w}^{(i,j)})^{V}\|{\sqrt{k}}\epsilon}{\|\mathbf{w}^{(i,j)}\|}\leq K^{2}{\sqrt{k}}\epsilon\;.$
+Fix $i,j\in[K]$ with $i\neq j$.
+Notice that for a sign change to occur, i.e., $\mathrm{sign}(\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j})\neq\mathrm{sign}((\mathbf{w}^{(i,j)})^{\perp V}\cdot\mathbf{x}+t_{i,j}+t^{\prime}_{i,j})$ it is necessary that
+$\Bigl{|}\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\Bigr{|}\leq\Bigl{|}\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-t_{i,j}^{\prime}\Bigr{|}.$
+To see that, note that we can add and subtract $\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}$, and we get that if for some $\mathbf{x}\in S$ we have $\mathrm{sign}(\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j})\neq\mathrm{sign}((\mathbf{w}^{(i,j)})^{\perp V}\cdot\mathbf{x}+t_{i,j}+t^{\prime}_{i,j})$ we also have, $\mathrm{sign}(\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j})\neq\mathrm{sign}((\mathbf{w}^{(i,j)})\cdot\mathbf{x}+t_{i,j}+t^{\prime}_{i,j}-\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V})$, hence it should hold that $\Bigl{|}\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\Bigr{|}\leq\Bigl{|}\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-t_{i,j}^{\prime}\Bigr{|}$.
+Therefore, it suffices to bound the probability that $\operatorname*{\mathbf{Pr}}[|\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}|\leq|\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-t_{i,j}^{\prime}|\mid\mathbf{x}\in S]$.
+By the definition of $f_{\mathcal{G}_{S}}$, if $\mathbf{x}\in S$ then for any $i,j\in[K]$ we have
+$\left|\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-t_{i,j}^{\prime}\right|=\left|\sum_{l=1}^{k}(\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)})(\mathbf{x}\cdot\mathbf{v}^{(l)}-z_{l})\right|.$
+This equality holds because by definition the projection of $\mathbf{x}$ onto the subspace $V$ with orthonormal basis $\{\mathbf{v}^{(1)},\ldots,\mathbf{v}^{(k)}\}$ is given by $\mathbf{x}^{V}=\sum_{l=1}^{k}(\mathbf{x}\cdot\mathbf{v}^{(l)})\mathbf{v}^{(l)}$, and the term $t_{i,j}^{\prime}$ is defined as $t_{i,j}^{\prime}=\sum_{l=1}^{k}z_{l}\,(\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)})$.
+Next, by applying the triangle inequality we have
+$\left|\sum_{l=1}^{k}(\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)})(\mathbf{x}\cdot\mathbf{v}^{(l)}-z_{l})\right|\leq\sum_{l=1}^{k}\Bigl{|}\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)}\Bigr{|}\Bigl{|}\mathbf{x}\cdot\mathbf{v}^{(l)}-z_{l}\Bigr{|}.$
+By the definition of the partition $S$, if $\mathbf{x}\in S$ then for every $l\in\{1,\ldots,k\}$ the projection of $\mathbf{x}$ onto the direction $\mathbf{v}^{(l)}$ is $\epsilon^{\prime}$-close to the corresponding value $z_{l}$, i.e., $\Big{|}\mathbf{x}\cdot\mathbf{v}^{(l)}-z_{l}\Big{|}\leq\epsilon^{\prime}$.
+Therefore,
+$\sum_{l=1}^{k}\Bigl{|}\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)}\Bigr{|}\Bigl{|}\mathbf{x}\cdot\mathbf{v}^{(l)}-z_{l}\Bigr{|}\leq\sum_{l=1}^{k}\Bigl{|}\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)}\Bigr{|}\epsilon^{\prime}.$
+Finally, by applying the Cauchy-Schwarz inequality, we obtain
+$\sum_{l=1}^{k}\Bigl{|}\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)}\Bigr{|}\leq\sqrt{k}\,\sqrt{\sum_{l=1}^{k}\Bigl{(}\mathbf{w}^{(i,j)}\cdot\mathbf{v}^{(l)}\Bigr{)}^{2}}=\sqrt{k}\,\|(\mathbf{w}^{(i,j)})^{V}\|\;.$
+Combining the above, we have $\left|\mathbf{w}^{(i,j)}\cdot\mathbf{x}^{V}-t_{i,j}^{\prime}\right|\leq\|(\mathbf{w}^{(i,j)})^{V}\|\sqrt{k}\,\epsilon^{\prime}$. Averaging over $S\in\mathcal{S}$ gives us
+$\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]$ $\leq\sum_{i\neq j\in[K]}\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}[\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\rvert\leq\|(\mathbf{w}^{(i,j)})^{V}\|\sqrt{k}\epsilon^{\prime}\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]$
+$\leq\sum_{i\neq j\in[K]}\operatorname*{\mathbf{Pr}}[\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\rvert\leq\|(\mathbf{w}^{(i,j)})^{V}\|\sqrt{k}\epsilon^{\prime}]$
+$\leq\sum_{i\neq j\in[K]}\operatorname*{\mathbf{Pr}}[\lvert\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j}\rvert\leq\|\mathbf{w}^{(i,j)}\|\sqrt{k}\epsilon^{\prime}]$
+$\leq\sum_{i\neq j\in[K]}\sqrt{k}\epsilon^{\prime}\leq K^{2}\sqrt{k}\epsilon^{\prime}\;,$
+where we used the anti-concentration of the Gaussian distribution (see A.1).
+Finally, since $\epsilon^{\prime}\leq\epsilon^{2}/(K^{2}\sqrt{k})$ it holds that $\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]\leq\epsilon^{2}$.
+To conclude the proof, let ${\mathcal{S}}_{bad}$ be the set of $S\in\mathcal{S}$ such that if $S\in{\mathcal{S}}_{bad}$ then $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]>\epsilon$.
+Then, we have that
+$\epsilon^{2}\geq\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]\geq\sum_{S\in{\mathcal{S}}_{bad}}\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}[S]>\epsilon\sum_{S\in{\mathcal{S}}_{bad}}\operatorname*{\mathbf{Pr}}[S]\;.$
+Hence, $\sum_{S\in{\mathcal{S}}_{bad}}\operatorname*{\mathbf{Pr}}[S]\leq\epsilon$. By taking $\mathcal{T}=\mathcal{S}/{\mathcal{S}}_{bad}$ we conclude the proof. ∎
 We now proceed with the proof of Proposition 2.8.
@@
 Let $\mathbf{w}^{(1)},\ldots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$ and $t_{1},\ldots,t_{K}\in\mathbb{R}$ be the parameters defining $f$, i.e., $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$.
-Denote by $\mathbf{w}^{(i,j)}=\mathbf{w}^{(i)}-\mathbf{w}^{(j)}$, by $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}=(\mathbf{w}^{(i,j)})^{\perp V}/\left\|(\mathbf{w}^{(i,j)})^{\perp V}\right\|$, and let $C$ be a constant that satisfies the statement of Proposition 2.4.
-Let $\mathcal{G}_{S}$ be the set of linear equations associated with each cube $S\in\mathcal{S}$ as defined in Definition 2.2, and let $f_{\mathcal{G}_{S}}$ be the projection of $f$ onto $\mathcal{G}_{S}$ as defined in Definition 2.9.
-Furthermore, denote by $\mathbf{u}^{(S,i)}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(y=i)\mid\mathbf{x}\in S],S\in\mathcal{S},i\in[K]$ the population moments of each discretization set and by $\widehat{\mathbf{u}}^{(S,i)}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\widehat{D}}[\mathbf{x}\mathds{1}(y=i)\mid\mathbf{x}\in S],S\in\mathcal{S},i\in[K]$ its empirical counterpart, computed in Line 4.
-Also, let $\widehat{\mathbf{U}}$ be the matrix computed at Line 5 of Algorithm 2.
-We first show that if the current subspace $V$ does not lead to a good approximation of $f$, then by splitting the space into small cubes with respect to $V$, there will be a nontrivial fraction of them for which $f$ is far from constant.
+For brevity, define the difference vectors $\mathbf{w}^{(i,j)}=\mathbf{w}^{(i)}-\mathbf{w}^{(j)}$, and let $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}=(\mathbf{w}^{(i,j)})^{\perp V}/\left\|(\mathbf{w}^{(i,j)})^{\perp V}\right\|$ denote the normalized projection of $\mathbf{w}^{(i,j)}$ onto the orthogonal complement of a given subspace $V$.
+Also, define $W=\mathrm{span}(\mathbf{w}_{1},\dots,\mathbf{w}_{K})$, and let $C$ be a constant as specified in Proposition 2.4.
+For each cube $S\in\mathcal{S}$ (as defined in Definition 2.2), let $\mathcal{G}_{S}$ be the corresponding set of linear equations, and let $f_{\mathcal{G}_{S}}$ be the projection of $f$ onto $\mathcal{G}_{S}$ (see Definition 2.9).
+Furthermore, for every $S\in\mathcal{S}$ and $i\in[K]$, define the population moment $\mathbf{u}^{(S,i)}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}\mathds{1}(y=i)\mid\mathbf{x}\in S]$,and let $\widehat{\mathbf{u}}^{(S,i)}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\widehat{D}}[\mathbf{x}\mathds{1}(y=i)\mid\mathbf{x}\in S]$ its empirical counterpart (see Line 4).
+Also, denote by $\widehat{\mathbf{U}}$ the matrix computed at Line 5 of Algorithm 2.
+We first show that if the current subspace $V$ does not lead to a good approximation of $f$, meaning that any function defined only on $V$ gets error at least $\Omega(\mathrm{OPT}+\epsilon)$ with the labels $y$, then when we partition the space into small cubes with respect to $V$, on a nontrivial fraction of them $f$ is far from being constant.
 Claim 2.11 (Large Fraction of Good Cubes).
-If $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{{\mathcal{S}}}(\mathbf{x})\neq y]>(C+1)\mathrm{OPT}+5\epsilon$, then there exists a subset $\mathcal{T}\subseteq\mathcal{S}$ such that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]\geq\epsilon$; and for all $i\in[K],S\in\mathcal{T}$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]+\epsilon$.
+If $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>(C+1)\mathrm{OPT}+5\epsilon$, then there exists a subcollection $\mathcal{T}\subseteq\mathcal{S}$ satisfying $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]\geq\epsilon$; and for every $S\in\mathcal{T}$ we have $\min_{i\in[K]}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]+\epsilon$.
 Proof of Claim 2.11.
-Define $\mathcal{T}$ to be all sets $S\subseteq\mathcal{S}$ such that $\operatorname*{\mathbf{Pr}}[i\neq y\mid\mathbf{x}\in S]>(C+1)\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y\mid\mathbf{x}\in S]+2\epsilon$, for all $i\in[K],S\in\mathcal{T}$.
+Define $\mathcal{T}$ to be the collection of $S\subseteq\mathcal{S}$ such that $\min_{i\in[K]}\operatorname*{\mathbf{Pr}}[i\neq y\mid\mathbf{x}\in S]>(C+1)\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y\mid\mathbf{x}\in S]+2\epsilon$.
 Assume for the sake of contradiction that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]\leq 2\epsilon$.
@@
 $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]$ $\leq\sum_{S\in\mathcal{S}}\min_{\widehat{y}\in[K]}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[\widehat{y}\neq y\mid\mathbf{x}\in S]\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]+\epsilon$
-$\leq\sum_{S\not\in\mathcal{T}}\min_{\widehat{y}\in[K]}\left((C+1)\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}\in S]+2\epsilon\right)\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]+3\epsilon$
+$\leq\sum_{S\not\in\mathcal{T}}\left((C+1)\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}\in S]+2\epsilon\right)\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]+3\epsilon$
 $\leq(C+1)\mathrm{OPT}+5\epsilon\;,$
-where in the first equality, we used the fact that at most an $\epsilon$ fraction of the mass is outside the discretization (see A.1).
-This leads to a contradiction and hence $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]>2\epsilon$.
-Furthermore, note that by Lemma 2.10 there exists $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ such that $\sum_{S\in\mathcal{T}^{\prime}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]>\epsilon$ and for all $i\in[K],S\in\mathcal{T}^{\prime}$, it holds that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[i\neq y\mid\mathbf{x}\in S]>(C+1)\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(x)\neq y\mid\mathbf{x}\in S]+\epsilon$.
-Since $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[i\neq y\mid\mathbf{x}\in S]\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]+\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]$ we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]+\epsilon$ for all $S\in\mathcal{T}^{\prime}$, which completes the proof.
+where in the first line we used the fact (see A.1) that the mass of points outside the discretization is at most $\epsilon$, in the second line we applied our definition of $\mathcal{T}$ and in the last inequality we used that the total error of $f$ is $\mathrm{OPT}$.
+This contradicts our hypothesis that $\operatorname*{\mathbf{Pr}}[h_{\mathcal{S}}(\mathbf{x})\neq y]>(C+1)\mathrm{OPT}+5\epsilon$.
+Hence, it must be that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[S]>2\epsilon$.
+Next, recall that the cube width $\epsilon^{\prime}$ is chosen to be less than $c\epsilon^{2}/(K^{2}\sqrt{k})$ for a sufficiently small constant $c$.
+Taking $c<1/(C+1)$, we can apply Lemma 2.10 and get that there exists a collection $\mathcal{T}^{\prime}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}^{\prime}}\operatorname*{\mathbf{Pr}}[S]\geq 1-\epsilon$, such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq f_{\mathcal{G}_{S}}(\mathbf{x})\mid\mathbf{x}\in S]\leq\epsilon/(C+1)$.
+Note that for $\mathcal{T}^{\prime\prime}=\mathcal{T}^{\prime}\cap\mathcal{T}$, it must hold that $\sum_{S\in\mathcal{T}^{\prime\prime}}\operatorname*{\mathbf{Pr}}[s]\geq\epsilon$.
+For all $S\in\mathcal{T}^{\prime\prime}$, it holds that $\min_{i\in[K]}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[i\neq y\mid\mathbf{x}\in S]>(C+1)\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(x)\neq y\mid\mathbf{x}\in S]+\epsilon$.
+Since,
+$\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\Bigl{[}i\neq y\mid\mathbf{x}\in S\Bigr{]}\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\Bigl{[}f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S\Bigr{]}+\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\Bigl{[}f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S\Bigr{]}\;,$
+we get that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]+\epsilon$ for all $S\in\mathcal{T}^{\prime\prime}$, which completes the proof.
 ∎
@@
 Claim 2.12 (Correlating Moments).
-Let $S\in\mathcal{S}$ be a subset of the partition such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]+\epsilon$.
-Then there exists some $i\in[K]$ and a unit vector $\mathbf{w}\in W$ for which $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{u}^{(S,i)}\geq\mathrm{poly}(1/K,\epsilon)$.
+Let $S\in\mathcal{S}$ such that $\min_{i\in[K]}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i\mid\mathbf{x}\in S]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y\mid\mathbf{x}\in S]+\epsilon$.
+Then there exists $i,j,l\in[K]$ with $i\neq j$ such that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{u}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$.
 Proof of Claim 2.12.
-First, notice that the output label of $f_{\mathcal{G}_{S}}$ is only determined by the projection at the orthogonal complement of $V$, i.e.,
+First, notice that the output label of $f_{\mathcal{G}_{S}}$ is only determined by the projection onto the orthogonal complement of $V$, i.e.,
 $f_{\mathcal{G}_{S}}(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}((\mathbf{w}^{(i)})^{\perp V}\cdot\mathbf{x}+t_{i}+t_{\mathcal{G}_{S},i}^{\prime})=\operatorname*{argmax}_{i\in[K]}((\mathbf{w}^{(i)})^{\perp V}\cdot\mathbf{x}^{\perp V}+t_{i}+t_{\mathcal{G}_{S},i}^{\prime})=f_{\mathcal{G}_{S}}(\mathbf{x}^{\perp{V}}).$
-Furthermore, note that the condition whether $\mathbf{x}$ belongs to $S$ depends only on the projection of $\mathbf{x}^{V}$, hence without loss of generality it can written as $\mathds{1}(\mathbf{x}^{V}\in S)$.
+Furthermore, observe that whether $\mathbf{x}$ belongs to $S$ depends only on its projection onto $V$, hence without loss of generality it can written as $\mathds{1}(\mathbf{x}^{V}\in S)$.
 Thus, we have that
@@
 […]
-$=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y]\;,$
-where $D_{V^{\perp}}$ is defined to be the marginal obtained by averaging $D$ over $V$ conditioned on $S$.
-Similarly, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y]+\epsilon$ for all $i\in[K]$.
-Moreover, by the properties of the standard Gaussian distribution, we have that the $\mathbf{x}$-marginal of $D_{V^{\perp}}$ is a standard Gaussian.
-As a result, by applying Proposition 2.4 for $f_{\mathcal{G}_{S}}$, we have that there exist $i,j\in[K],i\neq j$ such that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[\mathbf{x}\mathds{1}[y=i]]\geq\mathrm{poly}(1/K,\epsilon)$.
-Furthermore, note that by averaging over $V$ conditioning on $S$ we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}^{\perp V}\mathds{1}[y=i]\mid\mathbf{x}\in S]=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[\mathbf{x}\mathds{1}[y=i]]$.
+$=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}^{S}}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y]\;,$
+where $D_{V^{\perp}}^{S}$ denotes the marginal distribution on $V^{\perp}$ obtained by averaging $D$ over $V$ while conditioning on the event $\mathbf{x}^{V}\in S$.
+Similarly, we have $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}^{S}}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq i]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}^{S}}[f_{\mathcal{G}_{S}}(\mathbf{x})\neq y]+\epsilon$ for all $i\in[K]$.
+Moreover, by the properties of the standard Gaussian distribution, the $\mathbf{x}$-marginal of $D_{V^{\perp}}^{S}$ is a standard Gaussian.
+As a result, by applying Proposition 2.4 for $f_{\mathcal{G}_{S}}$, we have that there exist $i,j\in[K],i\neq j$ such that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[\mathbf{x}\mathds{1}(y=i)]\geq\mathrm{poly}(1/K,\epsilon)$.
+Furthermore, note that by averaging over $V$ conditioning on $S$ we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}^{\perp V}\mathds{1}(y=i)\mid\mathbf{x}\in S]=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}^{S}}[\mathbf{x}\mathds{1}(y=i)]$.
 Therefore, there exist $i,j\in[K],i\neq j$, such that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{u}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$ for some $l\in[K]$.
@@
 Let $p(\epsilon,1/K)$ be a sufficiently small polynomial in $\epsilon$ and $1/K$.
-If $N=d(k/\epsilon^{\prime})^{Ck}(K/\epsilon)^{C}\log(1/\delta)$ for a sufficiently large universal constant $C>0$, then with probability $1-\delta$ we have that $\left\|\widehat{\mathbf{u}}^{(S,i)}-{\mathbf{u}}^{(S,i)}\right\|\leq{p(\epsilon,1/K)}$ and $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S},i\in[K]$.
+If $N=d(k/\epsilon^{\prime})^{Ck}(K/\epsilon)^{C}\log(1/\delta)$ for a sufficiently large universal constant $C>0$, then with probability $1-\delta$ we have that $\left\|(\widehat{\mathbf{u}}^{(S,i)}-{\mathbf{u}}^{(S,i)})^{\perp V}\right\|\leq p(\epsilon,1/K)$ and $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S},i\in[K]$.
 Proof of Claim 2.13.
 First note that by the definition of an approximating partition (see Definition 2.2) and A.1, for all $S\in\mathcal{S}$ we have that $\operatorname*{\mathbf{Pr}}_{D}[S]=(\epsilon^{\prime}/k)^{\Omega(k)}$ and $\lvert\mathcal{S}\rvert=(k/\epsilon^{\prime})^{O(k)}$.
-Therefore, by union bound and Hoeffding’s inequality, it holds that if $N\geq{(k/\epsilon^{\prime})^{Ck}}\log(\lvert\mathcal{S}\rvert/\delta)={(k/\epsilon^{\prime})^{O(k)}}\log(1/\delta)$ for a sufficiently large constant $C>0$, then with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]-\operatorname*{\mathbf{Pr}}_{D}[S]\rvert\leq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S}$.
-Hence, it is true that $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$, i.e., the number of samples that fall in each set $S\in\mathcal{S}$ is at least $N\operatorname*{\mathbf{Pr}}_{D}[S]/2{=N(\epsilon^{\prime}/k)^{\Omega(k)}}$.
-Furthermore, note that conditioning on any $S$, the random vector $\mathbf{x}\mathds{1}(y=i)$ with $(\mathbf{x},y)\sim D$ is subgaussian in any unit direction with subgaussian norm at most a constant.
-Because, we can decompose $\mathbf{v}\cdot\mathbf{x}=\mathbf{v}\cdot\mathbf{x}^{V}+\mathbf{v}\cdot\mathbf{x}^{\perp V}$ with $\mathbf{v}\cdot\mathbf{x}^{V},\mathbf{v}\cdot\mathbf{x}^{\perp V}$ being independent, where the zero mean centering $\mathbf{v}\cdot\mathbf{x}^{V}$ is bounded in any direction by $\epsilon^{\prime}\sqrt{k}\leq\epsilon$ since we have conditioned on $S$ and $\mathbf{v}\cdot\mathbf{x}^{\perp V}$ is a standard gaussian since $S$ depends only on $\mathbf{x}^{V}$.
-Therefore, by the use of a standard covering argument it follows that if the number of samples that falls in each discretization set $S$ is at least $d\mathrm{poly}(K,1/\epsilon)\log(K/\delta)$ for a sufficiently large polynomial, then $\left\|\widehat{\mathbf{u}}^{(S,i)}-{\mathbf{u}}^{(S,i)}\right\|\leq{p(\epsilon,1/K)}$ for all $i\in[K]$.
-Combining the above results with the allong with union bound gives us that there exists an $N=d(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}\log(1/\delta)$ that suffices.
-Applying Claims 2.11 and 2.12, we have that if $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{{\mathcal{S}}}(\mathbf{x})\neq y]>(C+1)\mathrm{OPT}+\epsilon$ then there exists a subset $\mathcal{T}\subseteq\mathcal{S}$ of the $\epsilon^{\prime}$-approximating partition with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{D}[S]=\Omega(\epsilon)$ such that for all $S\in\mathcal{T}$ we have that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{u}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$ for some $i,j,l\in[K],i\neq j$.
+Therefore, by the union bound and Hoeffding’s inequality, if $N\geq(k/\epsilon^{\prime})^{Ck}\log(\lvert\mathcal{S}\rvert/\delta)=(k/\epsilon^{\prime})^{O(k)}\log(1/\delta)$ for a sufficiently large absolute constant $C>0$, with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]-\operatorname*{\mathbf{Pr}}_{D}[S]\rvert\leq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S}$.
+Hence, $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$, i.e., the number of samples that fall in each set $S\in\mathcal{S}$ is at least $N\operatorname*{\mathbf{Pr}}_{D}[S]/2=N(\epsilon^{\prime}/k)^{\Omega(k)}$.
+Furthermore, note that by the definitions of ${\mathbf{u}}^{(S,i)}$ and $\widehat{\mathbf{u}}^{(S,i)}$ it holds that
+$(\widehat{\mathbf{u}}^{(S,i)}-{\mathbf{u}}^{(S,i)})^{\perp V}$ $=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\widehat{D}}[\mathbf{x}^{\perp V}\mathds{1}(y=i)\mid\mathbf{x}\in S]-\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{x}^{\perp V}\mathds{1}(y=i)\mid\mathbf{x}\in S]$
+Therefore, to prove our concentration result, it suffices to show that when conditioning on $\mathbf{x}\in S$ the random vector $\mathbf{x}^{\perp V}\mathds{1}(y=i)$ with $(\mathbf{x},y)\sim D$ is subgaussian in every unit direction with subgaussian norm bounded by a constant.
+To see this, note that whether $\mathbf{x}$ belongs in $S$ depends only on the projection of $\mathbf{x}$ onto $V$, and $\mathbf{x}^{\perp V}$ is independent of $\mathbf{x}^{V}$.
+Now fix a unit vector $\mathbf{v}\in\mathbb{R}^{d}$, we have that for any $p\geq 1$
+$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}\left[\left(\mathbf{v}\cdot\mathbf{x}^{\perp V}\mathds{1}(y=i)-\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{v}\cdot\mathbf{x}^{\perp V}\mathds{1}(y=i)\mid\mathbf{x}\in S]\right)^{p}\mid S\right]^{1/p}$ $\lesssim\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}\left[\left(\mathbf{v}\cdot\mathbf{x}^{\perp V}\mathds{1}(y=i)\right)^{p}\mid S\right]^{1/p}$
+$\leq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}\left[\left(\mathbf{v}\cdot\mathbf{x}^{\perp V}\right)^{p}\right]^{1/p}$
+$\leq\operatorname*{\mathbf{E}}_{z\sim\mathcal{N}(0,1)}\left[(z)^{p}\right]^{1/p}\lesssim\sqrt{p}\;,$
+where in the first inequality we used the triangle and Jensen’s inequalities, while in the second we exploited the independence of $\mathbf{x}^{V}$ and $\mathbf{x}^{\perp V}$ and the fact that the indicator function is bounded by 1.
+Therefore $\mathbf{x}^{\perp V}\mathds{1}(y=i)$ is subgaussian.
+Hence, by the use of a standard covering argument it follows that if the number of samples that falls in each discretization set $S$ is at least $d\mathrm{poly}(K,1/\epsilon)\log(K/\delta)$ for a sufficiently large polynomial, then $\left\|(\widehat{\mathbf{u}}^{(S,i)}-{\mathbf{u}}^{(S,i)})^{\perp V}\right\|\leq p(\epsilon,1/K)$ for all $i\in[K]$.
+Combining the above results along with the union bound gives us that there exists an $N=d(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}\log(1/\delta)$ that suffices. ∎
+Applying Claims 2.11 and 2.12, we have that if $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>(C+1)\mathrm{OPT}+\epsilon$ then there exists a subset $\mathcal{T}\subseteq\mathcal{S}$ of the $\epsilon^{\prime}$-approximating partition $\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{D}[S]=\Omega(\epsilon)$ such that for all $S\in\mathcal{T}$ we have $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{u}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$ for some $i,j,l\in[K],i\neq j$.
 Moreover, if $N=d(k/\epsilon^{\prime})^{C^{\prime}k}(K/\epsilon)^{C^{\prime}}\log(1/\delta)$ for a sufficiently large universal constant $C^{\prime}>0$, using the concentration result of Claim 2.13, we have that the empirical moments of each set satisfy the same guarantees, i.e., $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\widehat{\mathbf{u}}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$, for some $i,j,l\in[K],i\neq j,$ for all $S\in\mathcal{T}$.
 Furthermore, note that by Claim 2.13 we also have that $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]=\Omega(\epsilon)$.
-Therefore, there exists a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ and $i,j\in[K]$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\Omega(\epsilon/K^{2})$ such that for all $S\in\mathcal{T}$ we have that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\widehat{\mathbf{u}}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$ for some $l\in[K]$.
+Therefore, there exists a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ and $i,j\in[K]$ with $\sum_{S\in\mathcal{T}^{\prime}}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\Omega(\epsilon/K^{2})$ such that for all $S\in\mathcal{T}^{\prime}$ we have that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\widehat{\mathbf{u}}^{(S,l)}\geq\mathrm{poly}(1/K,\epsilon)$ for some $l\in[K]$.
 Hence, for some $i\neq j\in[K]$ we have that
@@
 where we used the fact that $\mathbf{x}\mathbf{x}^{\top}$ is PSD for any $\mathbf{x}\in\mathbb{R}^{d}$.
-Moreover, note that $(\mathbf{u}^{(S,i)})^{\perp V}=\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp V}}[\mathbf{x}^{\perp V}\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[\mathds{1}(y=i)\mid\mathbf{x}\in S]]$ as $S$ depends only on $\mathbf{x}^{V}$ and
+Since the set $S$ depends only on the projection $\mathbf{x}^{V}$, we have $(\mathbf{u}^{(S,i)})^{\perp V}=\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp V}}[\mathbf{x}^{\perp V}\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[\mathds{1}(y=i)\mid\mathbf{x}\in S]]$.
+Therefore, we can bound the norm of $(\mathbf{u}^{(S,i)})^{\perp V}$ as follows:
 $\left\|(\mathbf{u}^{(S,i)})^{\perp V}\right\|$ $\leq\max_{\left\|\mathbf{u}\right\|=1}\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp V}}[\mathbf{u}\cdot\mathbf{x}^{\perp V}\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[\mathds{1}(y=i)\mid\mathbf{x}\in S]]$
 $\leq\max_{\left\|\mathbf{u}\right\|=1}(\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp V}}[(\mathbf{u}\cdot\mathbf{x}^{\perp V})^{2}]\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathds{1}(y=i)\mid\mathbf{x}\in S]])^{1/2}\leq 1\;,$
-where we used the Cauchy-Schwarz inequality and the dual characterization of the $\ell_{2}$ norm. Therefore, by the triangle inequality, it follows that
+where we used the Cauchy-Schwarz inequality and the dual characterization of the $\ell_{2}$ norm.
+In the last inequality we used that $\mathbf{x}^{\perp V}$ is a standard Gaussian, and henceforth the first expectation is at most $1$, and the second expectation is at most $1$ as well because $\mathds{1}(y=i)$ is bounded by $1$.
+By the triangle inequality, it follows that
 $\left\|\widehat{\mathbf{U}}\right\|_{F}\leq\sum_{S\in\mathcal{S},i\in[K]}\|(\widehat{\mathbf{u}}^{(S,i)})^{\perp V}\|\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\leq\sum_{S\in\mathcal{S},i\in[K]}(1+\mathrm{poly}(\epsilon,1/K))\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\leq 2K\;.$
-Hence, by Claim B.1, we have that there exists a unit eigenvector $\mathbf{v}$ of $\widehat{\mathbf{U}}$ with eigenvalue at least ${\mathrm{poly}(\epsilon/K)}$ for a sufficiently small polynomial, such that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{v}\geq\mathrm{poly}(1/K,\epsilon)$, for some $i\neq j\in[K]$.
-Moreover, the number of such eigenvectors is at most $2K/\mathrm{poly}(1/K,\epsilon)=\mathrm{poly}(K,1/\epsilon)$, which completes the proof of Proposition 2.8.
-Before proving our main result of this section, Theorem 2.1, we need to show that if we have learned a direction of our target model to sufficiently high accuracy, then we can assume, without loss of generality, that the model is defined only on the remaining directions.
+Hence, by Claim B.1, there exists a unit eigenvector $\mathbf{v}$ of $\widehat{\mathbf{U}}$ with eigenvalue at least $\mathrm{poly}(\epsilon/K)$ (for a sufficiently small polynomial), such that $(\widehat{\mathbf{w}}^{(i,j)})^{\perp V}\cdot\mathbf{v}\geq\mathrm{poly}(1/K,\epsilon)$, for some $i\neq j\in[K]$.
+Moreover, the number of such eigenvectors is at most $2K/\mathrm{poly}(1/K,\epsilon)=\mathrm{poly}(K,1/\epsilon)$.
+This completes the proof of Proposition 2.8.
+∎
+Before proving our main result of this section, Theorem 2.1, we first show that if we have learned one direction of our target model to sufficiently high accuracy, then we may, without loss of generality, restrict the model to the remaining directions.
 Specifically, we establish the following:
 Lemma 2.14.
-Let $K\geq 2,\epsilon\in(0,1)$ and $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$ with $\mathbf{w}^{(i)}\in\mathbb{R}^{d}$ for $i\in[K]$ be a multiclass linear classifier.
-Let $V$ be a subspace such that for all $i,j\in[K]$ with $i\neq j$ it holds that $\|(\mathbf{w}^{(i)}-\mathbf{w}^{(j)})^{V}\|\leq(\epsilon^{2}/K^{4})\|\mathbf{w}^{(i)}-\mathbf{w}^{(j)}\|$.
+Let $K\geq 2,\epsilon\in(0,1)$ and $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i})$ with $\mathbf{w}^{(i)}\in\mathbb{R}^{d},t_{i}\in\mathbb{R}$ for $i\in[K]$ be a multiclass linear classifier.
+Let $V$ be a non-empty subspace of $\mathbb{R}^{d}$ such that for all $i,j\in[K]$ with $i\neq j$ it holds that $\|(\mathbf{w}^{(i)}-\mathbf{w}^{(j)})^{V}\|\leq(\epsilon^{2}/K^{4})\|\mathbf{w}^{(i)}-\mathbf{w}^{(j)}\|$.
 Then, it holds that
@@
 Proof.
-For simplicity, let $\mathbf{w}^{(i,j)}=\mathbf{w}^{(i)}-\mathbf{w}^{(j)}$ and $t_{i,j}=t_{i}-t_{j}$. We have that
+For simplicity, let $\mathbf{w}^{(i,j)}=\mathbf{w}^{(i)}-\mathbf{w}^{(j)}$ and $t_{i,j}=t_{i}-t_{j}$. Using the union bound inequality, we have that
 $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f^{\perp V}(\mathbf{x})]\leq\sum_{i,j\in[K]}\operatorname*{\mathbf{Pr}}[\mathrm{sign}(\mathbf{w}^{(i,j)}\cdot\mathbf{x}+t_{i,j})\neq\mathrm{sign}((\mathbf{w}^{(i,j)})^{\perp V}\cdot\mathbf{x}+t_{i,j})]\;.$
-We can write $\mathbf{w}^{(i,j)}=(\mathbf{w}^{(i,j)})^{V}+(\mathbf{w}^{(i,j)})^{\perp V}$, with $\|(\mathbf{w}^{(i,j)})^{V}\|\leq(\epsilon^{2}/K^{4})\|\mathbf{w}^{(i,j)}\|$ and $\|(\mathbf{w}^{(i,j)})^{\perp V}\|\geq\sqrt{1-(\epsilon^{4}/K^{8})}\|\mathbf{w}^{(i,j)}\|$.
+We can write $\mathbf{w}^{(i,j)}=(\mathbf{w}^{(i,j)})^{V}+(\mathbf{w}^{(i,j)})^{\perp V}$.
+According to the assumptions of Lemma 2.14, it holds that $\|(\mathbf{w}^{(i,j)})^{V}\|\leq(\epsilon^{2}/K^{4})\|\mathbf{w}^{(i,j)}\|$ and $\|(\mathbf{w}^{(i,j)})^{\perp V}\|\geq\sqrt{1-(\epsilon^{4}/K^{8})}\|\mathbf{w}^{(i,j)}\|$.
 Therefore, using the independence of any two orthogonal directions of a standard Gaussian, along with A.1, we have that
@@
 In this section, given Proposition 2.8, we proceed to the proof of Theorem 2.1.
+In particular, we show that Algorithm 1 returns, with probability at least $1-\delta$, a hypothesis $h$ whose 0-1 error is $O(\mathrm{OPT})+\epsilon$.
 Proof of Theorem 2.1.
-We show that Algorithm 1, with high probability, returns a hypothesis $h$ with 0-1 error $O(\mathrm{OPT})+\epsilon$.
 Let $\mathbf{w}^{*(1)},\ldots,\mathbf{w}^{*(K)}\in\mathbb{R}^{d}$ and $t_{1}^{*},\ldots,t_{K}^{*}\in\mathbb{R}$ be the parameters of the affine functions of an optimal classifier $f^{*}$, i.e., $f^{*}(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{*(i)}\cdot\mathbf{x}+t_{i}^{*})$ and denote by $W^{*}=\mathrm{span}(\mathbf{w}^{*(1)},\ldots,\mathbf{w}^{*(K)})$.
-Let $C>0$ be an absolute constant that satisfies the statement of Proposition 2.8.
-Let $L_{t}$ be the list of vectors updated by the algorithm (Line 3c) and $V_{t}=\mathrm{span}(L_{t})$, $\dim(V_{t})=k_{t}$.
-Also let $\epsilon^{\prime}=\mathrm{poly}(\epsilon/K)$ for a sufficiently small polynomial, and for $t\in[T]$ let $\mathcal{S}_{t}$ be arbitrary $\epsilon^{\prime}$-approximating partitions with respect to $V_{t}$ (see Definition 2.2).
-Let $h_{t}:\mathbb{R}^{d}\to[K]$ be piecewise constant classifiers, defined as $h_{t}=h_{\mathcal{S}_{t}}$ according to Definition 2.3 for the distribution $D$.
-To prove the correctness of Algorithm 1, we need to show that if $h_{t}$ has significant error, then the algorithm improves its approximation $V_{t}$ to $W^{*}$.
-For quantifying the improvement at every step, we consider the following potential function
+Let $C>0$ be the constant that satisfies the statement of Proposition 2.8.
+The algorithm maintains a list of vectors $L_{t}$ (updated at Line 3c), and define $V_{t}=\mathrm{span}(L_{t})$, $\dim(V_{t})=k_{t}$.
+We set $\epsilon^{\prime}=\mathrm{poly}(\epsilon,1/K)$, for a sufficiently small polynomial (Line 2) which will be quantified later, and for each $t\in[T]$ let $\mathcal{S}_{t}$ be an arbitrary $\epsilon^{\prime}$-approximating partitions with respect to $V_{t}$ (see Definition 2.2).
+Let $h_{t}:\mathbb{R}^{d}\to[K]$ be the piecewise constant classifier defined by $h_{t}=h_{\mathcal{S}_{t}}$ as in Definition 2.3, for the distribution $D$.
+To prove the correctness of Algorithm 1, we need to show that if $h_{t}$ has significant error, then the algorithm improves its current approximation $V_{t}$ to $W^{*}$.
+To quantify this improvement, we define the potential function
 $\Phi_{t}=\sum_{i=1}^{K}\left\|(\mathbf{b}^{(i)})^{\perp V_{t}}\right\|^{2}\;,$
@@
 Claim 2.15 (Large Error Implies Good Correlation).
-For all $k_{t}=\mathrm{poly}(K/\epsilon)$, if $\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+2\epsilon$ and $N\geq d\mathrm{poly}(K,k_{t},1/\epsilon)^{k_{t}}\log(1/\delta)$, then with probability at least $1-\delta$ there exists a unit vector $\mathbf{v}^{(t)}\in V_{t+1}$ such that $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\mathrm{poly}(1/K,\epsilon)$ for some unit vector $\mathbf{w}\in W^{*}$.
+Suppose that at iteration $t$ the classifier $h_{t}$ satisfies $\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+2\epsilon$, and that $\epsilon^{\prime}$ is bounded as $\epsilon^{\prime}\leq c\epsilon^{2}/(K^{2}\sqrt{k_{t}})$, for a sufficiently small absolute constant $c>0$.
+Then there exists sample size, $N\geq d(k_{t}/\epsilon^{\prime})^{O(k_{t})}(K/\epsilon)^{O(1)}\log(1/\delta)$, such that with probability at least $1-\delta$ there exists unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ and $\mathbf{w}\in W^{*}$ such that $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\mathrm{poly}(1/K,\epsilon)$.
 Proof of Claim 2.15.
-First, note that using a sufficiently small polynomial for $\epsilon^{\prime}$, we have that $\epsilon^{\prime}\leq c\epsilon^{2}/(K^{2}\sqrt{k_{t}})$ for all $k_{t}=\mathrm{poly}(K/\epsilon)$ for a sufficiently small constant $c$.
-Denote by $\mathbf{u}^{(1)},\dots,\mathbf{u}^{(k^{\prime})}$ the unit vectors return by Algorithm 2 at iteration $t$, i.e., the vectors in $\mathcal{E}_{t}$.
-For the sake of contradiction assume that $\lvert\mathbf{u}^{(i)}\cdot\mathbf{w}\rvert\leq\epsilon^{2}/(C^{2}K^{4})$, for all unit vectors $\mathbf{w}\in W^{*}$.
-Let $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}$ be vectors in $\mathbb{R}^{d}$.
-For simplicity, define $\mathbf{w}^{(i,j)}\coloneqq\mathbf{w}^{(i)}-\mathbf{w}^{(j)},\mathbf{w}^{*(i,j)}\coloneqq\mathbf{w}^{*(i)}-\mathbf{w}^{*(j)}$.
-By Lemma 2.14, there exists a classifier $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}^{*})$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq f^{*}(\mathbf{x})]\leq\epsilon/C$, and $\mathbf{w}^{(i)}=(\mathbf{w}^{*(i)})^{\perp U},$ for all $i\in[K]$, where $U=\mathrm{span}(\mathbf{u}^{(1)},\dots,\mathbf{u}^{(k^{\prime})})$.
-Since $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]\leq\mathrm{OPT}+\epsilon/C,$ we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]+\epsilon$.
-Therefore, by applying Proposition 2.8, for $f$, with probability $1-\delta$ there exists a vector $\mathbf{u}^{(l)}\in\mathcal{E}_{t}$ such that $\mathbf{u}^{(l)}\cdot\mathbf{w}^{(i,j)}\geq\mathrm{poly}(1/K,\epsilon)\left\|(\mathbf{w}^{(i,j)})^{\perp V_{t}}\right\|$ for some $i,j\in[K]$ with $(\mathbf{w}^{(i,j)})^{\perp V_{t}}\neq\mathbf{0}$.
-However, since $\mathbf{w}^{(i,j)}\in U^{\perp}$ we have a contradiction, which concludes the proof.
+Let $\mathcal{E}_{t+1}=\{\mathbf{u}^{(1)},\dots,\mathbf{u}^{(k^{\prime})}\}$ be the set of unit vectors returned by Algorithm 2 at iteration $t+1$.
+For the sake of contradiction assume that for every $\mathbf{u}\in\mathcal{E}_{t+1}$ and every unit vector $\mathbf{w}\in W^{*}$, it holds that $\lvert\mathbf{u}\cdot\mathbf{w}\rvert\leq\epsilon^{2}/(C^{2}K^{4})$.
+By Lemma 2.14, there exists a classifier $f(\mathbf{x})=\operatorname*{argmax}_{i\in[K]}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}^{*})$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq f^{*}(\mathbf{x})]\leq\epsilon/C$, and for every $i\in[K]$, the vector $\mathbf{w}^{(i)}$ is defined as $\mathbf{w}^{(i)}=(\mathbf{w}^{*(i)})^{\perp U},$ where $U=\mathrm{span}(\mathbf{u}^{(1)},\dots,\mathbf{u}^{(k^{\prime})})$.
+Since $f$ approximates $f^{*}$ up to error $\epsilon/C$, we have $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]\leq\mathrm{OPT}+\epsilon/C$.
+Thus, the assumption of Claim 2.15 that $\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+2\epsilon$ implies that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]+\epsilon$.
+Therefore, by applying Proposition 2.8 for $f$, we deduce that the returned set $\mathcal{E}_{t+1}$ should contain a vector $\mathbf{u}^{(l)}\in\mathcal{E}_{t+1}$ so that for some indices $i,j\in[K]$, $i\neq j$, it holds $\mathbf{u}^{(l)}\cdot\mathbf{w}^{(i,j)}\geq\mathrm{poly}(1/K,\epsilon)\left\|(\mathbf{w}^{(i,j)})^{\perp V_{t}}\right\|$, where $\mathbf{w}^{(i,j)}\coloneqq\mathbf{w}^{(i)}-\mathbf{w}^{(j)}$ with $(\mathbf{w}^{(i,j)})^{\perp V_{t}}\neq\mathbf{0}$.
+However, since $\mathbf{w}^{(i,j)}\in U^{\perp}$ it should hold that $\mathbf{w}^{(i,j)}\cdot\mathbf{u}^{(l)}=0$, which is a contradiction.
+Therefore, there exists $\mathbf{u}\in\mathcal{E}_{t+1}$ and unit vector $\mathbf{w}\in W^{*}$, such that $\lvert\mathbf{u}\cdot\mathbf{w}\rvert\geq\epsilon^{2}/(C^{2}K^{4})$, which concludes the proof.
 ∎
-Our next claim quantifies the decrease of the potential function $\Phi_{t}$ after adding a new weight vector.
+Our next claim quantifies the decrease of the potential function $\Phi_{t}$ between the iterations $t$ and $t+1$.
 Claim 2.16 (Potential Decrease).
-Let $\beta\geq 0$. If there exists a unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ and $\mathbf{w}\in W^{*}$ such that
+Let $\beta\geq 0$. If there exists unit vector $\mathbf{v}^{(t)}\in V_{t+1}$ and unit vector $\mathbf{w}\in W^{*}$ such that
 $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\beta\;,$
-then $\Phi_{t+1}\leq\Phi_{t}-\beta^{2}/2$.
+then $\Phi_{t+1}\leq\Phi_{t}-\beta^{2}$.
 Proof of Claim 2.16.
-First note that $\sum_{1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})(\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}})\geq\beta$ because we can write $\mathbf{w}$ as $\sum_{1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})\mathbf{b}^{(i)}$.
-Also note that $\lvert\mathbf{b}^{(i)}\cdot\mathbf{w}\rvert\leq\left\|\mathbf{w}\right\|=1$ for all $i\in[K]$.
-Hence, by the Pythagorean theorem and the dual characterization of norms, we have that
-$\Phi_{t+1}=\sum_{i=1}^{K}\|(\mathbf{b}^{(i)})^{\perp V_{t}}\|^{2}-\|(\mathbf{b}^{(i)})^{V_{t+1}\cap\perp V_{t}}\|^{2}$ $\leq\Phi_{t}-\sum_{i=1}^{K}\left\lvert\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\right\rvert^{2}$
-$\leq\Phi_{t}-\left\lvert\sum_{i=1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})(\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}})\right\rvert^{2}/2$
-$\leq\Phi_{t}-\beta^{2}/2\;.$
+First, note that since $\{\mathbf{b}^{(i)}\}_{i=1}^{K}$ is an orthonormal basis for $W^{*}$, any vector $\mathbf{w}\in W^{*}$ can be written as $\sum_{i=1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})\mathbf{b}^{(i)}$.
+Thus, we can write $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\beta$ as $\sum_{i=1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})(\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}})\geq\beta$.
+Recall that our potential function is defined as $\Phi_{t}=\sum_{i=1}^{K}\left\|(\mathbf{b}^{(i)})^{\perp V_{t}}\right\|^{2}$.
+Hence, by the Pythagorean theorem, we have that
+$\Phi_{t+1}=\sum_{i=1}^{K}\|(\mathbf{b}^{(i)})^{\perp V_{t}}\|^{2}-\|(\mathbf{b}^{(i)})^{V_{t+1}\cap\perp V_{t}}\|^{2}=\Phi_{t}-\sum_{i=1}^{K}\|(\mathbf{b}^{(i)})^{V_{t+1}\cap\perp V_{t}}\|^{2}\;.$
+Using the dual characterization of the norm, we can further bound
+$\sum_{i=1}^{K}\|(\mathbf{b}^{(i)})^{V_{t+1}\cap\perp V_{t}}\|^{2}\geq\sum_{i=1}^{K}\left\lvert\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\right\rvert^{2}$
+Note that from Cauchy-Schwarz inequality we have that
+$\sum_{i=1}^{K}\left|\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\right|^{2}\geq\left|\sum_{i=1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})(\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}})\right|^{2}\;,$
+where we used that $\sum_{i=1}^{K}\left|\mathbf{b}^{(i)}\cdot\mathbf{w}\right|^{2}=1$ from Pythagorean theorem. Therefore, we have
+$\Phi_{t+1}$ $\leq\Phi_{t}-\left\lvert\sum_{i=1}^{K}(\mathbf{b}^{(i)}\cdot\mathbf{w})(\mathbf{b}^{(i)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}})\right\rvert^{2}$
+$\leq\Phi_{t}-\beta^{2}\;.$
 This completes the proof of Claim 2.16. ∎
@@
 Proof of Claim 2.17.
-Since the statement of Proposition 2.8 holds for any $\epsilon^{\prime}$-partitions $\mathcal{S}_{t},t\in[T]$, independent of the choice of basis and threshold points, we can assume that all the classifiers $h_{t}$ are computed with respect to extensions of a common orthonormal basis and that all threshold points are aligned.
+Since Proposition 2.8 holds for any $\epsilon^{\prime}$-partitions $\mathcal{S}_{t},t\in[T]$, independent of the choice of basis and threshold points, we can assume that all the classifiers $h_{t}$ are computed with respect to extensions of a common orthonormal basis and that all threshold points are aligned.
 Hence, $\mathcal{S}_{t+1}$ is a subdivision of $\mathcal{S}_{t}$.
@@
 Therefore, by the definition of $h_{t}$, we have that
-$\operatorname*{\mathbf{Pr}}[h_{t+1}(\mathbf{x})\neq y,\mathbf{x}\in S]$ $=\sum_{i=1}^{l}\operatorname*{\mathbf{Pr}}[h_{t+1}(\mathbf{x})\neq y\mid\mathbf{x}\in S_{i}]\operatorname*{\mathbf{Pr}}[S_{i}]=\sum_{i=1}^{l}\min_{\widehat{y}\in[K]}\operatorname*{\mathbf{Pr}}[y\neq\widehat{y}\mid\mathbf{x}\in S_{i}]\operatorname*{\mathbf{Pr}}[S_{i}]$
-$\leq\min_{\widehat{y}\in[K]}\sum_{i=1}^{l}\operatorname*{\mathbf{Pr}}[y\neq\widehat{y}\mid\mathbf{x}\in S_{i}]\operatorname*{\mathbf{Pr}}[S_{i}]=\min_{\widehat{y}\in[K]}\operatorname*{\mathbf{Pr}}[\widehat{y}\neq y,\mathbf{x}\in\cup_{i=1}^{l}S_{i}]$
-$\leq\operatorname*{\mathbf{Pr}}[h_{t}(\mathbf{x})\neq y,\mathbf{x}\in S]\;.$
-Summing over all $S\in\mathcal{S}$ we conclude the statement. ∎
-In the following claim, we show that using a sufficient number of samples, we can obtain a good estimate for the population piecewise constant approximation function.
-Claim 2.18.
-Let $\epsilon,\epsilon^{\prime},\delta\in(0,1),k,K\in\mathbb{Z}_{+}$ with $\epsilon^{\prime}\leq\epsilon/2$.
-Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$, and let $h$ and $\widehat{h}$ be the piecewise constant approximating functions defined according to the same partition, with width $\epsilon^{\prime}$ and subspace $V$ with respect to distributions $D$ and $\widehat{D}$, respectively (Definition 2.3), where $\widehat{D}$ is the empirical distribution of $N$ i.i.d.
-samples from $D$.
-There exists $N={(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}}\log(1/\delta)$ such that with probability $1-\delta$ we have that $\operatorname*{\mathbf{Pr}}[\widehat{h}(\mathbf{x})\neq y]\leq\operatorname*{\mathbf{Pr}}[h(\mathbf{x})\neq y]+\epsilon\;.$
+$\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[h_{t+1}(\mathbf{x})\neq y,\mathbf{x}\in S]$ $=\sum_{i=1}^{l}\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[h_{t+1}(\mathbf{x})\neq y\mid\mathbf{x}\in S_{i}]\operatorname*{\mathbf{Pr}}[S_{i}]=\sum_{i=1}^{l}\min_{\widehat{y}\in[K]}\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[y\neq\widehat{y}\mid\mathbf{x}\in S_{i}]\operatorname*{\mathbf{Pr}}[S_{i}]$
+$\leq\min_{\widehat{y}\in[K]}\sum_{i=1}^{l}\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[y\neq\widehat{y}\mid\mathbf{x}\in S_{i}]\operatorname*{\mathbf{Pr}}[S_{i}]=\min_{\widehat{y}\in[K]}\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[\widehat{y}\neq y,\mathbf{x}\in\cup_{i=1}^{l}S_{i}]$
+$\leq\operatorname*{\mathbf{Pr}}_{(x,y)\sim D}[h_{t}(\mathbf{x})\neq y,\mathbf{x}\in S]\;.$
+Summing over all $S\in\mathcal{S}$ we conclude the proof. ∎
+In the following claim, we show that if we use a sufficiently large number of samples, then with high probability we obtain an accurate estimate of the population piecewise constant approximation function.
+Claim 2.18 (Concentration of the Piecewise Constant Approximation).
+Let $\epsilon,\epsilon^{\prime},\delta\in(0,1)$ and $k,K\in\mathbb{Z}_{+}$ with $\epsilon^{\prime}\leq\epsilon/2$.
+Let $V\subseteq\mathbb{R}^{d}$ be a $k$-dimensional subspace, and consider a piecewise constant approximation $h:\mathbb{R}^{d}\to[K]$ of $D$, defined with respect to an $\epsilon^{\prime}$-approximating partition of $V$ (see Definition 2.3).
+Let $\widehat{D}$ be the empirical distribution obtained from $N$ i.i.d.
+samples drawn from $D$, and let $\widehat{h}$ be a piecewise constant approximation of $\widehat{D}$ defined with respect to the same partition.
+Then there exists $N=(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}\log(1/\delta)$ such that, with probability at least $1-\delta$, we have $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[\widehat{h}(\mathbf{x})\neq y]\leq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]+\epsilon\;.$
 Proof.
-Denote by $\mathcal{S}$ the approximating partition for the functions $h,\widehat{h}$.
-First, note that by the definition of an approximating partition (Definition 2.2) and by A.1, we have that $\operatorname*{\mathbf{Pr}}_{D}[S]={(\epsilon^{\prime}/k)^{\Omega(k)}}$ and hence $\lvert\mathcal{S}\rvert={(k/\epsilon^{\prime})^{O(k)}}$.
-Therefore, by the union bound and Hoeffding’s inequality it holds that if $N\geq C\log(\lvert\mathcal{S}\rvert K/\delta)/(\epsilon/\lvert S\rvert)^{2}={(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}}\log(1/\delta)$ for a sufficiently large universal constant $C>0$, then with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[\mathbf{x}\in S,y=i]-\operatorname*{\mathbf{Pr}}_{D}[\mathbf{x}\in S,y=i]\rvert\leq\epsilon/(4\lvert\mathcal{S}\rvert)$ for all $S\in\mathcal{S},i\in[K]$.
-Hence, with probability $1-\delta$ we have that for all $S\in\mathcal{S}$ it holds $\operatorname*{\mathbf{Pr}}[\widehat{h}(\mathbf{x})\neq y]\leq\operatorname*{\mathbf{Pr}}[h(\mathbf{x})\neq y]+\epsilon/(2\lvert\mathcal{S}\rvert)$.
+Let $\mathcal{S}$ denote the approximating partition used to define both $h$ and $\widehat{h}$.
+By the definition of an approximating partition (Definition 2.2) and using A.1, we have that for any $S\in\mathcal{S}$ it holds $\operatorname*{\mathbf{Pr}}_{D}[S]=(\epsilon^{\prime}/k)^{\Omega(k)}$ which implies that $\lvert\mathcal{S}\rvert=(k/\epsilon^{\prime})^{O(k)}$.
+Therefore, by the union bound and Hoeffding’s inequality it holds that if $N\geq C\log(\lvert\mathcal{S}\rvert K/\delta)/(\epsilon/\lvert S\rvert)^{2}=(k/\epsilon^{\prime})^{O(k)}(K/\epsilon)^{O(1)}\log(1/\delta)$ for a sufficiently large universal constant $C>0$, then with probability at least $1-\delta$, it holds that for every $S\in\mathcal{S}$ and every $i\in[K]$, we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[\mathbf{x}\in S,y=i]-\operatorname*{\mathbf{Pr}}_{D}[\mathbf{x}\in S,y=i]\rvert\leq\epsilon/(4\lvert\mathcal{S}\rvert)$.
+Hence, with probability $1-\delta$ we have that for all $S\in\mathcal{S}$ it holds $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[\widehat{h}(\mathbf{x})\neq y]\leq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]+\epsilon/(2\lvert\mathcal{S}\rvert)$.
 The claim follows, as the total error is simply the summation of the error over cubes and the error outside the partition which is at most $\epsilon^{\prime}$ from Definition 2.2.
 ∎
-Note that from Lines 1, 3 of Algorithm 1 we perform at most $\mathrm{poly}(K,1/\epsilon)$ iterations.
-Furthermore, in each iteration, we update the vector set with at most $\mathrm{poly}(K,1/\epsilon)$ vectors.
-Hence, it follows that $k_{t}\leq\mathrm{poly}(K,1/\epsilon)$ for all $t=1,\dots,T$.
-Assume that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+\epsilon$ for all $t=1,\dots,T$.
-Using the fact that $N=d2^{\mathrm{poly}(K,1/\epsilon)}\log(1/\delta)$ for a sufficiently large polynomial (Line 2 of Algorithm 1), we can apply Claim 2.15 and conclude that with probability $1-\delta$ there exist unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ for $t=[T]$ such that $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\mathrm{poly}(1/K,\epsilon)$.
-Thus, from Claim 2.16, we have that with probability $1-\delta$, for all $t\in[T]$, $\Phi_{t}\leq\Phi_{t-1}-\mathrm{poly}(\epsilon,1/K)$.
-After $T$ iterations, it follows that $\Phi_{T}\leq\Phi_{0}-T\mathrm{poly}(\epsilon,1/K)=1-T\mathrm{poly}(\epsilon,1/K)$.
-However, if $T$ is set to be a sufficiently large polynomial of $\epsilon$ and $K$, we would arrive at a contradiction, since $\Phi_{T}\geq 0$.
-Hence, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]\leq C\cdot\mathrm{OPT}+\epsilon$ for some $t\in\{1,\dots,T\}$.
-Since the error of $h_{t}$ can only be decreasing by Claim 2.17 and $h_{t}$ is close to its sample variant by Claim 2.18, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]<C\cdot\mathrm{OPT}+\epsilon$.
-Sample and Computational Complexity: From the analysis above, we have that the algorithm terminates in $\mathrm{poly}(1/\epsilon,K)$ iterations and at each iteration we draw $d2^{\mathrm{poly}(K,1/\epsilon)}\log(1/\delta)$ samples.
+Note that from Lines 1, 3 of Algorithm 1 we perform at most ${T=\mathrm{poly}(K,1/\epsilon)}$ iterations.
+Moreover, in each iteration we add at most $\mathrm{poly}(K,1/\epsilon)$ vectors to our current vector set $L_{t}$, hence the dimension $k_{t}$ of $V_{t}$ satisfies $k_{t}=\mathrm{poly}(K,1/\epsilon)$ for all $t=1,\dots,T$.
+Assume that for every $t=1,\dots,T$ the error satisfies $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>C\cdot\mathrm{OPT}+\epsilon$.
+Note that since $k_{t}=\mathrm{poly}(K,1/\epsilon)$ for all $t\in[T]$, if we choose $\epsilon^{\prime}=\mathrm{poly}(\epsilon/K)$ for a sufficiently small polynomial, then for all $t\in[T]$ we have $\epsilon^{\prime}\leq c\epsilon^{2}/(K^{2}\sqrt{k_{t}})$ for a sufficiently small absolute constant $c>0$.
+Hence, using that $N=d2^{\mathrm{poly}(K,1/\epsilon)}\log(1/\delta)$ for a sufficiently large polynomial (Line 2 of Algorithm 1), we can apply Claim 2.15 and conclude that with probability $1-\delta$ there exist unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ and unit vectors $\mathbf{w}^{(t)}\in W^{*}$ for $t\in[T]$ such that $\mathbf{w}^{(t)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\mathrm{poly}(1/K,\epsilon)$.
+By Claim 2.16, it follows that with probability at least $1-\delta$, the potential function decreases by at least $\mathrm{poly}(\epsilon,1/K)$ at each iteration, that is $\Phi_{t}\leq\Phi_{t-1}-\mathrm{poly}(\epsilon,1/K)$.
+After $T$ iterations, it obtain that $\Phi_{T}\leq\Phi_{0}-T\mathrm{poly}(\epsilon,1/K)=1-T\mathrm{poly}(\epsilon,1/K)$.
+However, since $T$ is chosen to be a sufficiently large polynomial of $\epsilon$ and $K$, we would arrive at a contradiction, since $\Phi_{T}\geq 0$.
+Hence, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]\leq C\cdot\mathrm{OPT}+\epsilon$ for some $t\in[T]$.
+Since the error of $h_{t}$ can only decrease over the iterations (see Claim 2.17) and $h_{t}$ is close to its empirical version (Claim 2.18), we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]<C\cdot\mathrm{OPT}+\epsilon$.
+Sample and Computational Complexity: From the analysis above, we have that the algorithm terminates in $T=\mathrm{poly}(1/\epsilon,K)$ iterations and at each iteration we draw $d2^{\mathrm{poly}(K,1/\epsilon)}\log(1/\delta)$ samples.
 Hence, we have that the total number of samples is of the order of $d2^{\mathrm{poly}(K,1/\epsilon)}\log(1/\delta)$.
-Moreover, the algorithm runs in $\mathrm{poly}(N)$ time as all operations can be implemented in polynomial time.
+Moreover, the algorithm runs in $\mathrm{poly}(N)$ time as all the computational operations can be implemented in polynomial time.
 ∎
 2.4 Proof of Lemma 2.7
-This section proves Lemma 2.7, one of the supporting lemmas used in the proof of Proposition 2.4.
-For completeness we restate the lemma.
-See 2.7 The proof of the lemma immediately follows from Lemma 2.19 and Claim 2.20.
-First, we establish Lemma 2.19, which provides a bound on the tails of a piecewise affine function.
+This section proves Lemma 2.7, one of the supporting technical lemmas used in the proof of Proposition 2.4. For completeness we restate the lemma below. See 2.7
+Proof.
+The proof of the lemma immediately follows from Claim 2.19 and Claim 2.20.
+First, we establish Claim 2.19, which provides a bound on the tails of a piecewise affine function.
 Then, using Claim 2.20, we show that there is indeed a suitable $\gamma$ to which this tail bound can be applied.
-Lemma 2.19 (Tail Bounds).
+Claim 2.19 (Tail Bounds).
 Let $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d}$, $t_{1},\dots,t_{K}\in\mathbb{R}$ and $s:\mathbb{R}^{d}\to\mathbb{R}$ such that $s(\mathbf{x})=\max_{i\in[K]}\max(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i},0)$.
@@
 from $\mathcal{N}(\mathbf{0},\mathbf{I})$.
-Now if for some $i\in[n^{2}]$ we have that $s(\mathbf{z}^{(i)})\geq n\gamma$, then there exists a $j\in[K]$ such that $\mathbf{w}^{(j)}\cdot\mathbf{z}^{(i)}+t_{j}\geq n\gamma$ and, furthermore, we have that
-$\mathbf{w}^{(j)}\cdot\mathbf{x}+t_{j}$ $=\frac{1}{n}\mathbf{w}^{(j)}\cdot\mathbf{z}^{(i)}+t_{j}+\frac{1}{n}\sum_{k\neq i}\mathbf{w}^{(j)}\cdot\mathbf{z}^{(k)}\geq\gamma+\mathbf{w}^{(j)}\cdot\bigg{(}\frac{1}{n}\sum_{k\neq i}\mathbf{z}^{(k)}\bigg{)}.$
-Note that $\frac{1}{n}\sum_{k\neq i}\mathbf{z}^{(k)}$ is distributed according to $\mathcal{N}(\mathbf{0},\frac{n^{2}-1}{n^{2}}\mathbf{I})$; hence, by symmetry we have that $\operatorname*{\mathbf{Pr}}[\mathbf{w}^{(j)}\cdot(\frac{1}{n}\sum_{k\neq i}\mathbf{z}^{(k)})\geq 0]=1/2$.
+Now if for some $i\in[n^{2}]$ we have that $s(\mathbf{z}^{(i)})\geq n\gamma$, then there exists $j\in[K]$ such that $\mathbf{w}^{(j)}\cdot\mathbf{z}^{(i)}+t_{j}\geq n\gamma$ (by definition of $s(\mathbf{x})$) and, furthermore, we have that
+$\mathbf{w}^{(j)}\cdot\mathbf{x}+t_{j}$ $=\frac{1}{n}\mathbf{w}^{(j)}\cdot\mathbf{z}^{(i)}+t_{j}+\frac{1}{n}\sum_{k\neq i,k\in[n^{2}]}\mathbf{w}^{(j)}\cdot\mathbf{z}^{(k)}\geq\gamma+\mathbf{w}^{(j)}\cdot\bigg{(}\frac{1}{n}\sum_{k\neq i,k\in[n^{2}]}\mathbf{z}^{(k)}\bigg{)}.$
+Note that $\frac{1}{n}\sum_{k\neq i}\mathbf{z}^{(k)}$ is distributed according to $\mathcal{N}(\mathbf{0},\frac{n^{2}-1}{n^{2}}\mathbf{I})$.
+Hence, by symmetry we have that $\operatorname*{\mathbf{Pr}}[\mathbf{w}^{(j)}\cdot(\frac{1}{n}\sum_{k\neq i}\mathbf{z}^{(k)})\geq 0]=1/2$.
 As a result we have that $\operatorname*{\mathbf{Pr}}[s(\mathbf{x})\geq\gamma\mid s(\mathbf{z}^{(i)})\geq n\gamma]\geq 1/2$, for all $i\in[n^{2}]$.
 Denote by $p_{n}=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq n\gamma]$.
-To prove Lemma 2.19, it suffices to show that $p_{n}\leq 4p_{1}/n^{2}$, for all $n\geq 1$.
-Note that this holds trivially true for $n=1$.
+To prove Claim 2.19, it suffices to show that $p_{n}\leq 4p_{1}/n^{2}$, for all $n\geq 1$.
+We will prove it by induction.
+Note that for $n=1$ the hypothesis holds trivially.
 Assume as an inductive hypothesis that $p_{n-1}\leq 4p_{1}/(n-1)^{2}$ for some $n\geq 2$.
@@
 $\geq p_{n}/2-(n^{2}-1)p_{n}^{2}\;,$
-where in the first inequality we used the union bound, and in the last inequality we used the independence of the $\mathbf{z}^{(i)}$’s and that $\operatorname*{\mathbf{Pr}}[s(\mathbf{x})\geq\gamma\mid s(\mathbf{z}^{(i)})\geq n\gamma]\geq 1/2$, for all $i\in[n^{2}]$.
+where in the first inequality we used the union bound, and in the second inequality we used the independence of the $\mathbf{z}^{(i)}$’s and that $\operatorname*{\mathbf{Pr}}[s(\mathbf{x})\geq\gamma\mid s(\mathbf{z}^{(i)})\geq n\gamma]\geq 1/2$, for all $i\in[n^{2}]$.
 Note that it holds $p_{n}\leq p_{n-1}$, since $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq n\gamma]\leq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq(n-1)\gamma]$.
@@
 $p_{1}=\operatorname*{\mathbf{Pr}}[A]\geq\operatorname*{\mathbf{Pr}}\left[A,\bigcup_{i\in[n^{2}]}\left(B_{i}^{(n)},\bigcap_{j\neq i}B_{j}^{(n)}\right)\right]\geq n^{2}p_{n}/4\;.$
-Therefore, we have that $p_{n}\leq 4p_{1}/n^{2}$, which by the principle of mathematical induction completes the proof of Lemma 2.19. ∎
-In order to apply Lemma 2.19, we need to find a threshold such that the probability that $s$ is greater than this threshold is bounded from above by $1/{48}$. We show the following
+Therefore, we have that $p_{n}\leq 4p_{1}/n^{2}$, which completes the proof of Claim 2.19. ∎
+To apply Claim 2.19, we must choose a threshold $\gamma$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(0,I)}[s(\mathbf{x})\geq\gamma]\leq 1/48.$ We now show the following result.
 Claim 2.20.
-Fix $c\in(0,1)$, there exists $\gamma>0$ such that
+For any sufficiently small $c\in(0,1)$, there exists $\gamma>0$ such that
 $c\geq\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]\geq\frac{c}{2}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\neq 0]\;.$
 Proof of Claim 2.20.
-First notice that the quantity $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]$ is continuous with respect to $\gamma$ for any $\gamma>0$.
+Notice that the quantity $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[s(\mathbf{x})\geq\gamma]$ is continuous with respect to $\gamma$ for any $\gamma>0$.
 Indeed, $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\gamma\leq s(\mathbf{x})\leq\gamma+\epsilon]\leq\sum_{i=1}^{K}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[-t_{i}+\gamma\leq\mathbf{w}^{(i)}\cdot\mathbf{x}\leq-t_{i}+\gamma+\epsilon]$ where each term such that $\mathbf{w}^{(i)}\neq\mathbf{0}$ tends to 00 as $\epsilon$ tends to 00 and each term such that $\mathbf{w}^{(i)}=\mathbf{0}$ is equal to 00 from the fact that $-t_{i}+\gamma>0$.
@@
 ∎
-Combining the the above two statements completes the proof of Lemma 2.7.
+Combining the the above two statements completes the proof of Lemma 2.7. ∎
 3 Learning Multi-Index Models with Agnostic Noise
-In this section, we extend the algorithm of Section 2 for a general class of functions that satisfy specific regularity conditions (Definition 1.2), establishing Theorem 1.4.
-First, we state the formal versions of the conditions and theorem, and then provide pseudocode for the algorithm along with the necessary definitions.
-As we will discuss later, we need a continuity condition on the function in order to be able to approximate it over small regions.
+In this section, we extend the algorithm presented in Section 2 for a general class of MIMs that satisfy specific regularity conditions (Definition 1.2), establishing Theorem 1.5.
+We begin by stating the formal versions of the regularity conditions and the theorem.
+Following this, we provide the pseudocode for the extended algorithm along with the necessary definitions.
+As we will discuss later, we need a continuity condition on the target function in order to be able to accurately approximate it over small regions.
 To this end, we define the Gaussian Surface Area of a multiclass classifier.
@@
 For a class of multiclass concepts $\mathcal{C}$, we define $\Gamma(\mathcal{C}):=\sup_{f\in\mathcal{C}}\Gamma(f)$.
-Now we proceed to define the regularity conditions that we will assume in order to get efficient algorithms. Formally, we define a class of well-behaved functions as follows:
+Now we proceed to define the regularity conditions that we assume in order to obtain efficient algorithms.
 Definition 3.2 (Well-Behaved Multi-Index Model).
-Fix $d,K,m\in\mathbb{Z}_{+},\delta,\epsilon,\sigma,\tau\in(0,1)$ and $\mathcal{Y}\subseteq\mathbb{Z}$ of finite cardinality.
-We say that a class of functions $\mathcal{F}$ from $\mathbb{R}^{d}\mapsto\mathcal{Y}$ is a class of $(m,\epsilon,\delta,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models if, for every $f\in\mathcal{F}$, the following conditions hold:
+Let $d,K,m\in\mathbb{Z}_{+},\alpha,\zeta,\sigma,\tau\in(0,1)$ and $\mathcal{Y}\subseteq\mathbb{Z}$ of finite cardinality.
+We say that a class of functions $\mathcal{F}$ from $\mathbb{R}^{d}\mapsto\mathcal{Y}$ is a family of concept classes of $(m,\zeta,\alpha,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models if, for every $f\in\mathcal{F}$, the following conditions hold:
   1. 1.
@@
   3. 3.
-    Let $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, for any random variable $y$ supported on $\mathcal{Y}$ with $\operatorname*{\mathbf{Pr}}_{\mathbf{x},y}[f(\mathbf{x})\neq y]\leq\epsilon$ and any subspace $V\subseteq\mathbb{R}^{d}$,
+    Let $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
+For any random variable $y$ supported on $\mathcal{Y}$ with $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]\leq\zeta$ and any subspace $V\subseteq\mathbb{R}^{d}$,
     1. (a)
-      either there exists a function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x},y}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$.
+      either there exists a function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x},y}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$,
     2. (b)
-      or with probability at least $\delta$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, where $\mathbf{x}_{0}$ is independent of $\mathbf{x}$, there exists a polynomial $p:U\to\mathbb{R}$ of degree at most $m$ and a label $z\in\mathcal{Y}$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x}}[p(\mathbf{x}^{U})]=0$ and $\operatorname*{\mathbf{E}}_{\mathbf{x},y}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>\sigma\|p(\mathbf{x}^{U})\|_{2}$, where $U=(V+W)\cap V^{\perp}$.
-We can now state the main theorem of this section.
+      or with probability at least $\alpha$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, where $\mathbf{x}_{0}$ is independent of $\mathbf{x}$, there exists a polynomial $p:U\to\mathbb{R}$ of degree at most $m$ and a label $z\in\mathcal{Y}$ such that $\operatorname*{\mathbf{E}}[p(\mathbf{x}^{U})]=0$ and $\operatorname*{\mathbf{E}}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>\sigma\|p(\mathbf{x}^{U})\|_{2}$, where $U=(V+W)\cap V^{\perp}$.
+We can now state the formal version of Theorem 1.5.
 Theorem 3.3 (Approximately Learning Multi-Index Models with Agnostic Noise).
@@
 If $\zeta\geq\mathrm{OPT}+\epsilon$, then, Algorithm 3 draws $N={d}^{O(m)}2^{\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))}\log(1/\delta)$ i.i.d.
-samples from $D$, runs in time $\mathrm{poly}(d,N,\lvert\mathcal{Y}\rvert)$, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon\;.$
+samples from $D$, runs in time $\mathrm{poly}(N)$, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon\;.$
 Definition 3.2 outlines the crucial properties necessary for our algorithm to produce a satisfactory solution.
-We first assume that our concept class consists of Multi-Index Models with a bounded Gaussian surface area.
-This condition bounds the complexity of the decision boundary of the multi-index models or equivalently it bounds how fast the function changes labels.
-We use this condition in order to show that if we partition our space to sufficiently small cubes then the function should be almost constant.
+We first assume that our concept class consists of Multi-Index Models with finite Gaussian surface area.
+This condition bounds the complexity of the decision boundary of the multi-index models.
+We leverage this condition in order to show that if we partition our space to sufficiently small cubes then the function should be almost constant within each cube.
 This is a condition that holds for many classes of interest, such as multiclass linear classifiers and intersections of halfspaces.
-In addition, we impose a structural condition on the class to ensure progress at each step of the algorithm.
-Intuitively, if $V$ is the subspace we have currently identified, we require that either it already provides a good approximation Condition 3(a) or we can use moments to uncover further directions in the hidden subspace $W$ Condition 3(b).
-Specifically, we assume that for any perturbation $y$ of a function $f$ from the class that we are attempting to learn, and for any given subspace $V$, one of the following must hold: (i) there exists a good approximation to $y$ using a function defined over $V$ (Condition 3(a) of Definition 3.2), or (ii) with some probability over orthogonal cylinders to $V$, there exists a low-degree, zero-mean, variance-one polynomial along directions that correlate with the ground-truth subspace $W$, which non-trivially correlates with the level sets of $y$ over that orthogonal cylinder (Condition 3(b) of Definition 3.2).
+In addition, we impose a structural condition on the target class to guarantee progress at each step of the algorithm.
+Intuitively, if $V$ is the subspace we have currently identified, we require that either it already provides a good approximation (Condition 3(a)) or we can use moments to uncover further directions in the hidden subspace $W$ Condition 3(b).
+Specifically, we assume that for any perturbation $y$ of a function $f$ from the class that we are attempting to learn, and for any given subspace $V$, one of the following must hold: (i) there exists a good approximation to $y$ using a function defined over $V$ (Condition 3(a)), or (ii) with some probability over orthogonal cylinders to $V$, there exists a low-degree, zero-mean, variance-one polynomial along directions that correlate with the ground-truth subspace $W$, which non-trivially correlates with the level sets of $y$ over that orthogonal cylinder (Condition 3(b).
 In this manner, extracting these directions and adding them to $V$ enables an improved approximation of the hidden-subspace $W$.
-Although Condition 3(b) is defined over specific points with a measure of zero, we demonstrate in Lemma 3.5 that, by leveraging our assumptions, there exists a $\delta$ fraction of finite width cylinders that exhibit this property.
+Although Condition 3(b) is defined over specific points with a measure of zero, we demonstrate in Lemma 3.5 that, by leveraging our assumptions, there exists a $\alpha$ fraction of finite width cylinders that exhibit this property.
 Condition 3(a) of Definition 3.2 further provides an approximation factor that the optimal function within the subspace $V$ can achieve.
@@
 Consequently, the function that minimizes the disagreement over these cubes will also exhibit low error.
-Input: Accuracy $\epsilon\in(0,1)$, failure probability $\delta\in(0,1)$, and sample access to a distribution $D$ over $\mathbb{R}^{d}\times\mathcal{Y}$.
-Output: A hypothesis $h$ such that, with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$.
+Input: Accuracy $\epsilon\in(0,1)$, failure probability $\delta\in(0,1)$, and sample access to a distribution $D$ over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$.
+Output: A hypothesis $h$ such that, with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$, where $\mathrm{OPT}\coloneqq\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$.
   1. 1.
-    Let $T,p$ be sufficiently large polynomials in $1/\alpha,1/\epsilon,1/\sigma,K,\Gamma,\lvert\mathcal{Y}\rvert,m$, and let $C$ be a sufficiently large universal constant..
+    Let $T,p$ be sufficiently large polynomials in $1/\alpha,1/\epsilon,1/\sigma,K,\Gamma,\lvert\mathcal{Y}\rvert,m$, and let $C$ be a sufficiently large universal constant.
   2. 2.
@@
     1. (a)
-      Draw a set $S$ of $N$ i.i.d. samples from $D$.
+      Draw a set $S_{t}$ of $N$ i.i.d. samples from $D$.
     2. (b)
-      $\mathcal{E}_{t}\leftarrow$ Algorithm 4($\epsilon,\epsilon^{\prime},\delta,\mathrm{span}(L_{t}),S$).
+      $\mathcal{E}_{t}\leftarrow$ Algorithm 4($\epsilon,\epsilon^{\prime},\delta,\mathrm{span}(L_{t}),S_{t}$).
     3. (c)
@@
   5. 5.
-    Return a piecewise constant approximating classifier $h_{\mathcal{S}}$, as defined in Definition 2.3, using $N$ i.i.d.
-samples from $D$, by finding the most frequent label at any subset of $\mathcal{S}$.
+    Draw $N$ i.i.d.
+samples from $D$ and construct the piecewise constant classifier $h_{\mathcal{S}}$ as follows: For each $S\in\mathcal{S}$, assign a label that appears most frequently among the samples falling in $S$.
+Return $h_{\mathcal{S}}$.
 Algorithm 3 Learning Multi-Index Models with Agnostic Noise
-Input: $\epsilon>0$, $\epsilon^{\prime}>0$, $\delta>0$, a subspace $V$ of $\mathbb{R}^{d}$ and a set of samples $\{(\mathbf{x}^{(1)},y_{1}),\dots,(\mathbf{x}^{(N)},y_{N})\},\mathbf{x}^{(i)}\in\mathbb{R}^{d},y_{i}\in\mathcal{Y}$.
+Input: $\epsilon>0$, $\epsilon^{\prime}>0$, $\delta>0$, a subspace $V$ of $\mathbb{R}^{d}$ and a set of samples $\{(\mathbf{x}^{(1)},y_{1}),\dots,(\mathbf{x}^{(N)},y_{N})\}$ from a distribution $D$ over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$.
 Output: A set of unit vectors $\mathcal{E}$.
@@
   4. 4.
-    For each $S\in\mathcal{S}$ and each $i\in\mathcal{Y}$, find a polynomial $p_{S,i}(\mathbf{x})$ such that
+    For each $S\in\mathcal{S}$ and each $i\in\mathcal{Y}$, compute a polynomial $p_{S,i}(\mathbf{x})$ such that
     $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[$ $(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}^{\perp V}))^{2}\mid\mathbf{x}\in S]$
@@
 Lemma 3.4 (Piecewise Constant Approximation Suffices).
-Let $d\in\mathbb{Z}_{+},\epsilon,\epsilon^{\prime}\in(0,1)$ and let $\mathcal{Y}$ be a set of finite cardinality.
+Let $d,k\in\mathbb{Z}_{+},\epsilon,\epsilon^{\prime}\in(0,1)$ and let $\mathcal{Y}$ be a set of finite cardinality.
 Let $D$ be a distribution supported on $\mathbb{R}^{d}\times\mathcal{Y}$, whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, and let $V\subseteq\mathbb{R}^{d}$ be a $k$-dimensional subspace, with $k\geq 1$.
@@
 Additionally, assume that there exists a function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$.
-Let $\eta=\epsilon^{\prime 2}/(C\Gamma(f)\lvert\mathcal{Y}\rvert\sqrt{k})$, for $C>0$ a sufficiently large universal constant, and $\mathcal{S}$ be an $\eta$-piecewise approximating partition with respect to $V$, as defined in Definition 2.2.
-Then, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]\leq\tau+\epsilon+\epsilon^{\prime}$, where $h_{\mathcal{S}}$ is an $\eta$-piecewise constant approximation function as defined in Definition 2.3.
+Let $\eta=\epsilon^{\prime 2}/(C\Gamma(f)\lvert\mathcal{Y}\rvert\sqrt{k})$, for $C>0$ a sufficiently large universal constant, and $\mathcal{S}$ be an $\eta$-piecewise approximating partition with respect to $V$ (see Definition 2.2).
+Then, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]\leq\tau+\epsilon+\epsilon^{\prime}$, where $h_{\mathcal{S}}$ is an $\eta$-piecewise constant approximation function (see Definition 2.3).
 Proof.
-Fix $i\in\mathcal{Y}$ and consider the boolean functions $f_{i}(\mathbf{x})\coloneqq\mathds{1}(f(\mathbf{x})=i)$, $g_{i}(\mathbf{x})\coloneqq\mathds{1}(g(\mathbf{x}^{V})=i)$.
-By $T_{\rho}$ we denote the Gaussian noise operator defined in Definition A.3.
+Fix $i\in\mathcal{Y}$ and $\rho\in(0,1)$ and consider the boolean functions $f_{i}(\mathbf{x})\coloneqq\mathds{1}(f(\mathbf{x})=i)$, $g_{i}(\mathbf{x})\coloneqq\mathds{1}(g(\mathbf{x}^{V})=i)$.
+We denote the Gaussian noise operator by $T_{\rho}$, as defined in Definition A.3.
 By A.4, we have that $\left\|f_{i}-g_{i}\right\|_{L_{1}}\geq\left\|T_{\rho}f_{i}-T_{\rho}g_{i}\right\|_{L_{1}}$.
 Hence, from the triangle inequality we also have that $\left\|f_{i}-g_{i}\right\|_{L_{1}}+\left\|T_{\rho}f_{i}-f_{i}\right\|_{L_{1}}\geq\left\|f_{i}-T_{\rho}g_{i}\right\|_{L_{1}}$.
-Furthermore, by A.5, we have that $\left\|T_{\rho}f_{i}-f_{i}\right\|_{L_{1}}\leq\sqrt{\pi}\Gamma(f_{i})\rho$.
+Furthermore, by the Ledoux-Pisier inequality (A.5), we have that $\left\|T_{\rho}f_{i}-f_{i}\right\|_{L_{1}}\leq\sqrt{\pi}\Gamma(f_{i})\rho$.
 Therefore, we have that
 $\left\|f_{i}-T_{\rho}g_{i}\right\|_{L_{1}}\leq\left\|f_{i}-g_{i}\right\|_{L_{1}}+\sqrt{\pi}\Gamma(f_{i})\rho\;.$
-Moreover, note that for any continuously differentiable function $F:\mathbb{R}^{d}\mapsto\mathbb{R}$, it holds that $F(\mathbf{a})-F(\mathbf{b})=\int_{0}^{1}\nabla F(\mathbf{b}+t(\mathbf{a}-\mathbf{b}))\cdot(\mathbf{a}-\mathbf{b})\mathrm{d}t$ and from the Cauchy-Schwarz inequality, we have that $|F(\mathbf{a})-F(\mathbf{b})|\leq\max_{t\in(0,1)}\|\nabla F(\mathbf{b}+t(\mathbf{a}-\mathbf{b}))\|\|\mathbf{a}-\mathbf{b}\|$.
-By taking $F=T_{\rho}g_{i}$ and using A.4, we get that
-$|T_{\rho}g_{i}(\mathbf{x})-T_{\rho}g_{i}(\mathbf{y})|\leq\frac{\|\mathbf{x}-\mathbf{y}\|}{\rho}\;,$
-for any $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$. For each $S\in\mathcal{S}$, pick a point $\mathbf{x}_{S}\in S$. By averaging over the partition, we have that
+Moreover, note that for any continuously differentiable function $F:\mathbb{R}^{d}\mapsto\mathbb{R}$, it holds that $F(\mathbf{a})-F(\mathbf{b})=\int_{0}^{1}\nabla F(\mathbf{b}+t(\mathbf{a}-\mathbf{b}))\cdot(\mathbf{a}-\mathbf{b})\mathrm{d}t$.
+By the Cauchy-Schwarz inequality, we have that $|F(\mathbf{a})-F(\mathbf{b})|\leq\max_{t\in(0,1)}\|\nabla F(\mathbf{b}+t(\mathbf{a}-\mathbf{b}))\|\|\mathbf{a}-\mathbf{b}\|$.
+By taking $F=T_{\rho}g_{i}$ and using A.4, we have that for any $\mathbf{x},\mathbf{y}\in\mathbb{R}^{d}$, it holds
+$|T_{\rho}g_{i}(\mathbf{x})-T_{\rho}g_{i}(\mathbf{y})|\leq\frac{\|\mathbf{x}-\mathbf{y}\|}{\rho}\;.$
+For each $S\in\mathcal{S}$, pick a point $\mathbf{x}_{S}\in S$.
+Then for any $\mathbf{x}\in S$, the distance between $\mathbf{x}$ and $\mathbf{x}_{S}$ is at most the diameter of $S$.
+In our setting, since the partition is constructed along a $k$-dimensional subspace $V$ with edge length roughly $\eta$ (i.e., each coordinate is in an interval of length $\eta$), the distance between any $\mathbf{x}\in S$ and $\mathbf{x}_{S}$ is at most $\|\mathbf{x}-\mathbf{x}_{S}\|\leq\sqrt{k}\eta.$ Thus, for every $\mathbf{x}\in S$
+$|T_{\rho}g_{i}(\mathbf{x})-T_{\rho}g_{i}(\mathbf{x}_{S})|\leq\frac{\|\mathbf{x}-\mathbf{x}_{S}\|}{\rho}\leq\frac{\sqrt{k}\cdot\eta}{\rho}.$
+By averaging over the partition $\mathcal{S}$, we have that
 $\sum_{S\in\mathcal{S}}\left\|(T_{\rho}g_{i}(\mathbf{x}_{S})-T_{\rho}g_{i})\mathds{1}(\mathbf{x}\in S)\right\|_{L_{1}}\leq\sum_{S\in\mathcal{S}}\frac{\sqrt{k}\eta}{\rho}\operatorname*{\mathbf{Pr}}[S]\leq\frac{\sqrt{k}\eta}{\rho}\;.$
@@
 $\sum_{S\in\mathcal{S}}\left\|(T_{\rho}g_{i}(\mathbf{x}_{S})-f_{i})\mathds{1}(\mathbf{x}\in S)\right\|_{L_{1}}\leq\left\|f_{i}-g_{i}\right\|_{L_{1}}+\sqrt{\pi}\Gamma(f_{i})\rho+\frac{\sqrt{k}\eta}{\rho}\;.$
-Furthermore, observe that by the linearity of expectation and the fact that $\sum_{i\in\mathcal{Y}}g_{i}(\mathbf{x})=1$ we have that $\sum_{i\in\mathcal{Y}}T_{\rho}g_{i}(\mathbf{x})=1$ for $\mathbf{x}\in\mathbb{R}^{d}$.
+Furthermore, by linearity of expectation and the fact that $\sum_{i\in\mathcal{Y}}g_{i}(\mathbf{x})=1$ we have that $\sum_{i\in\mathcal{Y}}T_{\rho}g_{i}(\mathbf{x})=1$ for $\mathbf{x}\in\mathbb{R}^{d}$.
 Moreover, $T_{\rho}g_{i}(\mathbf{x})\in[0,1]$ for all $\mathbf{x}\in\mathbb{R}^{d},i\in\mathcal{Y}$.
-Therefore, we can consider the multiclass function $G:\mathbb{R}^{d}\to\mathcal{Y}$ such that $G(\mathbf{x})$ is a random label sampled from the probability distribution induced by $T_{\rho}g_{i}(\mathbf{x}_{S})$, where $S\in\mathcal{S}$ such that $\mathbf{x}\in S$.
-Note that for every probability vector $\mathbf{p}\in\mathbb{R}^{|\mathcal{Y}|}$ and label $z\in\mathcal{Y}$ we have that
+Now, we define a multiclass function $G:\mathbb{R}^{d}\to\mathcal{Y}$ in the following way: for each point $\mathbf{x}\in\mathbb{R}^{d}$, $G(\mathbf{x})$ is a random label chosen according to the probability distribution induced by $T_{\rho}g_{i}(\mathbf{x}_{S})$, where $S\in\mathcal{S}$ is the region such that $\mathbf{x}\in S$.
+Furthermore, for every probability vector $\mathbf{p}\in\mathbb{R}^{|\mathcal{Y}|}$ (where each element of $\mathbf{p}$ represents the probability of the corresponding label) and the unit vector $\mathbf{e}_{z}$, which represents the one-hot encoding of the label $z\in\mathcal{Y}$.
+We have that
 $\|\mathbf{p}-\mathbf{e}_{z}\|_{1}=\lvert\mathbf{p}_{z}-1\rvert+\sum_{i\neq z,i\in\mathcal{Y}}\lvert\mathbf{p}_{i}\rvert=2\operatorname*{\mathbf{Pr}}_{i\sim\mathbf{p}}[i\neq z]\;.$
@@
 $\leq\sum_{i\in\mathcal{Y}}\|f_{i}-g_{i}\|_{L_{1}}/2+\sqrt{\pi}\Gamma(f)\rho+\lvert\mathcal{Y}\rvert\sqrt{k}\eta/\rho\;.$
-Noting that $\sum_{i\in\mathcal{Y}}\|f_{i}-g_{i}\|=2\tau$ set $\rho=\epsilon^{\prime}/(C{\Gamma(f)})$ and $\eta={\epsilon^{\prime}\rho}/{(C|\mathcal{Y}|\sqrt{k}})$, for a sufficiently large constant $C>0$.
+Noting that $\sum_{i\in\mathcal{Y}}\|f_{i}-g_{i}\|=2\tau$, set $\rho=\epsilon^{\prime}/(C{\Gamma(f)})$ and $\eta={\epsilon^{\prime}\rho}/{(C|\mathcal{Y}|\sqrt{k}})$, for a sufficiently large constant $C>0$.
 With these choices, we obtain $\operatorname*{\mathbf{Pr}}[G(\mathbf{x})\neq f(\mathbf{x}),\mathbf{x}\in A]\leq\tau+\epsilon^{\prime}/2$.
 By the triangle inequality we have that $\operatorname*{\mathbf{Pr}}[G(\mathbf{x})\neq y,\mathbf{x}\in A]]\leq\tau+\epsilon+\epsilon^{\prime}/2$.
-Since $h_{\mathcal{S}}$ chooses the label that minimizes the error at each $S\in\mathcal{S}$, we have that $\operatorname*{\mathbf{Pr}}[h_{\mathcal{S}}(\mathbf{x})\neq y,\mathbf{x}\in A]\leq\operatorname*{\mathbf{Pr}}[G(\mathbf{x})\neq y,\mathbf{x}\in A]]$.
-Finally, considering that $\operatorname*{\mathbf{Pr}}[\mathbf{x}\not\in A]\leq\eta$ completes the proof of Lemma 3.4.
-In the following lemma, we show that, due to the bound on the Gaussian surface area of $f$, if there exist level sets of $f$ for which the conditional moments on a subspace $V$ are large with non-trivial probability over $V$, then the same property holds on sufficiently small cubes.
-That is, when $V$ is partitioned into sufficiently thin cubes, the conditional moments of the level sets of $f$ on these cubes will be large with non-trivial probability over the choice of cube.
+Since $h_{\mathcal{S}}$ chooses the label that minimizes the error at each $S\in\mathcal{S}$, we have that $\operatorname*{\mathbf{Pr}}[h_{\mathcal{S}}(\mathbf{x})\neq y,\mathbf{x}\in A]\leq\operatorname*{\mathbf{Pr}}[G(\mathbf{x})\neq y,\mathbf{x}\in A]$.
+Finally, considering that $\operatorname*{\mathbf{Pr}}[\mathbf{x}\not\in A]\leq\eta$ (recall that $\eta\leq\epsilon^{\prime 2}$) completes the proof of Lemma 3.4.
+∎
+In the following lemma, we show that, due to the bound on the Gaussian surface area of $f$, if there is a non-negligible probability that the conditional moments of $f$’s level sets along a subspace $V$ are large, the same property holds when $V$ is partitioned into sufficiently small cubes.
+That is, if we partition $V$ into thin cubes, then with non-trivial probability the conditional moments computed over each cube remain large.
 The intuition behind the proof is that, since the function is smooth (by the GSA bound), re-randomizing over sufficiently small cubes barely affects its output.
-However, by re-randomizing over cubes, the moments of the level-sets become independent of which point in the cube we condition on.
+However, by re-randomizing over cubes, the moments of the levelsets become independent of which point in the cube we condition on.
 As a result, if we observe distinguishing moments when conditioning on a point in the cube, then these moments persist when conditioning on the entire cube.
 Lemma 3.5 (Existence of Large Moments on Small Cubes).
-Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ and let $\mathcal{F}$ be a class of $(m,\epsilon,\delta,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models according to Definition 3.2.
-Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$, with $k\geq 1$, and let $\mathcal{S}$ be an $\eta$-approximating partition of $V$ according to Definition 2.2, with $\eta=\epsilon^{\prime 3}/(C\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2})$, for $C>0$ a sufficiently large universal constant.
-If for a function $f\in\mathcal{F}$ with $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\epsilon+\epsilon^{\prime}$ there exist no function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$, then there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\delta$ such that for each $S\in\mathcal{T}$ there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x},y}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}\in S]>\sigma\|p(\mathbf{x}^{U})\|_{2}$, where $U=(V+W)\cap V^{\perp}$.
+Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ and let $\mathcal{F}$ be a class of $(m,\zeta+O(\epsilon),\alpha,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models according to Definition 3.2.
+Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$, with $k\geq 1$, and let $\mathcal{S}$ be an $\eta$-approximating partition of $V$ according to Definition 2.2, with $\eta=\epsilon^{3}/(C\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2})$, for $C>0$ a sufficiently large universal constant.
+Suppose that for $f\in\mathcal{F}$ with $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\zeta$ there exists no function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$, then there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\alpha$ such that for each $S\in\mathcal{T}$ there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ and $z\in\mathcal{Y}$ such that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}\in S]>\sigma\|p(\mathbf{x}^{U})\|_{2}$, where $U=(V+W)\cap V^{\perp}$.
 Proof.
-Note that, since $\mathcal{F}$ satisfies Definition 3.2, with probability at least $\delta$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, independent of $\mathbf{x}$, there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that
+Since $\mathcal{F}$ satisfies Definition 3.2, with probability at least $\alpha$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, independent of $\mathbf{x}$, there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that
 $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim\mathcal{D}}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>\sigma\|p(\mathbf{x}^{U})\|_{2}\;,$
@@
 In order to prove the statement, we need to show that the above condition can be extended from a single point to encompass an entire cube $S$.
-To achieve this, we introduce a random variable $y^{\prime}$, which is $\epsilon$-close to $f$, such that, when conditioned on any $S\subset\mathcal{S}$, $y^{\prime}$ is distributed uniformly within $S$ along the directions defined by $V$, while the other directions remain consistent with those of $y$.
+To achieve this, we introduce a random variable $y^{\prime}$, which is $(\zeta+O(\epsilon))$-close to $f$, such that, when conditioned on any $S\in\mathcal{S}$, $y^{\prime}$ is distributed uniformly within $S$ along the directions defined by $V$, while the other directions remain consistent with those of $y$.
 For any $\mathbf{x}\in\mathbb{R}^{d}$, denote by $\mathcal{S}_{\mathbf{x}}$ the set $S\in\mathcal{S}$ such $\mathbf{x}\in S$.
@@
 Note that $\mathbf{x}^{\prime}$ also follows the standard normal distribution over $\mathbb{R}^{d}$ because its $V$-component is resampled from $\mathcal{N}(\mathbf{0},\mathbf{I})$.
-In the following claim we will prove that $f(\mathbf{x})$ and $f(\mathbf{x}^{\prime})$ are approximately the same.
+In the following claim we will prove that $f(\mathbf{x})$ and $f(\mathbf{x}^{\prime})$ are approximately close.
 Claim 3.6.
-It holds that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f(\mathbf{x}^{\prime})]=O(\epsilon^{\prime})$.
+It holds that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f(\mathbf{x}^{\prime})]=O(\epsilon)$.
 Proof.
-Define the boolean function $f_{i}(\mathbf{x})\coloneqq\mathds{1}(f(\mathbf{x})=i)$ and denote by $g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})\coloneqq T_{\rho}f_{i}(\mathbf{x}_{0}^{\perp V}+\mathbf{x}^{V})$, where $T_{\rho}$ denotes the Gaussian noise operator defined in Definition A.3.
-Fix $\mathbf{x}_{0}\in\mathbb{R}^{d},S\in\mathcal{S}$ we first show that $\lvert g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})-g_{\mathbf{x}_{0}}^{(i)}(\mathbf{z})\rvert$ is small for $\mathbf{x}^{V},\mathbf{z}^{V}\in S$.
-Note that for any continuously differentiable function $F:V\mapsto\mathbb{R},\mathbf{a},\mathbf{b}\in V$, it holds that $F(\mathbf{a})-F(\mathbf{b})=\int_{0}^{1}\nabla F(\mathbf{b}+t(\mathbf{a}-\mathbf{b}))\cdot(\mathbf{a}-\mathbf{b})\mathrm{d}t$ and from Cauchy-Schwarz inequality, we have that $|F(\mathbf{a})-F(\mathbf{b})|\leq\max_{t\in(0,1)}\|\nabla F(\mathbf{b}+t(\mathbf{a}-\mathbf{b}))\|\|\mathbf{a}-\mathbf{b}\|$.
-By taking $F(\mathbf{x})=g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})$ and using A.4, we get that
+For each $i\in\mathcal{Y}$, define the boolean functions $f_{i}(\mathbf{x})\coloneqq\mathds{1}(f(\mathbf{x})=i)$ and let $g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})\coloneqq T_{\rho}f_{i}(\mathbf{x}_{0}^{\perp V}+\mathbf{x}^{V})$, where $T_{\rho}$ is the Gaussian noise operator as given in Definition A.3.
+Fix a point $\mathbf{x}_{0}\in\mathbb{R}^{d}$ and a cube $S\in\mathcal{S}$.
+We first show that for any two points $\mathbf{x},\mathbf{z}\in S$, the difference $\lvert g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})-g_{\mathbf{x}_{0}}^{(i)}(\mathbf{z})\rvert$ is small.
+By using that the gradient of $g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})$ is bounded by $1/\rho$ and that the distance between any $\mathbf{x},\mathbf{z}\in S$ is at most $\|\mathbf{x}-\mathbf{x}_{S}\|\leq\sqrt{k}\eta$, we get that
 $|g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})-g_{\mathbf{x}_{0}}^{(i)}(\mathbf{z})|\leq\frac{\eta\sqrt{k}}{\rho}\;,$
-for any $\mathbf{x}^{V},\mathbf{\mathbf{z}}^{V}\in S$.
-Consequently by the triangle inequality and Claim A.7 we have
-$\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{E}}_{\mathbf{x}_{0},\mathbf{x}^{V}}[\lvert(f_{i}(\mathbf{x}_{0}^{\perp V}+\mathbf{x}^{V})g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x}))\mathds{1}(\mathbf{x}\in S)\rvert]$
-$\leq\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{E}}_{\mathbf{x}_{0},\mathbf{x}^{V}}[\lvert(f_{i}(\mathbf{x}_{0}^{\perp V}+\mathbf{x}^{V})-g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x}_{S}))\mathds{1}(\mathbf{x}\in S)\rvert]+\eta/\rho\lesssim\sqrt{\rho}\Gamma(f_{i})+\eta\sqrt{k}/\rho\;,$
+for any $\mathbf{x},\mathbf{\mathbf{z}}\in S$. Consequently by the triangle inequality and Claim A.7, we have
+$\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{E}}_{\mathbf{x}_{0},\mathbf{x}^{V}}[\lvert(f_{i}(\mathbf{x}_{0}^{\perp V}+\mathbf{x}^{V})-g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x}))\mathds{1}(\mathbf{x}\in S)\rvert]$
+$\leq\sum_{S\in\mathcal{S}}\operatorname*{\mathbf{E}}_{\mathbf{x}_{0},\mathbf{x}^{V}}[\lvert(f_{i}(\mathbf{x}_{0}^{\perp V}+\mathbf{x}^{V})-g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x}_{S}))\mathds{1}(\mathbf{x}\in S)\rvert]+\eta\sqrt{k}/\rho\lesssim\sqrt{\rho}\Gamma(f_{i})+\eta\sqrt{k}/\rho\;,$
 for any collection of points $\mathbf{x}_{S}\in S$.
-Furthermore, observe that by the linearity of expectation and the fact that $\sum_{i\in\mathcal{Y}}f_{i}(\mathbf{x})=1$ we have that $\sum_{i\in\mathcal{Y}}g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})=1$ and $g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})\in[0,1]$ for all $\mathbf{x},\mathbf{x}_{0}\in\mathbb{R}^{d},i\in\mathcal{Y}$.
+Furthermore, by linearity of expectation and the fact that $\sum_{i\in\mathcal{Y}}f_{i}(\mathbf{x})=1$, we have that $\sum_{i\in\mathcal{Y}}g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})=1$ and $g_{\mathbf{x}_{0}}^{(i)}(\mathbf{x})\in[0,1]$ for all $\mathbf{x},\mathbf{x}_{0}\in\mathbb{R}^{d},i\in\mathcal{Y}$.
 Therefore, we can consider the multiclass function $G:\mathbb{R}^{d}\to\mathcal{Y}$ such that $G(\mathbf{x})$ is a random label sampled from the probability distribution induced by $g_{\mathbf{x}}^{(i)}(\mathbf{x}_{S})$, where $S\in\mathcal{S}$ such that $\mathbf{x}\in S$.
-Note that for every probability vector $\mathbf{p}\in\mathbb{R}^{|\mathcal{Y}|}$ and label $z\in\mathcal{Y}$ we have that
-$\|\mathbf{p}-\mathbf{e}_{z}\|_{1}=\lvert\mathbf{p}_{z}-1\rvert+\sum_{i\neq z,i\in\mathcal{Y}}\lvert\mathbf{p}_{i}\rvert=2\operatorname*{\mathbf{Pr}}_{i\sim\mathbf{p}}[i\neq z]\;.$
-Therefore, we have that the $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq G(\mathbf{x})]\lesssim\sqrt{\rho}\Gamma(f)+\eta\sqrt{k}\lvert\mathcal{Y}\rvert/\rho\lesssim\epsilon^{\prime}$.
-Notice that since $\mathbf{x}^{\prime}$ is distributed as $\mathcal{N}(\mathbf{0},\mathbf{I})$, we have that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq G(\mathbf{x}^{\prime})]=O(\epsilon^{\prime})$.
-However, as $G(\mathbf{x})$ is independent $\mathbf{x}^{V}$, we have that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq G(\mathbf{x}^{\prime})]=\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq G(\mathbf{x})]=O(\epsilon^{\prime})$.
-Therefore, by the triangle inequality we have that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f(\mathbf{x}^{\prime})]=O(\epsilon^{\prime})$, which completes the proof of Claim 3.6.∎
+Note that for every probability vector $\mathbf{p}\in\mathbb{R}^{|\mathcal{Y}|}$ and label $z\in\mathcal{Y}$ we have that $\|\mathbf{p}-\mathbf{e}_{z}\|_{1}=2\operatorname*{\mathbf{Pr}}_{i\sim\mathbf{p}}[i\neq z]$.
+Therefore, we have that the $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq G(\mathbf{x})]\lesssim\sqrt{\rho}\Gamma(f)+\eta\sqrt{k}\lvert\mathcal{Y}\rvert/\rho\lesssim\epsilon$.
+Notice that since $\mathbf{x}^{\prime}$ is distributed as $\mathcal{N}(\mathbf{0},\mathbf{I})$, we have that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq G(\mathbf{x}^{\prime})]=O(\epsilon)$.
+However, as $G(\mathbf{x})$ is independent $\mathbf{x}^{V}$, we have that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq G(\mathbf{x}^{\prime})]=\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq G(\mathbf{x})]=O(\epsilon)$.
+Therefore, by the triangle inequality we have that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq f(\mathbf{x}^{\prime})]=O(\epsilon)$, which completes the proof of Claim 3.6.∎
 Define $y^{\prime}$ such that or all $\mathbf{z}\in\mathbb{R}^{d}$, the distribution of $y^{\prime}$ given $\mathbf{x}^{\prime}=\mathbf{z}$ is the same as the distribution of $y$ given $\mathbf{x}=\mathbf{z}$.
-By the definition of $y^{\prime}$, it follows that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}^{\prime},y^{\prime}}[f(\mathbf{x}^{\prime})\neq y^{\prime}]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}^{\prime}}\left[\operatorname*{\mathbf{E}}_{y^{\prime}}[f(\mathbf{x}^{\prime})\neq y^{\prime}\mid\mathbf{x}^{\prime}]\right]\leq\epsilon$.
-Therefore, it holds that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y^{\prime}]\leq\epsilon+O(\epsilon^{\prime})$.
-Therefore, with probability at least $\delta$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ independent of $\mathbf{x}$, there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that
+By the definition of $y^{\prime}$, it follows that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}^{\prime},y^{\prime}}[f(\mathbf{x}^{\prime})\neq y^{\prime}]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}^{\prime}}\left[\operatorname*{\mathbf{E}}_{y^{\prime}}[f(\mathbf{x}^{\prime})\neq y^{\prime}\mid\mathbf{x}^{\prime}]\right]\leq\zeta$.
+Therefore, it holds that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y^{\prime}]\leq\zeta+O(\epsilon)$.
+Therefore, with probability at least $\alpha$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ independent of $\mathbf{x}$, there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ and a label $y\in\mathcal{Y}$ such that
 $\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>\sigma\|p(\mathbf{x}^{U})\|_{2}\;.$
 However, since the moments of $y^{\prime}$ are identical when conditioned on each $\mathbf{x}_{0}^{V}$ that belongs to the same region in $S$, we deduce that if the moment condition holds for some polynomial at a particular point $\mathbf{x}_{0}^{V}$, then it must hold for the entire cube $S$, such that $\mathbf{x}_{0}^{V}\in S$.
-Therefore, there exists a subset $\mathcal{T}\subseteq\mathcal{S}$ of the partition with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\delta$ satisfying the following: for all $S\in\mathcal{T}$ there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}\in S]>\sigma\|p(\mathbf{x}^{U})\|_{2}\;.$
+Therefore, there exists a subset $\mathcal{T}\subseteq\mathcal{S}$ of the partition with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\alpha$ satisfying the following: for all $S\in\mathcal{T}$ there exists a zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}\in S]>\sigma\|p(\mathbf{x}^{U})\|_{2}\;.$
 Noticing that the moments with respect to $U$ on each $S\in\mathcal{S}$ are the same for $y$ and $y^{\prime}$, completes the proof of Lemma 3.5. ∎
 The remaining part of the section is structured as follows: In Section 3.1, we provide the analysis of Algorithm 4 with Proposition 3.7.
-In Section 3.2, we analyze Algorithm 3 providing the full proof of Theorem 3.3, which is the main result of this section.
+In Section 3.2, we analyze Algorithm 3 providing the full proof of Theorem 3.3.
 3.1 Finding a Relevant Direction
-In this section, we prove Proposition 3.7, which states that if the class satisfies Definition 3.2 and the current approximation of the ground-truth subspace provides a suboptimal piecewise approximation, then Algorithm 4 outputs a small set of vectors, at least one of which has a non-trivial projection onto the ground-truth subspace.
+In this section, we prove Proposition 3.7, which states that if the target function class satisfies Definition 3.2 and the current approximation of the ground-truth subspace provides a suboptimal piecewise approximation, then Algorithm 4 outputs a small set of vectors, at least one of which has a non-trivial projection onto the ground-truth subspace.
 Proposition 3.7.
@@
 Let $f\in\mathcal{F}$ with $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$ and let $W$ be a subspace such that $f(\mathbf{x})=f(\mathbf{x}^{W})$ for all $\mathbf{x}\in\mathbb{R}^{d}$, with $K=\dim(W)$.
-Additionally, let $h_{\mathcal{S}}$ be a piecewise constant approximation of $D$, with respect to an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$, according $V$, as defined in Definitions 2.2 and 2.3.
-There exists $N=(dm)^{O(m)}{(k/\epsilon^{\prime})^{O(k)}}\log(1/\delta){(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{O(1)}}$ such that, If $\zeta\geq\mathrm{OPT}$ and $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>{\tau+\mathrm{OPT}+3\epsilon}$, then Algorithm 4 given $N$ i.i.d.
-samples from $D$, runs in $\mathrm{poly}(N,d,\lvert\mathcal{Y}\rvert)$ time, and with probability at least $1-\delta$, returns a list of unit vectors $\mathcal{E}$ of size $|\mathcal{E}|=\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))$, such that: For some $\mathbf{v}\in\mathcal{E}$ and unit vector $\mathbf{w}\in W$ it holds that
+Additionally, let $h_{\mathcal{S}}$ be a piecewise constant approximation of $D$, with respect to an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$, according to $V$, as defined in Definitions 2.2 and 2.3.
+There exists $N=(dm)^{O(m)}(k/\epsilon^{\prime})^{O(k)}\log(1/\delta)(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{O(1)}$ such that: If $\zeta\geq\mathrm{OPT}$ and $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+3\epsilon$, then Algorithm 4 given $N$ i.i.d.
+samples from $D$, runs in $\mathrm{poly}(N)$ time, and with probability at least $1-\delta$, returns a list of unit vectors $\mathcal{E}$ of size $|\mathcal{E}|=\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))$ such that: For some $\mathbf{v}\in\mathcal{E}$ and unit vector $\mathbf{w}\in W$ it holds that
 $\mathbf{w}^{\perp V}\cdot\mathbf{v}\geq\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))\;.$
-Before proving Proposition 3.7, we need to show that if we have learned a direction of our target model to accuracy $\epsilon$, then we only observe moments that depend on the remaining directions.
-The intuition behind the proof is that re-randomizing over directions that have negligible projection to $W$ barely affects the function’s output.
-Hence, the function is close to a mixing $y^{\prime}$ of the noisy version $y$ that has these components re-randomized.
-As a result, if the function is far from a projection onto $V$, we have large moments for level-sets of $y^{\prime}$.
-Consequently, the re-randomized directions are averaged out, and we are left with distinguishing moments on the remaining directions.
-The proof is similar in structure to that of Lemma 3.5, the main difference being in what is being re-randomized.
+Before proving Proposition 3.7, we show that once a direction of our target model is learned to sufficiently small accuracy, the observed moments depend only on the remaining directions.
+The intuition is as follows.
+Re-randomizing over directions that have negligible projection onto $W$ barely affects the output of the function.
+Thus, the function is close to a version $y^{\prime}$ of the noisy label $y$ in which these components have been averaged.
+Consequently, if the function is far from being a function of the current subspace $V$ (i.e., it still depends significantly on directions outside $V$), then the level sets of $y^{\prime}$ will exhibit large distinguishing moments.
+In this way, the influence of the re-randomized (negligible) directions is eliminated, leaving only the significant, distinguishing moments on the remaining directions.
+The proof follows a structure similar to that of Lemma 3.5, with the main difference being the set of directions over which re-randomization occurs.
 Lemma 3.8.
-Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, let $\mathcal{F}$ be a class of $(m,\epsilon+O(\epsilon^{\prime}),\delta,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models according to Definition 3.2.
-Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$, with $k\geq 1$, and let $\mathcal{S}$ be an $\eta$-approximating partition of $V$ according to Definition 2.2, with $\eta=\epsilon^{\prime 3}/(C\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2})$, for $C>0$ a sufficiently large universal constant.
-Let $f\in\mathcal{F}$ be a function that depends only on a $K$-dimensional subspace $W$, and let $y$ be random variable over $\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\epsilon$.
-Suppose that there exists no function $g:V\to\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$.
-Then, there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\delta$ such that for each $S\in\mathcal{T}$ there exists a zero-mean, variance-one polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x},y}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}\in S]>\sigma$, where $U{=}(V{+}W)\cap V^{\perp}$.
-Moreover, for each such polynomial $p$, we have that $\nabla p(\mathbf{x}^{U})\cdot\mathbf{v}=0$ for any $\mathbf{v}\in E$ and $\mathbf{x}\in\mathbb{R}^{d}$, where $E\subseteq{V^{\perp}}$ is a subspace satisfying $\|\mathbf{v}^{W}\|\leq\epsilon^{\prime}/\Gamma(f)$ for every unit vector $\mathbf{v}\in E$.
+Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$.
+Let $\mathcal{F}$ be a class of $(m,\zeta+O(\epsilon),\alpha,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models as defined in Definition 3.2.
+Let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$, with $k\geq 1$, and let $\mathcal{S}$ be an $\eta$-approximating partition of $V$ (see Definition 2.2), where $\eta=\epsilon^{3}/(C\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2})$ and $C>0$ is a sufficiently large universal constant.
+Let $f\in\mathcal{F}$ be a function that depends only on a $K$-dimensional subspace $W$, and let $y$ be random variable over $\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\zeta$.
+Suppose that no function $g:V\to\mathcal{Y}$ exists with $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\leq\tau$.
+Then, there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\alpha$ such that for each $S\in\mathcal{T}$ the following hold:
+  1. 1.
+    There exists a label $z\in\mathcal{Y}$ and a zero-mean, variance-one polynomial $p:U\to\mathbb{R}$ of degree at most $m$ such that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[p(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}\in S]>\sigma$, where $U{=}(V{+}W)\cap V^{\perp}$.
+  2. 2.
+    For each such polynomial $p$ satisfying (1), it holds that $\nabla p(\mathbf{x}^{U})\cdot\mathbf{v}=0$ for every $\mathbf{v}\in E$ and all $\mathbf{x}\in\mathbb{R}^{d}$, where $E\subseteq{V^{\perp}}$ is a subspace satisfying $\|\mathbf{v}^{W}\|\leq\epsilon/\Gamma(f)$ for every unit vector $\mathbf{v}\in E$.
 Proof.
@@
 Notice that for any unit direction $\mathbf{w}\in W$, both $\mathbf{w}\cdot\mathbf{x}$ and $\mathbf{w}\cdot\mathbf{x}^{\prime}$ are distributed according to a standard Gaussian.
-Furthermore, we have that by independence
+Furthermore, by independence we have that
 $\operatorname*{\mathbf{E}}[(\mathbf{w}\cdot\mathbf{x})(\mathbf{w}\cdot\mathbf{x}^{\prime})]$ $=\operatorname*{\mathbf{E}}[(\mathbf{w}\cdot\mathbf{x}^{\perp E})^{2}]=\|\mathbf{w}^{\perp E}\|^{2}\;.$
-Given that $E$ is the subspace of small projections, i.e., $\left\|\mathbf{w}^{E}\right\|\leq\epsilon^{\prime}/\Gamma(f)$; it follows that $\operatorname*{\mathbf{E}}[(\mathbf{w}\cdot\mathbf{x}^{\perp E})^{2}]\geq 1-(\epsilon^{\prime}/\Gamma(f))^{2}$.
-As a result, $\mathbf{x}$ and $\mathbf{x}^{\prime}$ are at least $1-(\epsilon^{\prime}/\Gamma(f))^{2}$ correlated in $W$.
-Thus by the Gaussian noise sensitivity bound A.5, we have that
-$\operatorname*{\mathbf{Pr}}_{\mathbf{x},\mathbf{x}^{\prime}}[f(\mathbf{x})\neq f(\mathbf{x}^{\prime})]\leq O(\epsilon^{\prime})\;.$
-Consider the random variable $y^{\prime}$ supported on $\mathcal{Y}$ that is distributed like $y$ for $\mathbf{x}^{\prime}$, i.e., $D(y^{\prime}=i\mid\mathbf{x}^{\prime}=\mathbf{z})=D(y=i\mid\mathbf{x}=\mathbf{z})$, for all $\mathbf{z}\in\mathbb{R}^{d},i\in\mathcal{Y}$.
+Given that $E$ is the subspace of small projections, (i.e., for any unit vector $\mathbf{w}\in W$, we have $\|\mathbf{w}^{E}\|\leq\epsilon/\Gamma(f)$), it follows that $\operatorname*{\mathbf{E}}[(\mathbf{w}\cdot\mathbf{x}^{\perp E})^{2}]\geq 1-(\epsilon/\Gamma(f))^{2}$ for any unit vector $\mathbf{w}\in W$.
+This means that $\mathbf{x}$ and $\mathbf{x}^{\prime}$ are at least $1-(\epsilon/\Gamma(f))^{2}$ correlated in $W$.
+By the Gaussian noise sensitivity bound (see A.5), we have that
+$\operatorname*{\mathbf{Pr}}_{\mathbf{x},\mathbf{x}^{\prime}}[f(\mathbf{x})\neq f(\mathbf{x}^{\prime})]\leq O(\epsilon)\;.$
+Next, define a new random variable $y^{\prime}$ over $\mathcal{Y}$ which is distributed like $y$ but with respect to $\mathbf{x}^{\prime}$; that is, for every $\mathbf{z}\in\mathbb{R}^{d}$ and $i\in\mathcal{Y}$, $D(y^{\prime}=i\mid\mathbf{x}^{\prime}=\mathbf{z})=D(y=i\mid\mathbf{x}=\mathbf{z})$.
 We have that
-$\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y^{\prime}]\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq y^{\prime}]+O(\epsilon^{\prime})\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]+O(\epsilon^{\prime})\leq\epsilon+O(\epsilon^{\prime})\;.$
-As a result, by Lemma 3.5, we have that there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\delta$ such that for each $S\in\mathcal{T}$ there exists a zero-mean, variance-one polynomial $p:U\to\mathbb{R}$ of degree at most $m$ along with a label $z\in\mathcal{Y}$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}\in S]>\sigma$, where $U=(V+W)\cap V^{\perp}$.
+$\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y^{\prime}]\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x}^{\prime})\neq y^{\prime}]+O(\epsilon)\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]+O(\epsilon)\leq\zeta+O(\epsilon)\;.$
+As a result, by Lemma 3.5, we have that there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\alpha$ such that for each $S\in\mathcal{T}$ there exists a zero-mean, variance-one polynomial $p:U\to\mathbb{R}$ of degree at most $m$ and a label $z\in\mathcal{Y}$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}\in S]>\sigma$, where $U=(V+W)\cap V^{\perp}$.
 However we have that for all $S\in\mathcal{T}$
-$\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}\in S]=\operatorname*{\mathbf{E}}[\operatorname*{\mathbf{E}}_{\mathbf{x}^{E}}[p(\mathbf{x}^{U})\mid\mathbf{x}^{\perp E}]\operatorname*{\mathbf{E}}_{\mathbf{x}^{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp E}]\mid\mathbf{x}\in S]]\;.$
-Notice that $p^{\prime}(\mathbf{x}^{U})\coloneqq\operatorname*{\mathbf{E}}_{\mathbf{x}^{E}}[p(\mathbf{x}^{U})\mid\mathbf{x}^{\perp E}]$ is a mean-zero polynomial of degree at most $m$ with variance at most one by Jensen’s inequality.
+$\operatorname*{\mathbf{E}}_{\mathbf{x},y^{\prime}}[p(\mathbf{x}^{U})\mathds{1}(y^{\prime}=z)\mid\mathbf{x}\in S]=\operatorname*{\mathbf{E}}\left[\operatorname*{\mathbf{E}}_{\mathbf{x}^{E}}[p(\mathbf{x}^{U})\mid\mathbf{x}^{\perp E}]\operatorname*{\mathbf{E}}_{\mathbf{x}^{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp E}]\mid\mathbf{x}\in S\right]\;,$
+where we used that $y^{\prime}$ is independent of $\mathbf{x}^{E}$.
+Notice that $p^{\prime}(\mathbf{x}^{U})\coloneqq\operatorname*{\mathbf{E}}_{\mathbf{x}^{E}}[p(\mathbf{x}^{U})\mid\mathbf{x}^{\perp E}]$ is a zero-mean polynomial of degree at most $m$ with variance at most one by Jensen’s inequality.
 Furthermore, as $p^{\prime}$ is independent of $\mathbf{x}^{E}$ we have that $\nabla p^{\prime}(\mathbf{x}^{U})\cdot\mathbf{v}=0$ for any $\mathbf{v}\in E$ and $\mathbf{x}\in\mathbb{R}^{d}$.
@@
 Proof of Proposition 3.7.
-Denote by $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}$ an orthonormal basis of the subspace $W$.
-Also, let $\widehat{\mathbf{U}}$ be the matrix computed at Line 6 of Algorithm 4 and let $\eta^{2}$ be the $l_{2}^{2}$ error chosen in the polynomial-regression step, i.e., Line 4 of Algorithm 4.
-First note that as $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>{\tau+\mathrm{OPT}+3\epsilon}$ we have that form Lemma 3.4 for every function $g:V\to\mathcal{Y}$ it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\geq\tau$.
-Thus from Lemma 3.8 we have that there exists $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\alpha$, such that for each $S\in\mathcal{T}$ there exists a $z\in\mathcal{Y}$ and a zero-mean variance-one polynomial $q_{S}:U\to\mathbb{R}$ of degree at most $m$ such that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[q_{S}(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}\in S]>\sigma$, where $U=(V+W)\cap(V^{\perp})$.
-Furthermore, define $E_{S}\coloneqq\mathrm{span}(\{\mathbf{v}\in U_{S}:\left\|\mathbf{v}^{W}\right\|\leq\epsilon/(\sqrt{\lvert U_{S}\rvert}\Gamma)\})$.
+Denote by $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}$ an orthonormal basis for the subspace $W$.
+Let $\widehat{\mathbf{U}}$ be the matrix computed in Line 6 of Algorithm 4 and let $\eta^{2}$ be the $L_{2}^{2}$ error chosen in the polynomial-regression step (see Line 4 of Algorithm 4).
+First, observe that since $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+3\epsilon$, Lemma 3.4 for every implies that for every function $g:V\to\mathcal{Y}$ it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[g(\mathbf{x}^{V})\neq f(\mathbf{x})]\geq\tau$.
+Then, by Lemma 3.8 there exists a $\mathcal{T}\subseteq\mathcal{S}$ with $\sum_{S\in\mathcal{T}}\operatorname*{\mathbf{Pr}}[S]\geq\alpha$, such that for each $S\in\mathcal{T}$ there exists a label $z\in\mathcal{Y}$ and a zero-mean variance-one polynomial $q_{S}:U\to\mathbb{R}$ of degree at most $m$ satisfying $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[q_{S}(\mathbf{x}^{U})\mathds{1}(y=z)\mid\mathbf{x}\in S]>\sigma$, where $U=(V+W)\cap(V^{\perp})$.
+Furthermore, for each $S\in\mathcal{S}$, define $E_{S}\coloneqq\mathrm{span}(\{\mathbf{v}\in U_{S}:\left\|\mathbf{v}^{W}\right\|\leq\epsilon/(\sqrt{\lvert U_{S}\rvert}\Gamma)\})$.
 Then, for every unit vector $\mathbf{v}\in E_{S}$, we have $\nabla q_{S}(\mathbf{x}^{U})\cdot\mathbf{v}=0$ since $\left\|\mathbf{v}^{W}\right\|\leq\epsilon/\Gamma$ holds for every unit vector $\mathbf{v}\in E_{S}$.
-In the following claim, we leverage the existence of non-trivial moments over $U$ for regions $S\in\mathcal{T}$, to show that if the sample size is sufficiently large then for each region $S\in\mathcal{T}$ the associated set, $U_{S}$ (see Line 5 of Algorithm 4) contains a vector that correlates with $W$.
+In the following claim, we leverage the existence of non-trivial moments over $U$ for regions $S\in\mathcal{T}$.
+We show that if the sample size is sufficiently large then for each region $S\in\mathcal{T}$ the associated set $U_{S}$ (see Line 5 of Algorithm 4) contains a vector that correlates with $W$.
 Claim 3.9 (Existence of Correlating Vectors).
-Fix, $S\in\mathcal{T}$, if the number of samples that fall in $S$ is $N_{S}\geq(dm)^{Cm}\log(\lvert\mathcal{Y}\rvert/\delta)/{\eta}^{C}$, for a sufficiently large universal constant $C>0$, then with probability at least $1-\delta$ there exists a unit vector $\mathbf{v}\in U_{S}$ such that
-$(\mathbf{w}^{(i)})^{\perp V}\cdot\mathbf{v}\geq{\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)}\;,$
-for some $i\in[K]$. Moreover, it holds that $\lvert U_{S}\rvert=\mathrm{poly}(mK\lvert\mathcal{Y}\rvert/\sigma)$.
+Let $C>0$ be a sufficiently large universal constant.
+Fix $S\in\mathcal{T}$.
+If the number of samples that fall in $S$ is $N_{S}\geq(dm)^{Cm}\log(\lvert\mathcal{Y}\rvert/\delta)/\eta^{C}$, then with probability at least $1-\delta$ there exists a unit vector $\mathbf{v}\in U_{S}$ and $i\in[K]$ such that
+$(\mathbf{w}^{(i)})^{\perp V}\cdot\mathbf{v}\geq\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)\;.$
+Moreover, it holds that $\lvert U_{S}\rvert=\mathrm{poly}(mK\lvert\mathcal{Y}\rvert/\sigma)$.
 Proof of Claim 3.9.
-Denote by $p_{S,i}$ for $S\in\mathcal{S},i\in\mathcal{Y}$ the polynomials of degree at most $m$ computed at Line 4 of Algorithm 4.
-First, notice that since whether or not $\mathbf{x}\in S$ depends only in $\mathbf{x}^{V}$, we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}^{\perp V}))^{2}\mid\mathbf{x}\in S]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}^{\perp V}))^{2}\mid\mathbf{x}\in S]+\eta^{2}$ is equivalent to $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}))^{2}]+\eta^{2}$, where $D_{V^{\perp}}$ is defined to be the marginal obtained by averaging $D$ over $V$ conditioned on $S$, i.e., $D_{V^{\perp}}(\mathbf{x}^{\perp V},y)=\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[D(\mathbf{x},y)\mid\mathbf{x}\in S]$.
+Denote by $p_{S,i}$ (for $S\in\mathcal{S}$ and $i\in\mathcal{Y}$) the degree at most $m$ polynomial computed at Line 4 of Algorithm 4.
+Notice that membership of a point $\mathbf{x}$ in a cube $S$ depends solely on its projection onto $V$, i.e., on $\mathbf{x}^{V}$.
+Therefore, the error guarantee obtained in the regression step $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}^{\perp V}))^{2}\mid\mathbf{x}\in S]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}^{\perp V}))^{2}\mid\mathbf{x}\in S]+\eta^{2}$ is equivalent to $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}))^{2}]+\eta^{2}$, where $D_{V^{\perp}}^{S}$ is the marginal obtained by averaging $D$ over $V$ conditioned on $S$, i.e., $D^{S}_{V^{\perp}}(\mathbf{x}^{\perp V},y)=\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[D(\mathbf{x},y)\mid\mathbf{x}\in S]$.
 Moreover, by the properties of the Gaussian distribution, we have that the $\mathbf{x}$-marginal of $D_{V^{\perp}}$ is a standard Gaussian.
-Therefore, if $N_{S}\geq(dm)^{Cm}\log(\lvert\mathcal{Y}\rvert/\delta)/{\eta^{C}}$ for a sufficiently large universal constant $C>0$, then by the union bound and B.2, we have that with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}))^{2}]+\eta^{2}$.
-Moreover, by the orthogonality of Hermite polynomials we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]=\sum_{\beta\subseteq\mathbb{N}^{d}}(\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x})\mathds{1}(y=i)]-\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x})p_{S,i}(\mathbf{x})])^{2}$, it holds that $\sum_{t=1}^{m}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)^{[t]}-p_{S,i}^{[t]}(\mathbf{x}))^{2}]\leq\eta^{2}$.
-From the rotational invariance of the Gaussian without loss of generality, we can denote by $\mathbf{e}_{1},\dots,\mathbf{e}_{k^{\prime}}$ an orthonormal basis of the space of $U$ projected onto $E_{S}^{\perp}$.
-We have that $k^{\prime}\geq 1$ otherwise we would not have the existence of $q_{S}$ that is defined over this space.
-Furthermore, since $U$ is a subset of $V^{\perp}$ we similarly have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[q_{S}(\mathbf{x}^{U})\mathds{1}(y=z)]>\sigma$ for some $z\in\mathcal{Y}$.
-Decomposing $q_{S}(\mathbf{x}^{U})$ to the basis of Hermite polynomials we have that $q_{S}(\mathbf{x}^{U})=\sum_{\beta\in\mathbb{N}^{d^{\prime}},1\leq\left\|\beta\right\|_{1}\leq m}\hat{q}_{S}(\beta)H_{\beta}(\mathbf{x}^{U})$ where $d^{\prime}=\dim(U)$.
-Moreover, note that since $q_{S}$ has no component outside $\mathbf{e}_{1},\dots,\mathbf{e}_{k^{\prime}}$ we have that $q_{S}(\mathbf{x}^{U})=\sum_{\beta\in J}\hat{q}_{S}(\beta)H_{\beta}(\mathbf{x}^{U})$, where $J$ denotes the set of $\beta\in\mathbb{N}^{d^{\prime}}$ such that $\beta_{i}\geq 1$ for some $i\in[k^{\prime}]$.
-Considering the correlation of $q_{S}$ with the label region $z$ we have that
-$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[q_{S}(\mathbf{x}^{U})\mathds{1}(y=z)]$ $=\sum_{\beta\in J}\hat{q}_{S}(\beta)\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x}^{U})\mathds{1}(y=z)]$
-$\leq\left(\sum_{\beta\in J}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x}^{U})\mathds{1}(y=z)]^{2}\right)^{1/2}\;,$
+Therefore, if $N_{S}\geq(dm)^{Cm}\log(\lvert\mathcal{Y}\rvert/\delta)/\eta^{C}$ for a sufficiently large universal constant $C>0$, then by applying the union bound and B.2, we have that with probability at least $1-\delta$:
+$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}))^{2}]+\eta^{2}\;.$
+Furthermore, by the orthogonality of Hermite polynomials, we have that
+$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]=\sum_{\beta\subseteq\mathbb{N}^{d}}(\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[H_{\beta}(\mathbf{x})\mathds{1}(y=i)]-\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[H_{\beta}(\mathbf{x})p_{S,i}(\mathbf{x})])^{2}\;.$
+In particular, if we decompose the error into its Hermite polynomial components of degree $t$, we have that $\sum_{t=1}^{m}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[(\mathds{1}(y=i)^{[t]}-p_{S,i}^{[t]}(\mathbf{x}))^{2}]\leq\eta^{2}$.
+Next, by the rotational invariance of the Gaussian distribution, we may, without loss of generality, choose an orthonormal basis $\mathbf{e}_{1},\dots,\mathbf{e}_{k^{\prime}}$ for the subspace $U$ projected onto $E_{S}^{\perp}$.
+Note that $k^{\prime}\geq 1$, otherwise, the polynomial $q_{S}$ (which is defined over this space) could not exist.
+Furthermore, since $U$ is a subset of $V^{\perp}$, we similarly have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[q_{S}(\mathbf{x}^{U})\mathds{1}(y=z)]>\sigma$ for some $z\in\mathcal{Y}$.
+Decomposing $q_{S}(\mathbf{x}^{U})$ in the basis of Hermite polynomials we have that $q_{S}(\mathbf{x}^{U})=\sum_{\beta\in\mathbb{N}^{d^{\prime}},1\leq\left\|\beta\right\|_{1}\leq m}\widehat{q}_{S}(\beta)H_{\beta}(\mathbf{x}^{U})$, where $d^{\prime}=\dim(U)$.
+Moreover, note that since $q_{S}$ has no component outside $\mathbf{e}_{1},\dots,\mathbf{e}_{k^{\prime}}$, we have that $q_{S}(\mathbf{x}^{U})=\sum_{\beta\in J}\widehat{q}_{S}(\beta)H_{\beta}(\mathbf{x}^{U})$, where $J$ denotes the set of $\beta\in\mathbb{N}^{d^{\prime}}$ such that $\beta_{i}\geq 1$ for some $i\in[k^{\prime}]$.
+Considering the correlation of $q_{S}$ with the region of the label $z$, we have that
+$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[q_{S}(\mathbf{x}^{U})\mathds{1}(y=z)]$ $=\sum_{\beta\in J}\widehat{q}_{S}(\beta)\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[H_{\beta}(\mathbf{x}^{U})\mathds{1}(y=z)]$
+$\leq\left(\sum_{\beta\in J}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[H_{\beta}(\mathbf{x}^{U})\mathds{1}(y=z)]^{2}\right)^{1/2}\;,$
 where we used the fact that $\|q_{S}\|_{2}=1$ and the Cauchy-Schwarz inequality.
-Hence, we have that the sum of the squares of the Hermite coefficients for of the degree $m$ restriction the random variable $\mathds{1}(y=z)$ is at least $\sigma^{2}$.
+Hence, we deduce that the sum of the squares of the Hermite coefficients corresponding to degrees up to $m$ in the expansion of the random variable $\mathds{1}(y=z)$ is at least $\sigma^{2}$.
 Now evaluating the quadratic form of $\mathbf{M}=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim D_{\mathbf{x}}}[\nabla p_{S,z}(\mathbf{x}^{\perp V})\nabla p_{S,z}(\mathbf{x}^{\perp V})^{\top}\mid\mathbf{x}\in S]$ (Line 5 of Algorithm 4) using A.2 gives us
-$\sum_{i=1}^{k^{\prime}}\mathbf{e}_{i}^{\top}\mathbf{M}\mathbf{e}_{i}$ $=\sum_{i=1}^{k^{\prime}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\nabla p_{S,z}(\mathbf{x})\cdot\mathbf{e}_{i})^{2}]\geq\sum_{\beta\in\mathbb{N}^{d}}(\sum_{i=1}^{k^{\prime}}\beta_{i})(\hat{p}_{S,z}(\beta))^{2}$
-$\geq\sum_{\beta\in J}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x}^{U})\mathds{1}(y=z)]^{2}-2\eta$
+$\sum_{i=1}^{k^{\prime}}\mathbf{e}_{i}^{\top}\mathbf{M}\mathbf{e}_{i}$ $=\sum_{i=1}^{k^{\prime}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[(\nabla p_{S,z}(\mathbf{x})\cdot\mathbf{e}_{i})^{2}]\geq\sum_{\beta\in\mathbb{N}^{d}}\left(\sum_{i=1}^{k^{\prime}}\beta_{i})(\widehat{p}_{S,z}(\beta)\right)^{2}$
+$\geq\sum_{\beta\in J}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D^{S}_{V^{\perp}}}[H_{\beta}(\mathbf{x}^{U})\mathds{1}(y=z)]^{2}-2\eta$
 $\geq\sigma^{2}-2\eta\geq\sigma^{2}/2$
@@
 Additionally, let $U_{S,z}$ denote the set of eigenvectors of $\mathbf{M}$ with eigenvalues greater than ${\sigma^{2}}/{4K}$ (see Line 5 of Algorithm 4).
-Decomposing $\mathbf{e}_{j}$ to the basis of eigenvectors gives us
-$\frac{\sigma^{2}}{4K}+\sum_{\mathbf{u}^{(i)}\in U_{S,z}}\lambda_{i}(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\sum_{i=1}^{d-k}\lambda_{i}(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\frac{\sigma^{2}}{2K}\;.$
-Hence, there exists an eigenvector $\mathbf{u}^{(i)}\in U_{S,z}$ such that $\lambda_{i}(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\sigma^{2}/(4K\lvert U_{S,z}\rvert)$ thus $(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\sigma^{2}/(4\lvert U_{S,z}\rvert K\text{tr}(\mathbf{M}))$ as $\mathbf{M}$ is positive definite.
-Therefore, by B.3 we have that $\lvert U_{S,z}\rvert=\mathrm{poly}(mK/\sigma),\text{tr}(\mathbf{M})=O(m)$ which implies that $\lvert\mathbf{e}_{j}\cdot\mathbf{u}^{(i)}\rvert\geq\mathrm{poly}(\sigma/(mK))$.
+Decomposing $\mathbf{e}_{j}$ in the basis of eigenvectors gives us
+$\mathbf{e}_{j}^{\top}\mathbf{M}\,\mathbf{e}_{j}=\sum^{d-k}_{i=1}\lambda_{i}\,(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}.$
+Using that $U_{S,z}$ contains eigenvectors of $\mathbf{M}$ with eigenvalues greater than ${\sigma^{2}}/{4K}$, we get that
+$\frac{\sigma^{2}}{4K}+\sum_{\mathbf{u}^{(i)}\in U_{S,z}}\lambda_{i}(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\mathbf{e}_{j}^{\top}\mathbf{M}\,\mathbf{e}_{j}\geq\frac{\sigma^{2}}{2K}\;.$
+Hence, there exists an eigenvector $\mathbf{u}^{(i)}\in U_{S,z}$ such that $\lambda_{i}(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\sigma^{2}/(4K\lvert U_{S,z}\rvert)$.
+This implies that $(\mathbf{e}_{j}\cdot\mathbf{u}^{(i)})^{2}\geq\sigma^{2}/(4\lvert U_{S,z}\rvert K\text{tr}(\mathbf{M}))$ as $\mathbf{M}$ is positive definite.
+Therefore, by B.3, we have that $\lvert U_{S,z}\rvert=\mathrm{poly}(mK/\sigma),\text{tr}(\mathbf{M})=O(m)$ which implies that $\lvert\mathbf{e}_{j}\cdot\mathbf{u}^{(i)}\rvert\geq\mathrm{poly}(\sigma/(mK))$.
 Finally, note that since $\mathbf{e}_{j}\cdot\mathbf{u}^{(i)}\neq 0$, we have that $\mathbf{u}^{(i)}$ cannot belong in the space $E_{S}$.
@@
 ∎
-First note that by the definition of an approximating partition (see Definition 2.2) and A.1, for all $S\in\mathcal{S}$ we have that $\operatorname*{\mathbf{Pr}}_{D}[S]{=(\epsilon^{\prime}/k)^{\Omega(k)}}$ and $\lvert\mathcal{S}\rvert{=(k/\epsilon^{\prime})^{O(k)}}$.
-Therefore, by union bound and Hoeffding’s inequality, it holds that if $N\geq{(k/\epsilon^{\prime})^{Ck}}\log(\lvert\mathcal{S}\rvert/\delta)={(k/\epsilon^{\prime})^{O(k)}}\log(1/\delta)$ for a sufficiently large constant $C>0$, then with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]-\operatorname*{\mathbf{Pr}}_{D}[S]\rvert\leq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S}$.
-Hence, it is true that $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$, i.e., the number of samples that fall in each set $S\in\mathcal{S}$ is at least $N\operatorname*{\mathbf{Pr}}_{D}[S]/2=N{(\epsilon^{\prime}/k)^{\Omega(k)}}$.
-Therefore, for $N\geq(dm)^{Cm}{(k/\epsilon^{\prime})^{Ck}}\log(1/\delta){(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{C}}$ for a sufficiently large universal constant $C>0$, by applying Claim 3.9, we have that for all $S\in\mathcal{T}$ it holds that $\mathbf{v}\cdot(\mathbf{w}^{(i)})^{\perp V}\geq{\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)}$ for some $\mathbf{v}\in U_{S},i\in[K]$.
+By the definition of an approximating partition (see Definition 2.2) and A.1, for all $S\in\mathcal{S}$ we have that $\operatorname*{\mathbf{Pr}}_{D}[S]=(\epsilon^{\prime}/k)^{\Omega(k)}$ and $\lvert\mathcal{S}\rvert=(k/\epsilon^{\prime})^{O(k)}$.
+Therefore, by the union bound and Hoeffding’s inequality, it holds that if the sample size $N\geq(k/\epsilon^{\prime})^{Ck}\log(\lvert\mathcal{S}\rvert/\delta)=(k/\epsilon^{\prime})^{O(k)}\log(1/\delta)$ for a sufficiently large constant $C>0$, then with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]-\operatorname*{\mathbf{Pr}}_{D}[S]\rvert\leq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S}$.
+Hence, $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$, i.e., the number of samples that fall in each set $S\in\mathcal{S}$ is at least $N\operatorname*{\mathbf{Pr}}_{D}[S]/2=N(\epsilon^{\prime}/k)^{\Omega(k)}$.
+If we choose $N$ so that $N\geq(dm)^{Cm}(k/\epsilon^{\prime})^{Ck}\log(1/\delta)(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{C}$ for a sufficiently large universal constant $C>0$, by Claim 3.9, we have that for all $S\in\mathcal{T}$ it holds that $\mathbf{v}\cdot(\mathbf{w}^{(i)})^{\perp V}\geq\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)$ for some $\mathbf{v}\in U_{S},i\in[K]$.
 Hence, we have that there exists a subset $\mathcal{T}^{\prime}\subseteq\mathcal{T}$ and $i\in[K]$ with $\sum_{S\in\mathcal{T^{\prime}}}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\sum_{S\in\mathcal{T^{\prime}}}\operatorname*{\mathbf{Pr}}_{D}[S]/2=\Omega(\alpha/K)$ such that for all $S\in\mathcal{T^{\prime}}$ we have that $\mathbf{v}\cdot(\mathbf{w}^{(i)})^{\perp V}\geq\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)$ for some $\mathbf{v}\in U_{S}$.
-Denote by $\widehat{\mathbf{U}}$ the matrix computed at Line 6 of Algorithm 4. For some $i\in[K]$ we have that
+Denote by $\widehat{\mathbf{U}}$ the matrix computed in Line 6 of Algorithm 4. For some $i\in[K]$, we have that
 $\left((\mathbf{w}^{(i)})^{\perp V}\right)^{\top}\widehat{\mathbf{U}}(\mathbf{w}^{(i)})^{\perp V}\gtrsim\frac{\alpha}{K}\mathrm{poly}(\epsilon\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)\geq\mathrm{poly}(\epsilon\alpha\sigma/(mK\lvert\mathcal{Y}\rvert\Gamma)\;,$
-where we used the fact that $\widehat{\mathbf{U}}$ is psd.
+where we used the fact that $\widehat{\mathbf{U}}$ is a PSD matrix.
 Moreover, as $\left\|\mathbf{v}\right\|=1$ for all $\mathbf{v}\in U_{S},S\in\mathcal{S}$, by the triangle inequality we can see that
@@
 3.2 Proof of Theorem 3.3
-In this section, given Proposition 3.7, we proceed to the proof of Theorem 3.3.
+In this section, given Proposition 3.7, we prove Theorem 3.3.
 Recall that, in Proposition 3.7, we have shown that if our current approximation to the hidden subspace is not accurate enough to produce a classifier that has sufficiently small error, then Algorithm 4 efficiently finds a direction that has non-trivial correlation to the hidden subspace.
@@
 Let $W^{*}$ be a subspace of dimension $K$ such that $f^{*}(\mathbf{x})=f^{*}(\mathbf{x}^{W^{*}})$ and denote by $\mathbf{w}^{*(1)},\ldots,\mathbf{w}^{*(K)}\in\mathbb{R}^{d}$ an orthonormal basis of $W^{*}$.
-Let $L_{t}$ be the list of vectors updated by the algorithm (Line 3c of Algorithm 3) and $V_{t}=\mathrm{span}(L_{t})$, $\dim(V_{t})=k_{t}$.
-Also, let $\epsilon^{\prime}$ be the partition width parameter set at Line 2 of Algorithm 3, and for $t\in[T]$ let $\mathcal{S}_{t}$ be arbitrary $\epsilon^{\prime}$-approximating partitions with respect to $V_{t}$ (see Definition 2.2).
-Let $h_{t}:\mathbb{R}^{d}\to[K]$ be a piecewise constant classifiers, defined as $h_{t}=h_{\mathcal{S}_{t}}$ according to Definition 2.3 for the distribution $D$.
+Let $L_{t}$ be the list of vectors maintained by the algorithm (Line 3c of Algorithm 3) and $V_{t}=\mathrm{span}(L_{t})$, $\dim(V_{t})=k_{t}$.
+Let $\epsilon^{\prime}$ be the partition width parameter set at Line 2 of Algorithm 3.
+For $t\in[T]$, let $\mathcal{S}_{t}$ be arbitrary $\epsilon^{\prime}$-approximating partitions with respect to $V_{t}$ (see Definition 2.2).
+Let $h_{t}:\mathbb{R}^{d}\to[K]$ be piecewise constant classifiers, defined as $h_{t}=h_{\mathcal{S}_{t}}$ according to Definition 2.3 for the distribution $D$.
 To prove the correctness of Algorithm 3, we need to show that if $h_{t}$ has significant error, then the algorithm improves its approximation $V_{t}$ to $W^{*}$.
@@
 Note that from Lines 1 and 3 of Algorithm 3, we perform at most $\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))$ iterations.
-Furthermore, in each iteration, we update the vector set with at most $\mathrm{poly}(mK\Gamma{\lvert\mathcal{Y}\rvert}/(\epsilon\alpha\sigma))$ vectors.
+Furthermore, in each iteration, we update the vector set with at most $\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))$ vectors (see Proposition 2.8).
 Hence, it follows that $k_{t}\leq\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))$, for all $t=1,\dots,T$.
 Assume that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+\epsilon/2$ for all $t=1,\dots,T$.
-Using the fact that $N={d}^{Cm}2^{(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))^{C}}\log(1/\delta)$ for a sufficiently large universal constant $C>0$ (Line 2 of Algorithm 3), we can apply Proposition 3.7 and conclude that, with probability $1-\delta$, there exists unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ for $t=[T]$ such that $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))$.
-Thus, by Claim 2.16, we have that with probability $1-\delta$, for all $t\in[T]$, $\Phi_{t}\leq\Phi_{t-1}-\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))$.
+Using the fact that $N={d}^{Cm}2^{(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))^{C}}\log(1/\delta)$ for a sufficiently large universal constant $C>0$ (Line 2 of Algorithm 3), we can apply Proposition 3.7 and conclude that, with probability $1-\delta$, there exist unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ for $t=[T]$ such that $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))$.
+Thus, by Claim 2.16, we have that with probability $1-\delta$, for all $t\in[T]$,
+$\Phi_{t}\leq\Phi_{t-1}-\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))\;.$
 After $T$ iterations, it follows that $\Phi_{T}\leq\Phi_{0}-T\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))=K-T\mathrm{poly}(\epsilon\alpha\sigma/(mK\Gamma\lvert\mathcal{Y}\rvert))$.
@@
 Since the error of $h_{t}$ can only be decreasing by Claim 2.17 and $h_{t}$ is close to its sample variant by Claim 2.18, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$.
-Sample and Computational Complexity: From the analysis above we have that the algorithm terminates in $\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))$ iterations and at each iteration we draw of the order of ${d}^{O(m)}2^{\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))}\log(1/\delta)$ samples.
+Sample and Computational Complexity: From the analysis above we have that the algorithm terminates in $\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))$ iterations and at each iteration we draw of order ${d}^{O(m)}2^{\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))}\log(1/\delta)$ samples.
 Hence, we have that the total number of samples is ${d}^{O(m)}2^{\mathrm{poly}(mK\Gamma\lvert\mathcal{Y}\rvert/(\epsilon\alpha\sigma))}\log(1/\delta)$.
-Moreover, we use at most $\mathrm{poly}(N,d,\lvert\mathcal{Y}\rvert)$ time as all operations can be implemented in polynomial time.
+Moreover, we use at most $\mathrm{poly}(N)$ time as all operations can be implemented in sample polynomial time.
 ∎
 4 Algorithmic Applications of General MIM Learner
-In this section, we present applications of our main result.
+In this section, we highlight selected applications of our main result.
 First, in Section 4.1 we present an application of our algorithmic approach to the class of intersections of halfspaces, establishing Theorem 1.13.
-Second, in Section 4.2, we show that a slight modification of our algorithm can achieve improved performance in the RCN setting, establishing Theorem 1.8.
-Additionally, note that our result on learning Multiclass Linear Classifiers, as presented in Theorem 2.1, can also be derived as a corollary of Theorem 3.3, using similar arguments.
+Second, in Section 4.2, we show that a slight modification of our general algorithm can achieve improved complexity in the RCN setting, establishing Theorem 1.11.
+Additionally, we note that our result on learning Multiclass Linear Classifiers, as presented in Theorem 2.1, can also be derived as a corollary of Theorem 3.3, using similar arguments.
 4.1 Learning Intersections of Halfspaces with Adversarial Label Noise
@@
 Then, Algorithm 3 draws $N=d^{2}2^{\mathrm{poly}(K/\epsilon)}\log(1/\delta)$ i.i.d.
-samples from $D$, runs in time $\mathrm{poly}(d,N)$, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq K\,O(\mathrm{OPT}\log(1/\mathrm{OPT}))+\epsilon$, where $\mathrm{OPT}=\inf_{f\in\mathcal{F}_{K,d}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$.
-For this class of Multi-Index Models, we adapt the general algorithm Algorithm 4 to perform polynomial regression with polynomials of degree at most $2$ in Line 4 of Algorithm 4.
+samples from $D$, runs in time $\mathrm{poly}(N)$, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq K\,O(\mathrm{OPT}\log(1/\mathrm{OPT}))+\epsilon$, where $\mathrm{OPT}=\inf_{f\in\mathcal{F}_{K,d}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$.
+Proof.
+First, note that without loss of generality we can assume that $\mathrm{OPT}\geq\epsilon^{2}$ by randomly flipping an $\epsilon^{2}$ fraction of the labels with probability $1/2$.
+Therefore, by the mean value theorem we have that
+$(\mathrm{OPT}+\epsilon^{2})\log(1/(\mathrm{OPT}+\epsilon^{2})+1)$ $\leq(\mathrm{OPT})\log(1/(\mathrm{OPT})+1)+\epsilon^{2}\log(1/\mathrm{OPT}+1)$
+$\leq(\mathrm{OPT})\log(1/(\mathrm{OPT})+1)+\epsilon\;.$
+Hence, by applying Proposition 4.2 we have that the class of intersections of $K$-halfspaces is a class of $(2,\mathrm{OPT}+\epsilon,\epsilon,K,KO(\mathrm{OPT}\log(1/\mathrm{OPT}+1)+\epsilon),\mathrm{poly}(\epsilon/K),O(\sqrt{\log K}))$-well-behaved Multi-Index Models.
+Thus, by Theorem 3.3 we have that Algorithm 3 draws $N=d^{2}2^{\mathrm{poly}(K/\epsilon)}\log(1/\delta)$ and in $\mathrm{poly}(N)$ time returns a hypothesis $h$ with error $KO(\mathrm{OPT}\log(1/\mathrm{OPT}+1))+O(\epsilon)$.
+Finally, note that without loss of generality $\mathrm{OPT}\leq 1/2$, since one of the constant classifiers that predict 00 and $1$ can achieve error at most $1/2$.
+Consequently, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq K\,O(\mathrm{OPT}\log(1/\mathrm{OPT}))+O(\epsilon)$, since $x\log(1/x+1)=\Theta(x\log(1/x))$ for all $x\in(0,1/2]$.
+∎
+For the class of intersections of $K$ halfspaces, we adapt the general algorithm Algorithm 4 to perform polynomial regression with polynomials of degree at most $2$ in Line 4 of Algorithm 4.
 This degree suffices for learning intersections of halfspaces, as demonstrated in Proposition 4.2.
-This task can be efficiently reduced to covariance estimation with respect to the Frobenius norm, as the entries of the covariance matrix correspond to the Hermite coefficients.
-Therefore, we have that it can be performed in polynomial time using $O(d^{2})$ samples.
-Consequently, the final sample complexity of the algorithm scales with $d^{2}$ rather than $d^{O(1)}$.
-The following proposition shows that the class of intersections of halfspaces satisfies Definition 3.2.
-To prove this result, we leverage that if a function is far from any function defined over a given subspace, then it must also be far from any constant function on a non-trivial fraction of the cylinders restricted to that subspace.
+Moreover, the regression task can be efficiently reduced to covariance estimation with respect to the Frobenius norm because the entries of the covariance matrix correspond directly to the Hermite coefficients.
+Consequently, the regression can be executed in polynomial time using $O(d^{2})$ samples, so that the overall sample complexity of the algorithm scales as $d^{2}$ rather than $d^{O(1)}$.
+The following proposition shows that the class of intersections of halfspaces satisfies the regularity conditions in Definition 3.2.
 Proposition 4.2.
 Let $d,K\in\mathbb{Z}_{+},\zeta,\epsilon\in(0,1)$.
-$\mathcal{F}_{K,d}$ is a class of $(2,\zeta,\epsilon,K,KO(\zeta\log(1/\zeta))+O(\epsilon),\mathrm{poly}(\epsilon/K),O(\sqrt{\log K}))$-well-behaved Multi-Index Models according to Definition 3.2.
+$\mathcal{F}_{K,d}$ is a class of $(2,\zeta,\epsilon,K,KO(\zeta\log(1/\zeta+1)+\epsilon),\mathrm{poly}(\epsilon/K),O(\sqrt{\log K}))$-well-behaved Multi-Index Models according to Definition 3.2.
 Proof.
@@
 It remains to show that Condition 3 of Definition 3.2 is satisfied for the class $\mathcal{F}_{K,d}$.
-Let $D$ be a distribution over $\mathbb{R}^{d}\times\{0,1\}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\zeta$.
-Let $V$ be a subspace of $\mathbb{R}^{d}$, and let $\tau(x)=x\log(1/x)$.
-Assume that for all functions $g:V\to\{0,1\}$ it holds that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq g(\mathbf{x}^{V})]\geq CK\tau(\zeta)+2\epsilon$, for a sufficiently large constant $C$.
-Hence, the same holds true for the function $h$ defined as $h(\mathbf{x})=\operatorname*{argmin}_{i\in\{\pm 1\}}\mathbf{Pr}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}]$, since by the definition $h(\mathbf{x})=h(\mathbf{x}^{V})$.
-Thus, $\operatorname*{\mathbf{Pr}}[h(\mathbf{x})\neq y]\geq(C-1)K\tau(\zeta)+2\epsilon$.
-We first show that if the current subspace $V$ is far from optimal, then by splitting the space into cylinders with respect to each point of $V$, the target function is far from being constant on a nontrivial fraction of these cylinders.
+Let $D$ be a distribution over $\mathbb{R}^{d}\times\{0,1\}$ ,where the $\mathbf{x}$-marginals of $D$ are standard normal, such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]\leq\zeta$.
+Let $V$ be a subspace of $\mathbb{R}^{d}$, and define $\tau(x)=x\log(1/x+1)$.
+Assume that for every function $g:V\to\{0,1\}$ it holds that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq g(\mathbf{x}^{V})]>CK\tau(\zeta)+2\epsilon$, for a sufficiently large constant $C>0$.
+In particular, consider the function $h(\mathbf{x})=\operatorname*{argmin}_{i\in\{0,1\}}\mathbf{Pr}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}]$ which depends only on $\mathbf{x}^{V}$ (i.e., $h(\mathbf{x})=h(\mathbf{x}^{V})$).
+Then the same lower bound applies to $h$, hence $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]>(C-4)K\tau(\zeta)+2\epsilon$, since $\tau(x)\geq x/4$ for all $x\in(0,1].$
+We first show that if the subspace $V$ is far from optimal, then when we partition the space into cylinders determined by points in $V$, the target function is far from constant on a nontrivial fraction of these cylinders.
 Specifically, we show the following
 Claim 4.3 (Fraction of Good Points).
-With probability at least $\epsilon$ over $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, independent of $\mathbf{x}$, it holds that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>(C-2)K\tau\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\right)+\epsilon$ for all $i\in\{\pm 1\}$ and $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\leq 1/(C-2)$.
+With probability at least $\epsilon$ over a random draw of $\mathbf{x}_{0}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ (independent of $\mathbf{x}$), the following holds: for every $i\in\{0,1\}$,
+$\mathbf{Pr}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>(C-4)K\,\tau\left(\mathbf{Pr}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\right)+\epsilon\;,$
+and $\mathbf{Pr}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\leq 4/(C-4)$.
 Proof of Claim 4.3.
-For the sake of contradiction, assume that the set $S$ of $\mathbf{x}_{0}\in\mathbb{R}^{d}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>(C-1)K\tau\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\right)+\epsilon$ for all $i\in\{\pm 1\}$ has probability mass less than $\epsilon$.
-Then, we have that
-$\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]$ $=\operatorname*{\mathbf{E}}_{\mathbf{x}_{0}^{V}}\left[\min_{\hat{y}\in\{\pm 1\}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}\left[\hat{y}\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\right]\right]$ (3)
-$\leq(C-1)K\operatorname*{\mathbf{E}}_{\mathbf{x}_{0}^{V}}\left[\tau\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}\left[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\right]\right)\right]+2\epsilon$
-$\leq(C-1)K\tau(\zeta)+2\epsilon\;,$
-where in the last inequality we used Jensen’s inequality since $\tau$ is concave.
-This leads to a contradiction, and therefore we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{x}\in S]>\epsilon$.
-Observe that for each such $\mathbf{x}_{0}$ as $(C-1)K\tau(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}])+\epsilon\leq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\leq 1$, we have that, for a sufficiently large choice of $C$, the probability $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]$ is less than the constant specified in the statement for all $\mathbf{x}_{0}\in S$.
-Finally, noting that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\leq\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]+\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]$ and $\tau(x)\geq x$ for any $x\in(0,1)$ less than a constant gives us the statement.
-∎
+Assume for contradiction that the set $U$ that contains $\mathbf{x}_{0}\in\mathbb{R}^{d}$ such that for all $i\in\{0,1\}$ it holds $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]>(C-4)K\tau\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\right)+\epsilon$ has probability mass less than $\epsilon$, i.e., $\operatorname*{\mathbf{Pr}}[x\in U]<\epsilon$.
+Then, by the definition of the piecewise constant classifier we can write
+$\mathbf{Pr}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]=\operatorname*{\mathbf{E}}_{\mathbf{x}_{0}^{V}}\left[\min_{i\in\{0,1\}}\mathbf{Pr}_{(\mathbf{x},y)\sim D}\left[i\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\right]\right].$
+Note that for every $\mathbf{x}_{0}^{V}$ not in $U$, it holds that
+$\min_{i\in\{0,1\}}\mathbf{Pr}_{(\mathbf{x},y)\sim D}\bigl{[}i\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\bigr{]}\leq(C-4)K\,\tau\Bigl{(}\mathbf{Pr}_{(\mathbf{x},y)\sim D}\bigl{[}f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\bigr{]}\Bigr{)}+\epsilon.$
+Therefore, by taking the expectation over $\mathbf{x}_{0}^{V}$ and applying Jensen’s inequality (noting that $\tau(x)=x\log(1/x+1)$ is concave for all $x>0$), we obtain
+$\mathbf{Pr}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]$ $\leq(C-4)K\,\operatorname*{\mathbf{E}}_{\mathbf{x}_{0}^{V}}\Bigl{[}\tau\left(\mathbf{Pr}_{(\mathbf{x},y)\sim D}\left[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\bigr{]}\right)\right]+\epsilon+\operatorname*{\mathbf{Pr}}[\mathbf{x}^{V}\in U]$
+$\leq(C-4)K\,\tau\left(\operatorname*{\mathbf{E}}_{\mathbf{x}_{0}^{V}}\left[\mathbf{Pr}_{(\mathbf{x},y)\sim D}\left[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\right]\right]\right)+2\epsilon$
+$=(C-4)K\,\tau\left(\mathbf{Pr}_{(\mathbf{x},y)\sim D}\left[f(\mathbf{x})\neq y\right]\right)+2\epsilon\;,$
+which contradicts the assumed error lower bound for $h$.
+Therefore, it holds that $\mathbf{Pr}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{x}\in U]\geq\epsilon$.
+Finally, note that for every $\mathbf{x}_{0}$ in $U$ we have, by the definition of $U$,
+$(C-4)K\,\tau\left(\mathbf{Pr}_{(\mathbf{x},y)\sim D}\left[f(\mathbf{x})\neq y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\right]\right)+\epsilon\leq\mathbf{Pr}_{(\mathbf{x},y)\sim D}\left[y\neq i\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}\right]\leq 1.$
+By using that $\tau(x)\geq x/4$ for all $x\in(0,1]$ we deduce the desired claim. ∎
 Fix a point $\mathbf{x}_{0}\in\mathbb{R}^{d}$ that satisfies the statement of Claim 4.3.
-Let $D_{V^{\perp}}$ be a distribution supported on $V^{\perp}\times\{\pm 1\}$ defined as $D_{V^{\perp}}(\mathbf{x}^{\perp V},y)=D(\mathbf{x},y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V})$.
-By the properties of the Gaussian distribution, we have that the $\mathbf{x}$-marginal of $D_{V^{\perp}}$ is a standard Gaussian.
-Moreover, by the definition of $D_{V^{\perp}}$, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f^{\prime}(\mathbf{x})\neq i]>(C-2)K\tau(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f^{\prime}(\mathbf{x})\neq y])+\epsilon$ for all $i\in\{\pm 1\}$ and $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f^{\prime}(\mathbf{x})\neq y]\leq 1/(C-2)$, where $f^{\prime}(\mathbf{x})=f(\mathbf{x}^{\perp V}+\mathbf{x}_{0}^{V})$.
+Let $D_{V^{\perp}}$ be a distribution supported on $V^{\perp}\times\{0,1\}$ defined as $D_{V^{\perp}}(\mathbf{x}^{\perp V},y)=D(\mathbf{x},y\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V})$.
+By the properties of the Gaussian distribution, we have that the $\mathbf{x}$-marginal of $D_{V^{\perp}}$ is a standard Gaussian distribution.
+Moreover, by the definition of $D_{V^{\perp}}$, we have that for $f^{\prime}(\mathbf{x})=f(\mathbf{x}^{\perp V}+\mathbf{x}_{0}^{V})$ and all $i\in\{0,1\}$ it holds that
+$\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[y\neq i]$ $>(C-4)K\tau(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f^{\prime}(\mathbf{x})\neq y])+\epsilon$
+$\geq((C-4)K/4)\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f^{\prime}(\mathbf{x})\neq y]\log(1/\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[f^{\prime}(\mathbf{x})\neq y])+\epsilon$
+and $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f^{\prime}(\mathbf{x})\neq y]\leq 4/(C-4)$.
 Finally, applying Proposition 4.4 for every intersection of halfspaces $f^{\prime}(\mathbf{x})=f(\mathbf{x}^{\perp V}+\mathbf{x}_{0}^{V})$ for $\mathbf{x}_{0}^{V}$ that satisfies the statement of Claim 4.3 completes the proof of Proposition 4.2.
 ∎
-Below we state and prove a structural result concerning the intersections of halfspaces.
-This result asserts that if an intersection of halfspaces is far from being constant, then either the first or the second moment exhibits a non-trivial correlation with its weight vectors.
-This result is crucial for the execution of our algorithm, as it enables us to improve at every iteration.
+Below, we state and prove a structural result for intersections of halfspaces.
+In particular, we show that if an intersection of halfspaces is far from being a constant function, then either its first moment or its second moment exhibits a non-trivial correlation with the weight vectors that define it.
+This structural property is crucial for our algorithm, as it guarantees that at every iteration we can identify a direction along which to improve our approximation.
 Proposition 4.4 (Structural Result for Intersections of Halfspaces).
 Let $d\in\mathbb{Z}_{+}$, $\epsilon,\mathrm{OPT}\in(0,1)$, and let $D$ be a distribution over $\mathbb{R}^{d}\times\{0,1\}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$.
-Let $f$ be a classifier in $\mathcal{F}_{K,d}$ that satisfies $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$, and assume that $f$ can be expressed as $f(\mathbf{x})=\prod_{i\in[K]}\mathds{1}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}\geq 0)$ with $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d},t_{1},\dots,t_{K}\in\mathbb{R}$.
-If it holds that $\min(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq 0],\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq 1])\geq CK\mathrm{OPT}\log(1/\mathrm{OPT})+\epsilon$ and $\mathrm{OPT}\leq 1/C$ for some sufficiently large absolute constant $C$, then there exists some $i\in[K]$ such that
+Let $f\in\mathcal{F}_{K,d}$ be a classifier with error $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$, and assume that $f$ can be expressed in the form $f(\mathbf{x})=\prod_{i\in[K]}\mathds{1}(\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}\geq 0)$ with $\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)}\in\mathbb{R}^{d},t_{1},\dots,t_{K}\in\mathbb{R}$.
+Let $C>0$ be a sufficiently large universal constant.
+If it holds that $\min(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq 0],\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y\neq 1])\geq CK\mathrm{OPT}\log(1/\mathrm{OPT})+\epsilon$ and $\mathrm{OPT}\leq 1/C$, then there exists some $i\in[K]$ such that
 $\max\left(\left|\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}\left[\mathbf{w}^{(i)}\cdot\mathbf{x}y\right]\right|,\left|\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}\left[\left((\mathbf{w}^{(i)}\cdot\mathbf{x})^{2}-1\right)y\right]\right|\right)\geq\mathrm{poly}(1/K,\epsilon)\;.$
 Proof.
-Let $W=\mathrm{span}(\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)})$ and without loss of generality assume that $\left\|\mathbf{w}^{(i)}\right\|=1$ for all $i\in[K]$, since we can normalize each halfspace without changing the output of $f$.
-Assume that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]\rvert\leq p(\epsilon/K)$, for all $i\in[K]$, where $p$ is a sufficiently small universal polynomial (to be quantified later).
-We show that there exists $i\in[K]$ such that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[((\mathbf{w}^{(i)}\cdot\mathbf{x})^{2}-1)y]\rvert\geq\mathrm{poly}(\epsilon/K)$ for a sufficiently small universal polynomial.
+Let $W=\mathrm{span}(\mathbf{w}^{(1)},\dots,\mathbf{w}^{(K)})$ and without loss of generality assume that $\left\|\mathbf{w}^{(i)}\right\|=1$ for all $i\in[K]$, since normalizing the weight vectors does not change the output of $f$.
+Assume that for all $i\in[K]$ the first moments are small, i.e., $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]\rvert\leq\epsilon p(\epsilon/K)$, where $p$ is a sufficiently small universal polynomial (to be quantified later).
+We show that there exists $i\in[K]$ such that the second moment is large, i.e., $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[((\mathbf{w}^{(i)}\cdot\mathbf{x})^{2}-1)y]\rvert\geq\mathrm{poly}(\epsilon/K)$ for a sufficiently small universal polynomial.
 We can decompose the second-moment matrix as follows:
-$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})y]$ $=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})f(\mathbf{x})]+\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})(\mathds{1}(f(\mathbf{x})\neq y=1)-\mathds{1}(y\neq f(\mathbf{x})=1))]$
+$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})y]$ $=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})f(\mathbf{x})]+\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})(\mathds{1}(f(\mathbf{x})\neq y,y=1)-\mathds{1}(y\neq f(\mathbf{x}),f(\mathbf{x})=1))]$
 $=\underbrace{\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[({\mathbf{x}}{\mathbf{x}}^{\top}-\mathbf{I})f(\mathbf{x})]}_{\mathbf{I}_{1}}\underbrace{+\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[({\mathbf{x}}{\mathbf{x}}^{\top}-\mathbf{I})(T(\mathbf{x},y)-S(\mathbf{x},y))]}_{\mathbf{I}_{2}}\;,$
-where $S(\mathbf{x},y)\coloneqq\mathds{1}(y\neq f(\mathbf{x})=1),T(\mathbf{x},y)\coloneqq\mathds{1}(f(\mathbf{x})\neq y=1)$.
+where $S(\mathbf{x},y)\coloneqq\mathds{1}(y\neq f(\mathbf{x}),f(\mathbf{x})=1),T(\mathbf{x},y)\coloneqq\mathds{1}(f(\mathbf{x})\neq y,y=1)$.
 Note that $\mathbf{I}_{1}$ is the second moment of $f(\mathbf{x})$, i.e., the noiseless classifier, and $\mathbf{I}_{2}$ is the contribution corresponding to the noise.
@@
 Claim 4.5.
-Let $\epsilon\in(0,1)$ be less than a sufficiently small universal constant, and let $R:\mathbb{R}^{d}\to[0,1]$ be a function such that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]=\epsilon$.
-For all unit vectors $\mathbf{v}\in\mathbb{R}^{d}$ the following holds:
+Let $\epsilon\in(0,1)$ less than a sufficiently small universal constant and let $R:\mathbb{R}^{d}\to[0,1]$ be a function such that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]=\epsilon$.
+Then, for every unit vector $\mathbf{v}\in\mathbb{R}^{d}$, the following bounds hold:
   • i)
-    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[(\mathbf{v}\cdot\mathbf{x})^{2}R(\mathbf{x})]\lesssim\epsilon\log(1/\epsilon)$.
+    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})(\mathbf{v}\cdot\mathbf{x})^{2}]\lesssim\epsilon\log(1/\epsilon)$.
   • ii)
-    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{v}\cdot\mathbf{x}R(\mathbf{x})]\lesssim\epsilon\sqrt{\log(1/\epsilon)}$.
+    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})(\mathbf{v}\cdot\mathbf{x})]\lesssim\epsilon\sqrt{\log(1/\epsilon)}$.
 Proof.
-First, we will prove (i) and then we apply it to prove (ii).
-Note that the expectation $\operatorname*{\mathbf{E}}[(\mathbf{v}\cdot\mathbf{x})^{2}R(\mathbf{x})]$ is minimized if $R$ puts all of its mass at the tails of the function $(\mathbf{v}\cdot x)^{2}$, i.e., $R(\mathbf{x})=\mathds{1}((\mathbf{v}\cdot\mathbf{x})^{2}\geq y_{\epsilon})$, where $y_{\epsilon}$ is the smallest number such that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathds{1}((\mathbf{v}\cdot\mathbf{x})^{2}\geq y_{\epsilon})]\leq\epsilon$.
+First, we will prove (i) and then we will apply it to prove (ii).
+Note that the expectation $\operatorname*{\mathbf{E}}[(\mathbf{v}\cdot\mathbf{x})^{2}R(\mathbf{x})]$ is minimized if $R$ puts all of its mass at the tails of the function $(\mathbf{v}\cdot\mathbf{x})^{2}$, i.e., $R(\mathbf{x})=\mathds{1}((\mathbf{v}\cdot\mathbf{x})^{2}\geq y_{\epsilon})$, where $y_{\epsilon}$ is the smallest number such that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathds{1}((\mathbf{v}\cdot\mathbf{x})^{2}\geq y_{\epsilon})]\leq\epsilon$.
 Hence, we have that
@@
 […]
-$=\frac{1}{\sqrt{2\pi}}y_{\epsilon}e^{-{y_{\epsilon}}/2}-\epsilon$ (4)
+$=\frac{1}{\sqrt{2\pi}}y_{\epsilon}e^{-{y_{\epsilon}}/2}+\epsilon$ (4)
 $\leq\epsilon\log(1/\epsilon)+\epsilon\lesssim\epsilon\log(1/\epsilon)\;,$ (5)
-where in (4) we simply used partial integration.
-In (5) we used the fact that the function $xe^{-x/2}$ is decreasing when $x$ is larger than $2$, note that from A.1 we have that $y_{\epsilon}\gtrsim\log(1/\epsilon)\geq 2$, since $\epsilon$ is assumed to be less than some sufficiently small constant.
+where in (4) we simply used integration by parts.
+In (5) we used the fact that the function $xe^{-x/2}$ is decreasing for $x\geq 2$.
+Note that, by A.1, we have $y_{\epsilon}\gtrsim\log(1/\epsilon)\geq 2$, and since $\epsilon$ is assumed to be sufficiently small.
 Finally, we can prove (ii) by simply using Cauchy-Schwarz
-$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{v}\cdot\mathbf{x}R(\mathbf{x})]^{2}\leq\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[(\mathbf{v}\cdot\mathbf{x})^{2}R(\mathbf{x})]\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]\lesssim\epsilon^{2}\log(1/\epsilon)\;.$
+$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{v}\cdot\mathbf{x}R(\mathbf{x})]^{2}\leq\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[(\mathbf{v}\cdot\mathbf{x})^{2}R(\mathbf{x})]\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]\lesssim\epsilon^{2}\log(1/\epsilon)\;,$
 which concludes the proof of Claim 4.5. ∎
-We next show that there exists a direction within $W$ along which the second moment is small.
+We next show that there exists a direction within $W$ along which the second moment of $f(\mathbf{x})$ is small. More precisely, we have the following:
 Claim 4.6.
 There exists an $i\in[K]$ such that
-$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\mathbf{w}^{(i)}\lesssim-\frac{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]}{K}+p(\epsilon/K)+\mathrm{OPT}\log(1/\mathrm{OPT})\;.$
+$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\mathbf{w}^{(i)}\lesssim-\frac{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]}{K}+p\bigg{(}\frac{\epsilon}{K}\bigg{)}+\mathrm{OPT}\log(1/\mathrm{OPT})\;.$
 Proof.
-Denote by $\mu\coloneqq\operatorname*{\mathbf{E}}[\mathbf{x}\mid f(\mathbf{x})]$, by $\bar{\mathbf{x}}\coloneqq\mathbf{x}-\mathbf{\mu}$ and by $\mathbf{M}=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\bar{\mathbf{x}}\bar{\mathbf{x}}^{\top}f(x)]$.
-For a defining vector $\mathbf{w}^{(i)}$, by B.4 we have that $(\mathbf{w}^{(i)})^{\top}\mathbf{M}\mathbf{w}^{(i)}\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1](1-ce^{-(\max(0,t_{i}))^{2}/2}),$ for a sufficiently small universal constant $c>0$.
-Furthermore, we claim that there exists $i\in[K]$ for which the bias $t_{i}$ is at most $2\sqrt{\ln(\frac{K}{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]})}$.
-To see that, note that if for every $i\in[K]$ we have that $t_{i}>2\sqrt{\ln(\frac{K}{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]})},$ then by the union bound
+Denote by $\mu\coloneqq\operatorname*{\mathbf{E}}[\mathbf{x}f(\mathbf{x})]/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]$ and define $\bar{\mathbf{x}}\coloneqq\mathbf{x}-\mathbf{\mu}$.
+Let $\mathbf{M}=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\bar{\mathbf{x}}\bar{\mathbf{x}}^{\top}f(x)]$.
+For any defining vector $\mathbf{w}^{(i)}$, by B.4 we have
+$(\mathbf{w}^{(i)})^{\top}\mathbf{M}\mathbf{w}^{(i)}\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\big{(}1-ce^{-(\max(0,t_{i}))^{2}/2}\big{)}\;,$
+for a sufficiently small universal constant $c>0$.
+Next, we claim that there exists some $i\in[K]$ for which the bias $t_{i}$ is at most $2\sqrt{\ln({K}/{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]})}$.
+To see this, assume for contradiction that for every $i\in[K]$ we have $t_{i}>2\sqrt{\ln(\frac{K}{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]})}$ and note that if $t_{i}>2\sqrt{\ln(\frac{K}{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]})}$ then $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}\leq 0]<\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]/K$.
+By applying the union bound, we obtain
 $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})\neq 1]\leq\sum_{i=1}^{K}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{w}^{(i)}\cdot\mathbf{x}+t_{i}\leq 0]<\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]\;,$
 which is a contradiction.
-Hence, it holds that $(\mathbf{w}^{(i)})^{\top}\mathbf{M}\mathbf{w}^{(i)}\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1](1-c\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]/K)$ for some $i\in[K]$.
-Therefore, we have that
-$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\mathbf{w}^{(i)}$ $=(\mathbf{w}^{(i)})^{\top}\mathbf{M}\mathbf{w}^{(i)}-\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=1]+\frac{\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{w}^{(i)}\cdot\mathbf{x}f(\mathbf{x})]^{2}}{\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=1]}$
-$\leq-c\frac{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]}{K}+\frac{\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{w}^{(i)}\cdot\mathbf{x}f(\mathbf{x})]^{2}}{\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=1]}\;.$
+Hence, there must exist some index $i\in[K]$ with $(\mathbf{w}^{(i)})^{\top}\mathbf{M}\mathbf{w}^{(i)}\leq\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1](1-c\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]/K)$.
+Fix such an index $i\in[K]$.
+Then, by definition,
+$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\,\mathbf{w}^{(i)}=\;(\mathbf{w}^{(i)})^{\top}\mathbf{M}\,\mathbf{w}^{(i)}-\mathbf{Pr}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=1]+\frac{\left(\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\Big{[}\mathbf{w}^{(i)}\cdot\mathbf{x}\,f(\mathbf{x})\Big{]}\right)^{2}}{\mathbf{Pr}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=1]}.$
+Substituting the bound on $(\mathbf{w}^{(i)})^{\top}\mathbf{M}\,\mathbf{w}^{(i)}$, we obtain
+$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\,\mathbf{w}^{(i)}\leq-c\,\frac{\mathbf{Pr}[f(\mathbf{x})=1]\,\mathbf{Pr}[f(\mathbf{x})\neq 1]}{K}+\frac{\left(\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\Big{[}\mathbf{w}^{(i)}\cdot\mathbf{x}\,f(\mathbf{x})\Big{]}\right)^{2}}{\mathbf{Pr}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=1]}.$
 Note that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathbf{w}^{(i)}\cdot\mathbf{x}f(\mathbf{x})]=\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]+\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}(S(\mathbf{x},y)-T(\mathbf{x},y))]$.
-By the assumption that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]\rvert\leq\epsilon p(\epsilon/K)$ for all $i\in[K]$, and the fact that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\geq\operatorname*{\mathbf{Pr}}[y=1]-\mathrm{OPT}\geq\epsilon$, it holds that $\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]^{2}/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\leq p(\epsilon/K)$.
-Moreover, note that $\lvert\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}S(\mathbf{x},y)]\rvert=\lvert\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}R(\mathbf{x})]\rvert$, for $R(\mathbf{x})=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]$.
-Therefore, since $R(\mathbf{x})\in[0,1]$ for all $\mathbf{x}\in\mathbb{R}^{d}$ and $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]\leq\mathrm{OPT}$, by Claim 4.5 we have that $\lvert\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}(S(\mathbf{x},y)]\rvert\lesssim\mathrm{OPT}\sqrt{\log(1/\mathrm{OPT})}$.
-Consequently, since $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\geq\mathrm{OPT}$ we have that $\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}S(\mathbf{x},y)]^{2}/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\lesssim\mathrm{OPT}\log(1/\mathrm{OPT})$.
-Using similar arguments we have that $\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}T(\mathbf{x},y)]^{2}/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\lesssim\mathrm{OPT}\log(1/\mathrm{OPT})$.
-Combining the above results gives us
-$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\mathbf{w}^{(i)}\lesssim-\frac{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]}{K}+p(\epsilon/K)+\mathrm{OPT}\log(1/\mathrm{OPT})\;.$
+Recall that by assumption we have that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]\rvert\leq\epsilon p(\epsilon/K)$ for all $i\in[K]$.
+Moreover, using the triangle inequality $\mathbf{Pr}[f(\mathbf{x})=1]\geq\mathbf{Pr}[y=1]-\mathbf{Pr}[f(\mathbf{x})\neq y]$ together with the assumption in Proposition 4.4 that $\mathbf{Pr}[y=1]\geq CK\mathrm{OPT}\log(1/\mathrm{OPT})+\epsilon$, it follows that $\mathbf{Pr}[f(\mathbf{x})=1]\geq\mathrm{OPT}+\epsilon$.
+Thus,
+$\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}y]^{2}/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\leq p(\epsilon/K)\;.$
+Next, observe that $\lvert\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}S(\mathbf{x},y)]\rvert=\lvert\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}R(\mathbf{x})]\rvert$, where we define $R(\mathbf{x})=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[S(\mathbf{x},y)\mid\mathbf{x}]$.
+Since, for all $\mathbf{x}\in\mathbb{R}^{d}$ it holds that $R(\mathbf{x})\in[0,1]$, and $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[R(\mathbf{x})]\leq\mathrm{OPT}$, by Claim 4.5 we have that $\lvert\operatorname*{\mathbf{E}}[(\mathbf{w}^{(i)}\cdot\mathbf{x})S(\mathbf{x},y)]\rvert\lesssim\mathrm{OPT}\sqrt{\log(1/\mathrm{OPT})}$.
+Consequently, since $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\geq\mathrm{OPT}$, it follows that $\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}S(\mathbf{x},y)]^{2}/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\lesssim\mathrm{OPT}\log(1/\mathrm{OPT})$.
+A similar argument shows that $\operatorname*{\mathbf{E}}[\mathbf{w}^{(i)}\cdot\mathbf{x}T(\mathbf{x},y)]^{2}/\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\lesssim\mathrm{OPT}\log(1/\mathrm{OPT})$.
+Combining these bounds, we obtain
+$(\mathbf{w}^{(i)})^{\top}\mathbf{I}_{1}\mathbf{w}^{(i)}\lesssim-\frac{\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]}{K}+p\bigg{(}\frac{\epsilon}{K}\bigg{)}+\mathrm{OPT}\log(1/\mathrm{OPT})\;,$
 which concludes the proof of Claim 4.6. ∎
@@
 By using similar arguments we have that $\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{v}\cdot\mathbf{x})^{2}T(\mathbf{x},y)]\rvert=O(\log(1/\mathrm{OPT})\mathrm{OPT})$.
-Therefore, we have that
+Therefore, by triangle inequality, we have that
 $\lvert\mathbf{v}^{\top}\mathbf{I}_{2}\mathbf{v}\rvert$ $\leq\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{v}\cdot\mathbf{x})^{2}S(\mathbf{x},y)]\rvert+\lvert\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{v}\cdot\mathbf{x})^{2}T(\mathbf{x},y)]\rvert+\operatorname*{\mathbf{E}}[S(\mathbf{x},y)+T(\mathbf{x},y)]$
-$=O(\log(1/\mathrm{OPT})\mathrm{OPT})\;.$
+$=O(\mathrm{OPT}\log(1/\mathrm{OPT}))\;,$
 which concludes the proof of Claim 4.7. ∎
-Denote by $t$ the quantity $t\coloneqq C^{\prime}\log(1/\mathrm{OPT})\mathrm{OPT}$, for sufficiently large constant $C^{\prime}$. Combining the Claims 4.6 and 4.7 we have that
-$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})y]$ $\lesssim-\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]/K-t+p(\epsilon/K)$
+Denote by $t$ the quantity $t\coloneqq C^{\prime}\mathrm{OPT}\log(1/\mathrm{OPT})$, for sufficiently large constant $C^{\prime}$.
+Combining Claims 4.6 and 4.7, we have that there exists $i\in[K]$ such that
+$(\mathbf{w}^{(i)})^{\top}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathbf{x}\mathbf{x}^{\top}-\mathbf{I})y](\mathbf{w}^{(i)})$ $\lesssim-\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]/K-t+p(\epsilon/K)$
 $\leq-\left(\min_{i\in\{0,1\}}(\operatorname*{\mathbf{Pr}}[y=i])-\mathrm{OPT}\right)/(2K)-t+p(\epsilon/K)$
 $\leq-\epsilon/(2K)+p(\epsilon/K)\;,$
-where we used the fact that one of $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1],\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]$ is greater than $1/2$ and the assumption that for all labels $i\in\{0,1\}$, $\operatorname*{\mathbf{Pr}}[y=i]\geq CK\mathrm{OPT}\log(1/\mathrm{OPT})+\epsilon$, for a sufficiently large constant $C$.
-Setting $p(\epsilon/K)\leq c\epsilon/K$ for a sufficiently small universal constant $c>0$, concludes the proof of Proposition 4.4.
+where we used the fact that one of $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})=1]$ or $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq 1]$ must be at least $1/2$, and the assumption of Proposition 4.4 that for every label $i\in\{0,1\}$ it holds that $\operatorname*{\mathbf{Pr}}[y=i]\geq CK\mathrm{OPT}\log(1/\mathrm{OPT})+\epsilon$, for a sufficiently large constant $C$.
+Setting $p(\epsilon/K)\leq c\epsilon/K$ for a sufficiently small universal constant $c>0$, we ensure that the error term $p(\epsilon/K)$ is negligible compared to the main term.
+This concludes the proof of Proposition 4.4.
 ∎
@@
 We show that Algorithm 3 can be modified into Algorithm 5, which achieves improved complexity in the RCN regime compared to the agnostic noise setting.
-The main result of this section is Theorem 4.9.
+The main result of this section is Theorem 4.9 (Theorem 1.11).
 Definition 4.8 (RCN Distribution).
 Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$.
-We say that $D$ is a RCN distribution for the class $\mathcal{F}\subseteq\{f:\mathbb{R}^{d}\to\mathcal{Y}\}$ if there exists a function $f\in\mathcal{F}$ and a row-stochastic matrix $\mathbf{H}\in\mathbb{R}^{\lvert\mathcal{Y}\rvert\times\lvert\mathcal{Y}\rvert}$ with $\mathbf{H}_{i,i}>\mathbf{H}_{i,j}+\gamma$ for all $i,j\in\mathcal{Y}$, for some $\gamma>0$, such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y=i\mid\mathbf{x}]=\mathbf{H}_{f(\mathbf{x}),i}$.
+We say that $D$ is an RCN distribution for the class $\mathcal{F}\subseteq\{f:\mathbb{R}^{d}\to\mathcal{Y}\}$ if there exists a function $f\in\mathcal{F}$ and a row-stochastic matrix $\mathbf{H}\in\mathbb{R}^{\lvert\mathcal{Y}\rvert\times\lvert\mathcal{Y}\rvert}$ with $\mathbf{H}_{i,i}>\mathbf{H}_{i,j}+\gamma$ for all $i,j\in\mathcal{Y}$, for some $\gamma>0$, such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[y=i\mid\mathbf{x}]=\mathbf{H}_{f(\mathbf{x}),i}$.
 We refer to $\mathbf{H}$ as the confusion matrix of the distribution.
-The key distinction that allows for this improvement is that, in the RCN setting, the output label depends solely on the ground truth label.
-Therefore, the moments of the noisy class regions still rely only on the subspace that characterizes the classifier.
-Consequently, any non-zero moment results in vectors that approximately lie within that subspace (see Proposition 4.10).
-The main result of this section is Theorem 4.9.
-Input: Accuracy $\epsilon>0$, $\delta>0$ and sample access to a distribution $D$ over $\mathbb{R}^{d}\times\mathcal{Y}$ that satisfies Definition 4.8.
-Output: A hypothesis $h$ such that with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$.
+The key distinction that enables us to have an improved algorithm (in terms of complexity) in the RCN setting is that the output label is determined solely by the ground truth label.
+Consequently, the moments of the observed labels (noisy labels) depend only on the subspace defining the classifier.
+This implies that any nonzero moment yields vectors that lie approximately within that subspace (see Proposition 4.10).
+The main result of this section is stated in Theorem 4.9.
+Input: $\epsilon,\delta>0$ and sample access to a distribution $D$ over $\mathbb{R}^{d}\times\mathcal{Y}$ that satisfies Definition 4.8 and whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$.
+Output: A hypothesis $h$ such that with probability at least $1-\delta$ it holds that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$, where $\mathrm{OPT}\coloneqq\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x}){\neq}y]$.
   1. 1.
@@
     1. (a)
-      Draw a set $S$ of $N$ i.i.d. samples from $D$.
+      Draw a set $S_{t}$ of $N$ i.i.d. samples from $D$.
     2. (b)
-      $\mathcal{E}_{t}\leftarrow$ Algorithm 4($\epsilon,\epsilon^{\prime},\delta,\mathrm{span}(L_{t}),S$).
+      $\mathcal{E}_{t}\leftarrow$ Algorithm 4($\epsilon,\epsilon^{\prime},\delta,\mathrm{span}(L_{t}),S_{t}$).
     3. (c)
@@
   5. 5.
-    Return a piecewise constant approximating classifier $h_{\mathcal{S}}$, as described in Definition 2.3, using $N$ i.i.d.
-samples from $D$, by finding the most frequent label at any set in $\mathcal{S}$.
+    Draw $N$ i.i.d.
+samples from $D$ and construct the piecewise constant classifier $h_{\mathcal{S}}$ as follows: For each $S\in\mathcal{S}$, assign a label that appears most frequently among the samples falling in $S$.
+Return $h_{\mathcal{S}}$.
 Algorithm 5 Learning Multi-Index Models under RCN
-Theorem 4.9 (Learning with RCN).
-Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, and let $\epsilon,\delta\in(0,1)$.
-Let $\mathrm{OPT}=\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$, where $\mathcal{F}$ is a class of $(m,{\zeta+\epsilon},\alpha,K,\Gamma,\tau,\sigma)$-well-behaved Multi-Index Models, as in Definition 3.2.
+Theorem 4.9 (Learning MIMs with RCN).
+Let $D$ be a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, and let $\epsilon,\delta\in(0,1/2)$.
+Let $\mathrm{OPT}=\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]$, where $\mathcal{F}$ is a class of $(m,\zeta+\epsilon,\alpha,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models, as in Definition 3.2.
 If $\zeta\geq\mathrm{OPT}$ and $D$ satisfies Definition 4.8, then Algorithm 5 draws $N=(dm)^{O(m)}2^{\mathrm{poly}(K)}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{O(K)}\log(1/\delta)/(\sigma\alpha)^{O(1)}$ i.i.d.
-samples from $D$, runs in time $\mathrm{poly}(d,N,\lvert\mathcal{Y}\rvert)$, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$.
-We remark that for the case of multiclass linear classifiers, similarly to Proposition 4.2, using Proposition 2.4 we can prove that the class of multiclass linear classifiers is $(1,\zeta,\epsilon,K,O(\zeta)+O(\epsilon),\mathrm{poly}(\epsilon/K),O(K\sqrt{\log(K)}))$-well behaved.
-As a result, we obtain an algorithm for learning under RCN by applying Theorem 4.9.
-Moreover, note that in this case Algorithm 4 performs linear regression under Gaussian marginals (see Line 4).
+samples from $D$, runs in time $\mathrm{poly}(N)$, and returns a hypothesis $h$ such that, with probability at least $1-\delta$, it holds $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$.
+We remark that for the class of multiclass linear classifiers using Proposition 2.4, we can show that this class is $(1,\zeta,\epsilon,K,O(\zeta+\epsilon),\mathrm{poly}(\epsilon/K),O(K\sqrt{\log(K)}))$-well-behaved.
+As a result, by applying Theorem 4.9, we obtain an efficient algorithm for learning under Random Classification Noise (RCN).
+Moreover, for the class of multiclass linear classifiers Algorithm 4 performs linear regression under Gaussian marginals (see Line 4).
 This task that can be efficiently reduced to mean estimation, specially to estimating $\operatorname*{\mathbf{E}}[\mathds{1}(y=i)]$ and $\operatorname*{\mathbf{E}}[\mathbf{x}\mathds{1}(y=i)]$.
-Therefore, this step can be performed in polynomial time using $O(d)$ samples.
-Consequently, the final sample complexity of the algorithm scales with $d$ rather than $d^{O(1)}$, which establishes Theorem 1.8.
-Our proof of the Theorem 4.9 relies on the following proposition, which states that, given that our distribution is an RCN distribution with respect to a function in the class, all vectors returned by Algorithm 4 have an improved correlation with the ground truth subspace compared to the agnostic case (see Proposition 3.7).
-Proposition 4.10.
+Since these mean estimates can be computed in polynomial time using $O(d)$ samples, the overall sample complexity of the algorithm scales as $d$, rather than $d^{O(1)}$.
+This establishes Theorem 1.11.
+Our proof of the Theorem 4.9 relies on the following proposition, which states that, given that our distribution is an RCN distribution with respect to a function in the class, any vector returned by Algorithm 4 has an improved correlation with the ground truth subspace compared to the agnostic case (see Proposition 3.7).
+Proposition 4.10 (Structural Result for RCN).
 Let $d,k,K\in\mathbb{Z}_{+}$, $\epsilon,\delta\in(0,1)$ and $\epsilon^{\prime}\leq c\epsilon^{3}/(\lvert\mathcal{Y}\rvert\Gamma^{2}k^{2}))$, for $c$ a sufficiently small universal constant.
-Let $D$ be a distribution supported on $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$ and let $\mathcal{F}\subseteq\{f:\mathbb{R}^{d}\to\mathcal{Y}\}$ be a class of $(m,{\zeta+O(\epsilon)},\alpha,K,\Gamma,\tau,\sigma)$-well-behaved Multi-Index Models according to Definition 3.2.
-Let $f\in\mathcal{F}$ with $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$ and let $W$ be a subspace such that $f(\mathbf{x})=f(\mathbf{x}^{W})$ for all $\mathbf{x}\in\mathbb{R}^{d}$, with $K=\dim(W)$.
-Additionally, let $h_{\mathcal{S}}$ be a piecewise constant approximation of $D$, with respect to an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$, according to $V$, as defined in Definitions 2.2 and 2.3.
-There exists $N=(dm)^{O(m)}{(k/\epsilon^{\prime})^{O(k)}}\log(1/\delta){(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{O(1)}}$ such that, if $D$ satisfies Definition 4.8 for $f$, $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+\epsilon$ and $\zeta\geq\mathrm{OPT}$, then Algorithm 4 given $N$ i.i.d.
-samples from $D$, runs in $\mathrm{poly}(N,d,\lvert\mathcal{Y}\rvert)$ time, and with probability at least $1-\delta$, returns a list of unit vectors $\mathcal{E}$ of size $|\mathcal{E}|=\mathrm{poly}(mK\Gamma/(\epsilon\sigma\alpha))$, such that: for any vector $\mathbf{v}\in\mathcal{E}$, there exists a unit vector $\mathbf{w}\in W$ for which
+Let $D$ be a distribution supported on $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, let $V$ be a $k$-dimensional subspace of $\mathbb{R}^{d}$ and let $\mathcal{F}\subseteq\{f:\mathbb{R}^{d}\to\mathcal{Y}\}$ be a class of $(m,\zeta+O(\epsilon),\alpha,K,\tau,\sigma,\Gamma)$-well-behaved Multi-Index Models according to Definition 3.2.
+Consider a function $f\in\mathcal{F}$ with $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[f(\mathbf{x})\neq y]=\mathrm{OPT}$ and suppose that $f$ depends only on a $K$-dimensional subspace $W$, i.e., $f(\mathbf{x})=f(\mathbf{x}^{W})$ for all $\mathbf{x}\in\mathbb{R}^{d}$.
+In addition, let $h_{\mathcal{S}}$ be a piecewise constant approximation of $D$ with respect to an $\epsilon^{\prime}$-approximating partition $\mathcal{S}$ of $V$ (as defined in Definitions 2.2 and 2.3).
+Then, there exists a sample size $N=(dm)^{O(m)}(k/\epsilon^{\prime})^{O(k)}\log(1/\delta)(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{O(1)}$ such that if $D$ satisfies the RCN condition (see Definition 4.8) for $f$, and if $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{\mathcal{S}}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+\epsilon$ and $\zeta\geq\mathrm{OPT}$, then Algorithm 4, when given $N$ i.i.d.
+samples from $D$, runs in time $\mathrm{poly}(N)$ and, with probability at least $1-\delta$, returns a list of unit vectors $\mathcal{E}$ of size $|\mathcal{E}|=\mathrm{poly}(mK\Gamma/(\epsilon\sigma\alpha))$, such that for every vector $\mathbf{v}\in\mathcal{E}$, there exists a unit vector $\mathbf{w}\in W$ with
 $\mathbf{w}\cdot\mathbf{v}\geq 1-\epsilon\;.$
@@
 Denote by $\mathbf{H}$ the stochastic matrix that perturbs the labels of $f$.
-Also let $\widehat{\mathbf{U}}$ be the matrix computed at Line 6 of Algorithm 5 and let $\eta^{2}=\mathrm{poly}(\epsilon\sigma\alpha/(mK\Gamma))$ be the $l_{2}^{2}$ error chosen in the polynomial-regression step, i.e., Line 4 of Algorithm 5.
-We begin by showing that, for each subset of the partition, the computed vectors have only a small component in the subspace orthogonal to $W$.
-This, in turn, implies that any large vector must have a significant component in $W$.
+Also let $\widehat{\mathbf{U}}$ be the matrix computed in Line 6 of Algorithm 5 and let $\eta^{2}=\mathrm{poly}(\epsilon\sigma\alpha/(mK\Gamma))$ be the $L_{2}^{2}$ error chosen in the polynomial-regression step, i.e., Line 4 of Algorithm 5.
+We show that if the sample size is sufficiently large then for each region $S\in\mathcal{S}$ the associated set $U_{S}$ (see Line 5 of Algorithm 4) contains only vectors that have only a small component in the subspace orthogonal to $W$.
+This in turn implies that any large vector must have a significant component in $W$.
 Claim 4.11.
 Fix $S\in\mathcal{S}$.
-If the number of samples that fall in $S$ is $N_{S}\geq(dm)^{Cm}\log(\lvert\mathcal{Y}/\delta\rvert)/\eta^{C}$ for a sufficiently large universal constant $C>0$, then for any $\mathbf{u}\in U_{S}$ and unit vector $\mathbf{v}\in W^{\perp}$ we have that $\lvert\mathbf{v}\cdot\mathbf{u}\rvert\leq\sqrt{2Km}\eta/\sigma$.
+Suppose that the number of samples falling in $S$ satisfies $N_{S}\geq(dm)^{Cm}\log(\lvert\mathcal{Y}/\delta\rvert)/\eta^{C}$ for a sufficiently large universal constant $C>0$.
+Then, for any $\mathbf{u}\in U_{S}$ and any unit vector $\mathbf{v}\in W^{\perp}$ we have that $\lvert\mathbf{v}\cdot\mathbf{u}\rvert\leq\sqrt{2Km}\eta/\sigma$.
 Proof.
-Let $N_{S},S\in\mathcal{S}$, be the number of samples that fall in the set $S$.
+Let $N_{S},S\in\mathcal{S}$, be the number of samples that land in the set $S$.
 From the rotational invariance of the standard Gaussian, without loss of generality let $\mathbf{e}_{1},\dots,\mathbf{e}_{t}$ be a basis of $W^{\perp}$, $t\in\mathbb{Z}_{+}$.
 By B.2 and the union bound, we have that with $N_{S}=(dm)^{O(m)}\log(\lvert\mathcal{Y}\rvert/\delta)/\eta^{O(1)}$ i.i.d.
-samples and runtime $\mathrm{poly}(N_{S},d)$, we can find polynomials $p_{S,i}$ such that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}^{\perp V}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}^{\perp V}))^{2}]+\eta^{2}$.
-Hence, by averaging over $V$ conditioning on $S$, we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}))^{2}]+\eta^{2}$, where $D_{V^{\perp}}$ is defined to be the marginal obtained by averaging $D$ over $V$ conditioned on $S$, i.e., $D_{V^{\perp}}(\mathbf{x}^{\perp V},y)=\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[D(\mathbf{x},y)\mid\mathbf{x}\in S]$.
-Define $\hat{p}_{S,i}(\beta)\coloneqq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x})p_{S,i}(\mathbf{x})],\hat{g}_{i}(\beta)\coloneqq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x})\mathds{1}(y=i)]$ for each $\beta\in\mathbb{N}^{d}$.
-Note that $\hat{p}_{S,i}(\beta)$ and $\hat{g}_{i}(\beta)$ are the Hermite coefficients of $p_{S,i}$ and $\mathds{1}(y=i)$, respectively.
-Since by the orthogonality of Hermite polynomials we have that $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]=\sum_{\beta\in\mathbb{N}^{d}}(\hat{p}_{S,i}(\beta)-\hat{g}_{i}(\beta))^{2}$, it holds that $\sum_{\beta\in\mathbb{N}^{d},1\leq\|\beta\|_{1}\leq m}(\hat{p}_{S,i}(\beta)-\hat{g}_{i}(\beta))^{2}\leq\eta^{2}$.
-Moreover, if $\beta_{i}\neq 0$, then for some $i\in[t]$
-$\hat{g}_{i}(\beta)$ $=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[H_{\beta}(\mathbf{x}^{\perp V})\mathds{1}(y=i)\mid S]=\sum_{j=1}^{K}\mathbf{H}_{j,i}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[H_{\beta}(\mathbf{x}^{\perp V})\mathds{1}(f(\mathbf{x})=j)\mid S]$
+samples and runtime $\mathrm{poly}(N_{S},d)$, we can compute polynomials $p_{S,i}$ such that
+$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}^{\perp V}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}^{\perp V}))^{2}]+\eta^{2}\;.$
+Hence, by averaging over $V$ conditioned on the cube $S$, we obtain that
+$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]\leq\min_{p^{\prime}\in\mathcal{P}_{m}}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p^{\prime}(\mathbf{x}))^{2}]+\eta^{2}\;,$
+where $D_{V^{\perp}}$ is defined to be the marginal obtained by averaging $D$ over $V$ conditioned on $S$, i.e., $D_{V^{\perp}}(\mathbf{x}^{\perp V},y)=\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[D(\mathbf{x},y)\mid\mathbf{x}\in S]$.
+Define, for each multi-index $\beta\in\mathbb{N}^{d}$, the Hermite coefficients of $p_{S,i}$ and $\mathds{1}(y=i)$ by $\widehat{p}_{S,i}(\beta)\coloneqq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x})p_{S,i}(\mathbf{x})]$ and $\widehat{g}_{i}(\beta)\coloneqq\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[H_{\beta}(\mathbf{x})\mathds{1}(y=i)]$.
+By the orthogonality of Hermite polynomials, we have
+$\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\mathds{1}(y=i)-p_{S,i}(\mathbf{x}))^{2}]=\sum_{\beta\in\mathbb{N}^{d}}(\widehat{p}_{S,i}(\beta)-\widehat{g}_{i}(\beta))^{2}\;.$
+Thus, restricting the sum to multi-indices $\beta$ with $1\leq\|\beta\|_{1}\leq m$, it follows that $\sum_{\beta\in\mathbb{N}^{d},1\leq\|\beta\|_{1}\leq m}(\widehat{p}_{S,i}(\beta)-\widehat{g}_{i}(\beta))^{2}\leq\eta^{2}$.
+Moreover, if $\beta_{i}\neq 0$, then for some $i\in[t]$, it holds
+$\widehat{g}_{i}(\beta)$ $=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[H_{\beta}(\mathbf{x}^{\perp V})\mathds{1}(y=i)\mid S]=\sum_{j=1}^{K}\mathbf{H}_{j,i}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[H_{\beta}(\mathbf{x}^{\perp V})\mathds{1}(f(\mathbf{x})=j)\mid S]$
 $=\sum_{j=1}^{K}\mathbf{H}_{j,i}\operatorname*{\mathbf{E}}_{\mathbf{x}^{V}}[\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp V\cap W}}[H_{\beta}(\mathbf{x}^{\perp V\cap W})\mathds{1}(f(\mathbf{x}^{W})=j)]\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp V\cap\perp W}}[H_{\beta}(\mathbf{x}^{\perp V\cap\perp W})]\mid S]=0\;,$
-where we used the fact that the $\mathbf{x}\in S$ depends only on $\mathbf{x}^{V}$, and the fact that non-constant Hermite polynomials have mean zero under the Gaussian distribution.
-Therefore, we have that the sum of the Hermite coefficients $\hat{p}_{S,i}(\beta)$, for all $\beta\in\mathbb{N}^{d}$ such that $1\leq\|\beta\|_{1}\leq m$ and $\beta_{i}\neq 0$ for some $i\in[t]$, squared is less than $\eta^{2}$.
-By fact A.2, for any $S\in\mathcal{S},i\in[K]$, we have that
-$\mathbf{e}_{1}^{\top}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[\nabla p_{S,i}(\mathbf{x})\nabla p_{S,i}(\mathbf{x})^{\top}]\mathbf{e}_{1}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\nabla p_{S,i}(\mathbf{x})\cdot\mathbf{e}_{1})^{2}]$ $=\sum_{\beta\in\mathbb{N}^{d}}\beta_{1}(\hat{p}_{S,i}(\beta))^{2}$
+where we used the fact that each $\mathbf{x}\in S$ depends only on $\mathbf{x}^{V}$, and any non-constant Hermite polynomial has zero mean under the standard Gaussian distribution.
+Thus, let $E_{error}$ denote the set of multi-indices $\beta$ for which the corresponding Hermite coefficient of $\mathds{1}(y=i)$ is zero, i.e., $\widehat{g}_{i}(\beta)=0$.
+Then, we have that the sum of the Hermite coefficients, restricted to multi-indexes that in set $E_{error}$ is bounded by
+$\sum_{\beta\in E_{error}}\widehat{p}_{S,i}(\beta)^{2}\leq\eta^{2}\;.$
+By A.2, for any $S\in\mathcal{S},i\in\mathcal{Y}$, we have that
+$\mathbf{e}_{1}^{\top}\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[\nabla p_{S,i}(\mathbf{x})\nabla p_{S,i}(\mathbf{x})^{\top}]\mathbf{e}_{1}=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[(\nabla p_{S,i}(\mathbf{x})\cdot\mathbf{e}_{1})^{2}]$ $=\sum_{\beta\in\mathbb{N}^{d}}\beta_{1}(\widehat{p}_{S,i}(\beta))^{2}$
 $\leq m\eta^{2}\;.$
-Hence, as $U_{S,i}$ has the eigenvectors of $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[\nabla p_{S,i}(\mathbf{x})\nabla p_{S,i}(\mathbf{x})^{\top}]=\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D}[\nabla p_{S,i}(\mathbf{x}^{\perp V})\nabla p_{S,i}(\mathbf{x}^{\perp V})^{\top}\mid S]$ with eigenvalue greater than $\sigma^{2}/(2K)$, by Claim 4.12 we have the statement of the claim.
+Hence, since $U_{S,i}$ contains the eigenvectors of $\operatorname*{\mathbf{E}}_{(\mathbf{x},y)\sim D_{V^{\perp}}}[\nabla p_{S,i}(\mathbf{x})\nabla p_{S,i}(\mathbf{x})^{\top}]$ with eigenvalue greater than $\sigma^{2}/(2K)$, using Claim 4.12 we complete the proof.
 ∎
 Claim 4.12.
-Let $\mathbf{M}\in\mathbb{R}^{d\times d}$ be a symmetric, psd matrix and let $\mathbf{v}\in\mathbb{R}^{d}$ be a vector such that $\mathbf{v}^{\top}\mathbf{M}\mathbf{v}\leq\epsilon$.
-Let $U$ denote the set of eigenvectors of $\mathbf{M}$ with eigenvalue greater than $\lambda$.
+Let $\mathbf{M}\in\mathbb{R}^{d\times d}$ be a symmetric, PSD matrix and let $\mathbf{v}\in\mathbb{R}^{d}$ be a unit vector such that $\mathbf{v}^{\top}\mathbf{M}\mathbf{v}\leq\epsilon$.
+Let $U$ denote the set of eigenvectors of $\mathbf{M}$ with eigenvalue at least $\lambda$.
 Then, for every $\mathbf{u}\in U$, it holds that $\lvert\mathbf{u}\cdot\mathbf{v}\rvert\leq\sqrt{\epsilon/\lambda}$.
@@
 $\epsilon\geq\mathbf{v}^{\top}\mathbf{M}\mathbf{v}=\sum_{i=1}^{d}\lambda_{i}(\mathbf{u}^{(i)}\cdot\mathbf{v})^{2}\geq\sum_{\mathbf{u}\in U}\lambda(\mathbf{u}\cdot\mathbf{v})^{2}\;,$
-where we used the fact that $\mathbf{M}$ is psd. Which completes the proof of Claim 4.12. ∎
-First note that by the definition of an approximating partition (see Definition 2.2) and A.1, for all $S\in\mathcal{S}$ we have that $\operatorname*{\mathbf{Pr}}_{D}[S]{=(\epsilon^{\prime}/k)^{\Omega(k)}}$ and $\lvert\mathcal{S}\rvert{=(k/\epsilon^{\prime})^{O(k)}}$.
-Therefore, by union bound and Hoeffding’s inequality, it holds that if $N\geq{(k/\epsilon^{\prime})^{Ck}}\log(\lvert\mathcal{S}\rvert/\delta)={(k/\epsilon^{\prime})^{O(k)}}\log(1/\delta)$ for a sufficiently large constant $C>0$, then with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]-\operatorname*{\mathbf{Pr}}_{D}[S]\rvert\leq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S}$.
-Hence, it is true that $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$, i.e., the number of samples that fall in each set $S\in\mathcal{S}$ is at least $N\operatorname*{\mathbf{Pr}}_{D}[S]/2=N{(\epsilon^{\prime}/k)^{\Omega(k)}}$.
-Therefore, if $N\geq(dm)^{Cm}{(k/\epsilon^{\prime})^{Ck}}\log(1/\delta){(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{C}}$ for a sufficiently large universal constant $C>0$, then by applying Claim 4.11 we have that with probability $1-\delta$ for any unit vector $\mathbf{v}\in W^{\perp}$
-$\mathbf{v}^{\top}\widehat{\mathbf{U}}\mathbf{v}\leq\sum_{S\in\mathcal{S},\mathbf{u}\in U_{S}}(\mathbf{v}\cdot\mathbf{u})^{2}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\leq\eta^{2}\mathrm{poly}(mK\lvert\mathcal{Y}\rvert/\sigma)\;.$
+where we used the fact that $\mathbf{M}$ is a PSD matrix. This completes the proof of Claim 4.12. ∎
+First note that by the definition of an approximating partition (Definition 2.2) and A.1, for all $S\in\mathcal{S}$ we have that $\operatorname*{\mathbf{Pr}}_{D}[S]=(\epsilon^{\prime}/k)^{\Omega(k)}$ and $\lvert\mathcal{S}\rvert=(k/\epsilon^{\prime})^{O(k)}$.
+Therefore, by the union bound and Hoeffding’s inequality, it holds that if $N\geq(k/\epsilon^{\prime})^{Ck}\log(\lvert\mathcal{S}\rvert/\delta)=(k/\epsilon^{\prime})^{O(k)}\log(1/\delta)$ for a sufficiently large constant $C>0$, then with probability at least $1-\delta$ we have that $\lvert\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]-\operatorname*{\mathbf{Pr}}_{D}[S]\rvert\leq\operatorname*{\mathbf{Pr}}_{D}[S]/2$ for all $S\in\mathcal{S}$.
+Hence, $\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\geq\operatorname*{\mathbf{Pr}}_{D}[S]/2$, i.e., the number of samples that fall in each set $S\in\mathcal{S}$ is at least $N\operatorname*{\mathbf{Pr}}_{D}[S]/2=N(\epsilon^{\prime}/k)^{\Omega(k)}$.
+Therefore, if $N\geq(dm)^{Cm}(k/\epsilon^{\prime})^{Ck}\log(1/\delta)(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{C}$ for a sufficiently large universal constant $C>0$, then by applying Claim 4.11 we have that the following: with probability $1-\delta$, for any unit vector $\mathbf{v}\in W^{\perp}$
+$\mathbf{v}^{\top}\widehat{\mathbf{U}}\mathbf{v}\leq\sum_{S\in\mathcal{S},\mathbf{u}\in U_{S}}(\mathbf{v}\cdot\mathbf{u})^{2}\operatorname*{\mathbf{Pr}}_{\widehat{D}}[S]\leq\eta^{2}\mathrm{poly}(mK\lvert\mathcal{Y}\rvert/\sigma)\;,$
 where we used the fact that $\lvert U_{S}\rvert=\mathrm{poly}(mK\lvert\mathcal{Y}\rvert/\sigma)$ by Claim 3.9.
@@
 As a result, we have that if $\eta$ is chosen to be a sufficiently small polynomial of $\epsilon,\sigma,\alpha,1/\Gamma,1/m,1/K,1/\lvert\mathcal{Y}\rvert$, then for all $\mathbf{u}^{(i)}\in\mathcal{E}$ and any $\mathbf{v}\in W^{\perp}$ we have that $\lvert\mathbf{u}^{(i)}\cdot\mathbf{v}\rvert\leq\epsilon$.
-Moreover, from Proposition 3.7 we have that if $N\geq(dm)^{Cm}{(k/\epsilon^{\prime})^{Ck}}\log(1/\delta){(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{C}}$ for a sufficiently large universal constant $C>0$, then there exists a $\mathbf{v}\in\mathcal{E}$.
+Moreover, by Proposition 3.7 we have that if $N\geq(dm)^{Cm}(k/\epsilon^{\prime})^{Ck}\log(1/\delta)(K\Gamma\lvert\mathcal{Y}\rvert/(\sigma\epsilon\alpha))^{C}$ for a sufficiently large universal constant $C>0$, then there exists a $\mathbf{v}\in\mathcal{E}$.
 Moreover, as $\mathbf{v}$ is a unit vector and $\left\|\mathbf{v}^{\perp W}\right\|\leq\epsilon$, we have that $\left\|\mathbf{v}^{W}\right\|\geq\sqrt{1-\epsilon^{2}}\geq 1-\epsilon$ which proves Proposition 4.10.
 ∎
-Now we prove our main result for learning with RCN.
-Our proof similar to the proof of Theorem 3.3.
-The distinction is that we are going to use Proposition 4.10 which provides improved correlation when compared to Proposition 3.7.
+Now we are ready to prove Theorem 4.9.
+Our proof is similar to the proof of Theorem 3.3.
+The difference is that we are going to use Proposition 4.10 which provides improved correlation when compared to Proposition 3.7.
 Proof of Theorem 4.9.
@@
 Let $f^{*}$ be an optimal classifier and denote by $W^{*}$ the subspace of its weight vectors.
-Let $L_{t}$ be the list of vectors updated by the algorithm (Line 3d) and $V_{t}=\mathrm{span}(L_{t})$, $\dim(V_{t})=k_{t}$.
+Let $L_{t}$ be the set of vectors maintained by the algorithm (Line 3d) and $V_{t}=\mathrm{span}(L_{t})$, $\dim(V_{t})=k_{t}$.
 Denote by $T=2K$ the total number of iterations.
@@
 Assume that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]>\tau+\mathrm{OPT}+\epsilon$ for all $t=1,\dots,T$.
-Using the fact that $N={(dm)^{Cm}2^{K^{C}}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{CK}\log(1/\delta)/(\sigma\alpha)^{C}}$ for a sufficiently large universal constant $C>0$ (Line 2 of Algorithm 5), we can apply Proposition 4.10 and conclude that, with probability $1-\delta$, there exist unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ for $t\in[T]$ such that $\mathbf{w}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq 1-\epsilon\geq 1/2$.
+Using the fact that $N=(dm)^{Cm}2^{K^{C}}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{CK}\log(1/\delta)/(\sigma\alpha)^{C}$ for a sufficiently large universal constant $C>0$ (Line 2 of Algorithm 5), we can apply Proposition 4.10 and conclude that, with probability $1-\delta$, there exist unit vectors $\mathbf{v}^{(t)}\in V_{t+1}$ and unit vectors $\mathbf{w}^{(t)}\in W^{*}$ for $t\in[T]$ such that $\mathbf{w}^{(t)}\cdot(\mathbf{v}^{(t)})^{\perp V_{t}}\geq 1-\epsilon\geq 1/2$.
 Thus, from Claim 2.16, we have that with probability $1-\delta$, for all $t\in[T]$, $\Phi_{t}\leq\Phi_{t-1}-1/2$.
@@
 Hence, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h_{t}(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+\epsilon$ for some $t\in\{1,\dots,T\}$.
-Since the error of $h_{t}$ can only be decreasing by Claim 2.17, and $h_{t}$ is close to its sample variant by Claim 2.18, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+2\epsilon$.
+Since the error of $h_{t}$ can only be decreasing (see Claim 2.17), and $h_{t}$ is close to its sample variant by Claim 2.18, we have that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}[h(\mathbf{x})\neq y]\leq\tau+\mathrm{OPT}+2\epsilon$.
 Sample and Computational Complexity: Note that the algorithm terminates in $O(K)$ iterations and at each iteration we draw $N=(dm)^{O(m)}2^{\mathrm{poly}(K)}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{O(K)}\log(1/\delta)/(\sigma\alpha)^{O(1)}$ samples.
 Hence, the total sample size is $(dm)^{O(m)}2^{\mathrm{poly}(K)}(\Gamma\lvert\mathcal{Y}\rvert/\epsilon)^{O(K)}\log(1/\delta)/(\sigma\alpha)^{O(1)}$.
-Moreover we use at most $\mathrm{poly}(N,d,\lvert\mathcal{Y}\rvert)$ time as all operations can be implemented in polynomial time in $N,d$ and $\lvert\mathcal{Y}\rvert$.
+Moreover we use at most $\mathrm{poly}(N)$ time as all operations can be implemented in polynomial time in $N,d$ and $\lvert\mathcal{Y}\rvert$.
 ∎
 5 Statistical Query Lower Bounds
-Here we prove Theorems 1.5 and 1.7.
-First, in Section 5.1 we present SQ lower bounds about general classes of Multi-Index Models that do not satisfy Definition 1.2.
+In this section we prove Theorems 1.6 and 1.9.
+First, in Section 5.1 we present SQ lower bounds for general classes of Multi-Index Models that do not satisfy Definition 1.2.
 Essentially showing that the algorithm of Section 3 is quantitatively optimal in the SQ model.
 Second, in Section 5.2, we present an SQ lower bound for the problem of learning multiclass linear classifiers under RCN.
+Before presenting our hardness results, we first provide a short introduction to the SQ model.
+Basics on SQ Model
+SQ algorithms are a broad class of algorithms that, instead of having direct access to samples, are allowed to query expectations of bounded functions of the distribution.
+Definition 5.1 (SQ algorithms).
+Let $D$ be a distribution on $\mathcal{X}$.
+A statistical query is a bounded function $q:\mathcal{X}\to[-1,1]$.
+For $u>0$, the $\mathrm{VSTAT}(u)$ oracle responds to the query $q$ with a value $v$ such that $|v-\operatorname*{\mathbf{E}}_{\mathbf{x}\sim D}[q(\mathbf{x})]|\leq\tau$, where $\tau=\max(1/u,\sqrt{\mathrm{Var}_{\mathbf{x}\sim D}[q(\mathbf{x})]/u})$.
+We call $\tau$ the tolerance of the statistical query.
+A Statistical Query algorithm is an algorithm whose objective is to learn some information about an unknown distribution $D$ by making adaptive calls to the corresponding oracle.
+The SQ model was introduced in [Kea98] as a natural restriction of the PAC model [Val84].
+Subsequently, the model has been extensively studied in a range of contexts, see, e.g., [Fel16].
+The class of SQ algorithms is broad and captures a range of known supervised learning algorithms.
+More broadly, several known algorithmic techniques in machine learning are known to be implementable using SQs (see, e.g., [FGR+13, FGV17]).
+Throughout this section, we leverage a number of previously known facts about the SQ model that allow us to establish hardness. We refer the reader to Appendix C for their formal statements.
 5.1 SQ Lower Bounds for Learning Multi-Index Models
-Here we show that if a Multi-Index Model class $\mathcal{F}$ does not satisfy Definition 3.2, any SQ algorithm that learns $\mathcal{F}$ to an accuracy comparable to that of Theorem 3.3 requires super-polynomial complexity.
+We show that if a Multi-Index Model class $\mathcal{F}$ does not satisfy Definition 3.2, any SQ algorithm that learns $\mathcal{F}$ to an accuracy comparable to that of Theorem 3.3 requires super-polynomial complexity.
 Specifically, we establish the following theorem:
-Theorem 5.1.
-(SQ Optimality of General Algorithm) Let $d,K,m\in\mathbb{Z}_{+}$ with $d^{c}\geq\max(m,K)$ for a sufficiently small absolute constant $c>0$, and let $\mathcal{Y}$ be a set of finite cardinality.
+Theorem 5.2 (SQ Optimality of General Algorithm).
+Let $d,K,m\in\mathbb{Z}_{+}$ with $d^{c}\geq\max(m,K)$ for a sufficiently small absolute constant $c>0$, and let $\mathcal{Y}$ be a set of finite cardinality.
 Additionally, let $\tau,\mathrm{OPT}\in(0,1)$ with $\tau,\mathrm{OPT}\geq d^{-O(m)}$.
-Assume the existence of (i) a function $f:\mathbb{R}^{d}\mapsto\mathcal{Y}$ (ii) a random variable $y$ over $\mathcal{Y}$ such that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]=\mathrm{OPT}$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$, (iii) a subspace $W$ of dimension at most $K$, and iv) a subspace $V$ such that:
+Assume the existence of (i) a function $f:\mathbb{R}^{d}\mapsto\mathcal{Y}$ (ii) a random variable $(\mathbf{x},y)$ over $\mathbb{R}^{d}\times\mathcal{Y}$ with $\mathbf{x}$-marginal the standard Gaussian such that $\operatorname*{\mathbf{Pr}}[f(\mathbf{x})\neq y]=\mathrm{OPT}$ (iii) a subspace $W$ of dimension at most $K$, and iv) a subspace $V$ such that:
   1. 1.
@@
     For any $\mathbf{x}_{0}\in\mathbb{R}^{d},i\in[K]$, and for any zero-mean polynomial $p:U\to\mathbb{R}$ of degree at most $m$, it holds that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[p(\mathbf{x}^{U})\mathds{1}(y=i)\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]=0$, where $U=(V+W)\cap V^{\perp}$.
-Let $\mathcal{F}$ denote the class of functions defined as =$\mathcal{F}=\{f(\mathbf{x}^{\perp U},\mathbf{U}\mathbf{x}^{U}):\mathrm{for\ all\ }\mathbf{U}\in\mathbb{R}^{k\times d}:\mathbf{U}\mathbf{U}^{\top}=\mathbf{I}_{k}\}$, where $k=\dim(U)$.
-Then, any SQ algorithm that learns $\mathcal{F}$ up to an error of smaller than $\tau-2\mathrm{OPT}$ with probability at least $2/3$ requires either $2^{d^{\Omega(1)}}$ queries or at least one query to $\text{VSTAT}\left(d^{\Omega(m)}/|\mathcal{Y}|\right)$.
+Let $\mathcal{F}$ denote the class of MIMs defined as =$\mathcal{F}=\{f(\mathbf{x}^{\perp U},\mathbf{U}\mathbf{x}^{U}):\mathrm{for\ all\ }\mathbf{U}\in\mathbb{R}^{k\times d}:\mathbf{U}\mathbf{U}^{\top}=\mathbf{I}_{k}\}$, where $k=\dim(U)$.
+Then, any SQ algorithm that learns $\mathcal{F}$ in the presence of $\mathrm{OPT}$ adversarial label noise up to an error of smaller than $\tau-2\mathrm{OPT}$ with probability at least $2/3$ requires either $2^{d^{\Omega(1)}}$ queries or at least one query to $\text{VSTAT}\left(d^{\Omega(m)}/|\mathcal{Y}|\right)$.
 Proof.
@@
 For each $\mathbf{Q}\in\mathcal{Q}$, we define the function $f_{\mathbf{Q}}(\mathbf{x}):=f(\mathbf{x}^{\perp U},\mathbf{Q}\mathbf{x}^{U})$ along with the coupled random variables $y_{\mathbf{Q}}$ satisfying $\operatorname*{\mathbf{Pr}}[y_{\mathbf{Q}}=i|\mathbf{x}=\mathbf{z}]=\operatorname*{\mathbf{Pr}}[y=i|\mathbf{x}=\mathbf{z}^{\perp U}+\mathbf{Q}\,\mathbf{z}^{U}]$, for all $i\in\mathcal{Y},\mathbf{z}\in\mathbb{R}^{d}$.
-We consider the following class of functions $\mathcal{F}_{\mathcal{Q}}=\{f_{\mathbf{Q}}(\mathbf{x}):\mathbf{Q}\in\mathcal{Q}\}$ and note that by definition $\mathcal{F}_{\mathcal{Q}}\subseteq\mathcal{F}$.
+We consider the class of functions $\mathcal{F}_{\mathcal{Q}}=\{f_{\mathbf{Q}}(\mathbf{x}):\mathbf{Q}\in\mathcal{Q}\}$ and note that by definition $\mathcal{F}_{\mathcal{Q}}\subseteq\mathcal{F}$.
 For each $\mathbf{Q}\in\mathcal{Q}$, let $D_{\mathbf{Q}}$ be the joint distribution of $(\mathbf{x},y_{\mathbf{Q}})$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
 Denote by $\mathcal{D}=\{D_{\mathbf{Q}}:\mathbf{Q}\in\mathcal{Q}\}$ the class of these distributions.
-Denote by $D_{0}$ the distribution of pairs $(\mathbf{x},y_{0})$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $y_{0}$ a random variable supported on $\mathcal{Y}$ that depends only on $W^{V}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{0}}[y_{0}=i\mid\mathbf{x}^{\perp U}=\mathbf{z}^{\perp U}]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[y=i\mid\mathbf{x}^{\perp U}=\mathbf{z}^{\perp U}]$, for all $i\in\mathcal{Y},\mathbf{z}\in\mathbb{R}^{d}$.
+Denote by $D_{0}$ the distribution of pairs $(\mathbf{x},y_{0})$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $y_{0}$ a random variable supported on $\mathcal{Y}$ that depends only on $W^{V}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y_{0})\sim D_{0}}[y_{0}=i\mid\mathbf{x}^{\perp U}=\mathbf{z}^{\perp U}]=\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[y=i\mid\mathbf{x}^{\perp U}=\mathbf{z}^{\perp U}]$, for all $i\in\mathcal{Y},\mathbf{z}\in\mathbb{R}^{d}$.
 Denote by $y_{i}(\mathbf{x}^{\perp U})=\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{0}}[y_{0}=i\mid\mathbf{x}^{\perp U}]$.
-We first show that for all $D_{\mathbf{P}},D_{\mathbf{Q}}\in\mathcal{D}$ with $\mathbf{P}\neq\mathbf{Q}$, the correlation $\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{Q}})$ is small.
+We show that is SQ-Hard to determinate whether we have access to distribution $D_{0}$ or to a distribution $D\in\mathcal{D}$.
+To this end, we show that for all $D_{\mathbf{P}},D_{\mathbf{Q}}\in\mathcal{D}$ with $\mathbf{P}\neq\mathbf{Q}$, the correlation $\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{Q}})$ (see Definition C.1) is small.
 Using the law of total expectation, we have that
@@
 $=\sum_{i\in\mathcal{Y}}\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp U}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\bigg{[}\frac{1}{y_{i}(\mathbf{x}^{\perp U})}\operatorname*{\mathbf{E}}_{\mathbf{x}^{U}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[\mathds{1}(y_{\mathbf{P}}=i)-y_{i}(\mathbf{x}^{\perp U}))(\mathds{1}(y_{\mathbf{Q}}=i)-y_{i}(\mathbf{x}^{\perp U}))\mid\mathbf{x}^{\perp U}]\bigg{]}\;.$
-Considering each term separately, by C.5, we have that
+Considering each term separately, by C.5 and Assumption (3) of the theorem statement, we have that
 $\operatorname*{\mathbf{E}}_{\mathbf{x}^{U}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[(\mathds{1}(y_{\mathbf{P}}=i)-$ $y_{i}(\mathbf{x}^{\perp U}))(\mathds{1}(y_{\mathbf{Q}}=i)-y_{i}(\mathbf{x}^{\perp U}))\mid\mathbf{x}^{\perp U}]$
-$\leq\sum_{t=m}^{\infty}\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{t}\operatorname*{\mathbf{E}}[\mathds{1}(y=i)^{2}\mid\mathbf{x}^{\perp U}]\leq\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{m}\operatorname*{\mathbf{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp U}]\;,$
-where we used Parseval’s identity and Assumption (3) of the theorem statement. Therefore, we have that
-$\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{Q}})$ $\leq\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{m}\sum_{i\in\mathcal{Y}}\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp U}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\frac{\operatorname*{\mathbf{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp U}]}{y_{i}(\mathbf{x}^{\perp U})}\right]\leq\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{m}|\mathcal{Y}|\;.$
+$\leq\sum_{t=m}^{\infty}\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{t}(y^{[t]})^{2}\leq\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{m}\sum_{t=m}^{\infty}(y^{[t]})^{2}\;,$
+where $y^{[t]}$ is the degree-$t$ Hermite part of the random variable $\mathds{1}(y=i)\mid\mathbf{x}^{\perp U}$ and that $\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}\leq 1$.
+Furthermore, using Parseval’s identity we get that $\sum_{t=m}^{\infty}(y^{[t]})^{2}\leq\operatorname*{\mathbf{E}}[\mathds{1}(y=i)^{2}\mid\mathbf{x}^{\perp U}]=\operatorname*{\mathbf{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp U}]$.
+Therefore, we have that
+$\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{Q}})$ $\leq\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{m}\sum_{i\in\mathcal{Y}}\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp U}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\frac{\operatorname*{\mathbf{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp U}]}{y_{i}(\mathbf{x}^{\perp U})}\right]=\|\mathbf{P}\mathbf{Q}^{\top}\|_{2}^{m}|\mathcal{Y}|\;,$
+where we used that $\operatorname*{\mathbf{E}}_{\mathbf{x}^{\perp U}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[({\operatorname*{\mathbf{E}}[\mathds{1}(y=i)\mid\mathbf{x}^{\perp U}]})/{y_{i}(\mathbf{x}^{\perp U})}\right]=1$, by the definition of $y_{i}(\mathbf{x}^{\perp U})$.
 Hence, $\chi_{D_{0}}(D_{\mathbf{P}},D_{\mathbf{Q}})=d^{-\Omega(m)}\lvert\mathcal{Y}\rvert$.
@@
 Proof.
-Assume the existence of an SQ algorithm $\mathcal{A}$ that given access to a distribution $D^{\prime}$ over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, returns a hypothesis $h$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D^{\prime}}[h(\mathbf{x})\neq y]<\tau-2\mathrm{OPT}$, with $\mathrm{OPT}=\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D^{\prime}}[f(\mathbf{x})\neq y]$.
-Using $\mathcal{A}$, we can reduce the testing problem to the learning problem.
-Note that by running algorithm $\mathcal{A}$ to distribution $D_{\mathbf{P}}\in\mathcal{D}$, $\mathcal{A}$ returns a hypothesis $h$ with error $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{\mathbf{P}}}[h(\mathbf{x})\neq y]<\tau-2\mathrm{OPT}$.
-However, running a $\mathcal{A}$ with SQ oracle access to $D_{0}$ would result in a hypothesis $h$ that cannot perform better than $\tau-\mathrm{OPT}$, as $D_{0}$ is a product distribution for each fixed projection of $\mathbf{x}$ in $V$.
-As $\tau-\mathrm{OPT}>\tau-2\mathrm{OPT}$, we have that with an additional query of accuracy less than $\mathrm{OPT}$, we can solve $\mathcal{B}(\mathcal{D},D_{0})$.
+Assume there exists an SQ algorithm $\mathcal{A}$ that, when given SQ access to any distribution $D^{\prime}$ over $\mathbb{R}^{d}\times\mathcal{Y}$ with standard Gaussian $\mathbf{x}$-marginals, outputs a hypothesis $h$ satisfying $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D^{\prime}}[h(\mathbf{x})\neq y]<\tau-2\mathrm{OPT}$, where $\mathrm{OPT}=\inf_{f\in\mathcal{F}}\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D^{\prime}}[f(\mathbf{x})\neq y]$.
+Using such an algorithm $\mathcal{A}$, we show that we can solve the testing problem.
+Suppose we run $\mathcal{A}$ on a distribution $D_{\mathbf{P}}\in\mathcal{D}$.
+Then, by definition, $\mathcal{A}$ returns a hypothesis $h$ with error $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{\mathbf{P}}}[h(\mathbf{x})\neq y]<\tau-2\mathrm{OPT}$.
+In contrast, when we run $\mathcal{A}$ with SQ oracle access to $D_{0}$ no hypothesis can achieve error lower than $\tau-\mathrm{OPT}$, as $D_{0}$ is a product distribution for each fixed projection of $\mathbf{x}$ in $V$.
+Since $\tau-\mathrm{OPT}>\tau-2\mathrm{OPT}$, an additional query of tolerance less than $\mathrm{OPT}$, would distinguish between $D_{\mathbf{P}}$ and $D_{0}$, thus solving $\mathcal{B}(\mathcal{D},D_{0})$.
 ∎
 5.2 SQ Lower Bounds for Learning Multiclass Linear Classifiers with RCN
-In this section, we present SQ hardness results for learning Multiclass Linear Classifiers in the presence of RCN under the Gaussian distribution.
-In particular, we show that achieving $\mathrm{OPT}+\epsilon$ error in the SQ model requires, either a super-polynomial number of queries, or queries with accuracy that depends on the spectral properties of the confusion matrix (see Definition 4.8), rather than on the amount of noise.
+In this section, we present an SQ hardness result for learning Multiclass Linear Classifiers in the presence of RCN under the Gaussian distribution.
+In particular, we show that achieving $0-1$ error of $\mathrm{OPT}+\epsilon$ in the SQ model must either make a super-polynomial (with respect the dimension $d$) number of queries or queries with accuracy that depends on the spectral properties of the confusion matrix (see Definition 4.8), rather than on the amount of noise.
 The main result of this section is Theorem 1.9.
-Theorem 5.3 (SQ Lower Bound for Learning under RCN).
+Theorem 5.4 (SQ Lower Bound for Learning under MLC RCN).
 Any SQ algorithm that learns $K$-Multiclass Linear Classifiers under the $d$-dimensional standard normal in the presence of RCN to error $\mathrm{OPT}+\mathrm{poly}(1/K)$, requires either $2^{d^{\Omega(1)}}$ queries or at least one query to $\text{VSTAT}\left(d^{\Omega(K)}\right)$.
 SQ-Hard Instances
-To construct our family of hard instances, we first define a two-dimensional model $f:\mathbb{R}^{2}\to[K]$ as follows:
+To construct our family of hard instances, we first define a two-dimensional classifier $f:\mathbb{R}^{2}\to[K]$ as follows:
 $\mathbf{w}^{(k)}=\begin{bmatrix}\cos\left(\frac{2\pi k}{K}\right)\\ \sin\left(\frac{2\pi k}{K}\right)\\ \end{bmatrix},k\in[K],\text{ and }f(\mathbf{x})=\operatorname*{argmax}_{k\in[K]}(\mathbf{w}^{(k)}\cdot\mathbf{x})\;.$ (6)
 Geometrically, the classifier $f$ partitions the plane into $K$ equal angular sectors, assigning each sector to a different class.
-Now, given $\mathcal{Q}$, a set of $2\times d$ matrices, we can define a class of functions from $\mathbb{R}^{d}$ to $[K]$ as
+Given a set of $2\times d$ matrices $\mathcal{Q}$, we can define a class of functions from $\mathbb{R}^{d}$ to $[K]$ as follows
 $\mathcal{F}_{\mathcal{Q}}=\{f(\mathbf{Q}\mathbf{x}):\mathbf{Q}\in\mathcal{Q}\}\;.$ (7)
+We show that it is hard to distinguish a distribution labeled by a classifier in $\mathcal{F}_{\mathcal{Q}}$ corrupted with RCN from a distribution with random labels independent of $\mathbf{x}$.
 Before proving Theorem 5.4, we first need to show that we can match the appropriate number of moments with the standard Gaussian.
 Lemma 5.5 (Moment matching under RCN).
-Let $K\in\mathbb{Z}_{+}$ with $K\geq 8$ such that $K$ is divisible by $4$. There exists a $K\times K$ doubly stochastic matrix $\mathbf{H}$ such that the following hold:
+Let $K\in\mathbb{Z}_{+}$ be such that $K\geq 8$ and $K$ is divisible by $4$.
+Then there exists a $K\times K$ doubly stochastic matrix $\mathbf{H}$ and a distribution $D$ over $\mathbb{R}^{2}\times[K]$, whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$, satisfying Definition 4.8 for the confusion matrix $\mathbf{H}$ and the function $f$ defined in Equation 6, such that:
   1. i)
@@
     For every polynomial $p:\mathbb{R}^{2}\to\mathbb{R}$ of degree at most $K/2-2$, it holds that
-    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}\left[y=i\mid\mathbf{x}\right]-{1}/{K}\right)p(\mathbf{x})\right]=0$ for all $i\in[K]$ ,
-where $D$ is the distribution over $\mathbb{R}^{2}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ that satisfies Definition 4.8 for the confusion matrix $\mathbf{H}$ and the function $f$ as defined in Equation 6.
+    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D}\left[y=i\mid\mathbf{x}\right]-{1}/{K}\right)p(\mathbf{x})\right]=0$ for all $i\in[K]$
 Proof.
@@
 Also denote by $\boldsymbol{\omega}^{(k)}$ the root of unity vector of rate $k$, i.e.
-${\boldsymbol{\omega}}^{(k)}\in\mathbb{C}^{K}$ such that $\boldsymbol{\omega}_{j}^{(k)}=e^{\frac{2\pi kji}{K}}$.
-First, we show that a function $p$ has zero correlation with $\tilde{f}_{i}-1/K$ for all $i\in[K]$ under the standard Gaussian if and only if the vector of expectations of $p$ over the level sets belongs in the kernel of $\mathbf{H}^{\top}$.
+${\boldsymbol{\omega}}^{(k)}\in\mathbb{C}^{K}$ such that $\boldsymbol{\omega}_{j}^{(k)}=e^{\frac{2\pi kji}{K}}$, where $i$ is the imaginary unit.
+We start by showing that a function $p$ has zero correlation with $\tilde{f}_{i}-1/K$ for all $i\in[K]$ under the standard Gaussian if and only if the vector of expectations of $p$ over the level sets belongs in the kernel of $\mathbf{H}^{\top}$.
 Claim 5.6.
-Let $p:\mathbb{R}^{2}\to\mathbb{R}$, then we have that $\mathbf{v}^{(p)}\in\ker(\mathbf{H}^{\top})\text{ if and only if }\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}[(\tilde{f}_{i}(\mathbf{x})-1/K)p(\mathbf{x})]=0\text{ for all $i\in[K]$}\;.$
+Let $p:\mathbb{R}^{2}\to\mathbb{R}$.
+We have that $\mathbf{v}^{(p)}\in\ker(\mathbf{H}^{\top})$ if and only if for all $i\in[K]$ it holds $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}[(\tilde{f}_{i}(\mathbf{x})-1/K)p(\mathbf{x})]=0$.
 Proof of Claim 5.6.
-First note that from the rotational invariance of the standard Gaussian distribution we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}^{2}}[f(\mathbf{x})=i]=1/K$, for all $i\in[K]$.
-Hence, we have that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}[\tilde{f}_{i}(\mathbf{x})]=\sum_{j=1}^{K}\mathbf{H}_{ji}\operatorname*{\mathbf{E}}[\mathds{1}(f(\mathbf{x})=i)]=1/K$, from the assumption that $\mathbf{H}$ is column stochastic.
+By the rotational invariance of the standard Gaussian we have that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}^{2}}[f(\mathbf{x})=i]=1/K$, for all $i\in[K]$.
+From the assumption that $\mathbf{H}$ is column stochastic, we have that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}[\tilde{f}_{i}(\mathbf{x})]=\sum_{j=1}^{K}\mathbf{H}_{ji}\operatorname*{\mathbf{E}}[\mathds{1}(f(\mathbf{x})=i)]=1/K$.
 Thus, for all $i\in[K]$ we have that
@@
 $=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\tilde{f}_{i}(\mathbf{x})\bar{p}(\mathbf{x})\right]\;.$
-Therefore, substituting the definition of $\tilde{f}_{i},i\in[K],$ we have that
+Therefore, substituting the definition of $\tilde{f}_{i}$ we have that
 $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\tilde{f}_{i}(\mathbf{x})\bar{p}(\mathbf{x})\right]=\sum_{j=1}^{K}\mathbf{H}_{j,i}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\bar{p}(\mathbf{x})\mathds{1}(f(\mathbf{x})=j)\right]=(\mathbf{H}^{\top})_{i}\mathbf{v}^{(p)}\;,$
 which concludes the proof. ∎
-We next note that that all polynomials of degree at most $m$ can be expressed as a linear combination of polynomials of the form $(x+iy)^{a}(x-iy)^{b}$, where $a,b>0,a+b\leq m$.
-In the following claim, we show that the vectors associated with each polynomial of this basis are precisely the root of unity vectors.
+We next observe that all polynomials $\mathbb{R}^{2}\mapsto\mathbb{R}$ of degree at most $m$ can be expressed as a linear combination of polynomials of the form $(x+iy)^{a}(x-iy)^{b}$, where $a,b>0,a+b\leq m$ and $i$ is the imaginary unit.
+Specifically, we show that the vectors associated with each polynomial of this basis are precisely the root of unity vectors.
 Claim 5.7.
-For $p(x,y)=(x+iy)^{a}(x-iy)^{b},a,b\in\mathbb{Z}_{+}$.
+Let $p(x,y)=(x+iy)^{a}(x-iy)^{b},a,b\in\mathbb{Z}_{+}$.
 If $a\neq b$, then there exists a constant $c\in\mathbb{C}$ such that $\mathbf{v}^{(p)}=c\boldsymbol{\omega}^{(a-b)}$; and if $a=b$, then $\mathbf{v}=\mathbf{0}$.
 Proof of Claim 5.7.
-First, note that by the definition of $f$ the region $\mathds{1}(f(x,y)=j)$ is equivalent to $\theta\in\left[{2\pi j}/{K}-{\pi}/{K},{2\pi j}/{K}+{\pi}/{K}\right]$, where $\theta=\arctan\left({y}/{x}\right)$.
-Fix $a,b\in\mathbb{Z}_{+},a\neq b$, and let $\theta_{1,j}={2\pi j}/{K}-{\pi}/{K}$, $\theta_{2,j}={2\pi j}/{K}+{\pi}/{K}$, for $j\in[K]$, and $p(x,y)=(x+iy)^{a}(x+iy)^{b}$.
+By the definition of $f$, the region $\{(x,y)\in\mathbb{R}^{2}:f(x,y)=j\}$ is equivalent to $\theta\in\left[{2\pi j}/{K}-{\pi}/{K},{2\pi j}/{K}+{\pi}/{K}\right]$, where $\theta=\arctan\left({y}/{x}\right)$.
+Fix $a,b\in\mathbb{Z}_{+},a\neq b$, and let $\theta_{1,j}={2\pi j}/{K}-{\pi}/{K}$, $\theta_{2,j}={2\pi j}/{K}+{\pi}/{K}$, $j\in[K]$.
 We have that
@@
 […]
-Therefore, note that $\mathbf{v}_{(j+1)\mod K}^{(p)}=\mathbf{v}_{j}^{(p)}e^{\frac{2\pi i}{K}(a-b)}$ for all $j\in\{0,\dots,K-1\}$, since $\theta_{1,(j+1)}=\theta_{1,j}+{2\pi}/{K}$ and $\theta_{2,(j+1)}=\theta_{2,j}+{2\pi}/{K}$.
-Moreover, if $a\neq b$, we observe that the integral of $p(x,y)$ over the entire region is zero, which concludes the proof.
+Therefore, it holds that $\mathbf{v}_{(j+1)\mod K}^{(p)}=\mathbf{v}_{j}^{(p)}e^{\frac{2\pi i}{K}(a-b)}$ for all $j\in\{0,\dots,K-1\}$, since $\theta_{1,(j+1)}=\theta_{1,j}+{2\pi}/{K}$ and $\theta_{2,(j+1)}=\theta_{2,j}+{2\pi}/{K}$.
+Moreover, if $a\neq b$, we observe that the integral of $p(x,y)$ over $\mathbb{R}^{2}$ is zero, which concludes the proof of the claim.
 ∎
-Hence, by Claim 5.6, in order to have 00 correlation with all polynomials of degree at most $m$, we need to have that $\boldsymbol{\omega}^{(k)}\in\ker(\mathbf{H}^{\top})$ for all $k\in\{1,\dots,m,K-m,\dots,K-1\}$, as the roots of unity are symmetric, i.e., $\boldsymbol{\omega}^{(-k)}=\boldsymbol{\omega}^{(K-k)}$.
+By Claim 5.7, in order to have 00 correlation with all polynomials of degree at most $m$, we need to have that $\boldsymbol{\omega}^{(k)}\in\ker(\mathbf{H}^{\top})$ for all $k\in\{1,\dots,m,K-m,\dots,K-1\}$, as the roots of unity are symmetric, i.e., $\boldsymbol{\omega}^{(-k)}=\boldsymbol{\omega}^{(K-k)}$.
 We first construct a stochastic vector $\mathbf{h}\in\mathbb{R}^{K}$ with $\mathbf{h}_{0}-\mathbf{h}_{j}=\Omega(1/K^{3})$, for all $j\in\{1,\dots,K-1\}$ that is also orthogonal to the aforementioned root of unity vectors for $m=\Omega(K)$.
@@
 Hence, we have that all the coordinates of $\mathbf{h}^{\prime}$ are nonnegative.
-Moreover, $\mathbf{h}^{\prime}$ attains its maximum value at $\mathbf{h}_{0}^{\prime}={4}/{K}$, since $\mathbf{h}_{K/2}^{\prime}=0$, because $(-1)^{j}\cos\left({2\pi j}/{K}\right)=-1$ when $j=K/2$ and $K$ is divisible by $4$.
-Furthermore, it holds that $\max_{j\in[K-1]}\mathbf{h}_{j}^{\prime}\leq 2/K+\max_{j\in[K-1]}\lvert\mathbf{h}_{j}^{(2)}\rvert=2/K+2/K\cos\left({2\pi}/{K}\right)$.
-As a result
+Moreover, $\mathbf{h}^{\prime}$ attains its maximum value at $\mathbf{h}_{0}^{\prime}={4}/{K}$, since $\mathbf{h}_{K/2}^{\prime}=0$ (because $(-1)^{j}\cos\left({2\pi j}/{K}\right)=-1$ when $j=K/2$ and $K$ is divisible by $4$).
+Furthermore, it holds that $\max_{j\in[K-1]}\mathbf{h}_{j}^{\prime}\leq 2/K+\max_{j\in[K-1]}\lvert\mathbf{h}_{j}^{(2)}\rvert=2/K+2/K\cos\left({2\pi}/{K}\right)$, and hence we have
 $\mathbf{h}^{\prime}_{0}-\max_{j\in[K-1]}\mathbf{h}_{j}^{\prime}\geq\frac{2}{K}\left(1-\cos\left(\frac{2\pi}{K}\right)\right)\geq\frac{2}{K}\left(\left(\frac{2\pi}{K}\right)^{2}/2-\left(\frac{2\pi}{K}\right)^{4}/24\right)\gtrsim\frac{1}{K^{3}}\;,$
-where the second inequality follows from the Taylor series $\cos(x)=1-\frac{x^{2}}{2!}+\frac{x^{4}}{4!}-\frac{x^{6}}{6!}+\cdots$, which is an alternating decreasing series when $x={2\pi}/{K}\leq 1$, thus it can be upper bounded by the first three terms.
+where the second inequality follows from the Taylor series of $\cos(x)=1-\frac{x^{2}}{2!}+\frac{x^{4}}{4!}-\frac{x^{6}}{6!}+\cdots$, which is an alternating decreasing series when $x={2\pi}/{K}\leq 1$ (thus it can be upper bounded by the first three terms).
 Let $\mathbf{h}$ be the normalization of $\mathbf{h}^{\prime}$ such that the sum of its coordinates equals $1$.
@@
 Now consider $\mathbf{H}$, the circulant matrix generated by row shifts of the vector $h$.
-By the definition of $\mathbf{H}$, we have that $\mathbf{H}$ is a row stochastic matrix and that $\min_{i\neq j}\mathbf{H}_{ii}-\mathbf{H}_{i,j}=\Omega(1/K^{3})$.
+By definition, we have that $\mathbf{H}$ is a row stochastic matrix and that $\min_{i\neq j}\mathbf{H}_{ii}-\mathbf{H}_{i,j}=\Omega(1/K^{3})$.
 Moreover, $\mathbf{H}$ is symmetric, because it holds that $\mathbf{H}_{i,j}=\mathbf{h}_{j-i\mod K}=\mathbf{h}_{i-j\mod K}=\mathbf{H}_{j,i}$ and $\mathbf{h}_{j}=\mathbf{h}_{K-j},j\in[K-1]$.
@@
 Hence $\mathbf{H}$ is also column stochastic.
-Finally, by A.6 we have that the eigenvectors of any circulant matrix are exactly $\boldsymbol{\omega}^{(k)}$ with eigenvalues $\mathbf{h}\cdot\boldsymbol{\omega}^{(k)},k\in\{0,\dots,K-1\}$.
+By A.6, the eigenvectors of any circulant matrix are exactly $\boldsymbol{\omega}^{(k)}$ with eigenvalues $\mathbf{h}\cdot\boldsymbol{\omega}^{(k)},k\in\{0,\dots,K-1\}$.
 Hence, $\boldsymbol{\omega}^{(k)}\in\ker(\mathbf{H})=\ker(\mathbf{H}^{\top})$ for all $k\in\{1,\dots,K/2-2,K/2,K/2+2,\dots,K-1\}$.
@@
 Proof of Theorem 5.4.
-Consider the case where $K$ is divisible by $4$ and let $\mathbf{H}$ be a matrix that satisfies the statement of Lemma 5.4 Let $f:\mathbb{R}^{2}\to[K]$ be the linear classifier defined in Equation 6, and let $\mathcal{F}_{\mathcal{Q}}$ be the class of function defined in Equation 7, where $\mathcal{Q}\subseteq\mathbb{R}^{d\times 2}$ the family that satisfies the statement of C.4.
+Consider the case where $K$ is divisible by $4$ and let $\mathbf{H}$ be a matrix that satisfies the statement of Lemma 5.5 Let $f:\mathbb{R}^{2}\to[K]$ be the linear classifier defined in Equation 6, and let $\mathcal{F}_{\mathcal{Q}}$ be the class of functions defined in Equation 7, where $\mathcal{Q}\subseteq\mathbb{R}^{d\times 2}$ is the family that satisfies the statement of C.4.
 We define the family of distributions $\mathcal{D}$ such that $D_{\mathbf{Q}}\in\mathcal{D}$ if and only if $D_{\mathbf{Q}}$ is a joint distribution $(\mathbf{x},y_{\mathbf{Q}})$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $y_{\mathbf{Q}}$ is a random variable supported on $[K]$ defined as $\operatorname*{\mathbf{Pr}}[y_{\mathbf{Q}}=i\mid\mathbf{x}]=\mathbf{H}_{f(\mathbf{Q}\mathbf{x}),i}$ for some $\mathbf{Q}\in\mathbf{\mathcal{missing}}{Q}$.
 Note that all distributions in $\mathcal{D}$ satisfy Definition 4.8 for confusion matrix $\mathbf{H}$ and some function in $\mathcal{F}_{\mathcal{Q}}$.
-Denote by $D$ the product distribution $\mathcal{N}^{d}\times\text{Unif}([K])$. For $D_{\mathbf{Q}},D_{\mathbf{P}}\in\mathcal{D}$ with $\mathbf{Q}\neq\mathbf{P}$, we have that
+Denote by $D$ the product distribution with $\mathbf{x}$-marginals distributed as standard normal and the labels $y$ are distributed uniformly on $[K]$.
+For $D_{\mathbf{Q}},D_{\mathbf{P}}\in\mathcal{D}$ with $\mathbf{Q}\neq\mathbf{P}$, we have that
 $\chi_{D}(D_{\mathbf{Q}},D_{\mathbf{P}})$ $=K\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{d}}\left[(\operatorname*{\mathbf{Pr}}[y_{\mathbf{Q}}=i\mid\mathbf{x}]-\frac{1}{K})(\operatorname*{\mathbf{Pr}}[y_{\mathbf{P}}=i\mid\mathbf{x}]-\frac{1}{K})\right]$
@@
 Consequently, we have that there exists a family of distributions of size $2^{\Omega(d^{c})}$ that is $(d^{\Omega(K)(c-1/2)}K,K)$ correlated.
-Hence, by C.3 we have that any SQ algorithm that solves the decision problem $\mathcal{B}(\mathcal{D},D)$ requires $2^{\Omega(d^{c})}d^{\Omega(K)(c-1/2)}$ queries to the $\text{VSTAT}(1/(d^{\Omega(K)(c-1/2)}))$.
-In order to complete the statement of the theorem, we need to show the following standard reduction from the testing problem to the learning problem.
+Hence, by C.3 we have that any SQ algorithm that solves the decision problem $\mathcal{B}(\mathcal{D},D)$ requires $2^{\Omega(d^{c})}d^{\Omega(K)(c-1/2)}=2^{d^{\Omega(1)}}$ (where we used the assumption that $d/\log(d)$ is much larger that $K$) queries to the $\text{VSTAT}(d^{\Omega(K)}$.
+In order to complete the proof of the theorem, we need to show the following standard reduction from the testing problem to the learning problem.
 Claim 5.8.
@@
 Proof.
-Assume that there exists an SQ algorithm $\mathcal{A}$ that if given access to a distribution $D^{\prime}$ over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is standard normal returns a hypothesis $h$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D^{\prime}}[h(\mathbf{x})\neq y]\leq\mathrm{OPT}+\gamma/4$.
-Using $\mathcal{A}$, we can reduce the testing problem to the learning problem.
+Assume that there exists an SQ algorithm $\mathcal{A}$ that if given access to a distribution $D^{\prime}$ over $\mathbb{R}^{d}\times\mathcal{Y}$ whose $\mathbf{x}$-marginal is the standard normal returns a hypothesis $h$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D^{\prime}}[h(\mathbf{x})\neq y]\leq\mathrm{OPT}+\gamma/4$.
+Using $\mathcal{A}$, we can solve the testing problem.
 Note that by running $\mathcal{A}$ to distribution $D_{\mathbf{P}}\in\mathcal{D}$, $\mathcal{A}$ returns a hypothesis $h$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},y)\sim D_{\mathbf{P}}}[h(\mathbf{x})\neq y]\leq\mathrm{OPT}+\gamma/4$.
-However, running $\mathcal{A}$ with SQ oracle access to $D$ would result in a hypothesis $h$ that cannot perform better than $1-1/K,$ as $D$ is a product distribution with the uniform distribution over the labels.
-Since $\mathrm{OPT}<1-1/K-\gamma/2$, because $\mathbf{H}_{i,i}\geq 1/K+\gamma/2$ as $\sum_{j}\mathbf{H}_{i,j}=1,i\in[K]$, we have that with an additional query of accuracy less than $\gamma/4$, we can solve $\mathcal{B}(\mathcal{D},D)$.
+Moreover, by running $\mathcal{A}$ with SQ oracle access to $D$ results in a hypothesis $h$ that cannot perform better than $1-1/K,$ as $D$ is a product distribution with the uniform distribution over the labels.
+Since $\mathrm{OPT}<1-1/K-\gamma/2$, because $\mathbf{H}_{i,i}\geq 1/K+\gamma/2$ as $\sum_{j}\mathbf{H}_{i,j}=1,i\in[K]$, we have that with an additional query of tolerance less than $\gamma/4$, we can solve $\mathcal{B}(\mathcal{D},D)$.
 ∎
-For $K$ not divisible by $4$, we add $r=4\mod K$ additional classes whose weight vectors are identical to $\mathbf{w}^{(1)}$, and with no noise, i.e., for these added classes $j$, we have that $\mathbf{H}_{jj}=1$ and $\mathbf{H}_{j,i}=\mathbf{H}_{i,j}=0$ for all $j\neq i$.
+For $K$ not divisible by $4$, we add $r=4\mod K$ additional classes whose weight vectors are identical to $\mathbf{w}^{(1)}$, and with no noise, i.e., for these added classes $j$, we have that $\mathbf{H}_{j,j}=1$ and $\mathbf{H}_{j,i}=\mathbf{H}_{i,j}=0$ for all $j\neq i$.
 For the remaining classes, i.e., for $i,j\in[K-r]$, we consider a matrix $\mathbf{H}_{i,j}$ that satisfies the conditions of Lemma 5.5.
 Note that since these additional classes have zero probability (as ties are broken by the smallest index), our SQ hardness result extends to the general case of any $K$.
+This completes the proof.
 ∎
@@
 Noting that the final expression is the Gaussian sensitivity, the result follows. ∎
+Legendre Polynomials
+Legendre polynomials are orthogonal on the interval $[-1,1]$. Their key properties are summarized in the following fact.
+Fact A.8 ([Sze67]).
+The Legendre polynomials $P_{k}$ for $k\in\mathbb{Z}$, satisfy the following properties:
+  1. 1.
+    $P_{k}$ is a $k$-degree polynomial and $P_{0}(x)=1$ and $P_{1}(x)=x$.
+  2. 2.
+    $\int_{-1}^{1}P_{i}(x)P_{j}(x)\mathrm{d}x=2/(2i+1)\mathds{1}(i=j)$, for all $i,j\in\mathbb{Z}$.
+  3. 3.
+    $|P_{k}(x)|\leq 1$ for all $|x|\leq 1$.
+  4. 4.
+    $P_{k}(x)=(-1)^{k}P_{k}(-x)$.
+  5. 5.
+    $P_{k}(x)=2^{-k}\sum_{i=1}^{\lceil k/2\rceil}\binom{k}{i}\binom{2k-2i}{k}x^{k-2i}$.
 Appendix B Omitted Content from Sections 2, 3 and 4
@@
 Moreover, if for some $T\in\mathbb{R}$ the halfspace $\{\mathbf{x}:\mathbf{u}\cdot\mathbf{x}+T\geq 0\}$ is one of the defining halfspaces of the intersection then, we have variance reduction along $u$, i.e., $\operatorname{Var}_{\mathbf{x}\sim\mathcal{N}_{d}|_{\mathcal{K}}}(\mathbf{u}\cdot\mathbf{x})\leq 1-\frac{1}{2}e^{-\frac{1}{2}\max\{0,T\}^{2}}$ for a sufficiently large universal constant $C>0$.
-Appendix C Basics on Statistical Query Model
+Appendix C Additional Background on the SQ Model
 In the following definition, we state the notion of pairwise correlation between two distributions with respect to a third distribution, and then we introduce the concept of a $(\gamma,\beta)$-correlated family.
@@
 Let $0<a,c<\frac{1}{2}$ and $m,n\in\mathbb{Z}_{+}$ such that $m\leq n^{a}$.
-There exists a set $S$ of $\Omega(n^{c})$ matrices in $\mathbb{R}^{m\times n}$ such that every $\mathbf{U}\in S$ satisfies $\mathbf{U}\mathbf{U}^{\top}=\mathbf{I}_{m}$ and every pair $\mathbf{U},\mathbf{V}\in S$ with $\mathbf{U}\neq\mathbf{V}$ satisfies $\|\mathbf{UV}^{\top}\|_{F}\leq O(n^{2c-1+2a})$.
+There exists a set $S$ of $2^{\Omega(n^{c})}$ matrices in $\mathbb{R}^{m\times n}$ such that every $\mathbf{U}\in S$ satisfies $\mathbf{U}\mathbf{U}^{\top}=\mathbf{I}_{m}$ and every pair $\mathbf{U},\mathbf{V}\in S$ with $\mathbf{U}\neq\mathbf{V}$ satisfies $\|\mathbf{UV}^{\top}\|_{F}\leq O(n^{2c-1+2a})$.
 Furthermore, to bound the correlation between two rotated versions of the same function, we will use the following fact, which relates this correlation to the Hermite coefficients of the function and the angle between the rotations.
@@
 where $g^{[t]}$ denotes the degree-$t$ Hermite part of $g$.
+Appendix D Relation to Other Complexity Measures
+D.1 Related Complexity Measures
+Several complexity measures have been proposed for learning real-valued concepts in the realizable setting.
+One measure closely related to our work is the generative exponent $k^{*}$, introduced in [DPVLB24].
+The generative exponent $k^{*}$ is defined as the smallest positive integer $k$ for which $\operatorname*{\mathbf{E}}_{y\sim D_{y}}[\zeta_{k}^{2}(y)]\neq 0$, where $\zeta_{k}(y)\coloneqq\operatorname*{\mathbf{E}}_{x\sim\mathcal{N}(0,1)}[He_{k}(x)\mid f(x)=y]$ and $f:\mathbb{R}\to\mathbb{R}$ is a single-index link function.
+This measure is used for the task of approximating the hidden direction $\mathbf{w}$ of a single-index model $f(\mathbf{w}\cdot\mathbf{x})$ from samples $(\mathbf{x},f(\mathbf{w}\cdot\mathbf{x}))$ with $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
+Using the measure $k^{*}$, [DPVLB24] derived nearly tight SQ bounds, with complexity of order $d^{k^{*}/2}$.
+(Technically speaking, their upper bound requires knowledge of $\zeta_{k^{*}}(y)$ or the existence of an orthonormal polynomial basis with respect to the marginal distribution over the labels $D_{y}$, where $\zeta_{k^{*}}$ has a large low-degree component.)
+Our definition of well-behaved MIMs (Definition 1.2) implicitly introduces a generalization of the generative exponent for PAC learning discrete-valued MIMs in the presence of noise.
+As we demonstrate in Section D.2, for the special case of realizable learning of Single-Index models, our complexity measure is essentially equivalent to the generative exponent.
+We establish this under two additional regularity assumptions: (i) a bound on the GSA of the ground truth concept, and (b) the existence of non-trivial distinguishing moments with respect to an already discovered subspace.
+(These regularity assumptions hold for several well-studied function classes such as multiclass linear classifiers and intersections of halfspaces.)
+Note that our learning algorithm does not require prior knowledge of the moments conditioned on each level-set for two reasons: (i) when the number of labels is finite, we can leverage the moments of all level-sets efficiently; and (ii) the bounded GSA assumption allows us to partition an already discovered subspace into finite-width regions to uncover additional directions using moments.
+Moreover, by considering the existence of distinguishing moments only when our approximation to the target ground truth subspace is insufficient, our notion allows for relaxed approximation guarantees that yield more efficient algorithms.
+Other relevant complexity measures that have gained attention recently are the information exponent [AGJ21] and the leap complexity [ABAB+21, AAM22, AAM23].
+The information exponent is defined as the smallest integer $k$ for which a single-index link function $f:\mathbb{R}\to\mathbb{R}$ has a non-zero Hermite coefficient, i.e., the minimum $k$ such that $\operatorname*{\mathbf{E}}_{x\sim\mathcal{N}(0,1)}[f(x)He_{k}(x)]\neq 0$.
+The leap complexity generalizes this notion to real-valued MIMs: we say that a function $f:\mathbb{R}^{K}\to\mathbb{R}$ has leap complexity $\mathrm{Leap}(f)=k$ if every Hermite decomposition of $f$ can be ordered such that each consecutive term introduces new directions via a polynomial of degree at most $k$.
+Since these measures rely solely on the non-negativity of the Hermite coefficients and do not incorporate label-based information (e.g., via level-sets), they are closely related to the CSQ complexity, which can be arbitrarily larger than the SQ complexity.
+D.2 Near-equivalence to Generative Exponent for Sindle-Index Case
+In this section, we compare our definition of well-behaved MIMs (Definition 1.2) to the generative exponent (Definition D.1) introduced in [DPVLB24].
+We first give a formal definition of the generative exponent:
+Definition D.1 (Generative exponent).
+Let $f:\mathbb{R}\to\mathbb{R}$ be a single index link function.
+Let $D$ be the joint distribution of $(x,f(x))$, where $x\sim\mathcal{N}(0,1)$.
+The generative exponent $k^{*}$ is defined as the smallest positive integer $k$ for which $\operatorname*{\mathbf{E}}_{y\sim D_{y}}[\zeta_{k}^{2}(y)]\neq 0$, where $\zeta_{k}(y)\coloneqq\operatorname*{\mathbf{E}}_{x\sim\mathcal{N}(0,1)}[He_{k}(x)\mid f(x)=y]$.
+As we have mentioned in the previous subsection, the generative exponent focuses on subspace recovery (in the realizable setting) rather than accurate label prediction.
+Hence, in order for the two notions to be comparable, we first slightly modify our definition of Definition 1.2 to this setting.
+Definition D.2 (Well-Behaved MIMs).
+Fix $m\in\mathbb{Z}_{+},\sigma\in(0,1)$ and $\mathcal{Y}\subseteq\mathbb{Z}$ of finite cardinality.
+Let $f:\mathbb{R}^{d}\to\mathcal{Y}$ be a $K$-MIM function with hidden subspace $W$.
+We say that a $K$-MIM function $f:\mathbb{R}^{d}\to\mathcal{Y}$ is $m$-well-behaved MIM if for any subspace $V\subseteq\mathbb{R}^{d}$ the following conditions holds:
+  1. a)
+    $V$ contains $W$, i.e., $W\subseteq V$, or
+  2. b)
+    there exists a point $\mathbf{x}_{0}\in\mathbb{R}^{d}$, a zero mean, unit-variance polynomial $p:U\to\mathbb{R}$ of degree at most $m$ and $z\in\mathcal{Y}$ such that
+    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[p(\mathbf{x}^{U})\mathds{1}(f(\mathbf{x})=z)\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\neq 0\;,$
+    where $U=(V+W)\cap V^{\perp}$.
+Compared to Definition 1.2, our modified definition of well-behaved MIMs in Definition D.2 introduces the following changes:
+  1. (i)
+    We omitted the technical assumption of bounded GSA to direct our attention on the existence of distinguishing moments.
+  2. (ii)
+    Instead of requiring that the function be approximately equal to one that depends solely on the projection onto $V$ (as specified in condition (a) of Definition 1.2), we now require that $V$ contain the hidden subspace $W$.
+This change emphasizes subspace recovery in the realizable setting and aligns our definition with that of the generative exponent.
+  3. (iii)
+    Since we are focusing on the realizable setting, we consider only the moments of $f$ rather than noisy version of $f$.
+In the following lemma, we show that for SIMs, the smallest degree $m$ for which Definition D.2 is satisfied is equal to the generative exponent.
+Lemma D.3.
+Let $\mathcal{Y}$ be a finite cardinality set.
+Let $f:\mathbb{R}^{d}\to\mathcal{Y}$ be a SIM, that is, there exists a unit vector $\mathbf{w}$ such that $f(\mathbf{x})=f(\mathbf{x}^{\mathbf{w}})$ for all $\mathbf{x}\in\mathbb{R}^{d}$.
+The smallest positive integer $m$ for which $f$ is an $m$-well-behaved MIM (as per Definition D.2) is equal to the generative exponent $k^{*}$ of $f$ (see Definition D.1).
+Proof.
+We first show that $f$ is not an $m$-well-behaved MIM for all positive integers $m<k^{*}$. Note that since $\mathcal{Y}$ is discrete for all $k\in\mathbb{Z}_{+}$ we have that
+$\operatorname*{\mathbf{E}}_{y\sim D_{y}}[\zeta_{k}^{2}(y)]=\sum_{y\in\mathcal{Y}}\zeta_{k}^{2}(y)\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(0,1)}[f(x\mathbf{w})=y]\;,$
+where $\zeta_{k}(y)\coloneqq\operatorname*{\mathbf{E}}_{x\sim\mathcal{N}(0,1)}[He_{k}(x)\mid f(x\mathbf{w})=y]$ (see Definition D.1).
+Hence for each positive integer $m<k^{*}$ by the definition of the generative exponent for all $y\in\mathcal{Y}$ it holds that
+$\operatorname*{\mathbf{E}}_{x\sim\mathcal{N}(0,1)}[He_{m}(x)\mid f(x\mathbf{w})=y]^{2}\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(0,1)}[f(x\mathbf{w})=y]$
+$=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[He_{m}(\mathbf{w}\cdot\mathbf{x})\mid f(\mathbf{x})=y]^{2}\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=y]=0\;.$
+Otherwise, the generative exponent would be less than $k^{*}$.
+Thus, for every $m\in[k^{*}-1]$ and $y\in\mathcal{Y}$ either $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[He_{m}(\mathbf{w}\cdot\mathbf{x})\mid f(\mathbf{x})=y]$ or $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[f(\mathbf{x})=y]=0$.
+In either case, this implies that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[He_{m}(\mathbf{w}\cdot\mathbf{x})\mathds{1}(f(\mathbf{x})=y)]=0$, for all $m\in[k^{*}-1]$ and $y\in\mathcal{Y}$.
+Since any zero-mean polynomial of degree less $k^{*}$ can be decomposed to a linear combination of non-constant Hermite polynomials of degree less than $k^{*}$, it follows that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[p(\mathbf{w}\cdot\mathbf{x})\mathds{1}(f(\mathbf{x})=y)]=0$, for any zero-mean polynomial $p$ of degree less than $k^{*}$.
+Furthermore, because $f$ is a SIM, for any polynomial $p:\mathbb{R}^{d}\to\mathbb{R}$ we have
+$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[p(\mathbf{x})\mathds{1}(f(\mathbf{x})=y)]=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[p((\mathbf{w}\cdot\mathbf{x})\mathbf{w})\mathds{1}(f(\mathbf{x})=y)]\;.$
+Therefore, for any zero-mean polynomial $p$ of degree less $k^{*}$ it holds that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[p(\mathbf{x})\mathds{1}(f(\mathbf{x})=y)]=0$, so $f$ does not satisfy the condition (b) of Definition D.2 even for $V=\{0\}$.
+This concludes the first part of the proof.
+Now to conclude the proof it suffices to show that $f$ is a $k^{*}$-well-behaved MIM.
+Note that since $\operatorname*{\mathbf{E}}_{y\sim D_{y}}[\zeta_{k^{*}}^{2}(y)]\neq 0$ and $\mathcal{Y}$ is discrete, we have that there exists a label $y\in\mathcal{Y}$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[He_{k^{*}}(\mathbf{w}\cdot\mathbf{x})\mathds{1}(f(\mathbf{x})=y)]\neq 0$.
+Consider a subspace $V$ of $\mathbb{R}^{d}$ that does not contain $\mathbf{w}$ and denote by $\mathbf{u}$ the projection of $\mathbf{w}$ onto $V^{\perp}$.
+Note that since $V$ does not contain $\mathbf{w}$ it holds that $\left\|\mathbf{u}\right\|=\left\|\mathbf{w}^{\perp V}\right\|=\mathbf{w}\cdot\mathbf{u}\neq 0$.
+Since $f$ is a SIM ,we have that for any $\lambda\in\mathbb{R}$
+$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[He_{k^{*}}(\lambda\mathbf{u}\cdot\mathbf{x})\mathds{1}(f(\mathbf{x})=y)]=\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[He_{k^{*}}(\lambda(\mathbf{w}\cdot\mathbf{u})(\mathbf{w}\cdot\mathbf{x}))\mathds{1}(f(\mathbf{x})=y)]\;.$
+Thus, setting $\lambda=1/(\mathbf{w}\cdot\mathbf{u})$ gives us that
+$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[He_{k^{*}}\left(\frac{\mathbf{u}\cdot\mathbf{x}}{\mathbf{w}\cdot\mathbf{u}}\right)\mathds{1}(f(\mathbf{x})=y)\right]\neq 0\;.$
+Note that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[He_{k^{*}}\left({\mathbf{u}\cdot\mathbf{x}}/{(\mathbf{w}\cdot\mathbf{u})}\right)]=0$ by a change of variables, since $\mathbf{u}/(\mathbf{w}\cdot\mathbf{u})$ is a unit vector.
+Thus, normalizing the above polynomial implies the existence of a mean-zero, variance-one polynomial $p:\mathbb{R}\to\mathbb{R}$ of degree $k^{*}$ such that
+$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[p(\mathbf{u}\cdot\mathbf{x})\mathds{1}(f(\mathbf{x})=y)]\neq 0\;.$
+Therefore, there has to exist an $\mathbf{x}_{0}^{V}$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[p(\mathbf{u}\cdot\mathbf{x})\mathds{1}(f(\mathbf{x})=y)\mid\mathbf{x}^{V}=\mathbf{x}_{0}^{V}]\neq 0\;$, otherwise averaging over $\mathbf{x}_{0}^{V}$ would result to the above moment being zero.
+Consequently, we have that $f$ is a $k^{*}$-well-behaved MIM, which completes the proof.
+∎
+Appendix E SQ Lower Bound for Multiclass Linear Classification from Partial Labels
+In this section, we will leverage the moment-matching construction developed in the previous section (for learning with RCN) in order to prove SQ hardness for learning multiclass linear classifiers with partial labels.
+First we give a formal definition of the model of corruption and then we present our construction for the hard instances along with our main theorem.
+Definition E.1 (Partial Label Distribution).
+Let $D$ be a distribution over $\mathbb{R}^{d}\times 2^{\mathcal{Y}}$.
+We say that $D$ is a partial label distribution for the class $\mathcal{F}\subseteq\{f:\mathbb{R}^{d}\to\mathcal{Y}\}$ if there exists a function $f\in\mathcal{F}$ and a row-stochastic matrix $\mathbf{H}\in\mathbb{R}^{\lvert\mathcal{Y}\rvert\times 2^{\lvert\mathcal{Y}\rvert}}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},S)\sim D}[S=T\mid\mathbf{x}]=\mathbf{H}_{f(\mathbf{x}),T}$ for all $\mathbf{x}\in\mathbb{R}^{d},T\subseteq S$.
+Moreover, $\mathbf{H}_{i,S}=0$ if $i\not\in S$ and $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},S)\sim D}[y\not\in S\mid\mathbf{x}]=\sum_{S\not\ni y}\mathbf{H}_{f(\mathbf{x}),S}\geq\gamma$, for all $y\neq f(\mathbf{x})$ for all $\mathbf{x}\in\mathbb{R}^{d}$, for some $\gamma>0$.
+We refer to $\mathbf{H}$ as the confusion matrix of the distribution.
+SQ-Hard Instances
+To construct our family of hard instances, we consider planted versions of a two-dimensional model.
+This part of the construction is the same as the one used in Section 5.2 and it is defined in Equations 6 and 7.
+As for the confusion matrix, we consider a subset of learning with partial labels problems called learning with contrastive labels [INHS17, ZLF+20].
+In this setting, we observe examples of the form $\{(\mathbf{x}^{(i)},\bar{y}_{i})\}_{1}^{m}$, where $\bar{y}_{i}$ is almost surely different to the ground truth label $y_{i}$.
+Without loss of generality we will refer to the example distribution, $D$, as a distribution over $\mathbb{R}^{d}\times\mathcal{Y}$ rather than a distribution over $\mathbb{R}^{d}\times 2^{\mathcal{Y}}$ as there are $\lvert\mathcal{Y}\rvert$ subsets $S\subseteq\mathcal{Y}$ where $\lvert S\rvert=\lvert\mathcal{Y}\rvert-1$.
+Moreover, by Definition E.1, we have that there exists a $\lvert\mathcal{Y}\rvert\times\lvert\mathcal{Y}\rvert$ stochastic matrix $\mathbf{H}$ such that $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},\bar{y})\sim D}[\bar{y}=j\mid\mathbf{x},f(\mathbf{x})=i]=\mathbf{H}_{i,j}$ for all $\mathbf{x}\in\mathbb{R}^{d},i,j\in\mathcal{Y}$ and the identifiability constraint corresponds to the existence of a constant $\gamma\in(0,1]$ satisfying $\mathbf{H}_{i,j}=0,\text{ if }i=j\text{ and }\mathbf{H}_{i,j}\geq\gamma,\text{if }i\neq j$, for all $i,j\in\mathcal{Y}$.
+Using this construction we prove the main result of this section:
+Theorem E.2 (SQ Lower Bound for Learning with Partial Labels).
+Any SQ algorithm that learns $K$-class Linear Classifiers given partial labels, with $\gamma=\mathrm{poly}(1/K)$ (see Definition E.1), under the $d$-dimensional standard normal to error $1/(2K)$, requires either $2^{d^{\Omega(1)}}$ queries or at least one query to $\text{VSTAT}\left(d^{\Omega(K)}\right)$.
+Before proving Theorem E.2, we first need to show that we can match the appropriate number of moments with the standard Gaussian. The proof of this lemma is similar to the proof of Lemma 5.5.
+Lemma E.3.
+Let $K\in\mathbb{Z}_{+}$ with $K\geq 8$ such that $K$ is divisible by $4$. There exists a $K\times K$ doubly stochastic matrix $\mathbf{H}$ such that the following hold:
+  1. i)
+    $\mathbf{H}_{i,i}=0$, for all $i\in[K]$ and $\min_{i\neq j}\mathbf{H}_{i,j}=\Omega(1/K^{3}).$
+  2. ii)
+    For every polynomial $p:\mathbb{R}^{2}\to\mathbb{R}$ of degree at most $K/2-2$ we have
+    $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[\left(\operatorname*{\mathbf{Pr}}_{(\mathbf{x},\bar{y})\sim D}\left[\bar{y}=i\mid\mathbf{x}\right]-\frac{1}{K}\right)p(\mathbf{x})\right]=0$ for all $i\in[K]$.
+where $D$ is a contrastive label distribution over $\mathbb{R}^{2}\times[K]$ whose $\mathbf{x}$-marginal is $\mathcal{N}(\mathbf{0},\mathbf{I})$ that satisfies Definition E.1 for the confusion matrix $\mathbf{H}$ and the function $f$ as defined in Equation 6.
+Proof.
+Let $\mathbf{H}\in\mathbb{R}^{K\times K}$ be a doubly stochastic matrix and define the functions $\tilde{f}_{i}$ for $i\in[K]$ as $\tilde{f}_{i}(\mathbf{x})=\sum_{j=1}^{K}\mathbf{H}_{j,i}\mathds{1}(f(\mathbf{x})=j)$.
+Also denote by $\boldsymbol{\omega}^{(k)}$ the root of unity vector of rate $k$, i.e.
+${\boldsymbol{\omega}}^{(k)}\in\mathbb{C}^{K}$ such that $\boldsymbol{\omega}_{j}^{(k)}=e^{\frac{2\pi kji}{K}}$.
+Exactly as in the proof of Lemma 5.5 we have that by Claims 5.6 and 5.7 in order to have 00 correlation with all polynomials of degree at most $m$, it suffices to have that $\boldsymbol{\omega}^{(k)}\in\ker(\mathbf{H}^{\top})$ for all $k\in\{1,\dots,m,K-m,\dots,K-1\}$, since the roots of unity are symmetric, i.e., $\boldsymbol{\omega}^{(-k)}=\boldsymbol{\omega}^{(K-k)}$.
+We first construct a stochastic vector $\mathbf{h}\in\mathbb{R}^{K}$ such that $\mathbf{h}_{0}=0$ and $\mathbf{h}_{j}=\Omega(1/K^{3}),\forall j\in\{1,\dots,K-1\}$ that is orthogonal to the aforementioned root of unity vectors for $m=\Omega(K)$.
+Next, we extent this vector to a matrix.
+Define the vectors $\mathbf{h}^{(1)}=\boldsymbol{\omega}^{(0)}/K$, $\mathbf{h}^{(2)}=(\boldsymbol{\omega}^{(K/2-1)}+\boldsymbol{\omega}^{(K/2+1)})/K$ and $\mathbf{h}^{\prime}=2\mathbf{h}^{(1)}-\mathbf{h}^{(2)}$.
+It is easy to see that $\mathbf{h}^{\prime}$ is a vector with non-negative entries and $\mathbf{h}_{0}^{\prime}=0$.
+Moreover, it is true that
+$\min_{j\in[K-1]}\mathbf{h}_{j}^{\prime}\geq 2/K-\max_{j\in[K-1]}\mathbf{h}_{j}^{(2)}=2/K-2/K\cos\left(\frac{2\pi}{K}\right)\gtrsim\frac{1}{K^{3}}\;,$
+where the last equality follows from the Taylor series $\cos(x)$ which is an alternating decreasing series when $x=\frac{2\pi}{K}\leq 1$.
+Let $\mathbf{h}$ be the normalization of $\mathbf{h}^{\prime}$ such that the sum of its coordinates equals $1$.
+Since each element of $\mathbf{h}^{\prime}$ is less than $4/K$ we have that the normalization factor is at most $4$ and hence $\min_{j\in[K-1]}\mathbf{h}_{j}\gtrsim\frac{1}{K^{3}}$.
+Moreover, from the well-known fact that $\boldsymbol{\omega}^{(0)},\dots,\boldsymbol{\omega}^{(K-1)}$ form an orthogonal basis of $\mathbb{C}^{K}$, we have that $\mathbf{h}$ is orthogonal to $\boldsymbol{\omega}^{(k)}$ for all $k\in\{1,\dots,K/2-2,K/2,K/2+2,\dots,K-1\}$, since it is a linear combination of $\boldsymbol{\omega}^{(0)},\boldsymbol{\omega}^{(K/2-1)},\boldsymbol{\omega}^{(K/2+1)}$.
+Now consider $\mathbf{H}$, the circulant matrix generated by row shifts of the vector $h$.
+By construction, it is true that $\mathbf{H}$ is row stochastic with $\mathbf{H}_{i,i}=0$ for all $i\in[K]$ such that $\min_{i\neq j}\mathbf{H}_{i,j}=\Omega(1/K^{3})$.
+Moreover, $\mathbf{H}$ is symmetric, because it holds that $\mathbf{H}_{i,j}=\mathbf{h}_{j-i\mod K}=\mathbf{h}_{i-j\mod K}=\mathbf{H}_{j,i}$ and $\mathbf{h}_{j}=\mathbf{h}_{K-j},j\in[K-1]$.
+This equality is justified by the identity $(-1)^{j}\cos\left({2\pi j}/{K}\right)=(-1)^{K-j}\cos\left({2\pi(K-j)}/{K}\right)$, which holds since $K$ is an even integer.
+Hence $\mathbf{H}$ is also column stochastic.
+Hence, $\boldsymbol{\omega}^{(k)}\in\ker(\mathbf{H})=\ker(\mathbf{H}^{\top})$, for all $k\in\{1,\dots,K/2-2,K/2,K/2+2,\dots,K-1\}$.
+Therefore, applying Claim 5.6 and A.6 concludes the proof of Lemma E.3.
+∎
+E.1 Proof of Theorem E.2
+Now, given Lemma E.3, we proceed to the proof of Theorem E.2.
+The proof is essentially the same to that of Theorem 5.4, except that it employs the construction from Lemma E.3 rather than Lemma 5.5.
+Despite that we include the complete proof here for completeness.
+Proof of Theorem E.2.
+Consider the case where $K$ is divisible by $4$ and let $\mathbf{H}\subseteq\mathbb{R}^{K\times K}$ be a matrix that satisfies the statement of Lemma E.3 Let $f:\mathbb{R}^{2}\to[K]$ be the linear classifier defined in Equation 6, and let $\mathcal{F}_{\mathcal{Q}}$ be the class of function defined in Equation 7, where $\mathcal{Q}\subseteq\mathbb{R}^{d\times 2}$ the family that satisfies the statement of C.4.
+We define the family of distributions $\mathcal{D}$ such that $D_{\mathbf{Q}}\in\mathcal{D}$ if and only if $D_{\mathbf{Q}}$ is a joint distribution $(\mathbf{x},y_{\mathbf{Q}})$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $y_{\mathbf{Q}}$ is a random variable supported on $[K]$ defined as $\operatorname*{\mathbf{Pr}}[y_{\mathbf{Q}}=i\mid\mathbf{x}]=\mathbf{H}_{f(\mathbf{Q}\mathbf{x}),i}$ for some $\mathbf{Q}\in\mathbf{\mathcal{missing}}{Q}$.
+Note that all distributions in $\mathcal{D}$ satisfy Definition E.1 for confusion matrix $\mathbf{H}$ and some function in $\mathcal{F}_{\mathcal{Q}}$.
+Denote by $D$ the product distribution $\mathcal{N}^{d}\times\text{Unif}([K])$. For $D_{\mathbf{Q}},D_{\mathbf{P}}\in\mathcal{D}$ with $\mathbf{Q}\neq\mathbf{P}$, we have that
+$\chi_{D}(D_{\mathbf{Q}},D_{\mathbf{P}})$ $=K\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{d}}\left[(\operatorname*{\mathbf{Pr}}[y_{\mathbf{Q}}=i\mid\mathbf{x}]-\frac{1}{K})(\operatorname*{\mathbf{Pr}}[y_{\mathbf{P}}=i\mid\mathbf{x}]-\frac{1}{K})\right]$
+$=K\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{d}}\left[(\mathbf{H}_{f(\mathbf{Q}\mathbf{x}),i}-\frac{1}{K})(\mathbf{H}_{f(\mathbf{P}\mathbf{x}),i}-\frac{1}{K})\right]$
+$=K\sum_{i=1}^{K}\sum_{j=1}^{\infty}\left\|\mathbf{P}\mathbf{Q}^{\top}\right\|_{2}^{j}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[((\mathbf{H}_{f(\mathbf{x}),i}-\frac{1}{K})^{[j]})^{2}\right]$
+$=\left\|\mathbf{P}\mathbf{Q}^{\top}\right\|_{2}^{K/2-2}K\sum_{i=1}^{K}\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}^{2}}\left[(\mathbf{H}_{f(\mathbf{x}),i}-\frac{1}{K})^{2}\right]\leq d^{\Omega(K)(c-1/2)}K\;,$
+where we used C.5, Parseval’s identity and the fact that $\mathbf{H}$ satisfies the statement of Lemma E.3.
+Similarly, we have that $\chi_{D}(D_{\mathbf{Q}},D_{\mathbf{Q}})\leq K$ for all $D_{\mathbf{Q}}\in\mathcal{D}$.
+Consequently, we have that there exists a family of distributions of size $2^{\Omega(d^{c})}$ that is $(d^{\Omega(K)(c-1/2)}K,K)$ correlated.
+Hence, by C.3 we have that any SQ algorithm that solves the decision problem $\mathcal{B}(\mathcal{D},D)$ requires $2^{\Omega(d^{c})}d^{\Omega(K)(c-1/2)}$ queries to the $\text{VSTAT}(1/(d^{\Omega(K)(c-1/2)}))$.
+In order to complete the statement of the theorem, we need to show the following standard reduction from the testing problem to the learning problem.
+Claim E.4.
+Let $\epsilon\leq 1/K$.
+Any SQ algorithm that learns multiclass linear classifiers with partial labels under the standard normal with error better than $\epsilon$ can be used to solve the decision problem $\mathcal{B}(\mathcal{D},D)$ using one more query of accuracy $1/K-\epsilon$.
+Proof.
+Assume that there exists an SQ algorithm $\mathcal{A}$ that if given access to a partial label distribution $D^{\prime}$ (see Definition E.1) over $\mathbb{R}^{d}\times 2^{\mathcal{Y}}$ whose $\mathbf{x}$-marginal is standard normal returns a hypothesis $h$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[h(\mathbf{x})\neq f(\mathbf{x})]\leq\epsilon$.
+Using $\mathcal{A}$, we can reduce the testing problem to the learning problem.
+Note that by running $\mathcal{A}$ to distribution $D_{\mathbf{P}}\in\mathcal{D}$ that corresponds to the concept $f(\mathbf{P}\mathbf{x})\in\mathcal{F}_{\mathcal{Q}}$, $\mathcal{A}$ returns a hypothesis $h$ such that $\operatorname*{\mathbf{Pr}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[h(\mathbf{x})\neq f(\mathbf{P}\mathbf{x})]\leq\epsilon$.
+As a result, by the reverse triangle inequality, we have that testing the returned hypothesis disagreement with the contrastive labels gives us $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},\bar{y})\sim D}[h(\mathbf{x})\neq\bar{y}]\geq 1-\epsilon$.
+However, running $\mathcal{A}$ with SQ oracle access to $D$ would result in a hypothesis $h$ that has $\operatorname*{\mathbf{Pr}}_{(\mathbf{x},\bar{y})\sim D}[h(\mathbf{x})\neq\bar{y}]=1-1/K,$ since $D$ is a product distribution with the uniform distribution over the labels.
+Since, $\epsilon\leq 1/K$ we have that we can solve $\mathcal{B}(\mathcal{D},D)$ with an additional query of accuracy less than $1/K-\epsilon$.
+∎
+For $K$ not divisible by $4$, we add $r=4\mod K$ additional classes whose weight vectors are identical to $\mathbf{w}^{(1)}$.
+We construct the new confusion matrix as follows: first let $\mathbf{H}$ be the matrix that satisfies the conditions of Lemma E.3 for $i,j\in[K-r]$.
+For the added classes $j\in\{K-r+1,\dots,K\}$, we set $\mathbf{H}_{j,j}=0$, $\mathbf{H}_{j,i}=1/K$ for all $j\neq i$ and $\mathbf{H}_{i,j}=1/K$ for $i\in[K-r]$.
+Finally we normalize the elements of the rows such that the matrix is row stochastic.
+Note that since these additional classes have zero probability (since ties are broken by the smallest index) and for each of the classes $i\in[K-r]$ the added labels are observed with the same rate, our SQ hardness result extends to the general case of any $K$.
+Finally, note since the normalization factor is at most $1+4/K$ we have that $\min_{i\neq j}\mathbf{H}_{i,j}\gtrsim(1/(1+4/K))(1/K^{3})\geq\mathrm{poly}(1/K)$, which completes the proof of Theorem E.2.
+∎
+Appendix F Lower Bound For Single-Shot Optimization
+In this section, we present an example in which a single-shot approach fails to learn the target function to sufficiently small error even if it is given access to exactly the first $O(1)$-moments.
+On the other hand, we show that if we take an iterative approach then we can learn the function in polynomial time using just $O_{\epsilon}(d^{2})$ complexity.
+Our results build on moment-matching constructions and properties of Legendre polynomials.
+Theorem F.1.
+Let $d\in\mathbb{Z}_{+}$ and $C_{1},C_{2}>0$ be sufficiently large universal constants. There exists a family of functions
+=$\mathcal{F}=\{f(\mathbf{U}\mathbf{x}):\mathbf{U}\in\mathbb{R}^{2\times d},\;\mathbf{U}\mathbf{U}^{\top}=\mathbf{I}_{2}\},$
+such that:
+  • No algorithm given the first $C_{1}$ exact moments of $f(\mathbf{U}\mathbf{x})\in\mathcal{F}$ can output another function $f^{\prime}\in\mathcal{F}$ such that $\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[(f(\mathbf{U}\mathbf{x})-f^{\prime}(\mathbf{x}))^{2}]\leq 1/C_{2}$.
+  • There exists an SQ algorithm that for any $\epsilon>0$, given SQ oracle access to a distribution $D$ over $\mathbb{R}^{d}\times\{0,1\}$ (with $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $y=f(\mathbf{U}\mathbf{x})$ for some $f\in\mathcal{F}$), outputs an estimate $\widehat{\mathbf{U}}$ satisfying $\|\widehat{\mathbf{U}}-\mathbf{U}\|_{2}\leq\epsilon$ using $O(d^{2})$ queries to $\mathrm{VSTAT}(C_{2}d^{2}/\epsilon^{2})$.
+Proof.
+The proof is based on a careful construction of a two-dimensional distribution and a corresponding function family.
+Assume that the algorithm is given $C_{1}=O(1)$ moments.
+Let $k=2C_{1}$.
+Define the set $A$ (we use $A$ as the indicator function of this set) on $\mathbb{R}^{2}$ as in Proposition F.2 and let
+$B(x,y)=\mathds{1}(|x+y|\geq 2b)\;,$
+where $b=\Theta(k)$ and is the smallest value such that $A(x,y)=0$ whenever $|x|\geq b$ or $|y|\geq b$. We then define the boolean function
+$f(x,y)=A(x,y)+B(x,y)\;.$
+Let $\mathcal{Q}\subseteq\mathbb{R}^{d\times 2}$ be a set of matrices that satisfy C.4, i.e.,
+$|\mathcal{Q}|=2^{d^{\Omega(1)}},\quad\mathbf{Q}\mathbf{Q}^{\top}=\mathbf{I}_{2},\quad\text{and}\quad\|\mathbf{Q}\mathbf{P}^{\top}\|_{F}\leq O\Big{(}d^{-\Omega(1)}\Big{)}\;.$
+For each $[\mathbf{v},\mathbf{u}]=\mathbf{Q}\in\mathcal{Q}$, we define $f_{\mathbf{Q}}(\mathbf{x})$ as follows: Let $\mathbf{w}^{(1)}=\mathbf{v}/\sqrt{2}+\mathbf{u}/\sqrt{2}$ and $\mathbf{w}^{(2)}=\mathbf{v}/\sqrt{2}-\mathbf{u}/\sqrt{2}$ and we let $f_{\mathbf{Q}}(\mathbf{x})=f(\mathbf{w}^{(1)}\cdot\mathbf{x},\mathbf{w}^{(2)}\cdot\mathbf{x})$.
+Let $\mathcal{F}_{\mathcal{Q}}=\{f_{\mathbf{Q}}(\mathbf{x}):\mathbf{Q}\in\mathcal{Q}\}\subseteq\mathcal{F}$.
+For each $\mathbf{Q}\in\mathcal{Q}$, let $D_{\mathbf{Q}}$ denote the joint distribution of $(\mathbf{x},f_{\mathbf{Q}}(\mathbf{x}))$, where $\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
+Define the class of distributions $\mathcal{D}=\{D_{\mathbf{Q}}:\mathbf{Q}\in\mathcal{Q}\}$.
+Now, suppose we have the first $k/2$-moments of a distribution $D$ drawn from $\mathcal{D}$.
+Note that because $\operatorname*{\mathbf{E}}[p(\mathbf{x})A(\mathbf{w}^{(1)}\cdot\mathbf{x},\mathbf{w}^{(2)}\cdot\mathbf{x})]=0$ for any zero mean mean polynomial $p$ of degree at most $k$, we have that $\operatorname*{\mathbf{E}}[p(\mathbf{x})f(\mathbf{x})]=\operatorname*{\mathbf{E}}[p(\mathbf{x})B(\mathbf{x})]=\operatorname*{\mathbf{E}}[p(\mathbf{x})B(\mathbf{w}^{(1)}\cdot\mathbf{x}+\mathbf{w}^{(2)}\cdot\mathbf{x})]=\operatorname*{\mathbf{E}}[p(\mathbf{x})B(\mathbf{v}\cdot\mathbf{x}\sqrt{2})]$.
+Therefore, we can exactly extract from the moments only $\mathbf{w}^{(1)}+\mathbf{w}^{(2)}=\mathbf{v}\cdot\mathbf{x}\sqrt{2}$.
+Therefore, with access to $O(1)$-moments, we can only extract the direction $\mathbf{v}$.
+We show that for any vector $\mathbf{w}$ so that $\mathbf{w}\cdot\mathbf{u}\leq c_{1}$ where $c_{1}>0$ is a sufficiently small absolute constant, then the hypothesis $A(\mathbf{w}^{(3)}\cdot\mathbf{x},\mathbf{w}^{(4)}\cdot\mathbf{x})$ achieves large $L_{2}^{2}$ error where $\mathbf{w}^{(3)}=\mathbf{v}/\sqrt{2}+\mathbf{w}/\sqrt{2}$ and $\mathbf{w}^{(4)}=\mathbf{v}/\sqrt{2}-\mathbf{w}/\sqrt{2}$.
+First, we choose $\mathbf{w}$ so that $\mathbf{w}\cdot\mathbf{u}=0$.
+Note that from Jensen’s inequality it holds that
+$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}[$ $(A(\mathbf{w}^{(1)}\cdot\mathbf{x},\mathbf{w}^{(2)}\cdot\mathbf{x})-A(\mathbf{w}^{(3)}\cdot\mathbf{x},\mathbf{w}^{(4)}\cdot\mathbf{x}))^{2}]$
+$\geq\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\left(\operatorname*{\mathbf{E}}_{\mathbf{v}\cdot\mathbf{x}\sim\mathcal{N}(0,1)}[A(\mathbf{w}^{(1)}\cdot\mathbf{x},\mathbf{w}^{(2)}\cdot\mathbf{x})-A(\mathbf{w}^{(3)}\cdot\mathbf{x},\mathbf{w}^{(4)}\cdot\mathbf{x})]\right)^{2}\right]\;,$ (8)
+where the last term corresponds to the variance over the direction $\mathbf{u}$.
+Note that $A$ is a $2$-dimensional instance that depends only on matching the first $C_{1}$ moments where $C_{1}$ is an absolute constant therefore Equation 8 (as by construction is non zero) is bounded from below by an absolute constant.
+Therefore, there exists a sufficiently small absolute constant $c_{2}>0$ such that
+$\operatorname*{\mathbf{E}}_{\mathbf{x}\sim\mathcal{N}(\mathbf{0},\mathbf{I})}\left[\left(\operatorname*{\mathbf{E}}_{\mathbf{v}\cdot\mathbf{x}\sim\mathcal{N}(0,1)}[A(\mathbf{w}^{(1)}\cdot\mathbf{x},\mathbf{w}^{(2)}\cdot\mathbf{x})-A(\mathbf{w}^{(3)}\cdot\mathbf{x},\mathbf{w}^{(4)}\cdot\mathbf{x})]\right)^{2}\right]\geq c_{2}\;.$
+Furthermore, note that the $L_{2}^{2}$ of $A(\mathbf{w}^{(1)}\cdot\mathbf{x},\mathbf{w}^{(2)}\cdot\mathbf{x})$ and $A(\mathbf{w}^{(3)}\cdot\mathbf{x},\mathbf{w}^{(4)}\cdot\mathbf{x})$ distance is a decreasing function with respect to $\theta$, where $\theta=\theta(\mathbf{w},\mathbf{u})$.
+Using the fact that the set $\theta\in[0,\pi/2]$ is compact, there exists a sufficiently small absolute constant $c_{1}>0$ such that if $\theta(\mathbf{u},\mathbf{w})\geq\pi/2-c_{1}$, then the $L_{2}^{2}$ of $A(\mathbf{w}^{(1)}\cdot\mathbf{x},\mathbf{w}^{(2)}\cdot\mathbf{x})$ and $A(\mathbf{w}^{(3)}\cdot\mathbf{x},\mathbf{w}^{(4)}\cdot\mathbf{x})$ is at least $c_{2}/2$.
+Using the fact that $d$ is sufficiently large the probability of finding $\mathbf{w}$ so that $\theta(\mathbf{u},\mathbf{w})\leq\pi/2-c_{1}$ is at least $\exp(-\Omega(d))$.
+On the other hand, there exists an SQ Algorithm with only $O(d^{2})$ queries can learn the two directions of $f$.
+With $O(d)$ queries (with accuracy $O(2^{-k}\epsilon)=O(\epsilon)$) recovers an $\epsilon$ approximation to the direction $\mathbf{w}=(\mathbf{v}+\mathbf{u})/\sqrt{2}$.
+Next if we conditioning on the event $\{\mathbf{w}\cdot\mathbf{x}\geq\delta\}$ where $\delta>0$ is chosen so that $A(\mathbf{w}^{(}1)\cdot\mathbf{x},\mathbf{w}^{(2)}\cdot\mathbf{x})$ does not match moments with respect $\mathbf{u}$, which such exists and it is an absolute constant as it reweighs the probability mass of the intervals, therefore
+$\operatorname*{\mathbf{E}}\big{[}y\,\mathbf{x}\mathbf{x}^{\top}\mid\mathbf{w}\cdot\mathbf{x}\geq\delta\big{]}$
+reveals an $\epsilon$-approximation to the direction $\mathbf{u}$. This completes the proof. ∎
+F.1 Moment-Matching Construction
+We now describe the constructions used in the proof. These results build piecewise constant functions and moment-matching distributions that are central to our lower bound.
+Proposition F.2.
+Let $c>1$ be a universal constant and $k\in\mathbb{Z}_{+}$. There exists a distribution $A$ on $\mathbb{R}^{2}$, for which the following are true:
+  1. 1.
+    $A$ matches its first $k$ moments with the 2-dimensional standard Gaussian.
+  2. 2.
+    for all $\mathbf{x}\in\operatorname{supp}(A)$ it holds that $\lvert\mathbf{x}_{1}\rvert,\lvert\mathbf{x}_{2}\rvert\leq ck$ and $A$ is a union of $k+3$ intervals (in each orthogonal direction).
+Proof.
+Let $x,y$ be two orthogonal directions. Define $A(x,y)=U(x)U(y)/\operatorname*{\mathbf{Pr}}[x\in U]^{2}$ where $U$ is the set from Proposition F.3. Then, we have that $A$ satisfies Items 1 and 2. ∎
+Proposition F.3.
+For any $k\in\mathbb{Z}_{+}$ there exists a set $U\subseteq\mathbb{R}$ with $\operatorname*{\mathbf{Pr}}_{x\sim\mathcal{N}(0,1)}[x\in U]=1/2$ such that $U$ is a union of $k$ intervals with $|\max(U)|=O(k)$ and
+$\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}\big{[}y^{t}\mid y\in U\big{]}=\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}]$
+for all integers $t=0,\ldots,k$.
+Proof.
+We first show that there exists a piecewise constant function $f:\mathbb{R}\to\{0,1\}$ that matches the moments of the standard normal distribution and vanishes for $|x|\geq O(k)$.
+This is formalized in the following lemma.
+Lemma F.4.
+Fix $k\in Z_{+}$ and $\eta\in(0,1)$. There exists a $O(k)$-piecewise-constant function $f:\mathbb{R}\to\{0,2\}$ such that $f(x)=0$ for $|x|\geq\Theta(k)$ and for all $t\leq k$ it holds that
+$\left|\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}f(y)]-\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}]\right|=0\;.$
+Note that because $f(x)$ takes values in $\{0,2\}$, it holds that $\operatorname*{\mathbf{Pr}}[f(x)=2]=1/2$ (which holds from Lemma F.4 for $t=0$). It also holds that
+$\left|\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}\mid f(y)=2]-\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}]\right|=\left|\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}f(y)]\frac{1}{2\operatorname*{\mathbf{Pr}}[f(y)=2]}-\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}]\right|=0\;.$
+Taking $U=\mathds{1}(f(x)=2)$, we complete the proof of Proposition F.3. ∎
+F.2 Proof of Lemma F.4
+To prove Lemma F.4, we first show that there is a function that satisfies our properties up to an error and has exponential pieces.
+Lemma F.5.
+Fix $k\in Z_{+}$ and $\delta\gtrsim k$ and $\eta\in(0,1)$.
+There exists a $(1/\eta+1)^{\mathrm{poly}(k)}$-piecewise-constant function $f_{\eta}:\mathbb{R}\to\{0,2\}$ such that $f(x)=0$ for $|x|\geq\delta$ and for all $t\leq k$ it holds that
+$\left|\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}f(y)]-\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}]\right|\leq\eta\;.$
+Proof.
+We first show that there exists a continue function that satisfies are properties.
+Claim F.6.
+Fix $k\in Z_{+}$ and $\delta\gtrsim k$. There exists an $f:\mathbb{R}\to[0,2]$ such that $f(x)=0$ for $|x|\geq\delta$ and for all $t\leq k$ it holds that
+$\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}f(y)]=\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}]\;.$
+We fix the following parameters throughout the proof (where $C$ denotes a sufficiently large absolute constant) $s=(\eta/(\delta k))^{\mathrm{poly}(k)}$ and we choose $s$ so that there exists $i^{\prime}\in\mathbb{Z}$ so that $-\delta+i^{\prime}s=0$.
+We partition the subset $[-\delta,\delta]$ in pieces $[is,(i+1)s)$ for $i\in\mathbb{Z}$.
+We define $\tilde{f}$ to be the following random piecewise-constant function: For each $i\in\{-1/\delta,\ldots,1/\delta\}$ we let $\tilde{f}(x)$ be constant in the interval $x\in[is,(i+1)s)$, taking the following value:
+$\tilde{f}(x)=\begin{cases}2,&\text{with probability $p_{i}:=\int_{is}^{(i+1)s}\phi(x)f(x)\mathrm{d}x/(2\int_{is}^{(i+1)s}\phi(x)\mathrm{d}x$})\\ 0,&\text{with probability $1-p_{i}$}\end{cases}$ (9)
+and we define $\tilde{f}(x)=0$ with probability 1 in the entire $(-\infty,-\delta]\cup[\delta,+\infty)$.
+Our goal is to show that $\left|\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}\tilde{f}(y)]-\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}]\right|\leq\eta$.
+First note that for any $i\in[-\delta,-\delta+1,\ldots,\delta]$ and $t\leq k$, we have on expectation that
+$\operatorname*{\mathbf{E}}_{\tilde{f}}\left[\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}\tilde{f}(y)\mathds{1}(y\in(is,(i+1)s))]\right]=\int_{is}^{(i+1)s}\phi(y)f(y)\mathrm{d}y\frac{\int_{is}^{(i+1)s}y^{t}\phi(y)\mathrm{d}y}{\int_{is}^{(i+1)s}\phi(y)\mathrm{d}y}\;.$ (10)
+Because we have choose $s$ so that no interval $(is,(i+1)s)$ contains 00, that means that $y^{t}$ has constant sign inside the integration.
+Therefore, from mean valued theorem for integrals, we have that that there exists $\xi_{1},\xi_{2}\in(is,(i+1)s)$ so that the following are true:
+$\int_{is}^{(i+1)s}y^{t}\phi(y)\mathrm{d}y=\xi_{1}^{t}\int_{is}^{(i+1)s}\phi(y)\mathrm{d}y\quad\mathrm{and}\quad\int_{is}^{(i+1)s}\phi(y)f(y)y^{t}\mathrm{d}y=\xi_{2}^{t}\int_{is}^{(i+1)s}\phi(y)f(y)\mathrm{d}y\;.$
+Combining above with Equation 10, we have that
+$\left|\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}f(y)\{y\in(is,(i+1)s)\}]-\operatorname*{\mathbf{E}}_{\tilde{f}}\left[\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}\tilde{f}(y)\{y\in(is,(i+1)s)\}]\right]\right|$ $=|\xi_{1}^{t}-\xi_{2}^{t}|\int_{is}^{(i+1)s}\phi(y)f(y)\mathrm{d}y$
+$\leq 2s|\xi_{1}^{t}-\xi_{2}^{t}|$
+$\leq 2s^{2}t\delta^{t-1}\;,$
+where in the first inequality we used that $|f(y)\phi(y)|\leq 2$ and in the second we use that $|\xi_{1}^{t}-\xi_{2}^{t}|=\int_{\xi_{1}}^{\xi_{2}}tx^{t-1}\mathrm{d}x$ and $|\xi_{1}|,|\xi_{2}|\leq\delta$.
+We start using a probabilistic argument, by taking the expectation and using Fubini’s theorem, we have that
+$\left|\operatorname*{\mathbf{E}}_{\tilde{f}}\left[\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}\tilde{f}(y)-y^{t}]\right]\right|$ $=\left|\operatorname*{\mathbf{E}}_{\tilde{f}}\left[\sum_{i=-\delta}^{\delta-1}\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[(y^{t}\tilde{f}(y)-y^{t})\{y\in(is,(i+1)s)\}]\right]-\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}\{y\not\in(-\delta,\delta\}]\right|$
+$\leq\left|\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}f(y)-y^{t}]\right|+\sum_{i=-\delta}^{\delta-1}2s^{2}t\delta^{t-1}=4s^{2}t\delta^{t}=O(\eta)\;,$
+where we used that by construction $\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}f(y)-y^{t}]=0$.
+To complete the proof note that from Lemma F.8, we have that there exists $f:\mathbb{R}\mapsto\{0,2\}$ that satisfies the properties of Lemma F.4.
+F.3 Proof of Claim F.6
+Proof of Claim F.6.
+Let
+$f(x)=\left(1+\frac{p(x)\mathds{1}(x\in[-1,1])}{\phi(x)}\right)\mathds{1}(|x|\leq\delta)\;,$
+for some appropriate chosen $k$-degree polynomial $p$. We have that
+$\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}f(y)]=\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}\mathds{1}(|y|\leq\delta)]+\int_{-1}^{1}p(y)y^{t}\mathrm{d}y\;.$
+We want to chose $p$ so that $\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}f(y)]=\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}]$, therefore, we want
+$\int_{-1}^{1}p(y)y^{t}\mathrm{d}y=\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t}\mathds{1}(|y|\geq\delta)]\;.$
+Equivalently because Legendre polynomials form a complete basis, it suffices to find a polynomial $p$ so that
+$\int_{-1}^{1}p(y)P_{t}(y)\mathrm{d}y=\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[P_{t}(y)\mathds{1}(|y|\geq\delta)]\;,$
+for all $t\leq k$. Therefore, we can write $p(y)=\sum_{i=0}^{k}a_{i}P_{i}(y)$ and $a_{i}=\frac{2i+1}{2}\int_{-1}^{1}P_{i}(y)p(y)\mathrm{d}y$. Hence, we have that
+$a_{i}=\frac{2}{2i+1}\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[P_{t}(y)\mathds{1}(|y|\geq\delta)]\;.$
+Because for the Legendre polynomials it holds that $|P_{t}(y)|\leq 1$ for all $|y|\leq 1$ in order to have $|p(x)|\leq\phi(1)$ it suffices to show that $\sum_{i=0}^{k}|a_{i}|\leq\phi(1)$.
+Using the property that $P_{k}(x)=2^{-k}\sum_{i=1}^{\lceil k/2\rceil}\binom{k}{i}\binom{2k-2i}{k}x^{k-2i}$, we have that
+$a_{t}$ $=\frac{2}{2t+1}\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[P_{t}(y)\mathds{1}(|y|\geq\delta)]=\frac{2}{2t+1}2^{-t}\sum_{i=1}^{\lceil t/2\rceil}\binom{t}{i}\binom{2t-2i}{k}\operatorname*{\mathbf{E}}_{y\sim\mathcal{N}(0,1)}[y^{t-2i}\mathds{1}(|y|\geq\delta)]$
+$\lesssim\frac{\exp(-\delta^{2}/2)}{2t+1}2^{-t}\sum_{i=1}^{\lceil t/2\rceil}\binom{t}{i}\binom{2t-2i}{k}\delta^{t-2i}\lesssim\frac{\exp(-\delta^{2}/2)}{2t+1}P_{t}(\delta)\;.$
+Furthermore, using the Corollary 5.4 in [DKS17], we have that $|P_{t}(\delta)|\leq O(\delta^{t})$ Hence, we have that $|a_{i}|\lesssim\frac{\exp(-\delta^{2}/2)\delta^{k}}{2i+1}$.
+Choosing $\delta\gtrsim k$, we have that $\sum_{i=0}^{k}|a_{i}|\leq\phi(1)$ and that completes the proof.
+F.4 Reducing the Number of Intervals
+The proof is very similar to the proof of [DKZ20] but with the extra condition that we bound the maximum breakpoint of the piecewise constant function.
+For completeness, we state and add the extra details of the statements.
+Proposition F.7.
+Let $k,\ell$ be positive integers with $\ell\geq k+3$ and $a,b\in\mathbb{R}$ with $b>a$.
+Let $D$ be a continuous distribution over $\mathbb{R}$ and let $\nu_{0},\ldots,\nu_{k-1}\in\mathbb{R}$.
+Fix $b>0$.
+If for any $\eta>0$ there exists an at most $\ell$-piecewise constant function $g_{\eta}:\mathbb{R}\to\{a,b\}$ with $\{b_{1}^{\eta},b_{2}^{\eta},\ldots,b_{\ell-1}^{\eta}\}$ breakpoints such that $|\operatorname*{\mathbf{E}}_{z\sim D}[g_{\eta}(z)z^{t}]-\nu_{t}|\leq\eta$ for every non-negative integer $t<k$ and $|b_{1}^{\eta}|,|b_{\ell-1}^{\eta}|\leq b$, then there exists an at most $(k+3)$-piecewise constant function $f:\mathbb{R}\to\{a,b\}$ with $\{b_{1},b_{2},\ldots,b_{k+2}\}$ breakpoints such that $\operatorname*{\mathbf{E}}_{z\sim D}[f(z)z^{t}]=\nu_{t}$, for every non-negative integer $t<k$ and $|b_{1}|,|b_{k+2}|\leq b$.
+Proof.
+Note that, we can always transform the function $g_{\eta}:\mathbb{R}\mapsto\{a,b\}$ to a $g_{\eta}^{\prime}:\mathbb{R}\mapsto\{\pm 1\}$ that satisfies similar properties.
+We define $g_{\eta}^{\prime}(z)\coloneqq(2g_{\eta}(z)-a-b)/(b-a)$ and let $\nu_{t}^{\prime}=2\nu_{t}/(b-a)+(a+b)/(b-a)\operatorname*{\mathbf{E}}_{z\sim D}[z^{t}]$ and $\eta^{\prime}=\eta(2/(b-a))$.
+Hence, we have that for any $\eta^{\prime}>0$, there exists an at most $\ell$-piecewise constant function $g^{\prime}_{\eta^{\prime}}:\mathbb{R}\to\{\pm 1\}$ such that $|\operatorname*{\mathbf{E}}_{z\sim D}[g^{\prime}_{\eta}(z)z^{t}]-\nu_{t}^{\prime}|\leq\eta^{\prime}$ for every non-negative integer $t<k$.
+By applying Lemma F.9 and Lemma F.8, we obtain that there exists an at most $(k+1)$-piecewise constant function $f^{\prime}:\mathbb{R}\to\{\pm 1\}$ such that $\operatorname*{\mathbf{E}}_{z\sim D}[f^{\prime}(z)z^{t}]=\nu^{\prime}_{t}$ which keeps the first and last breakpoints the same as the original function, for every non-negative integer $t<k$.
+By setting $f(z)=(f^{\prime}(z)(b-a)+a+b)/2$, we complete the proof of Proposition F.7.
+∎
+Lemma F.8.
+Let $k$ be a positive integer and $b>0$.
+Let $D$ be a continuous distribution over $\mathbb{R}$ and let $\nu_{0},\ldots,\nu_{k-1}\in\mathbb{R}$.
+If for any $\eta>0$ there exists an at most $(k+3)$-piecewise constant function $g_{\eta}:\mathbb{R}\to\{\pm 1\}$ with $\{b_{1}^{\eta},b_{2}^{\eta},\ldots,b_{k+2}^{\eta}\}$ breakpoints such that $|\operatorname*{\mathbf{E}}_{z\sim D}[g_{\eta}(z)z^{t}]-\nu_{t}|\leq\eta$ and $|b_{1}^{\eta}|,|b_{k+2}^{\eta}|\leq b$, for every non-negative integer $t<k$, then there exists an at most $(k+3)$-piecewise constant function $f:\mathbb{R}\to\{\pm 1\}$ with $\{b_{1},b_{2},\ldots,b_{k+2}\}$ breakpoints such that $\operatorname*{\mathbf{E}}_{z\sim D}[f(z)z^{t}]=\nu_{t}$, for every non-negative integer $t<k$ and $|b_{1}|,|b_{k+2}|\leq b$.
+Lemma F.8 follows from the above using a compactness argument.
+Proof.
+[…]
+The following lemma is analogous to Lemma 3.8 in [DKZ20]. We include its proof for completeness, as our work considers more general distributions and requires specifying exact moment values.
+Lemma F.9.
+Let $m$ and $k$ be positive integers such that $m>k+3$ and $\eta>0$.
+Let $D$ be a continuous distribution over $\mathbb{R}$ and let $\nu_{0},\ldots,\nu_{k-1}\in\mathbb{R}$.
+If there exists an $m$-piecewise constant $f:\mathbb{R}\mapsto\{\pm 1\}$ such that $|\operatorname*{\mathbf{E}}_{z\sim D}[f(z)z^{t}]-\nu_{t}|<\eta$ for all non-negative integers $t<k$, then there exists an at most $(m-1)$-piecewise constant $g:\mathbb{R}\mapsto\{\pm 1\}$ such that $|\operatorname*{\mathbf{E}}_{z\sim D}[g(z)z^{t}]-\nu_{t}|<\eta$ for all non-negative integers $t<k$ and if $\{b_{1},b_{2},\ldots,b_{m-1}\}$ are the breakpoints of $f$ and $\{b_{1}^{\prime},b_{2}^{\prime},\ldots,b_{m-2}^{\prime}\}$ are the breakpoints of $g$, it holds that $b_{1}^{\prime}=b_{1}$ and $b_{m-2}^{\prime}=b_{m-1}$.
+Proof.
+[…]
+$\frac{\partial}{\partial z_{j}}M_{i}(\mathbf{z})=2a_{j-1}z_{j}^{i}p(z_{j})\quad\text{and}\quad\frac{\partial}{\partial z_{j}}\mathbf{M}(\mathbf{z})=2a_{j-1}p(z_{j})[1,z_{j}^{1},\ldots,z_{j}^{k-1}]^{T}\;.$
+We now argue that for any $\mathbf{z}$ with distinct coordinates that there exists a vector $\mathbf{u}\in\mathbb{R}^{m-1}$ such that $\mathbf{u}=(0,\mathbf{u}_{2},\ldots,\mathbf{u}_{k+1},0,0,\ldots,1,0)$ and the directional derivative of $\mathbf{M}$ in the $\mathbf{u}$ direction is zero.
+To prove this, we construct a system of linear equations such that $\nabla_{\mathbf{u}}M_{i}(\mathbf{z})=0$, for all $0<i<k$.
+Indeed, we have $\sum_{j=2}^{k}\frac{\partial}{\partial z_{j}}M_{i}(\mathbf{z})\mathbf{u}_{j}=-\frac{\partial}{\partial z_{m-2}}M_{i}(\mathbf{z})$ or $\sum_{j=2}^{k}a_{j-1}z_{j}^{i}p(z_{j})\mathbf{u}_{j}=-a_{m-3}z_{m-2}^{i}p(z_{m-2})$, which is linear in the variables $\mathbf{u}_{j}$.
+Let $\widehat{\mathbf{u}}$ be the vector with the first $k$ variables and let $\mathbf{w}$ be the vector of the right hand side of the system, i.e., $\mathbf{w}_{i}=-a_{m-3}z_{m-2}^{i}p(z_{m-2})$.
+Then this system can be written in matrix form as $\mathbf{V}\mathbf{D}\widehat{\mathbf{u}}=\mathbf{w}$, where $\mathbf{V}$ is the Vandermonde matrix, i.e., the matrix that is $\mathbf{V}_{i,j}=\alpha_{i}^{j-1}$, for some values $\alpha_{i}$ and $\mathbf{D}$ is a diagonal matrix.
+In our case, $\mathbf{V}_{i,j}=z_{i}^{j-1}$ and $\mathbf{D}_{j,j}=2a_{j-1}p(z_{j})$.
+It is known that the Vandermonde matrix has full rank iff for all $i\neq j$ we have $\alpha_{i}\neq\alpha_{j}$, which holds in our setting.
+Thus, the matrix $\mathbf{V}\mathbf{D}$ is nonsingular and there exists a solution to the equation.
+Thus, there exists a vector $\mathbf{u}$ with our desired properties and, moreover, any vector in this direction is a solution of this system of linear equations.
+Note that the vector $\mathbf{u}$ depends on the value of $\mathbf{z}$, thus we consider $\mathbf{u}(\mathbf{z})$ be the (continuous) function that returns a vector $\mathbf{u}$ given $\mathbf{z}$.
+[…]
+$\frac{\mathrm{d}}{\mathrm{d}T}\mathbf{M}(\mathbf{v}(T))=\frac{\mathrm{d}}{\mathrm{d}\mathbf{v}(T)}\mathbf{M}(\mathbf{v}(T))\frac{\mathrm{d}}{\mathrm{d}T}\mathbf{v}(T)=\frac{\mathrm{d}}{\mathrm{d}\mathbf{v}(T)}\mathbf{M}(\mathbf{v}(T))\mathbf{u}(\mathbf{v}(T))=\mathbf{0}\;,$
+where we used the chain rule and that the directional derivative in $\mathbf{u}(\mathbf{v}(T))$ direction is zero.
+This means that the function $\mathbf{M}(\mathbf{v}(t))$ is constant, and for all $0\leq j<k$, we have $|M_{j}-\nu_{j}|<\eta$, because we have that $|\operatorname*{\mathbf{E}}_{z\sim D}[F(z_{1},\ldots,z_{m-1},z)z^{t}]-\nu_{t}|<\eta$.
+Furthermore, since $\mathbf{u}(\mathbf{v}(T))$ is continuous in $\mathbf{v}(T)$, this differential equation will be well founded and have a solution up until the point where either two of the $z_{i}$ approach each other or one of the $z_{i}$ approaches plus or minus infinity (the solution cannot oscillate, since $\mathbf{v}_{m-2}^{\prime}(T)=1$ for all $T$).
+Running the differential equation until we reach such a limit, we find a limiting value $\mathbf{v}^{\ast}$ of $\mathbf{v}(T)$ so that: There is an $i$ such that $\mathbf{v}_{i}^{\ast}=\mathbf{v}_{i+1}^{\ast}$, which gives us a function that is at most $(m-2)$-piecewise constant, i.e., taking $F(\mathbf{v}^{\ast},z)$.
+Which is guaranteed to exist since $\mathbf{v}_{m-2}^{\prime}(T)=1$ for all $T$, thus there exists $T^{\prime}$ such that $\mathbf{v}_{m-2}^{\ast}=\mathbf{v}_{m-1}^{\ast}$.
+Thus, we have a function with at most $m-1$ breakpoints and the same moments.
+Furthermore, because $\mathbf{u}_{0}=\mathbf{u}_{m}=0$ the differential equation keeps the first and the last breakpoint constant we have that the first and last breakpoints of the new function are the same as in the original function.
+This completes the proof.
+∎